A Complete Bibliography of Publications in Synthese, 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.14

Title word cross-reference

5 [779]. _{e} [1050, 1160]. A [650, 461]. ℵ_{0} [1070]. B [650, 461]. p [607, 1224]. S5 [9]. × [431]. Sein [721]. Seinsollen [721].

-chains [1160, 1050]. -theoretic [461]. -year-olds [779].

2nd [1729].

36th [104]. 37th [243].

4- [779].

A-symmetric [1852]. A. [476]. Abduction [727, 925, 230]. abilities [1841]. ability [702, 699, 1358]. Aboutness [1746]. absence [1450, 649]. absence-based [649]. absences [692]. absent [1930]. absentminded [1093]. Abstract [631, 2023, 206]. abstraction [1820]. abstractionism [228]. abuses [1990, 1997]. Academic [1546]. Acceptance [451, 121, 1722, 1996]. access [203]. Accessibility [1440]. Accommodation [2032]. accord [1752]. Account [2085, 111, 1477, 347, 1320, 594, 1937, 1436, 1749, 861, 195, 1407, 1382, 1514, 1837, 1446, 1418, 342, 982, 806, 1107, 1842, 866, 1742, 1646, 1811, 944, 1965, 2068, 117, 1655, 1549, 1087]. accounting [120]. accounts [711, 522, 22, 689, 1266, 1556, 1639, 1736, 1293, 25]. accuracy [1822]. achievement [1109]. achievements [2044]. Achilles [1086]. Acknowledgements [174]. acquainted [1092]. acquire [1363]. acquisition [947, 1823]. across [1600, 1591, 1143]. Act [1899, 1786, 1740, 229]. Acting [708, 1730]. Action [1818, 1730, 1264, 1323, 1424, 172, 708, 1713, 2082, 1923, 825, 1961, 1839]. action-prediction [1961]. actions [1937, 516, 1263, 1006, 1149, 1186, 254]. Active [1699, 1091]. Acts [1895, 413]. Actual [82, 1615, 1950, 1512, 1004]. actualism [65]. actuality [603]. actually [306, 335]. ad [1066, 932]. additivity [782, 1033, 1018]. address [570]. addresses [99]. adequacy [1098, 658, 898, 929]. adequate [851, 1340]. Adjectival [1513]. adjoint [1056]. adjointness [1796]. adjunction [390]. adults [1362]. Advances [36]. advantage [1853]. adverbial [1986, 2075]. Advice [1751, 587]. advisor [1973]. aesthetic [177, 862]. aesthetics [1457, 862]. affairs [422, 487]. affect [2]. affection [1802]. afford [655]. Affordances [1544, 2074, 1818]. afraid [889, 1768]. after [902, 1931]. again [1894, 271]. Against [1268, 1464, 493, 1395, 1972, 808, 81, 817, 1270, 1088, 1907, 1985, 1651, 1687, 2079, 1522, 867, 1639, 2027, 2010, 1901, 1713, 2082, 1415, 933, 589, 1486]. age [1455, 1363]. Agency [1466, 1461, 1012, 1786, 823, 1840, 1108, 611, 2059, 1676]. Agent [1784, 1480, 1238, 17, 99, 1595]. agent-based [17]. Agent-causal [1784]. agent-causation [1480]. Agents [1833, 1239, 1038, 1369, 1958]. agglomeration [78]. Aggregating [673, 1958]. aggregation [412, 411, 2020, 931, 1745]. aggregators [1971]. AGM [651]. Agnostic [981]. agnosticism [439]. agreed [707]. Agreeing [754]. agreement [583, 1601, 922, 547]. Agrippa [573]. Aharonov [1922, 1796]. aid [130, 1717]. aim [528, 1234, 807, 1737]. aiming [141]. Aimless [794]. air [213]. akrasia [1672]. alethic [1410, 1649]. alethically [910]. algebras [1046, 400, 1054]. Alice [1585]. aliens [2004]. alike [1376]. alleged [37]. alpha [998]. also [1819]. Alternative [1836, 1372]. alternatives [2012, 2015, 2014, 2013]. always [329]. am [154]. Amalgamating [1977, 639]. amalgamation [1981, 1975, 1974]. Ambiguity [3, 1210]. Ambiguous [261, 1960]. Amsterdam [189]. Analog [1187]. analogical [1158]. analogue [757]. analogy [1467]. analyses [1661, 308]. Analysis [669, 1392, 685, 1248, 871, 939, 1554, 112, 688, 969, 1936, 805, 491, 1505, 1659, 1548, 1976, 1153, 1383, 1834, 1812, 853, 140, 1884, 1559, 1698, 433, 1215, 1599]. analytic [1334, 1148, 73, 59]. analytic-synthetic [59]. analyticity [1723]. anaphora [1000]. Anatomy [1892]. ancestral [1955]. Anchoring [1647]. angle [475]. angry [1505, 1296]. Angsty [1531]. animal [1982]. animals [184, 751]. annotation [278]. Announcement [1966]. annual [104, 243]. anorexia [1818]. Anselm [1084]. answer [1032, 1500]. ante [1678]. antecedent [363]. anthropic [1852]. anthropology [180]. Anti [1719, 2016, 1859, 1273, 773, 1672, 553, 987, 783, 1999, 74]. anti-akrasia [1672]. anti-intellectualism [1273]. Anti-luck [1719, 773]. anti-mathematicism [1999]. anti-psychologism [783]. anti-realism [2016, 1799, 605, 74]. anti-realist [1859, 553]. anti-representational [987]. anticipated [578]. anticipating [1691]. antinomy [1066]. Antirealism [30]. antirealist [689]. anxiety [1281]. anything [1761, 1210]. apologetics [1065]. aposteriori [1082]. apparent [1223]. applicability [1494, 236, 1565]. applications [392]. applied [943, 134, 569, 863, 1994, 631, 580]. apply [1914]. Applying [1617]. appraisal [634]. apprehension [1940]. approach [1952, 81, 864, 1615, 2038, 2007, 440, 658, 1694, 1492, 531, 1148, 1821, 992, 1152, 967, 842, 1535, 592, 946, 304, 1816, 1668, 384, 894, 560, 906, 1652]. approaches [1313, 2067, 841, 602, 1832, 1601, 589]. appropriate [1635]. appropriation [1605]. approximation [856, 1280, 898]. approximations [1912]. apriori [1082]. apriori/aposteriori [1082]. apriority [236, 117]. apt [708]. Arabic [371]. archaeological [1594]. Arguing [1533]. Argument [640, 1310, 716, 124, 590, 498, 1589, 130, 426, 1270, 302, 1568, 1765, 1985, 636, 377, 1084, 1172, 1659, 1791, 340, 1227, 1800, 1390, 218, 809, 1009, 993, 579, 1639, 1173, 1414, 917, 1802, 1291, 1192, 678, 2027, 2010, 1901, 1713, 2082, 1345, 234, 1680, 1278, 775, 2015, 1511, 1415, 933, 1222, 132, 1844, 1174]. argumentation [89, 740, 696, 24, 1227, 1009, 631]. arguments [718, 1530, 505, 1993, 978, 1888, 867, 1384, 1591, 1171, 675, 748, 589]. Aristotelian [1455, 1063]. Aristotle [58, 1226, 1832, 1301, 1020]. Arithmetic [1807, 520, 1881, 1190, 1951, 946]. Armchair [722, 203, 931]. arrangement [1014]. arrow [1015, 40, 827]. arrows [1571]. Art [176, 1406]. artefacts [574]. artefactualness [348]. Arthur [1315, 477]. article [2071]. articulacy [1235]. artifacts [978]. artifactual [758]. arts [177]. ascription [39, 517]. asked [136]. aspects [289]. assertibility [1603]. Assertion [1401, 798, 1801, 448, 705, 1279, 1220, 301, 1621, 648, 2049, 973]. Assertions [1149, 1011]. assertoric [493, 1020]. assessment [1140, 1391]. associations [1322]. assumptions [936, 1140, 561, 452]. asymmetric [2053]. asymmetries [1703]. Asymmetry [1612, 1365, 1836, 1575]. asymptotic [1910]. atheism [165]. ATLAS [1344]. atomism [977]. attempt [2047]. Attention [1324, 744, 514, 1523]. attentional [1965]. attitude [1305, 835, 1363]. attitudes [985, 1261, 1899]. Attitudinal [1630, 1465]. attributing [1370]. attribution [1269, 1749]. attributions [1260, 23, 255, 660]. Aufbau [224, 1025]. authoritative [923]. authorship [740]. automated [1971]. Automatically [1539]. Autonomy [345, 1575, 27, 1667]. Autopoiesis [346, 1695, 1693]. Autopsy [1344]. auxiliary [125]. aversion [3]. avoid [885]. Awareness [559, 1144, 1324]. away [691, 1223, 221]. axiom [908]. axiomatic [623, 940, 1045, 290, 1148, 294, 1601]. axiomatics [877, 293]. axiomatizability [1059]. Axiomatizations [1951]. Axiomatizing [1055]. axioms [404, 235, 385, 381].

B [712, 190]. B-theory [712]. back [271, 454]. backtrackers [1571]. Bad [1986, 1260, 678, 902, 1278, 2075]. badly [1023]. ball [227]. Banach [914]. banker [320]. Barcan [1304, 1307]. Bargaining [1182]. base [1765, 1390]. based [918, 939, 1185, 17, 1791, 1888, 2025, 2076, 1823, 1152, 1468, 583, 1535, 649, 597, 1434, 1134]. Basic [1608, 1370, 1818]. Basing [2008, 665]. Basis [1299, 1934, 923, 1963]. Battaly [473]. battles [341]. Bayes [1199, 1677]. Bayesian [1024, 1388, 112, 1968, 398, 1795, 1348, 556, 195, 2008, 651, 1227, 387, 1221, 119, 1513, 7, 595, 1654, 1521, 1884, 589, 837, 1594, 148, 1329]. Bayesianism [148]. be [426, 1169, 1021, 1388, 1270, 888, 915, 1825, 87, 321, 1597, 1859, 414, 1499, 1286, 801, 1902, 201, 1964, 1100, 1070, 1532, 896, 694, 1429, 306, 335, 1865]. beable [1101]. Bealer [27]. bearer [44]. beauties [1088]. beauty [896, 69, 253, 593]. Because [952]. begging [1425]. Begriffsschrift [380, 61]. behalf [1237]. behavior [1536, 2]. behavioral [842]. behaviour [1476]. behind [391]. Being [153, 304]. Belief [39, 1246, 121, 1234, 1551, 517, 464, 1631, 1451, 864, 1269, 718, 683, 856, 1801, 749, 1281, 556, 1987, 1405, 1988, 1408, 557, 779, 1466, 404, 24, 1432, 750, 1512, 866, 1776, 36, 857, 674, 1664, 1590, 693, 935, 1471, 1650, 862, 1463, 542, 2020, 694, 804, 1184, 332, 1141, 1467, 984, 469, 642, 1640, 1060]. belief-attribution [1269]. beliefs [1496, 1145, 2053, 1261, 938, 1147, 741, 118, 1960, 1434]. believe [427, 189, 2071, 699, 1752, 19, 504, 822]. Believing [1904, 838, 1462, 1269, 2011, 501]. Bell [1285]. belong [358]. Benacerraf [598]. benchmark [966]. Berkeley [1307]. Berlin [241]. best [1163, 695, 1554, 230, 1904, 365, 566, 1164, 1278, 1655]. beta [999]. Beth [189, 192, 188, 187]. bets [1496]. better [696, 587, 136, 504]. betting [557]. Between [445, 1260, 230, 421, 1065, 846, 1694, 1238, 530, 1148, 1586, 157, 1724, 194, 178, 747, 1040, 1603, 869, 633, 455]. Betwixt [1694, 547]. beware [536]. Beyond [1821, 70, 1927, 1116, 1907, 758, 1881, 127, 1980, 1109, 1793, 28, 1044]. beyond-Turing [1044]. Bhaskar [45]. bias [1874, 1987, 170]. Biased [1944, 2009]. biases [2, 562]. bifurcated [1963]. Big [399, 1374, 2013]. biocentrism [1380]. biological [724, 594, 359, 1661, 129, 1583, 1384, 1381, 1560, 1781]. biology [358, 356, 1759, 530, 1661, 1138, 2043, 1575, 184]. biomedical [139, 2035]. bird [736]. Birkhoff [293]. Bisimulation [1451]. bitter [2006]. Black [159]. blame [1260, 1681]. blind [83]. blindness [994]. blindspots [1355]. blood [1454]. bodies [1739, 1698]. bodily [1324]. body [1818, 1619, 2083, 1930, 126, 1927]. Bogen [266]. Bohm [1922, 1796]. Bohmian [1103]. Bohrification [400]. Bois [1682]. Boltzmann [1106]. Bolyai [1052]. Bolzanian [292]. Bolzano [60, 1967, 59]. book [1088, 2065]. books [253]. Boolean [1185]. Bootstrap [1692, 508, 210]. Bootstrapping [497, 733, 524]. border [753, 344]. Borel [1458, 989]. boson [1342, 1343, 1352]. boundaries [344, 1013]. bounded [39, 1287, 99]. bounds [928, 96]. Bourbaki [940]. box [1240, 408]. Bradley [88]. brain [1691, 511, 1202, 1962, 1212]. brains [1795, 1202, 1688, 1698]. Branching [459, 729, 478, 457, 461, 458, 424, 1058, 460]. Brandom [1256, 1637, 1828]. breaking [1527]. bridge [1178, 1956]. Bridging [1148]. Briggs [689]. bringing [619]. broad [1476]. Brogaard [2085]. brutes [1828]. Building [974, 530, 331, 2051]. bundle [2028, 777]. bundles [1249]. Burge [1213]. Buridan [488]. bush [736]. buyer [536]. bystander [860].

Caenorhabditis [724]. Calabrese [140]. calculi [1744, 43]. calculus [1566, 1015]. Calibrated [571]. calibration [1345]. Californian [1318]. calling [860]. camouflage [1406]. Can [426, 1388, 1270, 1145, 888, 2066, 1622, 1210, 1223, 1100, 1799, 303, 694, 953, 208, 1907, 189, 915, 1038, 414, 823, 1933, 1070, 853, 655, 254, 1698]. Canberra [1797]. cancellation [1896, 1897]. cancer [1981]. cannot [801]. Canon [1415]. canonical [742]. Can’t [157, 122, 2084, 1476]. Cantorian [974]. capacity [32]. capital [886]. capture [1000]. Cardano [947]. Carnap [1333, 1683, 223, 224, 877, 1331, 1341, 1332, 1330, 1337, 1025]. Carnapian [1338, 1340, 1335]. carriers [1996]. Cartesian [1789, 1793, 1788, 1794]. Cartwright [45]. case [509, 977, 947, 1584, 1651, 1140, 843, 1522, 1933, 1153, 773, 1346, 213, 1326, 2037, 634, 322, 2049, 370, 140, 64, 210, 1347, 533, 1368, 532, 1653, 783, 1915]. case-study [532]. cases [930, 1417]. Cassirer [180, 177, 179, 185, 175, 181, 182, 178, 184]. catastrophe [1973]. categorematic [1064]. categorial [627]. categorical [1483, 1371, 1302, 452]. Categoricalism [791]. categorisation [1182]. Category [600, 1326]. Category-theoretic [600]. cathedral [2014]. Cats [1806]. Causal [949, 1569, 568, 1199, 1969, 942, 1645, 509, 1589, 540, 1593, 37, 550, 816, 360, 1195, 1837, 1571, 1840, 1579, 1888, 832, 1096, 1202, 1138, 409, 1642, 2003, 1703, 842, 1200, 592, 1444, 1784, 1277, 1612, 1197, 1198, 408, 1413, 1351, 282]. Causality [256, 354, 1197, 1198, 440, 965, 1107, 1285]. Causation [880, 1426, 1480, 2067, 540, 1615, 357, 1443, 1747, 1571, 82, 568, 1888, 246, 1441, 993, 1229, 1633, 1641, 1812, 1128, 1241, 1442, 284, 656, 1205, 1193]. Cause [928, 661, 336, 2066, 2026]. caused [306, 335]. Causes [48, 1157, 367]. Cavendish [1994, 1995]. cells [1206, 1188]. centenary [189]. central [1670]. centuries [379]. certain [1050, 1160]. certainty [1551, 2049]. Ceteris [1107, 1657, 33]. chains [1050, 1160]. challenge [1952, 1365, 1716, 1666, 1655, 987, 1903]. challenges [130, 1080, 135]. Challenging [1655]. chambers [864, 862]. Chan [642]. chance [1941, 387, 286, 599]. Chandler [330]. change [120, 683, 216, 569, 651, 911, 870, 857, 841, 198, 1905, 109, 143, 117, 1478]. changed [190]. changing [1860]. character [1940, 1686]. characteristica [62, 61]. characterization [887]. Characterizing [168]. charge [1587]. charity [1267]. Charles [997]. cheap [869]. Chemical [977, 347, 1645, 1564]. Cheshire [1806]. childhood [1716]. children [1363, 750, 752]. Chimps [1239]. China [73]. chip [1672]. choice [97, 1868, 1145, 431, 564, 844, 827, 10, 542, 1156, 1196]. choose [1726]. Choosing [225, 633, 406]. chronic [1725]. Chunk [891]. Church [653, 656]. circular [89]. Circularity [677, 472, 733, 823, 508, 1639, 1656, 471, 546]. circumstances [329, 494]. claims [111, 260, 1143, 659]. Clarity [255, 51]. Clark [1193, 1205]. clashing [1188]. classes [553, 1430]. Classical [531, 60, 56, 398, 544, 1262, 1529, 1356, 1826, 48, 532, 59]. classification [1050, 1160]. classifications [783]. classify [2060]. climate [1973, 1140, 1635, 1585, 841]. clinical [1874, 2039]. cloak [1832]. Close [2085, 1297, 269]. closer [296]. closure [466, 297, 2006, 936, 784, 1791, 1550, 867, 1023, 389, 521, 655, 467]. Clouds [1454]. Co [359, 740]. co-authorship [740]. Co-emergences [359]. coaster [545]. code [1115]. coders [1696]. coding [1175, 1816]. Coffa [22]. Cogency [1722, 1659]. Cognition [1536, 1741, 1534, 1287, 2074, 1531, 1506, 1989, 726, 1694, 1818, 1699, 1361, 739, 1370, 1688, 1816, 183, 1532, 1676, 262, 1252]. Cognitive [2, 1506, 1895, 132, 1851, 1561, 1214, 1276, 1531, 702, 1795, 955, 1819, 835, 957, 1774, 1821, 1138, 1210, 579, 2044, 299, 1535, 1340, 833, 1547, 747, 589, 1669, 1329, 229, 1989]. cognitivism [1693]. Coherence [1524, 922, 417, 646, 1791, 769, 1483, 1842, 1268, 1391, 1585, 566, 1526, 1845, 1099, 929, 1293, 432, 848]. Coherent [10]. coherentism [787, 769]. coherentist [894]. coincidences [1784]. coins [1093]. collaborative [935]. collapse [721, 1075]. Collective [2061, 413, 1263, 734, 854, 737]. Collectively [1933]. collectives [858]. Collider [1346, 1345, 1823]. Color [52, 1417]. comb [1416]. Combining [1911, 1583]. comfort [130]. command [955]. commensurability [1983]. commentary [318, 533]. Comments [1164]. commit [154]. commitment [1952, 24, 1176, 712, 844, 1631]. commitments [1954, 2025, 2076, 1071, 1373]. Committing [1176]. Commodious [1403]. Common [1404, 661, 336, 587, 1288, 1871, 908, 1182, 2026, 304, 76]. communication [1602, 614, 959, 1982, 648, 837, 102]. communities [1851, 137, 738]. community [219, 1555, 1556, 857]. commutative [1566, 521]. company [1986, 2075]. Comparative [523, 799]. Comparing [1601, 651]. comparison [1238, 841, 43]. comparisons [914, 793]. compatibilism [829]. compatibility [602, 1147]. compatible [1170]. compete [341]. competent [1546]. Competing [1293, 522, 1288]. complement [1044]. complete [1283, 1634]. Completely [1006]. completeness [1648, 1888, 1212, 642, 1230]. completeness-based [1888]. Complex [356, 947, 160, 527, 1838, 422, 1580, 345, 184]. Complexity [1054, 1661]. components [1272]. composite [2011]. Composition [48, 1761, 1605, 1607, 772]. compositional [1272, 493]. Compositionality [964, 904, 52]. comprehension [552]. compulsion [720]. computation [256, 354, 160, 806, 1738, 1068]. Computational [1321, 1693, 742, 1043, 805, 355, 310, 797, 1328, 1110, 1667]. computations [742]. computer [918, 690, 340, 271, 590, 1044]. computer-based [918]. computer-simulated [271]. computers [537]. Conceivability [1702, 1530, 1794, 1158, 1619, 2083]. Concept [1590, 120, 177, 1821, 1629, 1067, 1823, 879, 182, 1463, 1197, 1198, 907, 526]. Conception [1001, 1978, 17, 1827, 1081, 632, 1068, 1963, 291]. conceptions [1069]. concepts [290, 947, 1339, 1066, 1272, 523, 1361, 1152, 1739, 1327, 612]. Conceptual [1248, 1153, 669, 140, 569, 1522, 272, 344, 727, 630]. Conceptualism [79]. conceptualization [1126]. concern [2016]. concerning [1835]. Concessive [660, 23, 255]. Conciliationism [1275, 958]. conclusion [451]. Concrete [113, 1808, 1738]. condensed [530]. condition [2011, 584, 282]. Conditional [1156, 1451, 925, 336, 1686, 1548, 1458, 564, 392, 938, 453]. conditionalisation [782, 148]. conditionalization [285, 693]. Conditionals [550, 675, 1407, 113, 563, 1232]. Conditioning [1458, 1201, 341]. conditions [1704, 1107]. Conducive [1411, 786, 1925, 432]. Confabulation [1929]. confidence [1664, 1035]. configuration [1100]. Confirmation [318, 195, 935, 314, 1884, 1883, 977, 522, 1852, 321, 1245, 119, 322, 1654, 316, 315, 596]. confirmers [872]. confirming [319]. conflicts [344]. conformism [1713, 2082]. conformity [20]. confusing [322]. Confusion [23, 333, 336]. conglomerability [1632]. conjecture [1503]. conjunction [323, 320, 670, 1842, 628, 316, 317, 315]. conjunctive [683]. Connection [1817, 1586]. conscious [847, 469]. Consciousness [1815, 1677, 744, 1223, 747]. consensus [583]. consequence [339, 1800, 444, 602, 448, 450, 68, 25, 453, 454, 912]. consequence-accounts [25]. consequences [934, 1635, 1609, 547]. consequentialism [1459]. conservatism [900, 651, 1717, 2013, 1712]. considerations [1561]. considered [427, 1037, 1329]. consistency [1618, 1433, 1017]. conspiracism [1585]. Conspiracy [1934]. constants [364, 454]. Constituting [1621]. constitution [770]. constitutive [311, 442]. constrained [510]. Constraining [1997]. constraint [1099]. constraints [929]. construal [1637]. Constructing [843, 492, 2045, 1634]. construction [1561, 1998, 511, 2032, 1258, 254]. Constructive [373, 984, 641, 1929, 1917, 301, 229]. Constructivism [1854, 1074]. constructivist [2038, 818]. Constructor [732]. Consuming [1143]. contemporary [1793, 2035, 1855, 117]. Content [1690, 1440, 1151, 1476, 1613, 944, 327, 1822, 288, 1367, 602, 1174, 1523, 493, 1437, 1197, 1198, 423, 1663, 1960, 1220, 1134]. contents [1370, 1115, 646]. Context [912, 237, 714, 1801, 134, 1150, 52, 1006, 637, 1225, 298, 894]. context-sensitivity [298]. context-shifting [1150]. contexts [668, 959, 329, 1143]. Contextual [1371, 121, 969, 50, 204, 1918, 627, 657]. Contextualism [470, 1792, 915, 609, 1375, 972, 637]. contextualist [46]. contingencies [1831]. contingents [1313]. Continuity [1905, 1106, 1695, 1003, 1872]. Continuum [90, 2050]. contra [1637]. contraction [38, 684]. contradiction [1674]. contrapiction [1256]. contrast [1417]. Contrastive [522, 692, 246, 1293]. contribute [136, 1799]. contributions [1205, 135]. contributors [1004, 103]. control [934, 1819, 873, 1468, 1465, 1672, 945, 464, 878, 1469]. controversial [47]. controversy [746]. convention [1680, 547]. Conventionalism [1017]. Conventions [672]. convergence [194]. convexity [393, 1522]. conviction [420]. Convolution [1215]. convolutional [1820]. cooperation [407]. coordination [235, 560]. Copenhagen [698]. copula [1298]. core [990, 563, 1434]. Cornerstones [504]. correct [1145, 1224]. Correction [1989, 2054, 2080, 2084, 2079, 2076, 2083, 2075, 2078, 2082, 1947, 1948]. corrective [1575, 1073]. correctness [1400]. Correlated [1184]. Correlated-belief [1184]. Correlations [661, 949, 2003]. correspondence [194, 214, 241]. Corroboration [125, 591]. cosmological [2002]. cost [1007]. cotenability [551]. could [1825, 1597, 1902, 306, 335]. couldn’ta [1039]. count [1454, 1154]. countably [785]. counter [37]. counter-example [37]. counterexamples [434]. Counterfactual [363, 1657, 1703, 113, 799]. counterfactuals [548, 197, 491, 1569, 845]. counterintuitive [1459]. Counterlegal [1571]. counterlegals [1571]. counterpart [1247]. Counterpossibles [2085, 549]. covering [1174]. cracks [529]. Craft [1300]. crashes [1780]. created [1070]. Creating [978]. creationists [158]. creative [293]. creativity [1215]. credal [851, 535]. Credence [91, 1518, 1246, 286, 1022, 1640]. credences [1433]. credit [262]. Criteria [941, 1678, 663, 1349, 2077, 70]. criterion [530, 1202, 1472, 1701]. Critical [849, 971, 1780, 1834, 1812, 1559, 1215]. criticism [182]. Critique [2085, 966, 1683, 2040, 440, 1108, 1621, 185, 1529, 988, 452]. cross [707, 618]. cross-disciplinary [618]. cross-purposes [707]. crossing [58]. Crowdsourced [1611]. cubical [233]. Cultural [1326, 181, 741]. culture [185]. current [1708]. customs [341]. cut [698]. cycles [816]. cylindric [1046].

damaging [472]. dare [1131]. Darwin [93, 1879]. Darwinian [83, 344]. Darwinism [346, 962]. Data [272, 269, 274, 1919, 1043, 1346, 1823, 263, 270, 268, 266, 271, 532]. data-phenomena [266]. Dauer [724]. David [1604]. Davos [178]. Davos-debate [178]. De-idealization [533]. dead [1766]. deal [1038, 1887]. dealing [1661]. debate [1761, 1795, 1492, 1831, 1816, 1799, 1855, 178, 869, 1594, 605]. debates [136, 631]. debiasing [863]. debunking [1956]. decade [130]. deception [116]. Deceptive [410, 1936]. decide [766]. deciding [766]. Decision [949, 416, 8, 1589, 97, 37, 550, 1088, 403, 1038, 1772, 1201, 409, 2003, 1703, 764, 408, 1528, 1141, 2039]. decision-making [2039]. decisions [415, 64]. Declarations [1520]. declaratives [1031]. Decoherence [698]. decoherent [1026]. decomposability [1542]. Dedekind [293]. dedisciplined [617]. Deduction [1916, 447, 2080]. Deductive [1722, 389, 655]. deeds [406]. deep [1820, 616]. Default [701, 1846]. default-free [1846]. defeasibility [467]. Defeasible [199, 108]. defeasibly [251]. defeat [1710, 1719, 1712]. defeater [598]. Defeaters [1709, 1708, 1711, 872]. Defeating [1715, 1718, 287]. defects [1073]. defence [829, 1682, 937, 200, 116, 484, 902, 612, 1486, 498]. defended [1814, 972, 1453]. Defending [81, 1367, 1394]. defense [916, 961, 1862, 667, 226, 1276, 1714, 1976, 98, 1099, 899, 1363, 1681, 417, 418, 2068, 1779, 274]. deference [1576, 1848]. definability [1043, 1916, 2080, 188]. define [1439]. defined [1046]. defining [1534, 1615, 1759, 1982, 1283]. definitely [1648]. definition [58, 5, 1285, 1394, 351, 344, 1222]. definitions [1759, 928, 350, 2036, 1700]. deflated [1943]. Deflationary [1142, 1749]. Deflationism [44, 1881, 1943, 682]. Degree [671, 317]. Degrees [1512, 1451, 347, 51, 557, 694]. Deliberation [579, 1636, 1703, 581]. delta [891]. Demarcation [157, 530, 364]. Dembski [160]. Democracy [582, 580]. democratic [1682, 1555, 579]. democratisation [1556]. demon [1832, 1791]. demonstratives [1724]. Denial [1903, 1220]. denialism [1848]. density [1104]. denying [1423]. deontic [721, 1006, 1727, 433, 433]. deontological [632]. deontologism [1907]. Deontology [1814, 878]. depend [1157]. dependence [615, 1121, 1527, 1571, 796, 333, 1854, 737]. dephlogisticated [213]. derivation [811]. Derrida [179]. Descartes [1993, 1992, 367]. describe [1819]. description [957, 310]. descriptive [952, 1464]. Design [161, 164, 609, 1383, 162, 163, 574]. Desirability [1232, 938]. Desire [1072, 1630, 2000]. Desires [938]. destabilized [1956]. Destructive [1712]. detach [122, 2084]. detailed [433]. detection [116]. detectivism [923]. detector [1344]. determinable [303]. determination [1840]. Determinism [355, 457, 1832, 282]. Deterministic [286, 930, 1593, 633]. deterrence [101]. Developing [825]. Development [760, 1980, 1809, 1684, 296]. developmental [749, 1363]. Deviant [616]. deviations [661]. diachronic [1723, 1640]. diagnosis [746, 1930]. diagnostic [1931, 1351]. diagram [370]. Diagrammatic [380, 1996, 374, 376, 998, 384, 375]. Diagrams [383, 369, 379, 2019, 371, 382, 373, 372, 1004, 385, 381]. dialectic [1009, 349, 1712]. dialectical [89, 695]. dialectics [769]. dialogue [1870, 479, 283]. Dicisigns [1002]. Diderot [1999]. difference [568, 24, 179, 975, 157, 1724]. difference-making [568]. differences [1762]. Different [1555, 362]. differential [379]. differently [179]. difficult [1609]. difficulty [491]. diffusiveness [1452]. digital [757, 806]. dilemma [1906, 66, 1038, 1221, 1241, 921]. dilemmas [1702, 895]. Dilemmic [2021]. dimension [734]. dimensional [608, 569, 1777, 853, 433]. Dimensionalism [1552]. dimensionality [968, 1764]. Dirac [891]. direct [228, 1468, 1367]. directed [1778]. direction [901]. directions [105, 1721, 1855]. dis [1853]. disagree [754, 1958]. disagreeing [707, 1414]. Disagreement [439, 958, 1541, 590, 1275, 875, 855, 856, 955, 647, 578, 340, 571, 2053, 1414, 1713, 2082, 546, 2020]. disagreements [2064]. Discarded [1860]. disciplinary [614, 618, 1617]. discipline [277]. disconfirmed [1568]. discontents [161]. Discontinuities [1919]. discourse [776, 1000, 588, 384].

Discovering [685]. Discovery [1286, 198, 237, 1593, 1941, 1349, 2077, 1348, 1342, 1343, 1049, 1809, 1684, 1085, 293, 1351]. discovery/justification [237]. discrete [1086]. discretion [1219]. Discursive [507]. discussion [484, 1834, 1004, 747, 408]. disease [2036]. diseases [1725]. disguise [385, 381]. disgust [1824]. disjunctivism [941, 1963, 1077]. disjunctivist [1591]. disorders [1933]. Dispensability [1166, 1165]. Display [43]. dispositionalism [791]. dispositionalist [2067, 1412]. dispositions [1103, 1841, 822]. Disproportional [284]. dissent [1635]. dissenters [874]. dissolution [767]. Dissolving [706]. distance [1065]. distinct [1600]. distinction [237, 1525, 230, 1080, 1454, 1637, 421, 1082, 1666, 1154, 826, 1284, 266, 1358, 59]. distinctions [1475]. distinctiveness [810]. distinctness [916, 742, 127]. Distinguishing [1263, 583]. distortions [1706]. Distributed [1482, 1882, 739, 1209]. distributional [909]. distributions [393]. Diversifying [1583]. diversity [579, 1547]. divide [1951]. divided [235]. division [341, 1995]. DIVORCE [1751]. DNA [1809, 1684]. Do [1459, 1496, 1260, 341, 2003, 1646, 1762, 1559, 590, 1376, 1911, 635, 1750, 1157, 1696, 1023, 1797, 830, 268, 1914, 254, 1652]. doctoral [240]. Dodd [255]. Does [802, 87, 447, 1245, 1673, 1154, 1964, 1501, 288, 699, 1058, 573, 704, 365, 334]. doesn’t [297, 786, 644]. dogma [1363, 1395]. dogmatism [1125, 119, 1521, 1712]. dogmatist [2041, 810]. Doing [1269, 756, 1786]. Domain [326, 1537, 1609, 627]. Domain-sensitivity [326]. dome [1912, 664]. donkey [1000]. don’t [2071]. Doomsday [716]. Doppelt [1489]. double [359]. doubt [1664, 471]. down [1767, 201]. Downward [357, 993]. Doxastic [1630, 1567, 1461, 1463, 865, 829, 855, 1907, 1904, 1468, 1400, 506, 1464, 699, 1681, 1373, 1737, 878, 1469]. drawing [366]. dream [1698]. dreaming [1698]. Dreams [747, 1698]. Drift [1116, 1427]. Driftability [710]. driver [1093]. dropped [227]. drugs [2034]. ‘DSD’ [2037]. dual [749, 8]. dual-system [749]. Dualism [897, 1927, 1662]. dualities [541]. duality [2072]. Dubins [1587]. Due [1848]. Duhem [435, 125]. Duhem-Quine [435]. Duhemian [1488]. Dummett [1255]. dust [974]. Dutch [1088]. duties [427, 719]. dynamic [193, 401, 277, 1030, 754, 1029]. Dynamical [1445, 957, 248, 353]. dynamicism [987]. Dynamics [189, 771, 1948, 1977, 796, 569, 1146, 1144, 1775, 11, 1962, 36, 857, 1811, 1905, 1055, 860, 854].

e-research [1611]. E-type [1000]. E. [192]. each [1004]. Early [1005, 1990, 1190, 807, 877, 1331, 1332, 1991]. Earman [424]. earn [303]. earning [137]. earth [438]. ease [779]. easy [1244, 704, 1921, 1077, 1092]. echo [864, 862]. Echoes [1793]. ecological [1691, 2074, 1382, 1601]. ecology [1277]. economic [1240]. economics [861, 1973, 555, 634, 727]. economy [1094]. econophysics [1780]. ecosystem [1382]. ecosystemic [1647]. ecumenical [1417]. Editor [72]. Editorial [585, 996, 212, 463]. Editors [854]. education [1871]. Edward [140]. effability [953]. effect [1922, 1326, 641, 860, 1796, 1654]. Effective [431, 24]. effectiveness [1878]. effectivity [1833]. effects [1833, 1323, 1385, 20]. efficacy [1559]. efficient [1816]. effort [1966]. EG [998]. EG-alpha [998]. Egan [1038, 1038]. egocentric [1987, 170]. egocentricity [1309]. eighteenth [379]. Einstein [707, 1052]. either [757, 761]. elegans [724]. Elementary [48]. elements [1911, 461, 1054]. elimination [1928]. eliminativism [1274]. eliminativist [1810]. else [500]. Elusive [63]. embedded [1032]. Embedding [619]. embeddings [1056]. Embodied [1354, 1676, 1693, 1688, 1697, 2078]. embrace [1622]. embracing [1927]. Emergence [353, 337, 352, 356, 256, 354, 1604, 1043, 360, 1527, 355]. emergences [359]. Emergent [358, 360, 1444]. emergentism [359]. emerging [1999]. emotion [84, 895, 1697, 2078]. Emotional [2058, 1505]. emotions [1544, 991, 1367]. empathy [84, 1364]. Empirical [659, 1946, 111, 1119, 1480, 81, 918, 1098, 658, 690, 569, 5, 1813, 2002, 149, 898, 1341, 268, 1099, 1799, 1617, 1197, 1198, 747, 1493, 1011, 217, 402]. empirically [946, 1075]. Empiricism [145, 1820, 144, 1345, 154, 1867]. empiricists [707]. Empress [1994]. empty [1761]. Enactive [1667, 1691, 1688]. enactivism [1699]. encroachment [470, 711, 1264, 1360, 819, 1718, 1402]. enculturated [1694]. encyclopedia [277]. encyclopedic [281]. end [1937]. ended [355]. ends [216]. energy [1691, 1349, 2077, 1695, 1347]. engage [874]. Engagement [134, 2074, 143]. Engel [1400, 1399, 1402]. engineering [1911, 849, 971, 1329]. engines [1914]. Enhanced [1168, 1167, 1844]. enhancing [1382]. enough [1435, 1257, 1279]. enrichment [328]. entail [1501, 334, 699]. Entailment [795, 1018]. entails [654]. entangled [1697, 2078]. entanglement [1200]. Enthymematic [169, 718]. entirely [1804]. entities [441, 421, 258]. entitlement [503]. Entity [1857, 1192]. entropy [803]. environment [594]. epiphenomenalism [847]. episodic [1208, 743, 1211]. episodic-like [1211]. Epistemic [615, 528, 1525, 465, 1686, 936, 915, 584, 1238, 919, 520, 1483, 669, 809, 1461, 1557, 604, 1485, 972, 643, 429, 1497, 994, 68, 521, 894, 1752, 655, 1755, 1037, 467, 1754, 737, 470, 1459, 711, 1596, 472, 1748, 1753, 733, 120, 2007, 726, 817, 1401, 1710, 146, 1030, 260, 1907, 937, 955, 516, 647, 753, 754, 440, 1124, 608, 137, 1743, 1029, 1775, 508, 1742, 1556, 2036, 579, 570, 2037, 641, 1149, 632, 2010, 1471, 581, 171, 449, 1402, 1719, 902, 1186, 1073, 1728, 2035, 1798, 865, 2056, 317, 2048, 1756, 719, 114]. epistemic [2068, 1867, 580, 582, 469, 905, 596]. epistemically [1264, 168]. Epistemicism [986]. Epistemological [1077, 265, 939, 356, 706, 1683, 219, 941, 805, 1047, 63, 722, 1591, 989, 1963]. Epistemology [515, 555, 2021, 1413, 577, 802, 1638, 1435, 818, 556, 435, 164, 812, 1789, 610, 585, 773, 1616, 1176, 781, 200, 863, 571, 1679, 1183, 858, 116, 1708, 946, 649, 239, 764, 899, 576, 1719, 1793, 1788, 862, 1794, 1757, 1830, 791, 589, 1611, 905, 596, 229]. EPR [949, 2003]. equal [2027]. equational [1054, 1053]. equations [379]. equilibria [416]. equilibrium [559, 1184].

Equivalence [2047, 1883, 217]. Equivalent [1168]. Ernst [177, 181, 178]. erotetic [1028]. Erratum [305, 455, 1161, 442, 2077, 815, 204, 1095, 1096, 840, 620, 1160, 927, 911, 307, 1396, 1397, 1809, 1198, 385, 306, 443]. Error [1118, 805, 1731, 394, 1962, 1236, 680, 1005, 1984]. essay [82]. essence [1455, 1803, 1369, 2070]. essential [1796]. essentialism [1455]. established [1848]. Eternalism [498, 109]. Ethical [1298, 1235]. Ethics [1973, 1299, 1378, 141]. ethnobiology [1558]. ethotic [677]. etiological [1384]. Etiology [1650, 1281]. Euclid [1052, 86, 372]. Euclidean [426, 371, 1957]. evaluate [548, 1439]. Evaluating [1288, 739, 1952, 1362, 1971, 1211]. evaluation [329, 70, 1060]. Event [178, 1837, 107]. event-causal [1837]. Events [1253, 1845]. Everettian [460]. Evert [189, 187]. every [321]. Evidence [1981, 1975, 1273, 203, 1167, 1168, 1141, 624, 336, 2007, 1968, 1977, 439, 1934, 1574, 87, 1199, 2018, 659, 1752, 513, 719, 639, 1011, 1974, 170, 1204, 633, 1971, 931]. evidential [1088, 1038, 47, 593, 118]. Evidentialism [505, 1712]. evil [480, 479, 1791]. Evolution [165, 1455, 1605, 360, 1382, 1124, 22, 355, 279, 1378, 345]. Evolutionary [11, 1956, 160, 869, 741]. Exact [104, 243]. exactly [267]. exam [1481]. example [37, 1535]. example-based [1535]. examples [371, 1835]. excellence [711]. except [1378]. exceptions [1575]. exchange [1944]. Excluded [1250]. excuse [1525]. excuses [1783]. executable [1006]. executive [1670]. exercising [1786]. exhibit [262]. Exhibiting [800, 1008]. Existence [728, 106, 788, 1735]. existential [1954, 997]. Existentialism [2004]. existentialist [1492]. existentials [1502]. existing [1314]. exists [1761]. Exotic [1610]. Expanding [2067]. expect [1481]. expectations [828, 256, 354, 1458, 661, 1652]. expected [1512]. Experience [232, 241, 1866, 1978, 1078, 1997, 607, 415, 1079, 1367, 1733, 1787, 1040, 1734, 254]. experiences [287, 2056]. experientialism [810]. Experiment [331, 587, 1798]. Experimental [537, 1604, 1769, 609, 585, 1346]. experimentation [1823, 1969]. Experimenter [1345]. experiments [1817, 818, 690, 1361, 1774, 944, 1055, 1085, 2015, 296, 438]. expertise [774, 976, 582, 1235, 138]. Experts [735]. explain [208, 1382, 1223, 853, 1058, 588]. Explaining [1737, 1848, 1416, 906]. Explanation [968, 543, 1780, 310, 1168, 945, 509, 1163, 756, 871, 540, 695, 230, 742, 1439, 1807, 536, 1188, 1582, 365, 1209, 1441, 1138, 341, 2018, 1639, 692, 566, 1226, 1167, 1162, 671, 1959, 1164, 595, 1278, 1668, 1444, 1612, 1110, 631, 1796, 126, 2042, 283, 312, 1938, 309, 906, 1087, 1623]. explanationism [1479, 1388]. Explanationist [1717]. Explanations [1575, 1289, 1584, 1445, 1825, 1819, 129, 1583, 1170, 1266, 1577, 2015, 1328, 1168, 1350, 625, 1915, 425, 1599]. explanatorily [1444]. Explanatory [1212, 1126, 2051, 820, 783, 1419, 3, 1165, 1920, 1136, 1166, 1130, 957, 575, 1639, 566, 128, 257, 1197, 1198, 1258]. explication [1340, 1946]. explicit [679]. exploration [1968, 1346]. exploratory [1346]. Exploring [2019, 753, 999]. explosion [1574]. exponential [731]. expounding [219]. expression [1505]. expressions [1142]. Expressive [603, 1777]. expressivism [2040, 809, 1562]. expressivity [1451]. Extended [726, 1129, 1161, 661, 1412, 1538, 1482, 2058, 1501, 1676, 262, 1252]. Extending [2015, 2069]. extension [1538, 132]. extensional [629]. extensionality [1949]. extent [1914]. External [1133, 725, 1972, 233]. Externalism [1016, 1151, 288, 1091, 423]. externalist [289]. extra [1007]. Extrapolation [1979]. Eye [923].

F. [296]. fable [485, 486]. face [578, 1713, 2082, 18]. faces [1157, 1147, 1194]. facsimile [1875]. fact [1761, 1177]. factive [1141]. Factivity [1618, 831]. facts [1119, 1890, 1317, 1444]. factualism [1191]. factualist [667]. factually [1761]. fail [638, 447]. failed [1652]. fails [962, 390, 1259]. failure [500, 330, 595, 298]. failures [1968]. Fair [518, 821, 785]. faire [1354]. fairer [1499]. fairness [1243]. Faith [479, 503, 480]. faithful [528]. Faithfulness [2052, 1194]. Fallacies [314, 46, 199]. fallacious [867]. fallacy [1765, 320, 1390, 670, 1842, 316, 317, 315]. Fallibilism [790, 660]. fallibility [470]. False [863, 749, 750, 1216, 1603]. false-belief [750]. far [414, 678]. Fast [558]. Fate [948]. Faultless [955]. Faulty [1703]. favour [139]. feasible [946]. feature [1418]. feature-matching [1418]. fecundity [719]. Feel [1359]. feeling [1108, 1962, 152, 847]. Feminist [135]. Fermat [227]. Fiber [1249]. fiction [26, 758, 16, 1507, 1588, 1012]. fiction-making [1012]. fictional [14, 1012]. Fictionalism [1393, 1810]. fictions [529, 820]. fictions-for [820]. figures [261]. files [1739]. Finality [1449]. financial [533]. Financializing [2035]. Finding [529]. findings [1848]. fine [668, 2002, 494]. fine-grain [494]. fine-tuning [2002]. Finean [1191, 1803]. fineness [1894]. Finetti [1524, 1845]. Finite [782, 1911, 1566, 1033]. finitism [1067, 1063, 362]. finitude [362]. first [1309, 750, 1677, 1777, 327, 1951, 1059, 229]. first-order [1777, 1951, 1059]. first-order/second-order [1951]. first-person [229]. first-person-oriented [327]. Fisher [1116]. fit [1426]. Fitch [623, 653, 33, 656]. FL [1160, 1050]. flashback [1617]. flat [1325]. flawed [748]. flexibility [882]. flight [1562]. flow [1359, 539, 1058]. Fluency [779]. Fodor [962]. Foiling [159]. fold [184]. Folk [713, 1363]. follow [1409]. following [560]. foray [1138]. Force [1897, 1507, 1427, 1712, 1606]. forces [1372, 1277]. forecast [830]. forgotten [1396, 1336, 375, 2065]. form [177, 743, 377, 1805, 1900, 481, 178, 1220]. Formal [515, 1084, 1048, 585, 1300, 1392, 1936, 843, 1055, 764, 1283, 1634, 1012, 596, 979]. formalisms [1340]. formalization [1340, 663]. Formalizing [1265]. formation [83, 176]. forms [629, 266]. formula [1304, 1307]. Formulating [682, 1716]. foundation [1045, 1227, 1421, 1757]. foundation-oriented [1757]. foundational [1140, 292]. Foundationalism [1758, 228, 568, 1529]. foundationalist [239]. Foundations [276, 357, 403, 845, 1909, 1757, 1504, 1222]. founded [108]. Four [1080, 1552, 76, 485]. Four-Dimensionalism [1552]. fractional [1921]. fragment [479]. frame [2025, 2076, 1152, 1733]. frame-based [2025, 2076, 1152]. framework [1693, 356, 436, 846, 962, 1227, 687, 1542, 644, 225, 727, 1258, 979]. frameworks [1519, 1651]. framing [100]. Frankfurt [1835, 930]. Franssen [25]. Free [1868, 1682, 1691, 328, 1837, 812, 1840, 1864, 1838, 1695, 879, 1831, 746, 1846, 1467]. free-energy [1691]. Freedom [956, 584, 1834]. freely [812]. Frege [1735, 61, 380, 1507, 1877, 1337, 1701, 59]. Fregean [496]. frequentist [597, 1347]. friction [1256]. friendship [1904]. frontiers [1759]. frugal [558]. fruitfulness [2019]. Function [511, 891, 594, 1321, 2072, 1102, 1382, 1626, 1821, 512, 287, 1384, 1099, 1100, 960, 1097, 1737, 869]. Functional [871, 1137, 1738, 1383, 906, 347, 13, 353, 308]. Functionalism [1242, 1276, 1501, 1410, 903, 132, 1782]. functionality [1511]. functions [564, 1100, 10, 1022, 1560]. Fundamental [1099, 2064, 1601, 452]. fuss [1305]. Future [296, 1313, 1531, 1361, 462, 2061, 1543, 1237]. fuzzy [1511].

Galileo [227]. Galois [951]. Gambit [621]. game [1596, 101, 1238, 1692, 560, 1113]. game-theoretic [560]. games [102, 562, 1185, 1853, 431, 11, 752, 599]. Gamma [1744]. Gandalf [652]. Gao [2081]. gap [1148]. Garber [2018]. Garber-style [2018]. General [2074, 536, 249, 397, 80, 196, 1353, 644, 1901, 1277, 1415, 580, 1222, 1249, 1044]. general-relativistic [1044]. generalisation [1627]. Generalised [2026]. generality [967, 124]. generalization [1626]. generalizations [715, 1174]. generalized [1285, 1280, 1884]. Generalizing [898, 508]. generative [1690]. generis [1720]. genetic [1427]. genetics [1584, 136, 531, 532]. geo [392]. geo-logical [392]. Geodesic [1773]. geometrical [373]. geometry [376, 378, 1148, 1957, 372, 1877]. Gettier [468, 942, 1563]. Gettier-illusion [468]. Gettier-partialism [468]. Getting [1692, 1374, 1092]. GIRL [1598]. Girolamo [947]. Given [79, 1078, 321]. giving [1940]. Global [1887, 1562, 1385]. globality [950]. Glymour [1205, 1193]. go [1066]. Goal [1121, 120, 1678, 1778]. Goal-dependence [1121]. goal-directed [1778]. Gödel [75]. Going [1862]. gone [678, 745, 1318]. Good [1111, 888, 1257, 596]. Goodman [876]. Goodness [1310, 1378]. gossip [939]. gossip-based [939]. grades [2063]. gradualism [660]. grain [1894, 494]. grammar [668, 177, 627]. graph [434, 1744]. graph-theoretic [434]. graphical [832, 1096]. graphs [997, 1000]. gravitating [1914]. gravitational [1915]. great [1692, 1376]. Green [1195]. Griffin [165]. grip [1962]. Grothendieck [951]. ground [1955, 1613, 154]. ground-ancestral [1955]. ground-theoretic [1613]. grounded [1816]. Grounding [1747, 1619, 1229, 1173, 1174, 1283, 1634, 1967, 625, 2070, 301, 2083, 1174]. groundlessness [501]. Grounds [239, 418]. Group [969, 992, 2020, 1050, 1160, 1444, 2059, 580]. group-like [1050, 1160]. groups [615, 2060, 1933, 2055, 2056, 2057]. growth [216, 731]. grue [1195]. Grünbaum [1394]. GRW [1104, 1075]. guessing [1692]. Guest [463, 72]. Guidance [1672]. guide [1145]. guides [1115]. gullibility [116].

habit [2031]. Hacking [1192]. Hadron [1346, 1823, 1345]. haecceitism [1530]. Hall [64]. Hallucinating [924]. hallucination [1419, 168]. hand [736]. Handling [714]. Hans [238, 242, 232, 241]. Happily [1697, 2078]. Hard [1240]. harmony [1231]. harms [1850]. hath [321]. hats [4]. having [119]. hearing [1714]. Hearsay [1526]. heaven [1692, 1565]. Heidegger [178]. hell [1692]. help [1797, 288]. Henkin [1051, 1422]. hermeneutic [1678]. hermeneutics [1699]. heuristic [2072, 1473]. heuristics [558, 762, 1085, 425]. hexagon [1416]. hidden [219]. hierarchical [1690]. Hierarchies [115]. hierarchy [290, 1823, 1961]. Higgs [1348, 1342, 1343, 932, 1350, 1352, 1351]. high [1349, 2077, 1140, 1347]. high-resolution [1140]. higher [226, 2074, 1885, 2002, 1510, 1444, 1282]. higher-level [2002, 1444]. higher-order [226, 1885, 1510, 1282]. Hilbert [1148, 385, 381, 362, 1230]. Hinge [1071]. Hintikka [484, 1061, 554]. Historical [2017, 1489, 792, 1485, 628, 154, 117]. historically [1859]. histories [458]. History [1062, 2047, 616, 1646, 369, 1396, 1336, 135, 333]. hoc [932]. holds [25]. holism [769, 47, 1787, 1723]. Holistic [1418, 917, 1706]. Hope [1355]. hopefulness [624]. Horgan [593]. Horwich [1155, 1627, 1621]. hot [211]. HPS [723]. hues [1665]. Human [374, 96, 1211, 1834, 1258]. humanities [619]. Hume [928, 376, 1998, 876]. Humean [1639]. Humeans [1702]. Hungarian [1047]. hunger [920]. hunting [1111]. Husserl [1190, 1648, 75, 1230]. hybrid [1316]. hylomorphism [364]. hypergraphs [908]. Hyperintensional [983, 985, 1472, 981]. hyperspace [1806]. hypocrisy [1401]. hypodoxes [2073]. hypotheses [368, 932, 125, 532]. hypothesis [1108, 1357, 83, 319, 1493, 1993]. hypothetical [743, 452].

icons [687, 1669]. idea [2047, 1098, 614]. Ideal [1083, 1682]. Idealisations [586]. idealism [1721, 1893]. Idealization [905, 1920, 1780, 1918, 1212, 299, 1921, 1910, 533]. Idealizations [1166, 1796, 1165, 1911, 928, 1922, 1908, 885]. Idealized [1706, 826]. ideas [1834]. identicals [1843]. identifiability [1969]. identification [493]. identifying [529]. identities [1202]. Identity [434, 441, 811, 1482, 258, 1678, 1610, 2001, 777, 1448]. Ideological [709]. ideology [1952]. If [1891, 1208, 1599, 599, 190]. if’-explanations [1599]. Iffy [828]. Ignorance [1763, 1023, 859, 861, 1123, 902, 860, 1545]. ignoring [1850]. III [891, 1222]. ill [1933]. illocutionary [1886]. illusion [468, 1297, 1219]. illusions [582]. illustration [97]. I’m [1786]. image [1637]. imaginary [529]. Imagination [1993, 944, 1997, 1679, 1793, 1469]. imagining [2061]. immediate [1283, 1634, 1606]. immunity [1731]. impact [2013]. impenetrable [1734]. imperative [1222]. Imperfect [719]. imperfective [1846]. imperialism [1768]. impermissivist [1906]. implementation [1738, 797, 1044]. implication [1000, 298, 453]. implications [1600, 164, 919]. Implicature [1360, 1707]. implicatures [730]. Implicit [844, 2039, 831, 1261]. implies [1907]. imply [914]. import [858]. importance [24, 1224]. impossibility [645, 25, 639]. Impossible [2085, 645, 1988, 1092]. imprecise [1906, 1221, 896]. impressions [1367]. impulsivity [2031]. inaccessibility [2002]. inadequate [1130]. inappropriate [1635]. inbetweenness [1579]. incarnation [34]. incentivized [825]. including [433]. inclusion [582]. Inclusiveness [578, 577, 579]. incoherence [1410, 1022]. incoherent [1026, 1958]. incommensurability [1651, 759]. incommensurable [846]. Incomplete [947]. incompleteness [1393]. inconceivability [1794]. Inconsistency [676, 884, 892, 886, 889, 1483, 1302, 883, 882, 890]. Inconsistent [1013, 881, 2005, 888, 885]. incremental [321, 842]. indefeasible [941]. indefinitum} [1066]. independence [1185, 393, 1520]. independent [1845]. Indeterminacy [821, 951, 1945]. indeterminate [916, 3, 405]. Indeterminism [424, 457, 1838, 1831, 1834, 478]. indeterminist [1837]. indeterministic [633]. indexed [1523]. indexical [1150, 593, 612]. Indexicality [325]. Indexicals [298, 1036, 1724]. indication [998]. indirect [214, 633]. indiscernibility [951, 1843]. indiscernibles [777, 434]. Indispensability [640, 1844, 1170, 1162, 1168, 426, 1439, 1172, 1173, 917, 1167, 1171, 1174]. individual [1498, 1176, 375]. individualism [1379]. individualist [1865]. individuality [1392]. individuals [691, 458, 1303]. individuated [1891]. individuation [1738, 1782]. Indoctrination [1281]. induction [543, 678, 767, 731]. inductions [2017]. Inductive [959, 2018, 1825, 1973, 876, 6, 430, 597]. inductivism [238]. inequivalence [1262]. inexhaustible [154]. infallibilism [468, 1790]. Infallibilists [1751]. infallibility [247, 292]. infallible [1789]. infant [1540]. Inference [230, 365, 566, 1163, 695, 398, 1699, 388, 1677, 911, 870, 444, 391, 842, 1425, 448, 1164, 649, 449, 1278, 430, 451, 655, 597, 1118, 453, 1861, 502, 1351, 453]. inferences [2040]. Inferential [1028, 1686, 769, 995, 1397, 1953]. inferential-role [995, 1397, 1953]. inferentialist [1365, 1395]. inferentially [903, 269]. Infertility [2036]. Infinite [1920, 1908, 1438, 1918, 764, 1911, 1065, 382, 1916, 2080, 821, 785, 1758, 518]. Infinitesimal [1921, 1015]. Infinitesimals [785]. Infinitism [768, 171, 1686, 769, 2010, 573]. infinitum} [1066]. Infinity [1065, 1909, 1066, 1062, 1070, 1064]. influence [1476, 1762, 184]. inform [1245, 1698]. informal [437, 199]. Information [160, 784, 910, 2074, 688, 1592, 2054, 1127, 342, 806, 410, 1245, 1089, 539, 1739, 1233, 1964, 307, 641, 1944, 1044, 1134]. Informational [13, 1775, 1185, 1982, 860, 68]. informativeness [528]. injustice [2037, 1497]. Inner [897, 1677]. innocuous [329]. innovate [861]. innovation [2035]. inquiry [948, 936, 1675, 1029, 289, 362]. Inquisitive [1030, 1028]. Inscrutability [1474]. insight [944]. insights [344]. insolubilium} [488]. instability [409, 1764, 1352]. instance [1625]. institutional [825]. instructional [806]. instructions [1372]. instrumental [950]. instrumentalism [937, 604, 2064, 1752]. instruments [2015]. Integrating [308]. integration [1978, 614, 833]. integrative [1811]. intellectual [616, 927, 913]. Intellectualism [1269, 636, 1273]. intelligence [1770]. Intelligent [163, 164, 162]. Intensional [107]. intensionality [982, 325]. Intention [92, 2028, 1466, 1109, 611, 78]. Intention-sensitive [92]. Intentional [611, 1786, 1610, 2001]. Intentionality [804, 1449, 1781, 464]. intentionally [1462]. intentions [1656, 1595, 254]. inter [1128]. inter-level [1128]. Interaction [1252, 1185, 643, 561, 1617, 1771]. Interactions [1260]. interactive [1183]. intercontextual [912]. interdisciplinarity [616, 620, 613, 1609]. interdisciplinary [615, 614, 1609]. Interest [1554, 140, 117]. interesting [1356, 1578]. interests [361, 1860]. interface [1506, 1178, 657, 1989]. interference [557]. intermediate [736]. internal [2038, 872]. internalism [66, 865, 423]. Internalist [289, 861]. Internet [281]. interpersonal [793]. interpolation [188]. interpretation [668, 1090, 557, 1000, 1513, 1497, 1224, 597, 1984, 813]. interpretational [800]. Interpretations [1415]. Interpreting [718]. interrogative [1029]. interrogatives [1031, 1032]. intersection [1941]. Intersex [2037]. intersubjective [831]. Intersubstitutivity [1626]. Intervening [1685, 1201]. Intervention [282, 618, 860]. interventionism [823, 1564, 1120]. Interventionist [1528, 1569, 1107, 1668]. interventions [1289]. Intractability [425, 1599]. intransparency [883]. Intrinsicality [1247]. intrinsically [910]. Introduction [1179, 1, 343, 1729, 55, 1748, 156, 1869, 499, 1553, 12, 324, 1398, 1889, 1193, 2033, 734, 232, 1079, 980, 1461, 444, 36, 1641, 104, 314, 28, 1097, 95, 386, 186, 49, 290, 2012, 489, 403, 1990, 1508, 133, 1975, 1908, 1598, 480, 1688, 1577, 456, 1162, 1708, 1831, 1788, 212, 1927, 1620, 854, 1771, 1484, 443]. introductory [411]. introspection [1776]. introspectionism [923]. introspective [1292]. intuition [177, 1707, 378, 418]. intuitionism [875]. Intuitionistc [1521]. intuitionistic [191]. Intuitions [624, 1769, 666, 762, 609, 926, 1728, 215]. invariance [1333]. invariantism [1257]. Inverse [1782, 670, 167]. Investigating [1206, 1886]. Investigations [2022]. involuntarism [1907]. irreducibility [1223]. irreducible [2062]. irrelevance [1472]. Irrelevant [628]. isn’t [1128, 1631]. isomeric [1645]. Issue [314, 1179, 305, 275, 1748, 2012, 1193, 1205, 2033, 232, 1079, 1598, 1688, 1461, 263, 456, 1641, 1708, 2055, 1927, 1620, 1771, 1729]. issues [211, 1005]. Istanbul [242, 232, 241]. iterated [38]. Iterative [307, 1068]. I’ve [1991].

Jaakko [1061]. jade [1646]. Jean [488]. Jeffrey [693]. Jeffreys [2077, 1349]. ji [908]. John [941]. joint [1263, 1149, 611, 2056]. Jones [433]. Jones-Pörn [433]. Joongol [2075, 1986]. judge [1750]. judgement [862]. Judgment [412, 1745, 1477, 1514, 411]. judgments [2, 1298, 1769, 40]. Just [1750, 2071, 47, 1786, 41]. justice [584, 1854]. justification [318, 237, 768, 1748, 1753, 1525, 1686, 1789, 507, 769, 63, 670, 1980, 1387, 632, 1471, 1255, 894, 317, 1756, 1755, 289, 1415, 1092, 1606, 1159, 1754]. justification-excuse [1525]. justificational [719, 1712]. justifications [562, 1783, 1757]. justificatory [423]. Justified [502, 1147, 332]. justifies [903]. Justifying [695].

Kant [1066, 378, 22, 1735, 1905, 59, 291]. Kantian [270, 288]. keep [303]. Keeping [15]. Kepler [368]. key [1660, 1819]. Killer [1075]. Kim [1986, 1967, 1241]. kind [58, 1208, 755, 1646]. kind-crossing [58]. Kinds [1942, 1643, 1137, 247, 1153, 1122, 1642, 1644, 1645, 1641, 327, 760, 1647, 1931, 510]. kinematics [803]. KK [42]. Klein [573]. knew [1024]. Knight [159]. knots [1957]. know [1260, 635, 32, 1273, 406]. know-how [1273]. Knowability [106, 32, 815, 814, 685, 1618, 519, 30, 34, 653, 28]. Knowing [650, 1743, 254, 2044, 654, 1016, 1926, 1778, 292]. Knowing-how [1743]. Knowledge [420, 472, 730, 702, 21, 1749, 900, 1987, 1707, 1406, 1405, 1776, 1250, 621, 279, 833, 1923, 471, 473, 973, 262, 1358, 1709, 918, 704, 1260, 939, 1874, 1608, 1269, 2063, 1365, 1998, 941, 587, 23, 1355, 255, 31, 1408, 1597, 815, 814, 1424, 1570, 831, 137, 790, 990, 660, 1790, 2033, 2025, 2076, 1153, 1079, 908, 1403, 992, 867, 988, 60, 1123, 1354, 1558, 923, 708, 42, 506, 583, 1143, 482, 1091, 304, 593, 780, 35, 9, 113, 1963, 1279, 1077, 598, 332, 799, 1218, 1159, 737, 1365, 755]. knowledge-attributions [660]. knowledge-how [1269, 1570, 1354]. knowledge-infallibilism [1790]. knowledge-minimalism [1790]. known [1470]. knows [1129, 1161]. Kolmogorov [1458, 989]. Kornblith [2063]. Kralemann [2069]. Kripke [1238, 554]. Kuhn [846]. Kuhnian [723].

lab [1604]. labor [341, 299]. laden [1978]. ladenness [266]. Ladyman [1444]. landscape [459]. language [1729, 446, 1312, 1288, 1210, 175, 1331, 606, 188, 384]. language-transcendence [446]. languages [676, 1885]. Laplace [1832]. Large [1346, 1438, 1212, 1345, 1971, 1823]. large-scale [1971]. larva [724]. last [130]. lattices [293]. Lattman [2069]. law [689, 1639, 1845]. lawhood [1655]. Laws [1102, 510, 1554, 1412, 1178, 1784]. lawyer [1653]. Lay [1728, 137, 783]. layers [76]. laziness [902]. Leading [141]. learn [1210]. learned [1991]. Learning [147, 1600, 429]. least [1603]. leaving [629]. lecture [1404]. Lehrer [1653]. Leibniz [883]. lens [1526]. Leśniewski [62, 1700]. less [136]. lesson [510, 601]. Lessons [1823, 1558, 1242, 355]. letter [487]. level [1978, 2002, 1444, 1128]. level-laden [1978]. Levels [1602, 1439, 1942, 1302, 115]. Lévy [756]. Lewis [2079, 1604, 1687, 672, 1548, 11, 253, 113, 554, 1655, 789]. Lewisian [304, 1727, 553]. Lexical [1829, 1602]. Lexical-rule [1829]. LHC [1344]. liar [986]. liberal [1679, 1367]. liberal-content [1367]. libertarian [1839]. libertarianism [1784]. lies [1785, 1920]. Life [350, 347, 343, 1759, 359, 1090, 1939, 1297, 1695, 224, 351, 344, 345, 349, 1999]. life-mind [1695]. light [951, 653, 1302, 1378, 1917]. like [1050, 1160, 1211, 1787]. likeness [602]. limit [312]. limitations [1658, 84]. limited [1992]. limiting [248]. limits [558, 1815, 610, 874, 634, 239, 2064, 35, 1332, 1110, 1564, 486, 337]. Linda [319]. Lindley [2077, 1349]. line [462]. Lines [868]. lingua [61]. linguistic [636, 1893]. linguistics [1909, 54]. link [587, 1407]. Lipton [230]. lite [611]. Little [1595]. live [147]. Living [206, 356, 348]. Local [385, 381, 1862, 965, 1285, 1360, 1558, 1803]. localism [1862]. locality [1446, 950]. localized [1321]. location [285, 538, 91]. locavoracity [1113]. locus [209]. Logic [396, 544, 196, 77, 1300, 1035, 1793, 1041, 433, 1771, 1313, 397, 193, 401, 244, 1503, 925, 1593, 1042, 1316, 1185, 2007, 721, 338, 1030, 1031, 1955, 1311, 754, 776, 982, 608, 50, 204, 1238, 796, 1657, 447, 1954, 476, 400, 1616, 1286, 1777, 6, 392, 877, 1183, 108, 1302, 279, 1291, 484, 478, 1634, 483, 1826, 1036, 1299, 1330, 252, 1756, 1337, 1020, 1543, 1727, 477, 1064, 1222, 1778, 1469, 599, 1745, 1028, 905, 912, 402]. logic-based [1185]. Logical [2040, 857, 391, 1301, 1333, 645, 1334, 1891, 707, 1144, 801, 6, 392, 444, 1302, 353, 1803, 364, 475, 1834, 1900, 663, 481, 561, 1373, 1083, 1220, 395]. logically [888]. logician [483, 484]. logicians [1318]. logicism [807]. Logics [1027, 1186, 1952, 1774, 843, 1029, 1744]. look [1137, 1342]. looked [952]. Looking [1505, 991]. looks [1715]. LORI [1771]. LORI-5 [1771]. lossy [391]. Lost [1519, 1093, 873]. lot [1278]. Lotteries [1387, 821, 1438, 785, 518, 1640]. lottery [782, 392, 391]. low [750]. low-verbal [750]. Luck [361, 795, 114, 1218, 802, 1753, 726, 1436, 773, 641, 1719, 2068, 1754]. Lucky [780, 1546]. lying [771, 1948].

machine [1043]. Macro [245]. macrodata [527]. macroeconomics [1117]. made [1874, 1095, 970, 1565]. magic [1820, 1687, 2079]. main [748]. maintenance [1595]. make [975, 19, 751]. make-believe [19]. maker [1597]. makers [440]. makes [1665, 1609, 1381, 1378]. Making [1578, 679, 568, 826, 416, 1012, 2039]. Malfunction [1453]. malfunctioning [1010]. malfunctions [1658]. Mally [721]. man [1428, 1296, 1994]. Mandeville [1999]. Mandevillian [1770]. maneuver [817]. manifest [1637]. manifest-scientific [1637]. manifestationalism [961]. manifesto [1207]. manipulability [1579, 567]. Manipulation [1819, 1564]. manner [1284]. manuscript [371]. Many [441, 1026, 1952, 1503, 1209, 1157, 692, 789]. many-sorted [1503]. Mapping [2050, 1178]. Margaret [1994, 1995]. Margin [680]. mark [132]. Market [1780, 1830]. Martin [178]. Mass [53, 1018, 1454, 1154]. mass/count [1454, 1154]. match [1565]. matching [1418]. material [1808, 767, 675]. materialism [1995]. materialist [149]. Mathematical [516, 371, 629, 1163, 756, 236, 947, 1440, 714, 2023, 1998, 1997, 377, 805, 22, 1357, 1616, 1164, 1168]. mathematicism [1999]. mathematics [1494, 58, 2019, 105, 1065, 943, 1166, 1990, 1993, 1034, 1997, 1878, 2046, 1565, 629, 537, 369, 1992, 1674, 1994, 1793, 1617, 367, 375, 74, 892, 1106, 1917]. mathematization [366]. mathoids [701]. matter [1761, 1104, 1502, 786, 530]. matters [1177]. Matthew [1654, 321]. MaxCon [1412]. Maximal [684]. maximalist [921]. maximum [803]. may [975, 154]. maybe [864]. Mayr [1226]. maze [131]. McDowell [761, 941]. McGee [1155]. Me [1129, 1161]. meaning [265, 2028, 1604]. meanings [1714]. means [1937, 122, 2084]. means-end [1937]. measure [318, 670, 967, 995, 1397, 1953, 628, 417]. Measurement [54, 377, 80, 1847, 2081, 526, 1974]. measurement-precision [526]. measurements [1344]. measures [1483, 1268, 1391, 935, 929]. Measuring [1022]. mechanical [1356]. mechanics [1098, 778, 1566, 681, 1371, 1585, 1914, 214, 48, 460, 700, 979]. mechanism [1090, 932, 1209, 308, 1216, 1578]. Mechanisms [311, 1582, 128, 1217, 1938, 313, 816, 792, 440, 1583, 1811, 1128, 309, 442]. Mechanistic [419, 1584, 957, 1577, 312, 509, 1289, 871, 540, 1819, 962, 1209, 1212, 1738, 1110, 1938]. mediation [179]. Medical [2033, 1584, 440]. medicine [2038, 2029, 1874, 1584]. medieval [450, 477]. medium [102]. meet [1582, 269]. meeting [104, 243, 630]. meets [1238, 1029, 585, 69, 1873, 1563, 1118, 1028]. Meinongianism [2005, 1290]. Melia [85]. memoriam [1061]. memories [1093]. memory [1208, 743, 1815, 1482, 1253, 1211, 641, 1216, 1929, 131]. Meno [2029]. mensura} [1270]. Mental [1739, 1812, 1747, 1864, 1229, 2061, 745, 288, 1241, 284]. mentalistic [1369, 2081]. Mere [503, 957, 583]. merely [1275]. Mereological [1014, 777, 1231, 422, 1448, 808]. mereology [2005, 553]. merging [856, 1234, 1773]. message [102]. Meta [1710, 695, 1791, 1976, 1516, 991, 1712]. meta-analysis [1976]. meta-coherence [1791]. meta-emotions [991]. Meta-epistemic [1710]. meta-evidentialism [1712]. meta-syllogism [695]. meta-vagueness [1516]. metabolic [1434]. metalogic [1700]. metaphilosophy [280]. metaphors [762]. Metaphysical [220, 1662, 1392, 288, 1959]. Metaphysics [1243, 1643, 1683, 1443, 1993, 2060, 1687, 223, 127, 1562, 150, 1797, 1641, 1456, 1386, 144, 2079]. metarepresentational [2001]. Metasemantics [1656]. metatheory [1332]. meter [1945]. method [290, 1687, 2079, 294, 1284]. Methodological [895, 1114, 1862, 978, 1295, 1970, 911, 841, 124, 703, 1865]. Methodology [1120, 1112, 722, 1118, 971]. methods [410, 2015]. metrical [622]. Micro [245]. microphysical [1888]. middle [1955, 1092]. Might [2053, 136, 975, 1429]. Might-beliefs [2053]. millennia [56]. millennia-old [56]. Millian [139]. Mills [1249]. Milvian [1956]. Mind [923, 1693, 1729, 1538, 1202, 1619, 2083, 1695, 2058, 1697, 2078, 1501, 1178, 606, 252, 126, 1927, 1368, 1961, 751]. mind-body [1619, 2083, 126, 1927]. mind-brain [1202]. minded [1539]. Mindedness [1411, 1925]. minding [1370]. Mindless [1822]. Mindreading [1362, 779, 774, 1370, 1540, 1762, 1549]. minds [1370, 630, 1376]. mine [1114]. Minimal [742, 1912, 347, 410, 1183, 345, 1796]. Minimalism [1627, 1620, 328, 1790, 1622, 1624, 1623]. minimalist [1628]. minimality [282]. minimally [2009]. Minkowski [385, 381]. Minority [1853]. miracle [218, 1286]. Miracles [124, 1765, 197, 1390]. Mirror [513]. mirrored [84].

Mirroring [512, 245]. misidentification [1731]. misleading [1733]. misnomer [94]. misrepresentation [1690, 1115]. Missing [18, 1407, 1343, 1932]. missing-link [1407]. mistake [1673]. misunderstandings [2046]. mixed [1181]. Modal [65, 1158, 520, 552, 67, 1436, 1311, 1290, 447, 1777, 1679, 1384, 1744, 1114, 424, 43, 1794, 113, 813, 700, 1215]. Modal-Epistemic [520]. modalities [33]. modality [1494, 21, 2002, 1949]. modals [1775, 68]. mode [723]. Model [911, 870, 60, 56, 2067, 864, 1185, 1566, 80, 530, 127, 1029, 1823, 1266, 1513, 545, 877, 2032, 1824, 612, 2051, 1796, 597, 1884, 580, 727, 59, 1134]. model-based [1823, 597, 1134]. model-building [530]. model-explanations [1266]. Modeling [697, 207, 1240, 1130, 1418, 805, 2069, 687, 1118]. Modelling [816, 388, 741, 272, 19]. Models [16, 889, 2069, 1138, 687, 1280, 209, 532, 309, 399, 724, 717, 208, 742, 586, 14, 1166, 527, 1188, 1571, 15, 1238, 957, 1774, 17, 651, 1209, 1823, 1690, 1210, 967, 299, 1265, 206, 1706, 2015, 210, 533, 2045, 19, 633]. Moderately [1456]. modern [1990, 1020]. moderns [1991]. modes [1502, 143]. modest [961, 98, 1746]. modified [1394]. Modulation [1372]. Molecular [26, 724]. Monday [69]. Monism [778, 1630, 1531, 1886]. monistic [589]. Montagovian [686]. Monty [64]. mood [603]. Moore [41, 2071, 725, 469, 642]. Moore-paradoxical [469]. Moorean [1591]. Moorean-style [1591]. Moral [875, 2, 926, 1824, 1799]. morality [2038]. Morgan [1415]. Morris [131]. most [1291, 1603]. Motion [1802, 1057, 366]. motivated [1859]. motivation [1295, 809]. motivations [1555]. motor [1323, 1354, 510]. motoric [1040]. moved [153]. movement [1698]. moves [457, 829, 599]. Moving [1150, 127]. much [1024]. MUDdy [1389]. Muller [296]. multi [1238]. multi-agent [1238]. multidisciplinary [1978, 2039]. Multiple [38, 1460, 657, 1983, 1180, 130, 647, 451]. multiple-conclusion [451]. multiplicity [898]. multistable [1210]. Multitasking [498]. Multitude [446]. Multiverse [1069, 339]. music [2043]. Musschenbroek [1997]. Must [1532, 306, 335, 74]. mutable [1237]. Mutual [1579, 587]. my [923]. myopia [1821]. Myth [1529, 79, 182].

N [476]. Nachlass [1305]. Nagel [1884]. Naïve [1924, 2052]. name [90]. names [1586, 1829, 775]. Narratives [792, 1253]. Narrow [1076]. Narrow-scoping [1076]. nativism [1795]. Natural [1642, 1089, 1073, 676, 1643, 1208, 1312, 1879, 1153, 1122, 755, 1639, 1646, 1644, 1641, 760, 384, 1378]. naturalism [558, 1721, 722, 1378]. naturalist [418]. Naturalistic [1856, 1456, 2009]. Naturalized [280, 414]. Naturalizing [1172]. naturalness [1352]. nature [740, 1869, 1102, 1123, 1870, 57, 1664, 1822, 1995, 1217, 1850, 259, 1106, 366]. navigate [1119]. navigation [278]. Neat [1056]. necessarily [1032]. Necessary [35, 404, 1887, 1875, 1386]. Necessitarian [495]. necessitating [1833]. necessitation [490]. Necessity [1072, 811, 22, 1803, 1662]. need [1023, 1646, 1964, 1115, 1087]. negation [1298]. Negative [540, 1463, 1095, 970, 1746, 931]. Negotiating [344]. neither [404, 286]. Nell [1595]. neo [1414]. neo-Pyrrhonian [1414]. Nested [1226]. nets [1199, 1957, 1197, 1198]. Network [1661, 1580, 1584, 342, 1210, 837]. networks [1180, 1820, 1642, 648]. Neural [1327, 1206, 1820, 1321, 742, 1821, 1210, 1178, 1784, 1215]. neurobiological [1928]. neurobiology [1817, 1207]. neurons [513]. neuroprosthetics [1812]. neuropsychological [1666]. neuroscience [1214, 1321, 742, 310, 1210, 1325, 1811, 1326, 231, 308, 510, 1669]. neuroscientific [770, 746]. neutral [910]. neutrality [869]. never [707]. Newcomb [336, 1563, 99, 652, 601]. Newton [377, 883, 1996]. Newtonian [457, 1833, 214]. Niche [1258, 1561]. niches [1851]. Nieuwentijt [1997]. nihilism [1014, 808, 1259]. nineteenth [379]. No [1888, 258, 1109, 124, 1974, 441, 1761, 864, 2041, 285, 1765, 1007, 1390, 1009, 119, 1843]. No-Miracles [124]. nodes [1642]. Noesis [281]. nomic [683, 1280]. Nominalism [1858, 123, 1921]. nominalist [105]. Nominalistic [1174]. Non [667, 934, 715, 1808, 1713, 1459, 1392, 1804, 1632, 1819, 1566, 1446, 1046, 164, 957, 1138, 1191, 1949, 1009, 1360, 1485, 899, 529, 747, 921, 1282, 1413, 2082]. non-argumentation [1009]. non-causal [1138, 1413]. non-cognitive [747]. non-commutative [1566]. Non-concrete [1808]. non-conformism [1713, 2082]. non-conglomerability [1632]. non-epistemic [1485]. non-epistemology [164]. non-extensionality [1949]. non-factualism [1191]. Non-factualist [667]. non-fictions [529]. non-individuality [1392]. non-local [1360]. non-locality [1446]. non-maximalist [921]. non-mechanistic [1819, 957]. non-philosophers [1459]. non-probabilist [1282]. Non-psychological [934]. Non-rational [1713, 2082]. non-reductionism [899]. non-self-referential [1804]. non-standard [1046]. Non-zero [715]. nonassociative [521]. Nonconceptual [1940]. nonconceptualism [933]. nonmonotonic [1745]. nor [404, 286]. norm [2049, 1737, 973]. normality [1776]. Normative [696, 763, 1767, 1021, 586, 2011, 1227, 1629, 950, 1635, 1673, 547, 1087]. normatively [851]. Normativity [1767, 1516, 1525, 937, 1124, 672, 1544, 1378, 171, 1728, 1337]. Norms [648, 679, 101, 696, 1408, 1743, 1254, 1433, 738, 952, 2035, 1723, 1141]. Norton [1912, 664]. note [740, 1422]. Notes [721]. Nothing [1378, 1786, 1787, 761]. Nothingness [2005]. notion [265, 1065, 846, 2043, 424, 1283, 1634, 114, 1386, 1649, 979]. notions [951, 1646, 1302]. nouns [53, 76]. novelty [1962, 2043]. number [1019, 1572, 626, 979]. numbers [947, 572, 1986, 1559, 2075].

Object [2065]. objection [1480, 1418, 784, 962, 1986, 2075, 1521, 1511]. objections [1489, 1479]. Objective [148, 5, 1755, 1560]. Objectivity [841, 1025, 1306, 1553, 1555, 1858, 1556, 1558, 1559]. objects [714, 14, 2023, 1827, 1808, 1460, 2030, 1474, 1610, 1306]. objectual [985, 575]. obligation [484, 483]. obligations [874, 484, 483]. obscurantism [1240]. observation [1057]. observational [319]. observations [2015]. Ockhamism [868]. oddness [1407]. off [329]. offer [573, 74]. offs [1658, 530]. Okasha [330]. Old [1023, 2017, 300, 211, 712, 2018, 196, 1204, 56]. olds [779]. omission [1588]. Omissions [1652]. omitting [1786]. omniscience [645, 1373, 1083]. One [1101, 1209, 534, 321, 1016, 154, 907, 408]. one-box [408]. Online [1851]. only [358, 907]. ontic [600, 1446]. ontogeny [1327]. Ontological [2025, 1805, 278, 421, 1084, 843, 1117, 1176, 712, 2076]. ontologies [1099]. Ontology [1112, 977, 277, 1121, 1660, 1412, 1104, 1446, 15, 276, 538, 1644, 1142, 1115, 1118, 849, 2057, 19, 1999, 1120, 700]. Open [1411, 355, 1925]. open-ended [355]. Open-Mindedness [1411, 1925]. operational [982, 1759]. Operator [498, 400]. operators [1777]. opinion [24]. opportunistic [1021]. opportunities [141]. opportunity [1833]. optical [1996]. optics [368, 367]. Optimal [1844]. Optimality [1757]. Optimisation [756]. Optimistic [1490]. Optimus [626]. options [1254]. Optogenetics [1216]. order [290, 226, 1185, 843, 1954, 750, 1885, 1777, 57, 535, 1951, 1510, 1282, 1059]. order/second [1951]. Ordering [1385]. ordinal [1883]. ordinary [1848]. organisation [1377, 248, 1381]. organisms [1766]. organization [356]. oriented [327, 1757]. originalism [1590]. other [566, 1326, 1004, 43, 1058]. other-race [1326]. otherwise [699]. ought [1750, 122, 2084, 1907]. our [953, 96, 501, 497]. Outer [897]. Outline [80]. overall [1022]. overconfidence [837]. overdetermination [1747]. overlap [1268]. overlooked [1256]. overturned [783]. overview [1162]. own [1270].

P [886]. Pain [1274, 1866, 1978]. pannormism [1767]. papers [997]. Paraconsistency [1826, 2007]. Paraconsistent [302, 1146, 1313, 717, 1291]. paradigm [1295, 1733, 1611]. paradigms [1361, 151, 852]. Paradox [94, 106, 1458, 623, 1086, 244, 1596, 2029, 954, 1804, 41, 1349, 2077, 749, 1632, 2071, 1835, 782, 1587, 90, 34, 167, 392, 653, 1340, 481, 33, 989, 1373, 1913, 1944, 1846, 591, 1779, 642, 601, 1452, 1917, 41, 914]. Paradoxes [1180, 544, 1127, 391, 958]. Paradoxical [2073, 469]. parameters [527]. parametric [210]. paraphrase [1259]. paraphrasing [626]. paribus [1657, 1107, 33]. parity [132]. Parsimony [1163, 1164, 709, 169]. Parsing [824]. Part [1995, 166]. Part-whole [166]. Partial [1313, 1185, 658, 1988, 11, 1488, 214, 418, 804, 1960, 599]. Partial-order [1185]. partialism [468]. Partiality [840, 839, 1904]. partially [1006]. particle [1348, 1100, 1057, 1915, 979]. Particles [1443]. partitions [9]. partner [319]. partnership [1388]. Parts [553, 2066, 1150, 1808]. Pascal [1065]. past [788, 2061]. patchwork [1821]. path [1527, 1854]. path-dependence [1527, 1854]. patterns [268]. Peacocke [71]. peculiar [1653]. pedagogical [1566]. Peer [647, 856, 2013]. peer-review [2013]. peerhood [958]. peers [955]. Peirce [230, 1256, 2069, 998, 1000, 687, 1003, 1004, 997]. Peircean [1001]. penetration [1506, 1989]. people [1260, 1848]. perceptible [1267]. perception [706, 1506, 1989, 1940, 1406, 1210, 1367, 1213, 795, 680, 1368, 1924]. Perceptual [607, 1523, 612, 1965, 1078, 1505, 506, 1367, 1689, 1963, 1606]. perceptually [1600]. Pereboom [1480]. perils [527]. perimaximal [684]. permeate [891]. permissibility [1923]. permissive [1531, 1576]. Permissiveness [465, 1567]. permit [704]. Perseverance [927, 913]. Persistence [960, 1382, 1766]. person [1309, 1677, 327, 1367, 229]. personal [1482, 269]. perspectival [826]. perspectivalism [1157]. perspective [1691, 428, 843, 1540, 1212, 142, 561, 163, 2020, 837, 229]. perspectives [352, 852, 1348, 932, 1871, 395]. pessimistic [678, 731]. Peter [765]. Petersburg [765]. Petitio [89]. phase [1920, 1917]. Phenomena [263, 724, 1919, 358, 264, 273, 267, 1780, 1132, 272, 270, 266, 271, 1258, 274]. phenomenal [1717, 109, 1606, 254]. Phenomenalist [810]. phenomenological [1502, 1616]. phenomenology [1190, 75, 152]. phenomenon [265, 2059]. phi [1248]. philosopher [192]. Philosophers [2034, 1459, 105, 136, 157, 619]. Philosophical [352, 762, 932, 618, 1674, 183, 1085, 343, 2016, 852, 1342, 616, 75, 1834, 1225, 278, 976]. philosophically [1075]. Philosophy [617, 620, 613, 104, 243, 606, 1931, 1927, 305, 275, 1729, 277, 1207, 45, 1990, 134, 133, 211, 1193, 1205, 978, 1871, 276, 1769, 585, 73, 1210, 175, 1062, 196, 444, 57, 369, 764, 2009, 1793, 182, 135, 728, 27, 1877, 184, 931, 141, 138, 906, 74]. phlogiston [213, 214]. Phylogenetic [1278]. physical [396, 1494, 568, 545, 1099, 1100, 1265]. Physicalism [1827, 1493, 1242, 1888, 1447]. Physico [367]. Physico-mathematics [367]. Physics [890, 1349, 192, 1446, 273, 530, 1878, 1565, 1426, 176, 1347, 666, 395, 886, 2077]. picked [227]. picture [1583, 1374, 1809, 1684]. Pictures [1323, 86, 1225, 1674]. Picturing [398]. piece [1343]. piecemeal [1969]. piggyback [81]. pill [2006]. Pitfalls [97]. place [802, 1206, 1654]. placebos [1394]. Plain [1410]. plan [1797]. plane [372]. planets [1202]. planning [865]. plant [1207]. plasticity [1327]. Platitudes [1034]. Platonic [438]. Platonists [1169]. Plausibilistic [848]. plausibility [1981]. plausible [1837]. play [587]. players [1774]. playing [752]. plea [1375, 2009, 1839]. Plural [813]. pluralism [1531, 801, 1557, 929, 1649, 1927, 1779]. pluralist [70, 1760]. Pluralistic [860, 859, 861, 1742, 1302]. plus [1791]. PMI [1486]. Poincaré [1457, 1106]. Point [1915, 1371]. Point-particle [1915]. pointless [1759]. polarization [863, 862]. policies [852]. policy [1555, 164, 1973, 141]. political [2, 584, 582]. Pollock [1711]. polylithic [215]. polysemy [1622]. pooling [11]. Population [123, 1853, 531]. Pörn [433]. pose [2003]. position [470]. positive [302, 675]. positivism [1302]. possibilities [1836, 459, 1831]. Possibility [2043, 470, 1702, 1509, 1775, 1984, 486]. Possible [1950, 31, 1300, 1275, 491, 222, 729, 113, 25, 1603]. possible-world-accounts [25]. possibly [1314]. post [325]. post-semantics [325]. posteriori} [1080]. postulates [1280]. potential [815, 814, 1005]. power [1419, 3, 1833, 1136, 1130, 603, 957, 1777, 1946, 1452]. powers [1849, 622, 1840, 1645, 1449, 1782]. Practical [259, 1709, 695, 836]. practice [616, 1609, 1420, 1347, 18]. practices [1149, 619, 703]. Pragmatic [679, 711, 1264, 1347, 470, 41, 1257, 1492, 1360, 819, 1718, 2027, 1402, 118, 642]. pragmatics [900, 819, 657]. pragmatist [1621]. pre [926]. pre-theoreticality [926]. preachers [485]. precision [526]. predestination [476]. predicament [1839]. predicate [775]. Predicates [428, 1950, 1537, 1903]. predication [1150, 1900]. predicativism [1829]. predictability [355]. Prediction [232, 1353, 241, 1697, 2078, 2032, 1961]. predictions [828, 2015]. Predictive [1810, 1175, 1688, 1488, 1689, 1698, 1795, 1694, 1692, 1962, 1696, 1946]. preemptive [866]. Preface [1294, 205, 474, 173, 996, 525, 391, 1373]. prefer [1726]. Preference [651, 950, 1265, 1430, 77]. prejudice [840, 839]. prejudices [1502]. preliminary [1933]. prelude [476]. Prenective [1663]. presence [1965]. present [487]. Presentism [909, 109, 1985, 1421]. presentist [625]. preservation [995, 1953, 960, 1397]. presumptions [917]. presuppositions [1423, 1019]. primary [44]. prime [626]. Primitive [1104, 1446]. primitivism [1889]. Principia [377]. principii [89]. Principle [661, 81, 667, 1691, 1270, 323, 519, 1048, 803, 1843, 42, 1267, 48, 1282, 78, 434]. principled [1615]. principles [1626]. print [371]. Prior [1304, 1312, 1314, 556, 1315, 1316, 1311, 1305, 476, 1295, 1310, 1307, 1302, 484, 1834, 1308, 1301, 1303, 477, 488, 486]. Priorean [1308]. priori [237, 240, 247, 1079, 1081, 60, 537, 1018, 235, 659, 598, 1080]. Priorian [1306]. Prioritised [1657]. Priority [778, 175]. prisoner [481]. Pritchard [773]. probabilism [98, 748]. probabilist [1282]. Probabilistic [1391, 1433, 929, 768, 1519, 754, 1514, 1659, 1483, 1677, 1089, 1758, 316, 646, 315]. Probabilities [1046, 3, 323, 5, 405, 715, 571, 251, 1517]. Probability [6, 314, 64, 419, 393, 1508, 787, 240, 1566, 1193, 1205, 1516, 803, 286, 7, 535, 322, 914, 1521, 597, 2068]. probable [87]. probably [122, 2084, 1261]. probing [1075, 1564]. problem [1741, 300, 145, 2016, 285, 336, 1627, 93, 557, 665, 1847, 1889, 810, 823, 127, 978, 412, 476, 1888, 363, 1790, 520, 538, 1170, 1619, 2083, 2018, 2003, 1071, 692, 766, 803, 249, 942, 1200, 124, 767, 2081, 1875, 253, 1308, 2014, 1963, 1393, 1373, 233, 598, 625, 1431, 652, 1352, 2024, 1549, 1196, 1640, 1971, 789, 1703, 99, 64, 1595]. problematic [1840]. problems [343, 366, 1982, 728, 1701]. procedural [1570, 581, 1559]. procedure [1255, 580]. procedures [1891]. Proceedings [104]. process [1942, 2051, 8]. processes [1542, 1535, 1913]. processing [1810, 1694, 806, 1692]. Proclus [57]. product [1845]. productive [1388, 1241]. products [1180]. professionalization [2013]. program [75, 1211, 1306]. Progress [669, 792, 846, 134, 1490, 1863, 751]. prohibition [58]. Project [515, 618, 2065, 1047]. Projection [1122, 1054]. projections [1140]. projects [257]. prolegomena [616]. proliferation [1352]. prolonged [1594]. promise [1567]. pronoun [1000]. pronouns [1000].

Proof [1049, 445, 455, 653, 725, 656]. Proofs [86, 106, 437]. propensity [1090, 1382]. Proper [594, 524, 287, 828, 1276, 1637, 1537, 1586, 1738, 775, 903, 291]. properly [915]. properties [622, 717, 1950, 909, 1323, 321, 1954, 1428, 1803, 1267, 303, 791, 700]. property [552, 1733, 348]. proportionality [1633]. proposal [359, 2069, 139]. proposals [1555]. proposition [1889, 1893, 1892, 377]. Propositional [1753, 776, 498, 1902, 1363, 795, 1437, 1663, 1960, 1305]. Propositions [489, 1880, 1895, 1704, 1519, 1891, 1890, 983, 1894, 1123, 1317, 1900, 1901, 494, 495, 657, 1573, 1947]. propriety [260]. prospects [1388, 1599]. Prototypes [1272]. provided [130]. pruning [424]. Prussian [1587]. psychiatric [1933, 1928, 1930, 1927]. Psychiatry [1927, 1931, 1932]. psycho [1178]. psycho-neural [1178]. psychofunctionalist [933]. psychological [934, 1537, 1199, 1391, 1668, 309, 425]. psychologism [783]. psychologistic [1959]. psychology [1322, 415, 116, 634, 308, 1363, 713, 1217]. psychoneural [130]. psychotherapeutic [1559]. public [562, 164, 1854]. publish [736]. Pure [1098, 943, 1066, 1803, 1396, 1336, 1994]. Purposes [725, 707, 1781]. pursuit [894]. putative [434]. Putnam [117]. Putting [1540]. puzzle [4, 706, 146, 1343, 1813, 674, 1473, 1540, 1664, 2000, 974, 2048, 1040, 589]. Puzzles [1023]. puzzling [1617]. Pyrrhonian [1414]. Pythagorean [1473].

Qeauty [253]. qua [1973]. quadrant [999]. qualitative [1632, 1288, 127]. quality [1227]. quantification [668, 985]. quantificational [1433]. quantifiers [1949, 1603, 813]. quantifying [520]. quantities [58]. Quantum [193, 1287, 778, 323, 1566, 1392, 401, 548, 398, 1104, 1446, 681, 400, 1371, 253, 1921, 1543, 460, 1044, 700, 979]. quantum-information-theoretic [1044]. Quasi [197, 890]. Quasi-miracles [197]. quasi-truth [890]. quest [1873]. Question [1550, 1849, 1014, 1607, 1673, 1425, 162]. Questions [1592, 136, 1027, 744, 391, 1284, 1624, 1500, 2054]. quiet [15]. quietism [1856]. Quine [1503, 435, 781, 1177, 2065]. Quinean [1952, 338, 1723]. quirks [1305]. quo [1484, 212].

race [136, 1326]. Radical [1984, 600, 1375, 876, 1792]. rainbow [824]. ramifications [989]. Ramsey [517, 8, 110, 469]. Ramsification [430]. random [579, 1224]. Ranking [40]. rate [1765, 1390]. ratio [628]. Rational [1424, 407, 834, 1060, 1434, 1599, 146, 1145, 835, 1703, 1713, 542, 64, 1723, 158, 2082]. rationalism [300, 1666, 849, 971]. Rationality [404, 405, 1730, 558, 1589, 859, 1287, 120, 1338, 836, 414, 415, 1576, 950, 1742, 1673, 150, 96, 1601, 143, 142, 1330, 1083, 854, 56, 155, 1771]. Rationalizing [564, 118]. Ravens [94]. RCTs [1005]. Re [965, 112, 360, 1405, 486]. re-analysis [112]. re-emergence [360]. Re-thinking [965]. reactions [1564]. Reactive [191]. reading [180, 553, 1025]. real [136, 924, 992, 1831, 820]. real-world [992]. Realism [2016, 1450, 1203, 541, 70, 1489, 1862, 1849, 600, 2012, 1767, 1495, 1270, 1858, 1939, 1446, 1492, 1491, 213, 1488, 1918, 1485, 1873, 198, 1192, 1857, 1099, 1490, 1805, 1114, 1799, 206, 234, 1484, 1855, 884, 1487, 212, 1168, 703, 1861, 1413, 1856, 217, 605, 74, 85, 1924]. realism/anti [1799, 605]. realism/anti-realism [1799, 605]. realisms [45]. realist [228, 1320, 1859, 319, 225, 553, 1486]. realistic [97, 1772, 304, 1861]. realists [1169, 288]. reality [26, 1102, 1139, 207, 115, 825, 2045]. Realization [126, 130, 127, 1581, 128, 1983]. really [136, 1444, 1376]. Reason [948, 1940, 673, 2011, 1066, 119, 956, 1752, 1867, 1854]. reason-giving [1940]. reasonability [720]. reasonable [1595]. reasonableness [701]. Reasoning [251, 1935, 1136, 374, 376, 382, 779, 1657, 1199, 1158, 363, 750, 380, 373, 725, 417, 375, 694, 1282, 1429, 199]. Reasons [1730, 556, 1868, 772, 1124, 866, 708, 826, 118]. reassessed [731]. Reassessing [364]. reassessment [632]. received [1392]. reception [1396, 1336]. recipe [1487]. Reclaiming [781]. recognition [740, 1273, 1701]. Recollection [743]. reconciled [165, 149]. Reconciling [787, 423]. Reconsidered [1310, 920, 320, 793]. Reconsidering [218, 131]. reconstructions [1084]. recursion [1068]. Red [868, 462, 1787]. redefinition [1928]. reduced [38]. Reducing [724, 1640]. reductio [2010]. Reduction [846, 26, 1807, 195, 1190, 1928, 1884]. Reductionism [245, 1889, 1716, 899]. reductionists [130]. reductivism [851]. redundancy [1740, 1790]. Reed [473]. Reference [1588, 1001, 1176, 1460, 1590, 212, 214, 1430, 215]. reference-shift [1590]. referential [1804, 1573, 1947]. Referentialism [1919]. referentially [2004]. Refining [1552]. Reflecting [1033]. Reflection [667, 151]. Reflections [1495, 295, 873, 596, 614]. Reflective [472, 948, 473, 1769]. reform [139]. reformulated [1440]. reformulation [2081]. region [1075]. Regress [1890, 766, 88, 764, 1345, 1935]. regresses [1758]. Regret [409]. Regular [914]. regularities [1593]. Regularity [1568, 509]. Reichenbach [237, 236, 1316, 240, 238, 242, 232, 235, 239, 234, 233, 241]. Reichenbachian [2026]. Reid [2041]. reinterpretation [807]. Rejecting [1480]. rejection [954, 1848, 1585]. rejoinder [1172]. related [1323]. relation [665, 126]. relational [1054]. Relationalism [1267]. relations [1819, 128, 1040]. relative [237, 798, 1191, 1268]. relativism [1554, 610, 1557, 994, 1926]. relativist [817, 325]. Relativistic [1651, 1045, 1055, 1044]. relativity [397, 1042, 1048, 1049, 1394, 1353, 1041, 1249]. relativized [235]. Relativizing [235]. relevance [265, 102, 311, 1020, 442]. relevant [133, 348]. Reliabilism [733, 704, 1082, 942, 124, 1091, 903]. Reliability [240, 738, 1203, 870, 1541, 266, 922, 432, 417]. reliance [547]. religion [165, 157]. rem [1678]. Remarks [549, 1469, 519, 2081, 315]. remember [187]. Remembering [654, 743, 2061]. Reminiscing [2056]. renormalization [1444]. Repelling [1587]. replaced [1100]. Replacing [1487]. Replication [202, 2032]. replicators [202]. Replies [1408, 1876, 473]. Reply [1174, 255, 2081, 593, 642, 1166, 1168]. reports [1590, 984]. representable [1053]. Representation [920, 528, 277, 13, 1810, 1130, 1818, 1880, 17, 832, 1096, 394, 697, 279, 1736, 1580, 1533, 1854, 1594, 567]. Representation-hunger [920]. representational [800, 1937, 1491, 512, 1532, 1115, 987, 1008]. representationalism [1175, 1223, 261]. Representations [745, 1844, 1206, 714, 889, 2025, 1864, 1187, 1133, 1354, 826, 1215, 2076]. representatives [579]. representing [305, 275]. reputation [1526]. requirement [1235]. requirements [278, 1037]. requires [1354, 1843]. rereading [850]. resampling [210]. rescues [426]. research [724, 2050, 1981, 1502, 1199, 770, 841, 1211, 183, 139, 1930, 1927, 1559, 1611, 1329]. research- [1927]. resemblance [1500, 1669]. residuated [1053, 521]. resilience [1541]. Resisting [1489]. resolute [180, 1595]. resolution [1140, 24]. resolutions [366]. resolve [855]. resolving [344]. resource [99]. resource-bounded [99]. resourcefulness [1114]. resources [39, 278]. Response [296, 1414, 253, 1824, 1417, 573]. responses [1205]. Responsibility [1461, 1545, 1298, 1468, 506, 699, 902]. Responsible [1471, 1264]. rest [1673]. restatement [274]. restricting [519]. restriction [627]. restrictions [1843]. result [645, 248, 1072]. results [736, 1604]. resurgent [1529]. retention [1431]. Rethinking [1741, 1322]. retraction [1903]. Retractions [1732]. RETREAT [1751]. Retrocausality [1007]. return [1669]. revamped [1568]. reveal [1496]. Revelation [1447]. reversals [950]. reverse [1329, 1654]. reverse-engineering [1329]. reversible [1913]. Reversing [408]. review [411, 1004, 2013]. revisability [338]. revised [1548]. revision [565, 718, 651, 1060, 1434]. revisited [1436, 1749, 1847, 1548, 83, 77, 157, 1680, 125, 1189, 453, 1319, 599, 313, 402]. revisiting [1093]. revived [1449]. revolution [1214, 295, 1669]. revolutions [1445, 1574, 1809, 1684]. rhetoric [1203]. rich [1734]. Richness [1589, 102]. Rigged [1640]. rigid [436, 846]. Rigidity [1671, 1764, 249]. ripping [1217]. risk [1753, 1498, 1973, 959, 2036]. rivals [1450]. road [397, 1921, 241, 1092]. Robust [1939, 1435]. Robustness [1139, 299, 1021, 129, 1138]. role [1952, 228, 1136, 360, 1166, 1922, 770, 1331, 637, 372, 995, 1397, 1953, 293, 662, 1860]. roles [1079]. rollback [508]. roller [545]. Romeo [1766]. Rorty [1399]. Ross [1444]. Roush [87]. royal [1409]. Rule [560, 58, 1829, 1409]. Rule-following [560]. rules [1605, 391]. run [545]. Russell [807]. Russellian [1477]. Russo [440, 1979]. Ryoan [908]. Ryoan-ji [908].

s [1531, 673]. S. [2081]. Saatsi [85]. sad [1505]. safe [1451]. safely [1119]. safety [297, 1887, 705]. said [1614]. Saint [765]. Salerno [2085]. same [489, 211]. same-saying [489]. Sameness [496]. samesaying [497]. sandbag [904]. Satan [765]. savage [404]. save [297]. saved [426]. saves [1082]. Saving [773, 215, 264, 264]. savoir [1354]. savoir-faire [1354]. say [953]. Saying [2028, 489]. scalar [1352]. scale [2032, 1971]. scenario [841, 1052]. scenarios [363]. sceptical [784]. Scepticism [506, 1158, 1557, 1792, 1984]. Schaffner [1884]. schema [2051]. schemes [1227]. School [189]. Schröder [293]. Schupbach [315]. Science [2012, 1607, 1871, 60, 176, 210, 56, 679, 1978, 1874, 928, 1276, 893, 1869, 1531, 881, 818, 58, 45, 1553, 1795, 1941, 792, 359, 888, 134, 1555, 216, 133, 165, 1908, 136, 957, 127, 901, 1138, 1457, 1556, 196, 1585, 1870, 735, 1302, 895, 1641, 618, 1875, 764, 157, 1547, 135, 728, 713, 794, 1863, 8, 162, 662, 140, 2013, 141, 619, 589, 882, 1611, 1850, 138, 166, 731, 1329, 59, 291, 2013, 734]. sciences [290, 403, 1090, 690, 1819, 1939, 805, 1975, 129, 1583, 842, 1684, 1799, 1617, 619, 2062, 1999, 1809]. scientia [1270]. Scientific [14, 1574, 137, 1684, 1484, 1855, 1115, 1131, 265, 1862, 887, 352, 1937, 208, 740, 2016, 1136, 1169, 219, 13, 1495, 1270, 1637, 1274, 846, 1121, 1166, 1445, 1555, 1492, 889, 1768, 17, 531, 689, 911, 870, 1133, 213, 1848, 967, 1485, 1873, 738, 1490, 588, 1091, 1420, 1736, 234, 884, 212, 631, 241, 289, 296, 1932, 582, 117, 741, 1861, 783, 1856, 56, 737, 1809]. scientist [1691, 1973, 874, 181]. scientistic [149]. scientists [736]. SCOP [1601]. scope [1823, 1332, 1110]. scopers [1076]. scoping [1076]. scorekeeping [763]. se [1731, 542, 1145, 610]. sea [1119]. seams [1217]. search [1070, 367, 580]. Second [1954, 1627, 843, 750, 535, 1156]. Second-order [1954, 843, 750, 1951]. secret [1239]. sector [1350]. see [1787, 1762]. Seeing [384, 374, 607, 1738]. seeing-as [374]. seek [874]. Seemings [1720, 822]. seen [89]. selecting [363]. selection [969, 511, 1879, 1942, 1346, 579, 2032]. selective [1888, 1486]. selectivity [1485]. Self [562, 285, 248, 621, 42, 1573, 1320, 1608, 1804, 1365, 934, 1639, 91, 923, 1223, 154, 1914, 2056, 1796, 1932, 1698, 1947]. self- [1698]. self-adjointness [1796]. self-control [934]. self-explanation [1639]. self-gravitating [1914]. Self-Knowledge [621, 42, 1608, 923]. Self-location [285, 91]. Self-organisation [248]. Self-referential [1573, 1947]. self-representationalism [1223]. Self-serving [562]. Selfless [1011]. selflessness [610]. Sellars [1270, 1637, 1236]. Sellarsian [1529]. selves [1482, 1319, 1698]. Semantic [1943, 342, 1926, 599, 685, 120, 53, 1127, 531, 610, 1885, 910, 1964, 994, 1779, 76]. Semantics [630, 67, 1031, 1304, 1311, 758, 107, 608, 796, 998, 1602, 843, 1954, 1371, 744, 1074, 108, 1336, 981, 325, 452, 680, 1032, 92, 657, 904, 1335, 1028, 1396]. semantics/pragmatics [657]. semigroups [1053]. semiotic [687]. semiotics [1001]. semirealism [1857]. sensation [1213]. sense [1871, 1001, 1404, 1378, 496, 2056]. senses [362]. sensitive [1638, 92]. sensitivity [1021, 298, 326]. sensorimotor [2042]. Sensory [1658]. sentence [38, 1273]. sentences [67, 1305, 1019, 1017]. sentential [1805]. sentimentalism [1666]. sentimentalism/rationalism [1666]. SEP [594]. sequences [565, 1312, 1006]. sequent [451]. sequential [97]. serendipity [1941]. series [1066]. serious [1797]. serving [562]. set [940, 1069, 544, 1807, 527, 1046, 1068]. Sets [393, 1228, 1566, 1438, 535]. setting [40]. Settings [2046]. settle [747]. Shaky [1113]. shall [321]. Shannon [1233]. shape [893]. Shapiro [81]. shared [583, 611, 1984]. sharing [740, 137, 2058]. sheet [998]. shift [1590]. shifting [1150]. ship [1515]. shocks [1117]. Shogenji [318, 670, 417]. shores [629]. Should [1169, 1859, 1470, 736, 87, 351, 896, 408]. shouldn’ta [1039]. show [87]. showing [1743]. SI [734]. Sign [2065]. Signal [1703, 680]. signaling [1936, 11]. signals [1960]. signatures [271]. significance [1047, 2002, 950, 1341, 570, 449, 1946]. signs [2069, 687]. silent [1416]. similar [1555]. Similarity [551, 5]. Simple [1244, 184, 993, 1356, 545, 1649, 973]. simpler [1772]. simples [1412]. simplicity [662]. Simpson [244]. simulated [271]. simulating [1585]. Simulation [716, 793, 513, 527, 211, 340, 512, 1798, 170, 1698]. simulations [690, 638, 1212, 590]. Simultaneous [693]. single [38, 1100, 64, 1057]. single-case [64]. single-particle [1100, 1057]. single-sentence [38]. single-step [38]. Singular [1704, 1189, 2023, 1586, 1760, 337]. singularist [1857]. singularities [1919]. sitting [1786]. Situated [328]. situations [1600, 860, 1056]. Skeptical [1257, 1137, 505]. skepticism [1852, 507, 1888, 1385, 876, 1404, 1635, 628, 467]. sketches [383, 308]. skilful [1021]. skill [2031, 873]. Skills [1570]. Sleeping [896, 1088, 1698, 69, 253, 593]. slugfest [609]. small [1438, 785]. Smelling [2030]. Snow [2024]. Sober [81]. Social [101, 1254, 20, 1180, 1998, 2060, 1699, 1361, 1432, 538, 2033, 1370, 341, 863, 1556, 1074, 858, 774, 1149, 1367, 648, 862, 2059, 1762, 2057, 837, 2062, 1850, 854]. Socially [133]. Society [104, 243]. sociotechnical [1611]. Socratic [1873]. software [1010]. Solitary [1432]. solution [1489, 1086, 1627, 749, 93, 1813, 1587, 392, 1963, 1846, 1779, 652, 642]. solutions [2018]. solve [823, 1550]. Solving [1308]. Some [717, 519, 2081, 989, 315, 1702, 829, 522, 696, 434, 826, 452, 1011, 1059, 1609]. something [761]. sorites [1882, 1515, 1779]. sortalist [1827]. sorted [1503]. Sosa [472, 2063]. sounding [1505]. soup [82]. space [396, 457, 1833, 968, 376, 1100, 1097]. space-time [1097]. space-times [457]. spaces [1238, 569, 1522, 2043, 630, 1966]. spaceship [167]. spacetime [1045]. spacetimes [461, 424]. spatial [378, 131]. spatio [1059]. spatio-temporal [1059]. speak [1559]. Speakable [681]. Speaking [812]. Special [1729, 314, 1179, 305, 275, 397, 1748, 2012, 1014, 1607, 1193, 1205, 1048, 2033, 232, 1079, 1598, 1688, 1461, 263, 456, 1641, 1708, 2055, 140, 1927, 1620, 1771]. special-interest [140]. specialization [1809, 1684]. specialize [861]. species [93, 96, 1383]. specific [1430]. specification [283]. specificity [1609]. specified [160]. specter [1385]. speech [1677]. Spider [1994]. Spider-Man [1994]. splitting [188]. Spontaneous [1549]. spontaneousness [2031]. Spurious [1309]. Squares [1301]. St. [1084]. Stabilising [440]. stability [307, 1854]. stabilized [267]. staccato [545]. Stackelberg [1181]. Stance [152, 142, 1867, 852, 149, 270, 143, 574, 155]. Stances [151, 145, 146, 149]. Standard [433, 1046, 213, 1054, 452]. Stanford [295]. Stanfordite [1859]. Stars [1914]. Starting [1052]. state [1104, 487]. stated [915]. statement [1048]. Statements [1425]. states [422]. statistical [1916, 2080, 1914, 1784, 1118, 1347]. statistics [1921, 1005]. status [1117]. Steadfastness [1576]. steam [1914]. Stem [1188]. step [38, 1469]. Stereotypes [1961, 934]. stew [211]. Stewart [1215]. still [867, 822, 1230]. stochastic [1090]. stone [82]. stopped [351]. stopping [1005]. story [618]. strand [1396, 1336]. strategic [1185, 752, 1545]. strategically [1600]. Strategies [530, 1675, 2050, 1181, 1774, 1661, 307]. strategy [1980, 752, 1746, 1329]. streams [1838]. Street [1124]. strength [470]. striking [922]. string [541]. striving [1838]. strong [1713, 2082]. Structural [1320, 1491, 213, 70, 194, 214, 600, 1445, 1446, 1582, 1266, 222, 653, 198, 85]. Structuralism [1441, 1442, 111, 436, 223, 1678]. structuralist [846, 201, 1504]. structure [600, 668, 1443, 1357, 1888, 1902, 1685, 268, 422, 1809, 1684, 1910, 2042, 20]. Structured [1900, 1891, 1901]. Structures [1811, 494, 89, 695, 677, 658, 1416, 212, 521, 296]. Studies [73]. study [977, 947, 1687, 2079, 1295, 1483, 1346, 1391, 370, 746, 1012, 532, 979]. Studying [1774]. Sturgeon [1711]. style [1158, 2018, 1591]. sub [319]. sub-hypothesis [319]. subject [1502, 514]. subjective [1772, 1787, 1755]. subjunctive [563]. Sublating [1840]. subset [127]. Substantive [561]. substructural [912]. Success [605, 1488, 1861, 259]. success-to-truth [1861]. successive [693]. sufficient [404, 1875, 1606]. sui [1720]. superduperfunctionalism [1242]. supererogation [919]. superiority [1798]. superlative [1603]. supersets [1228]. superstars [1546]. supertask [1356]. Supervaluational [1437]. Supervenience [231, 1242, 770, 964, 1271]. support [112, 6, 593, 1758, 646, 590]. supports [2074]. supposed [1586]. supposition [218]. surprise [1550]. surprises [110]. surprising [1968, 1481]. suspended [1477]. Swampman [1276]. syllogism [695, 799]. syllogistic [1020]. symbolic [177, 179, 176]. symbolism [175]. symbols [516, 1669]. symmetric [1852, 431]. Symmetry [1527, 1122]. symptoms [1928]. syncategorematic [1064]. synchronic [1125]. synergy [360]. Syntactic [1144, 1190]. Synthese [1729, 305, 275, 1205, 734, 232, 263, 456, 1641, 2055]. synthetic [240, 1148, 59]. system [940, 1554, 1670, 749, 1655, 283, 433]. systematic [1136, 1873]. Systematicity [1874, 1873, 1875, 1872, 1869, 987]. Systematizing [1705]. Systems [832, 1096, 1362, 647, 1188, 1262, 1916, 1661, 1482, 11, 248, 2026, 1689, 1580, 1914, 18, 1667, 1549, 2080].

tableaux [191]. tacit [547]. takes [1276]. tale [1206, 257]. talk [1151, 626]. Talking [707]. tapestry [1838]. target [1635]. Tarski [914, 850, 1331, 907]. Tarskian [1335]. task [1273, 750]. tasks [753]. taste [1903]. taxonomies [1983]. Taxonomy [1644, 1727]. teams [2039]. Teleological [1377, 1379, 1381]. Teleology [1380, 1721, 1781]. teleosemantic [1540]. tell [268, 157]. telling [1291]. Temp [773]. temporal [1312, 1186, 1059]. temporal-epistemic [1186]. temporally [1676]. tenets [1485]. Tense [1317, 1313, 1316, 1311, 1660, 476, 1834, 1318, 478]. tense-logical [1834]. tense-logicians [1318]. term [755, 1448]. terms [223, 1982, 249, 1189]. Test [110, 219, 1550, 517, 976, 973, 469]. testimonial [1716, 738, 1650]. testimony [812, 1716, 116, 641, 899]. text [370]. Thagard [1215]. Thank [1310]. Thank-Goodness [1310]. Tharp [1386]. their [1439, 2066, 1187, 209, 303, 2060]. them [679, 2060, 952, 504]. theme [686]. themes [178]. themselves [1559]. theological [163]. theorem [447, 484, 827, 25, 639]. theorems [1601, 1386]. theoretic [600, 434, 461, 982, 1613, 967, 560, 1044]. Theoretical [1561, 67, 223, 1705, 1152, 319, 1863, 956, 1197, 1198, 19, 532]. theoreticality [926]. theories [887, 683, 13, 696, 1568, 544, 846, 1934, 536, 1288, 569, 531, 1879, 1199, 1885, 263, 1053, 8, 194, 747, 1511, 1899, 1854, 1059, 1860, 1903, 596, 402, 212]. theorist [1236]. theorists [1088, 1746, 408]. Theory [949, 436, 569, 513, 882, 940, 1069, 1589, 1596, 966, 97, 89, 2040, 37, 550, 951, 1807, 2023, 1247, 732, 1038, 787, 1104, 1566, 160, 216, 342, 758, 1772, 1238, 2069, 1357, 1209, 712, 409, 1980, 1986, 2075, 2003, 687, 844, 1613, 1384, 411, 1703, 850, 877, 1873, 653, 541, 88, 198, 1054, 1003, 634, 70, 1959, 1760, 331, 764, 963, 1085, 450, 797, 827, 1654, 793, 2001, 1612, 54, 266, 214, 1197, 1198, 777, 563, 1528, 1543, 1041, 379, 741, 269, 1258, 170, 1249, 1204, 1961, 1698, 1996, 1215, 1839, 301, 229, 751]. theory-ladenness [266]. There [1787, 1761, 1388, 874, 872, 952, 1654, 1128, 1277, 921, 694]. thermodynamics [1914, 1913]. these [1305]. Theseus [1515]. Thesis [339, 245, 795, 1968, 240, 1695, 1072, 1366, 125, 1883, 1979, 1872]. They’ve [873]. thick [1736]. Thin [868, 1736, 462]. things [427, 924, 196, 348, 1037, 1652]. think [1459, 953, 437, 1376]. Thinking [179, 2055, 1495, 743, 965, 123, 1484]. thinks [1016]. thought [818, 2023, 690, 1586, 944, 1055, 1760, 1085, 1040, 438]. thoughts [953, 1731, 610, 1677, 974, 1701]. threats [1485]. Three [1624, 1779, 1831, 958, 1194]. Three-valued [1779]. tied [316, 315]. tier [1156]. tiered [564]. ties [1920]. Time [482, 1806, 1478, 706, 1833, 1295, 2061, 478, 1097, 1058, 460]. timeless [880]. times [457, 250, 1319]. together [2061, 2056]. token [1358]. Tolerance [1882, 1510, 446, 50, 204]. tonk [1302]. too [1024, 1679, 678, 785]. Tool [1240, 1308]. Tool-box [1240]. Toolbox [618]. tools [1340]. topical [1003]. topics [936]. Topological [129, 858, 1271, 717, 1584, 1577, 1581, 1966]. topology [1583, 1210]. tortoise [1086]. toxicology [140]. toy [399, 1240]. toy-box [1240]. track [365]. tracking [412, 942]. tractability [753]. trade [1658, 530, 329]. trade-off [329]. trade-offs [1658, 530]. tradition [877, 370]. traditional [101, 1274, 1081]. Traditions [370]. transcendence [446]. Transcendental [45]. Transfinite [1068, 850]. transformational [1820]. transient [218]. Transition [796]. transitions [1920, 1917]. Transitivity [1639, 1633]. translation [1519]. translational [1304]. Transmission [867, 112, 500, 371, 1071, 330, 643, 595, 646, 1159]. transmits [1591, 644].

Transparency [1292, 1094, 1608, 1737]. Transparent [985, 1866]. travel [2061, 1806]. treatment [1150, 1930]. trees [71]. trenches [735]. trends [1729, 1117]. triangulation [1970]. trichotomous [864]. tricks [869]. tries [1832]. trilemma [1379, 573]. triviality [1985, 1671, 1072]. trivialization [249, 602]. trope [123, 88]. True [1785, 226, 1405, 1095, 970, 332]. Trumping [246]. Trust [866, 837, 138, 102, 562, 836, 835, 137, 838, 1526, 834, 2039]. trusting [840, 839]. Truth [2022, 856, 1411, 1623, 1333, 44, 1704, 445, 2007, 798, 440, 1567, 31, 1628, 786, 1193, 1880, 365, 1205, 1626, 412, 1636, 1629, 1191, 1622, 1280, 1488, 1873, 1925, 209, 334, 1410, 1142, 1399, 1898, 529, 194, 214, 907, 1511, 1012, 1603, 1386, 1620, 741, 1861, 605, 259, 292, 890, 1131, 1621, 455]. truth-conditions [1704]. truth-conducive [786, 1925]. truth-functionality [1511]. truth-makers [440]. truthful [1964]. truthlikeness [1234]. Truthmaker [640, 963, 1746, 921]. Truthmakers [257, 65, 909]. Truthmaking [490]. truths [978, 1095, 970, 1887, 35, 1746]. tube [219]. tuning [2002]. Turf [341]. Turing [1044]. turn [401]. tweaking [527]. Tweety [589]. Two [1479, 260, 1721, 658, 1127, 247, 1384, 327, 46, 997, 433, 457, 736, 1362, 65, 608, 564, 258, 1777, 1646, 1004, 1832, 494, 257, 178, 748, 1005, 184, 1147, 1549]. two-dimensional [608, 1777]. two-fold [184]. two-systems [1362, 1549]. two-tiered [564]. twofold [372]. Type [982, 1238, 1000, 850, 1358, 301, 229]. Type-theoretic [982]. type-token [1358]. types [1592, 658, 1774, 1572, 2054]. typicality [197].

ubiquity [1822]. Uexküll [184]. UKCP09 [1140]. unasked [136]. unbelief [480, 479]. uncertain [388, 391]. Uncertainty [535, 51, 1973, 10]. Unconceived [2014, 2013, 2012, 2015]. unconscious [1924]. Undead [1511]. undefined [979]. underdeterminacy [1864]. underdeterminacy-free [1864]. Underdetermination [219, 703, 217, 2028, 220, 218, 222, 541, 602, 212, 633]. underlie [1154]. undermines [1094]. Undermining [963, 546]. Understanding [1130, 1132, 2044, 1614, 1486, 567, 947, 749, 84, 1722, 1660, 957, 575, 1133, 1396, 1336, 1650, 780, 945, 1025, 1135, 1469, 1389, 1131]. unfathomable [154]. Unfolding [690]. ungrounded [1800]. unicepts [1540]. Unification [1980, 1500, 634, 1126, 775, 1967, 1292]. unificationist [1612]. unified [1842, 1542]. uniform [1032]. unique [1665]. uniquely [1211]. uniqueness [1401, 1000]. Unitary [1262]. Uniting [877]. units [421]. Unity [1898, 1890, 1889, 1893, 1325, 1302]. Univalent [1504]. universal [30, 990, 715, 1331]. universalis [62]. universalist [877]. universality [1042, 1780, 1706, 656]. universe [1357, 233, 1057]. universes [1357]. unknowability [534]. unknowable [1519, 35]. unknown [31, 1878]. unobservability [1366]. unreasonable [256, 354]. unresponsive [860]. unrestricted [2004]. unsatisfactory [1075]. unthinkable [1254]. unto [321]. unwelcome [339]. update [307]. updates [693]. updating [410, 1385, 674, 935]. upon [154]. use [1911, 527, 752, 425]. used [1270]. usefulness [1992]. useless [1093]. users [1187]. uses [1990, 333]. using [864, 718, 1458, 1055]. Utilitarian [576]. utilitarianism [1265]. Utility [100, 793]. utterance [1497].

vacuous [1156]. vadis [1484, 212]. Vague [1514, 1518, 1460]. Vagueness [565, 1508, 51, 50, 204, 226, 1251, 53, 1454, 302, 1509, 1516, 1513, 1154, 1510, 1511]. valid [449]. validity [800, 2052, 1972, 1649, 1222, 1008]. valuation [533]. Value [1339, 1547, 1682, 702, 1628, 790, 429, 581, 1402, 493, 348, 18, 837, 1218]. value-relevant [348]. valued [1779]. values [219, 901, 429, 1517, 1224, 2015, 596]. variability [858, 657]. variable [1196]. variables [1195, 1224]. variation [120, 53]. Variations [686]. Varieties [500, 65, 1172, 957, 1171]. variety [1968, 1350]. variety-of-evidence [1968]. vast [1119]. vector [2067]. vehicles [1491]. verbal [750]. Verballed [759]. verbs [107]. verdicts [1288]. verification [1049]. verificationist [1255]. Verisimilitude [683, 592, 602, 25]. Verities [1515]. veritistic [864]. vernacular [1305]. versa [1424]. version [623, 1394, 124, 598]. versus [2016, 2063, 13, 328, 587, 1445, 1889, 512, 769, 179, 213, 1736, 1575, 113, 1037, 453, 883, 1413]. very [1098, 614, 879]. via [1810, 240, 1288, 693, 1812]. viable [1489]. vice [1424]. vices [1420, 903]. victory [1810]. view [67, 1392, 1043, 189, 1885, 1371, 1325, 735, 1738, 1367, 775, 1110, 825, 1663]. viewed [1526]. views [1362, 1188]. Vindicating [1970, 953, 1302]. vindication [1464, 725]. virtual [1851]. virtue [1638, 1435, 435, 773, 200, 927, 913, 1091, 1719, 292]. Virtues [1420, 786, 1705, 911, 566, 1073]. Virtuous [1475]. vis [1071]. vis-à-vis [1071]. vision [281]. Visions [1051, 1422]. Visual [1734, 177, 1210, 1474]. visualisation [688]. Vital [1999]. volitional [1838]. Voluntarism [1463, 145, 146, 1464, 143, 144, 147, 148]. voluntary [1631]. Vortex [1993]. Voting [1251, 1636]. vs [11, 1399, 1994, 118, 789].

W [192]. wage [534]. wager [1568]. Wahrheitsbegriff [850]. Walk [756]. Wallace [1879]. Walter [320]. WAMs [2048]. want [1696]. WAR [1589]. Warrant [172, 334, 500, 1716, 1591, 644, 1197, 1198, 1347]. warrants [918]. wars [1810]. Was [1236, 2041, 1614]. water [131]. wave [1098, 1102, 1099, 1100, 1097]. wavefunction [1105]. waves [1915]. way [1767, 153, 201, 1738, 497, 747, 1058]. ways [1583, 494, 1755, 1926, 751]. weak [587, 964]. weak-link [587]. weakly [221]. Weakness [720, 934, 1813]. weasel [1111]. web [1114, 279]. weighted [1418]. weighting [2027]. Weisberg [1418]. Well [108, 351]. Well-founded [108]. were [707]. Weyl [236]. Where [872, 1556, 1318, 1118, 1135]. whether [650, 1761, 736, 786]. Which [1166, 84, 1610]. White [2024]. Who [1768, 1024, 1088, 889, 1958]. whole [166]. wholes [2066]. Whose [1781]. Wide [279, 1076]. wide-scopers [1076]. wiggled [190]. Wigner [1473]. wild [1784]. will [720, 934, 1837, 812, 1840, 1813, 1838, 879, 1831, 746, 1467, 847]. Willem [189]. Williamson [440, 113, 29, 1979]. Willusionism [847]. winning [978]. wisdom [1941]. witch [1610]. Wither [691]. Withering [221]. within [928, 2069, 241]. Without [928, 357, 1704, 490, 1269, 1820, 1246, 350, 1765, 368, 223, 1424, 832, 1096, 1000, 1428, 1370, 868, 697, 258, 1286, 202, 780, 1935, 1442, 9, 210, 1415, 262, 1306, 1974, 1478, 1938, 460]. witnesses [860]. Wittgenstein [943, 988, 1225, 1674, 501, 761, 897]. Wittgensteinian [344]. Woes [29]. Wonderland [1585]. won’t [1797]. Woodward [266]. Words [406, 72]. work [1331, 1540, 1733]. World [279, 880, 757, 548, 1101, 2033, 153, 992, 268, 1265, 725, 25, 252, 233, 1698]. world-simulation [1698]. Worlds [2085, 250, 1319, 645, 1026, 1988, 491, 492, 2004, 1300, 729, 113]. worry [220, 957]. worrying [136, 351]. worth [736]. would [1286]. wouldn’ta [1039]. would’ve [190]. Wright [1116, 77, 330].

x [1248]. x-phi [1248].

Ya [1039]. Yablo [1158]. Yablo-style [1158]. Yabloesque [1596, 1804]. Yang [1249]. year [779]. years [759, 408]. Yes [864]. You’d [504, 190]. young [1296].

Zeno [539, 1015]. zero [715]. Zwart [25].

References

Arlo-Costa:2010:I

[1] Horacio Arló-Costa and Jeffrey Helzner. Introduction. Synthese, 172 (1):1–6, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

      1573-0964 (electronic). URL http://link.springer.com/article/10.
      1007/s11229-009-9468-1;   http://link.springer.com/content/
      pdf/10.1007/s11229-009-9468-1.pdf.

Baron:2010:CBM

[2] Jonathan Baron. Cognitive biases in moral judgments that affect political behavior. Synthese, 172(1):7–35, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

      http://link.springer.com/article/10.1007/s11229-009-9478-z;
       http://link.springer.com/content/pdf/10.1007/s11229-009-9478-
      z.pdf.

Arlo-Costa:2010:AAE

[3] Horacio Arló-Costa and Jeffrey Helzner. Ambiguity aversion: the explanatory power of indeterminate probabilities. Synthese, 172(1):

37–55, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9475-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9475-2.pdf.

Bovens:2010:PH

[4] Luc Bovens and Wlodek Rabinowicz. The puzzle of the hats. Synthese, 172(1):57–78, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9476-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9476-1.pdf.

Gilboa:2010:DOP

[5] Itzhak Gilboa, Offer Lieberman, and David Schmeidler. On the definition of objective probabilities by empirical similarity. Synthese, 172 (1):79–95, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9473-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9473-4.pdf.

Levi:2010:PLL

[6] Isaac Levi. Probability logic, logical probability, and inductive support. Synthese, 172(1):97–118, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9474-3;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9474-3.pdf.

Maher:2010:BP

[7] Patrick Maher. Bayesian probability. Synthese, 172(1):119–127, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9471-6;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9471-6.pdf.

Sahlin:2010:DSR

[8] Nils-Eric Sahlin, Annika Wallin, and Johannes Persson. Decision science: from Ramsey to dual process theories. Synthese, 172(1):129–143, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9472-5;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9472-5.pdf.

Samet:2010:KP

[9] Dov Samet. S5 knowledge without partitions. Synthese, 172(1): 145–155, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9469-0;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9469-0.pdf.

Seidenfeld:2010:CCF

[10] Teddy Seidenfeld, Mark J. Schervish, and Joseph B. Kadane. Coherent choice functions under uncertainty. Synthese, 172(1):157–176, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9470-7;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9470-7.pdf.

Huttegger:2010:EDL

[11] Simon M. Huttegger, Brian Skyrms, Rory Smead, and Kevin J. S. Zollman. Evolutionary dynamics of Lewis signaling games: signaling systems vs. partial pooling. Synthese, 172(1):177–191, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9477-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9477-0.pdf.

Contessa:2010:I

[12] Gabriele Contessa. Introduction. Synthese, 172(2):193–195, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9501-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9501-4.pdf.

Chakravartty:2010:IVF

[13] Anjan Chakravartty. Informational versus functional theories of scientific representation. Synthese, 172(2):197–213, January 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9502-3;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9502-
3.pdf.

Contessa:2010:SMF

[14] Gabriele Contessa. Scientific models and fictional objects. Synthese, 172 (2):215–229, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9503-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9503-2.pdf.

French:2010:KQO

[15] Steven French. Keeping quiet on the ontology of models. Synthese, 172 (2):231–249, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9504-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9504-1.pdf.

Frigg:2010:MF

[16] Roman Frigg. Models and fiction. Synthese, 172(2):251–268, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9505-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9505-0.pdf.

Giere:2010:ABC

[17] Ronald N. Giere. An agent-based conception of models and scientific representation. Synthese, 172(2):269–281, January 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9506-z;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9506-
z.pdf.

Thomson-Jones:2010:MSF

[18] Martin Thomson-Jones. Missing systems and the face value practice. Synthese, 172(2):283–299, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9507-y;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9507-y.pdf.

Toon:2010:OTM

[19] Adam Toon. The ontology of theoretical modelling: models as make-believe. Synthese, 172(2):301–315, January 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9508-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9508-
x.pdf.

Zollman:2010:SSE

[20] Kevin James Spears Zollman. Social structure and the effects of conformity. Synthese, 172(3):317–340, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9393-8;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9393-
8.pdf.

Casullo:2010:KM

[21] A. Casullo. Knowledge and modality. Synthese, 172(3):341–359, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9396-5;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9396-5.pdf.

Goodwin:2010:CKE

[22] William Mark Goodwin. Coffa’s Kant and the evolution of accounts of mathematical necessity. Synthese, 172(3):361–379, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9397-4;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9397-4.pdf.

Dodd:2010:CAC

[23] Dylan Dodd. Confusion about concessive knowledge attributions. Synthese, 172(3):381–396, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9400-0;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9400-0.pdf.

Godden:2010:IBA

[24] David M. Godden. The importance of belief in argumentation: belief, commitment and the effective resolution of a difference of opinion. Synthese, 172(3):397–414, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9398-3;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9398-
3.pdf.

Schurz:2010:ZFI

[25] Gerhard Schurz and Paul Weingartner. Zwart and Franssen’s impossibility theorem holds for possible-world-accounts but not for consequence-accounts to verisimilitude. Synthese, 172(3):415–436, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9399-2;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9399-2.pdf.

Bregant:2010:MRR

[26] Janez Bregant, Andraz Stozer, and Marko Cerkvenik. Molecular reduction: reality or fiction? Synthese, 172(3):437–450, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9401-z;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9401-z.pdf.

Sarch:2010:BAP

[27] Alexander Sarch. Bealer and the autonomy of philosophy. Synthese, 172 (3):451–474, February 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9402-y;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9402-y.pdf.

Salerno:2010:IKB

[28] Joe Salerno. Introduction to knowability and beyond. Synthese, 173 (1):1–8, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9680-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9680-z.pdf.

Tennant:2010:WW

[29] Neil Tennant. Williamson’s woes. Synthese, 173(1):9–23, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9673-y;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9673-y.pdf.

Hand:2010:AUK

[30] Michael Hand. Antirealism and universal knowability. Synthese, 173 (1):25–39, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9674-x;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9674-x.pdf.

Edgington:2010:PKU

[31] Dorothy Edgington. Possible knowledge of unknown truth. Synthese, 173(1):41–52, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9675-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9675-9.pdf.

Fara:2010:KCK

[32] Michael Fara. Knowability and the capacity to know. Synthese, 173 (1):53–73, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9676-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9676-8.pdf.

Proietti:2010:FPC

[33] Carlo Proietti and Gabriel Sandu. Fitch’s paradox and ceteris paribus modalities. Synthese, 173(1):75–87, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9677-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9677-
7.pdf.

Kvanvig:2010:IKP

[34] Jonathan Kvanvig. The incarnation and the knowability paradox. Synthese, 173(1):89–105, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9678-6;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9678-6.pdf.

Routley:2010:NLK

[35] Richard Routley. Necessary limits to knowledge: unknowable truths. Synthese, 173(1):107–122, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9679-5;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9679-5.pdf.

Liu:2010:ABD

[36] F. Liu and O. Roy. Advances in belief dynamics: Introduction. Synthese, 173(2):123–126, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9710-x;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9710-x.pdf.

Cantwell:2010:ACE

[37] John Cantwell. On an alleged counter-example to causal decision theory. Synthese, 173(2):127–152, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9689-3;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9689-
3.pdf.

Hansson:2010:MIC

[38] Sven Ove Hansson. Multiple and iterated contraction reduced to single-step single-sentence contraction. Synthese, 173(2):153–177, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9688-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9688-4.pdf.

Alechina:2010:BAU

[39] Natasha Alechina and Brian Logan. Belief ascription under bounded resources. Synthese, 173(2):179–197, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-009-9706-6;   http://link.springer.
com/content/pdf/10.1007/s11229-009-9706-6.pdf.

Porello:2010:RJA

[40] Daniele Porello. Ranking judgments in Arrow’s setting. Synthese, 173 (2):199–210, March 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9568-y;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9568-y.pdf.

Chan:2010:MPJ

[41] Timothy Chan. Moore’s Paradox is not just another pragmatic paradox. Synthese, 173(3):211–229, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9403-x;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9403-x.pdf. See reply [642].

McHugh:2010:SKK

[42] Conor McHugh. Self-knowledge and the KK principle. Synthese, 173 (3):231–257, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9404-9;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9404-9.pdf.

Poggiolesi:2010:DCO

[43] Francesca Poggiolesi. Display calculi and other modal calculi: a comparison. Synthese, 173(3):259–279, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9425-4;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9425-
4.pdf.

Baave:2010:DPT

[44] Arvid Båve. Deflationism and the primary truth bearer. Synthese, 173 (3):281–297, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9428-1;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9428-1.pdf.

Clarke:2010:TRP

[45] Stephen Clarke. Transcendental realisms in the philosophy of science: on Bhaskar and Cartwright. Synthese, 173(3):299–315, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9427-2;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9427-2.pdf.

Montminy:2010:TCF

[46] Martin Montminy. Two contextualist fallacies. Synthese, 173(3):317–333, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9426-3;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9426-3.pdf.

Morrison:2010:JHC

[47] Joe Morrison. Just how controversial is evidential holism? Synthese, 173(3):335–352, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9440-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9440-5.pdf.

Smith:2010:ECM

[48] Sheldon R. Smith. Elementary classical mechanics and the principle of the composition of causes. Synthese, 173(3):353–373, April 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

008-9442-3;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9442-3.pdf.

vanRooij:2010:I

[49] Robert van Rooij.  Introduction.  Synthese, 174(1):1–3, May

2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9682-x;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9682-x.pdf.

Gaifman:2010:VTC

[50] Haim Gaifman. Vagueness, tolerance and contextual logic. Synthese, 174(1):5–46, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9683-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9683-9.pdf.

Egre:2010:VUD

[51] Paul Égré and Denis Bonnay. Vagueness, uncertainty and degrees of clarity. Synthese, 174(1):47–78, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9684-8;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9684-
8.pdf.

Kennedy:2010:CCC

[52] Christopher Kennedy and Louise McNally. Color, context, and compositionality. Synthese, 174(1):79–98, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9685-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9685-
7.pdf.

Chierchia:2010:MNV

[53] Gennaro Chierchia. Mass nouns, vagueness and semantic variation. Synthese, 174(1):99–149, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9686-6;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9686-6.pdf.

Sassoon:2010:MTL

[54] Galit Weidman Sassoon. Measurement theory in linguistics. Synthese, 174(1):151–180, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9687-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9687-5.pdf.

Betti:2010:I

[55] Arianna Betti and Willem R. de Jong. Introduction. Synthese, 174 (2):181–183, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9416-5;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9416-5.pdf.

deJong:2010:CMS

[56] Willem R. de Jong and Arianna Betti. The classical model of science: a millennia-old model of scientific rationality. Synthese, 174(2):185–203, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9417-4;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9417-4.pdf.

Martijn:2010:POP

[57] Marije Martijn. Proclus on the order of philosophy of nature. Synthese, 174(2):205–223, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9418-3;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9418-3.pdf.

Cantu:2010:APR

[58] Paola Cantù. Aristotle’s prohibition rule on kind-crossing and the definition of mathematics as a science of quantities. Synthese, 174(2):225–235, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9419-2;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9419-2.pdf.

deJong:2010:ASD

[59] Willem R. de Jong. The analytic-synthetic distinction and the classical model of science: Kant, Bolzano and Frege. Synthese, 174(2):237–261, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
008-9420-9;   http://link.springer.com/content/pdf/10.1007/
s11229-008-9420-9.pdf.

Lapointe:2010:BPK

[60] Sandra Lapointe. Bolzano, a priori knowledge, and the classical model of science. Synthese, 174(2):263–281, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9421-8;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9421-
8.pdf.

Korte:2010:FBL

[61] Tapio Korte. Frege’s Begriffsschrift as a lingua characteristica. Synthese, 174(2):283–294, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9422-7;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9422-7.pdf.

Betti:2010:LCU

[62] Arianna Betti. Leśniewski’s characteristica universalis. Synthese, 174 (2):295–314, May 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9423-6;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9423-6.pdf.

Hetherington:2010:EEJ

[63] Stephen Hetherington. Elusive epistemological justification. Synthese, 174(3):315–330, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-008-9451-2;   http://link.springer.com/content/
pdf/10.1007/s11229-008-9451-2.pdf.

Sprenger:2010:PRS

[64] Jan Sprenger. Probability, rational single-case decisions and the Monty Hall Problem. Synthese, 174(3):331–340, June 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-008-9455-y;
 http://link.springer.com/content/pdf/10.1007/s11229-008-9455-
y.pdf.

Contessa:2010:MTT

[65] Gabriele Contessa. Modal truthmakers and two varieties of actualism. Synthese, 174(3):341–353, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-008-9456-x;   http://link.springer.com/
content/pdf/10.1007/s11229-008-9456-x.pdf.

Crisp:2010:DI

[66] Thomas M. Crisp. A dilemma for internalism? Synthese, 174(3):355–366, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9457-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9457-4.pdf.

Andreas:2010:MVS

[67] Holger Andreas. A modal view of the semantics of theoretical sentences. Synthese, 174(3):367–383, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9458-3;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9458-
3.pdf.

Schulz:2010:EMI

[68] Moritz Schulz. Epistemic modals and informational consequence. Synthese, 174(3):385–395, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9461-8;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9461-8.pdf.

Karlander:2010:SBM

[69] Karl Karlander and Levi Spectre. Sleeping Beauty meets Monday. Synthese, 174(3):397–412, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9464-5;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9464-5.pdf.

Newman:2010:BSR

[70] Mark Newman. Beyond structural realism: pluralist criteria for theory evaluation. Synthese, 174(3):413–443, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9463-6;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9463-
6.pdf.

Millar:2010:PT

[71] Boyd Millar. Peacocke’s trees. Synthese, 174(3):445–461, June 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9465-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9465-4.pdf.

Bai:2010:GEW

[72] Tongdong Bai. Guest Editor’s words. Synthese, 175(1):1–2, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9536-6;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9536-6.pdf.

Jiang:2010:SAP

[73] Yi Jiang and Tongdong Bai. Studies in analytic philosophy in China. Synthese, 175(1):3–12, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9534-8;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9534-8.pdf.

Ye:2010:WAR

[74] Feng Ye. What anti-realism in philosophy of mathematics must offer. Synthese, 175(1):13–31, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9535-7;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9535-7.pdf.

Liu:2010:GPP

[75] Xiaoli Liu. Gödel’s philosophical program and Husserl’s phenomenology. Synthese, 175(1):33–45, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9532-x;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9532-x.pdf.

Zhou:2010:FSL

[76] Beihai Zhou and Yi Mao. Four semantic layers of common nouns. Synthese, 175(1):47–68, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9533-9;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9533-9.pdf.

Liu:2010:WLP

[77] Fenrong Liu. Von Wright’s “The Logic of Preference” revisited. Synthese, 175(1):69–88, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9530-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9530-z.pdf.

Zhu:2010:PIA

[78] Jing Zhu. On the principle of intention agglomeration. Synthese, 175 (1):89–99, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9531-y;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9531-y.pdf.

Tang:2010:CNM

[79] Refeng Tang. Conceptualism and the new myth of the given. Synthese, 175(1):101–122, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9529-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9529-5.pdf.

Frigerio:2010:OGM

[80] Aldo Frigerio, Alessandro Giordani, and Luca Mari. Outline of a general model of measurement. Synthese, 175(2):123–149, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9466-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9466-3.pdf.

Baltimore:2010:DPP

[81] Joseph A. Baltimore. Defending the piggyback principle against Shapiro and Sober’s empirical approach. Synthese, 175(2):151–168, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9467-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9467-2.pdf.

Glymour:2010:ACS

[82] Clark Glymour, David Danks, Bruce Glymour, Frederick Eberhardt, Joseph Ramsey, Richard Scheines, Peter Spirtes, Choh Man Teng, and Jiji Zhang. Actual causation: a stone soup essay. Synthese, 175 (2):169–192, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9497-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9497-9.pdf.

Kronfeldner:2010:DBH

[83] Maria E. Kronfeldner. Darwinian ‘blind’ hypothesis formation revisited. Synthese, 175(2):193–218, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9498-8;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9498-8.pdf.

Debes:2010:WEL

[84] Remy Debes. Which empathy? Limitations in the mirrored “understanding” of emotion. Synthese, 175(2):219–239, July 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9499-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9499-
7.pdf.

Yudell:2010:MSS

[85] Zanja Yudell. Melia and Saatsi on structural realism. Synthese, 175 (2):241–253, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9500-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9500-5.pdf.

Mumma:2010:PPE

[86] John Mumma. Proofs, pictures, and Euclid. Synthese, 175(2):255–287, July 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9509-9;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9509-9.pdf.

Fennell:2010:DRS

[87] Damien Fennell and Nancy Cartwright. Does Roush show that evidence should be probable? Synthese, 175(3):289–310, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9510-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9510-3.pdf.

Maurin:2010:TTB

[88] Anna-Sofia Maurin. Trope theory and the Bradley regress. Synthese, 175 (3):311–326, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9511-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9511-2.pdf.

Betz:2010:PPC

[89] Gregor Betz. Petitio principii and circular argumentation as seen from a theory of dialectical structures. Synthese, 175(3):327–349, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9512-1;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9512-1.pdf.

Kolman:2010:CNP

[90] Vojtech Kolman. Continuum, name and paradox. Synthese, 175(3): 351–367, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9527-7;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9527-7.pdf.

Lewis:2010:CSL

[91] Peter J. Lewis. Credence and self-location. Synthese, 175(3):369–382, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9528-6;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9528-6.pdf.

Stokke:2010:ISS

[92] A. Stokke. Intention-sensitive semantics. Synthese, 175(3):383–404, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9537-5;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9537-5.pdf.

Ereshefsky:2010:DSS

[93] Marc Ereshefsky. Darwin’s solution to the species problem. Synthese, 175 (3):405–425, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9538-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9538-4.pdf.

Clarke:2010:RPM

[94] Roger Clarke. “The ravens paradox” is a misnomer. Synthese, 175 (3):427–440, August 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9560-6;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9560-6.pdf.

Weirich:2010:I

[95] Paul Weirich. Introduction. Synthese, 176(1):1–3, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9479-y;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9479-y.pdf.

Morton:2010:HBR

[96] Adam Morton. Human bounds: rationality for our species. Synthese, 176 (1):5–21, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9481-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9481-4.pdf.

Bermudez:2010:PRD

[97] José Luis Bermúdez. Pitfalls for realistic decision theory: an illustration from sequential choice. Synthese, 176(1):23–40, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9482-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9482-3.pdf.

Kaplan:2010:DMP

[98] Mark Kaplan. In defense of modest probabilism. Synthese, 176(1): 41–55, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9483-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9483-2.pdf.

Pollock:2010:RBA

[99] John L. Pollock. A resource-bounded agent addresses the Newcomb Problem. Synthese, 176(1):57–82, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9484-1;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9484-
1.pdf.

Weirich:2010:UF

[100] Paul Weirich. Utility and framing. Synthese, 176(1):83–103, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9485-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9485-0.pdf.

Carlson:2010:SNT

[101] Lisa J. Carlson and Raymond Dacey. Social norms and the traditional deterrence game. Synthese, 176(1):105–123, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9486-z;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9486-z.pdf.

Bicchieri:2010:MMC

[102] Cristina Bicchieri, Azi Lev-On, and Alex Chavez. The medium or the message? Communication relevance and richness in trust games. Synthese, 176(1):125–147, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9487-y;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9487-y.pdf.

Weirich:2010:C

[103] Paul Weirich. The contributors. Synthese, 176(1):149–150, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9480-5;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9480-5.pdf.

Moffett:2010:IPA

[104] Marc A. Moffett. Introduction: Proceedings of the 36th annual meeting of the society for exact philosophy. Synthese, 176(2):151–152, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9488-x;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9488-x.pdf.

Chihara:2010:NDN

[105] Charles Chihara. New directions for nominalist philosophers of mathematics. Synthese, 176(2):153–175, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9489-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9489-
9.pdf.

Dean:2010:KPE

[106] W. Dean and H. Kurokawa. From the knowability paradox to the existence of proofs. Synthese, 176(2):177–225, September 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9490-3;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9490-
3.pdf.

Forbes:2010:IVE

[107] Graeme Forbes. Intensional verbs in event semantics. Synthese, 176(2): 227–242, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9491-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9491-2.pdf.

Maier:2010:WFS

[108] Frederick Maier and Donald Nute. Well-founded semantics for defeasible logic. Synthese, 176(2):243–274, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9492-1;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9492-
1.pdf.

Pelczar:2010:PEP

[109] Michael Pelczar. Presentism, eternalism, and phenomenal change. Synthese, 176(2):275–290, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9493-0;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9493-0.pdf.

Willer:2010:NSR

[110] Malte Willer. New surprises for the Ramsey test. Synthese, 176(2): 291–309, September 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9494-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9494-z.pdf.

Andreas:2010:NAE

[111] Holger Andreas. New account of empirical claims in structuralism. Synthese, 176(3):311–332, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9561-5;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9561-5.pdf.

Chandler:2010:TSB

[112] Jake Chandler. The transmission of support: a Bayesian re-analysis. Synthese, 176(3):333–343, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9570-4;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9570-4.pdf.

Sauchelli:2010:CPW

[113] Andrea Sauchelli. Concrete possible worlds and counterfactual conditionals: Lewis versus Williamson on modal knowledge. Synthese, 176 (3):345–359, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9571-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9571-3.pdf.

Steglich-Petersen:2010:LEN

[114] Asbjørn Steglich-Petersen. Luck as an epistemic notion. Synthese, 176 (3):361–377, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9569-x;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9569-x.pdf.

Rueger:2010:HLR

[115] Alexander Rueger and Patrick McGivern. Hierarchies and levels of reality. Synthese, 176(3):379–397, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9572-2;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9572-
2.pdf.

Michaelian:2010:DGE

[116] Kourken Michaelian. In defence of gullibility: the epistemology of testimony and the psychology of deception detection. Synthese, 176(3): 399–427, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9573-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9573-1.pdf.

Tsou:2010:PAA

[117] Jonathan Y. Tsou. Putnam’s account of apriority and scientific change: its historical and contemporary interest. Synthese, 176(3):429–445, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9574-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9574-0.pdf.

Vahid:2010:RBE

[118] Hamid Vahid. Rationalizing beliefs: evidential vs. pragmatic reasons. Synthese, 176(3):447–462, October 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9575-z;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9575-z.pdf.

Kung:2010:HNR

[119] Peter Kung. On having no reason: dogmatism and Bayesian confirmation. Synthese, 177(1):1–17, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9578-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9578-
9.pdf.

Brigandt:2010:EGC

[120] Ingo Brigandt. The epistemic goal of a concept: accounting for the rationality of semantic change and variation. Synthese, 177(1):19–40, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9623-8;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9623-8.pdf.

Cresto:2010:BCA

[121] Eleonora Cresto. Belief and contextual acceptance. Synthese, 177(1): 41–66, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9637-2;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9637-2.pdf.

Finlay:2010:WOP

[122] Stephen Finlay. What ought probably means, and why you can’t detach it. Synthese, 177(1):67–89, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9640-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9640-

7.pdf. See correction [2084].

Nanay:2010:PTT

[123] Bence Nanay. Population thinking as trope nominalism. Synthese, 177 (1):91–109, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9641-6;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9641-6.pdf.

Newman:2010:NMA

[124] Mark Newman. The no-miracles argument, reliabilism, and a methodological version of the generality problem. Synthese, 177(1):111– 138, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9642-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9642-5.pdf.

Rowbottom:2010:CAH

[125] Darrell P. Rowbottom. Corroboration and auxiliary hypotheses: Duhem’s thesis revisited. Synthese, 177(1):139–149, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9643-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9643-4.pdf.

Sullivan:2010:REM

[126] Jacqueline A. Sullivan. Realization, explanation and the mindbody relation. Synthese, 177(2):151–164, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9839-7;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9839-
7.pdf.

Gillett:2010:MBS

[127] Carl Gillett. Moving beyond the subset model of realization: The problem of qualitative distinctness in the metaphysics of science. Synthese, 177(2):165–192, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9840-1;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9840-1.pdf.

Polger:2010:MER

[128] Thomas W. Polger. Mechanisms and explanatory realization relations. Synthese, 177(2):193–212, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-010-9841-0;   http://link.springer.
com/content/pdf/10.1007/s11229-010-9841-0.pdf.

Huneman:2010:TER

[129] Philippe Huneman. Topological explanations and robustness in biological sciences. Synthese, 177(2):213–245, November 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9842-z;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9842-
z.pdf.

Bickle:2010:LDC

[130] John Bickle. Has the last decade of challenges to the multiple realization argument provided aid and comfort to psychoneural reductionists? Synthese, 177(2):247–260, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9843-y;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9843-
y.pdf.

Sullivan:2010:RSM

[131] Jacqueline A. Sullivan. Reconsidering ‘spatial memory’ and the Morris water maze. Synthese, 177(2):261–283, November 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9849-5;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9849-
5.pdf.

Walter:2010:CEP

[132] Sven Walter. Cognitive extension: the parity argument, functionalism, and the mark of the cognitive. Synthese, 177(2):285–300, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9844-x;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9844-x.pdf.

Fehr:2010:SRP

[133] Carla Fehr and Kathryn S. Plaisance. Socially relevant philosophy of science: an introduction. Synthese, 177(3):301–316, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9855-7;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9855-7.pdf.

Douglas:2010:EPA

[134] Heather Douglas. Engagement for progress: applied philosophy of science in context. Synthese, 177(3):317–335, December 2010. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9787-2;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9787-
2.pdf.

Richardson:2010:FPS

[135] Sarah S. Richardson. Feminist philosophy of science: history, contributions, and challenges. Synthese, 177(3):337–362, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9791-6;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9791-6.pdf.

Gannett:2010:QAU

[136] Lisa Gannett. Questions asked and unasked: how by worrying less about the ‘ really real’ philosophers of science might better contribute to debates about genetics and race. Synthese, 177(3):363–385, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9788-1;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9788-1.pdf.

Grasswick:2010:SLC

[137] Heidi E. Grasswick. Scientific and lay communities: earning epistemic trust through knowledge sharing. Synthese, 177(3):387–409, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9789-0;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9789-0.pdf.

Whyte:2010:TEP

[138] Kyle Powys Whyte and Robert P. Crease. Trust, expertise, and the philosophy of science. Synthese, 177(3):411–425, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9786-3;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9786-3.pdf.

Reiss:2010:FMP

[139] Julian Reiss. In favour of a Millian proposal to reform biomedical research. Synthese, 177(3):427–447, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9790-7;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9790-
7.pdf.

Shrader-Frechette:2010:CAS

[140] Kristin Shrader-Frechette. Conceptual analysis and special-interest science: toxicology and the case of Edward Calabrese. Synthese, 177(3): 449–469, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9792-5;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9792-5.pdf.

Tuana:2010:LEA

[141] Nancy Tuana. Leading with ethics, aiming for policy: new opportunities for philosophy of science. Synthese, 177(3):471–492, December 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9793-4;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9793-4.pdf.

Rowbottom:2011:SRP

[142] Darrell P. Rowbottom and Otávio Bueno. Stance and rationality: a perspective. Synthese, 178(1):1–5, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9526-8;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9526-
8.pdf.

Rowbottom:2011:HCI

[143] Darrell P. Rowbottom and Otávio Bueno. How to change it: modes of engagement, rationality, and stance voluntarism. Synthese, 178(1): 7–17, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9521-0;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9521-0.pdf.

Steup:2011:EMV

[144] Matthias Steup. Empiricism, metaphysics, and voluntarism. Synthese, 178(1):19–26, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9518-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9518-8.pdf.

Baumann:2011:ESP

[145] Peter Baumann. Empiricism, stances, and the problem of voluntarism. Synthese, 178(1):27–36, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9519-7;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9519-7.pdf.

Chakravartty:2011:PAV

[146] Anjan Chakravartty. A puzzle about voluntarism about rational epistemic stances. Synthese, 178(1):37–48, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9516-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9516-
x.pdf.

Teller:2011:LLV

[147] Paul Teller. Learning to live with voluntarism. Synthese, 178(1): 49–66, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9517-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9517-9.pdf.

Williamson:2011:OBB

[148] Jon Williamson. Objective Bayesianism, Bayesian conditionalisation and voluntarism. Synthese, 178(1):67–85, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9515-y;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9515-
y.pdf.

Ladyman:2011:SSE

[149] James Ladyman. The scientistic stance: the empirical and materialist stances reconciled. Synthese, 178(1):87–98, January 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9513-0;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9513-
0.pdf.

Lowe:2011:RM

[150] E. J. Lowe. The rationality of metaphysics. Synthese, 178(1):99–109, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9514-z;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9514-z.pdf.

Rowbottom:2011:SPR

[151] Darrell P. Rowbottom. Stances and paradigms: a reflection. Synthese, 178(1):111–119, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9524-x;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9524-x.pdf.

Ratcliffe:2011:SFP

[152] Matthew Ratcliffe. Stance, feeling and phenomenology. Synthese, 178 (1):121–130, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9525-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9525-9.pdf.

Jones:2011:BMW

[153] Ward E. Jones. Being moved by a way the world is not. Synthese, 178 (1):131–141, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9522-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9522-z.pdf.

Richardson:2011:WTI

[154] Alan Richardson. But what then am i, this inexhaustible, unfathomable historical self? Or, upon what ground may one commit empiricism? Synthese, 178(1):143–154, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9523-y;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9523-
y.pdf.

vanFraassen:2011:SR

[155] Bas C. van Fraassen. On stance and rationality. Synthese, 178(1): 155–169, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9520-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9520-1.pdf.

Branch:2011:I

[156] Glenn Branch. Introduction. Synthese, 178(2):171–176, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9548-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9548-2.pdf.

Pennock:2011:CPT

[157] Robert T. Pennock. Can’t philosophers tell the difference between science and religion?: Demarcation revisited. Synthese, 178(2):177–206, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9547-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9547-3.pdf.

Wilkins:2011:CR

[158] John S. Wilkins. Are creationists rational? Synthese, 178(2):207– 218, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9544-6;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9544-6.pdf.

Smith:2011:FBK

[159] Kelly C. Smith. Foiling the black Knight. Synthese, 178(2):219–235, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9545-5;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9545-5.pdf.

Elsberry:2011:ITE

[160] Wesley Elsberry and Jeffrey Shallit. Information theory, evolutionary computation, and Dembski’s “complex specified information”. Synthese, 178(2):237–270, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9542-8;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9542-8.pdf.

Weber:2011:DD

[161] Bruce H. Weber. Design and its discontents. Synthese, 178(2):271– 289, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9543-7;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9543-7.pdf.

Sarkar:2011:SQI

[162] Sahotra Sarkar. The science question in intelligent design. Synthese, 178 (2):291–305, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9540-x;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9540-x.pdf.

Shanks:2011:IDT

[163] Niall Shanks and Keith Green. Intelligent design in theological perspective. Synthese, 178(2):307–330, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9541-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9541-
9.pdf.

Forrest:2011:NEI

[164] Barbara Forrest. The non-epistemology of intelligent design: its implications for public policy. Synthese, 178(2):331–379, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9539-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9539-3.pdf.

Fetzer:2011:EAG

[165] James H. Fetzer. Evolution and atheism: Has Griffin reconciled science and religion? Synthese, 178(2):381–396, January 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9546-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9546-4.pdf.

Winther:2011:PWS

[166] Rasmus Grønfeldt Winther. Part-whole science. Synthese, 178(3): 397–427, February 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9647-0;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9647-0.pdf.

Laraudogoitia:2011:ISP

[167] J. P. Laraudogoitia. The inverse spaceship paradox. Synthese, 178(3): 429–435, February 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9649-y;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9649-y.pdf.

Pelling:2011:CHE

[168] Charlie Pelling. Characterizing hallucination epistemically. Synthese, 178 (3):437–459, February 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9651-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9651-4.pdf.

Paglieri:2011:EP

[169] Fabio Paglieri and John Woods. Enthymematic parsimony. Synthese, 178 (3):461–501, February 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9652-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9652-3.pdf.

Wallin:2011:EBE

[170] Annika Wallin. Is egocentric bias evidence for simulation theory? Synthese, 178(3):503–514, February 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9653-2;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9653-2.pdf.

Podlaskowski:2011:IEN

[171] Adam C. Podlaskowski and Joshua A. Smith. Infinitism and epistemic normativity. Synthese, 178(3):515–527, February 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9654-1;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9654-
1.pdf.

Gerken:2011:WA

[172] Mikkel Gerken. Warrant and action. Synthese, 178(3):529–547, February 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9655-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9655-0.pdf.

Pedersen:2011:P

[173] Esther Oluffa Pedersen, Claus Festersen, Steen Brock, and Stig Andur Pedersen. Preface. Synthese, 179(1):1–5, March 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9633-6;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9633-
6.pdf.

Pedersen:2011:A

[174] Esther Oluffa Pedersen, Claus Festersen, Steen Brock, and Stig Andur Pedersen. Acknowledgements. Synthese, 179(1):7, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9632-7;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9632-7.pdf.

Krois:2011:PSL

[175] John Michael Krois. The priority of “symbolism” over language in Cassirer’s philosophy. Synthese, 179(1):9–20, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9636-3;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9636-
3.pdf.

Schmitz-Rigal:2011:SAP

[176] Christiane Schmitz-Rigal. Science and art: physics as a symbolic formation. Synthese, 179(1):21–41, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9630-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9630-
9.pdf.

Bundgaard:2011:GAI

[177] Peer F. Bundgaard. The grammar of aesthetic intuition: on Ernst Cassirer’s concept of symbolic form in the visual arts. Synthese, 179 (1):43–57, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9631-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9631-8.pdf.

Schwemmer:2011:EFT

[178] Oswald Schwemmer. Event and form: two themes in the Davosdebate between Martin Heidegger and Ernst Cassirer. Synthese, 179 (1):59–73, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/s11229-009-9628-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9628-3.pdf.

Hoel:2011:TDD

[179] Aud Sissel Hoel. Thinking “difference” differently: Cassirer versus Derrida on symbolic mediation. Synthese, 179(1):75–91, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9629-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9629-2.pdf.

Brock:2011:RRC

[180] Steen Brock. A resolute reading of Cassirer’s anthropology. Synthese, 179(1):93–113, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9626-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9626-5.pdf.

Orth:2011:ECC

[181] Ernst Wolfgang Orth. Ernst Cassirer as cultural scientist. Synthese, 179 (1):115–134, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9627-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9627-4.pdf.

Renz:2011:PCM

[182] Ursula Renz. From philosophy to criticism of myth: Cassirer’s concept of myth. Synthese, 179(1):135–152, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9624-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9624-
7.pdf.

Plumacher:2011:PRC

[183] Martina Plümacher. Philosophical research on cognition. Synthese, 179 (1):153–167, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9625-6;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9625-6.pdf.

Stjernfelt:2011:SAC

[184] Frederik Stjernfelt. Simple animals and complex biology: Von Uexküll’s two-fold influence on Cassirer’s philosophy. Synthese, 179(1):169– 186, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9634-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9634-5.pdf.

Ikonen:2011:CCC

[185] Sirkku Ikonen. Cassirer’s critique of culture. Synthese, 179(1):187– 202, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9635-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9635-4.pdf.

vanBenthem:2011:I

[186] Johan van Benthem, Theo Kuipers, and Henk Visser. Introduction. Synthese, 179(2):203–206, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9774-7;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9774-7.pdf.

Seuren:2011:HRE

[187] Pieter A. M. Seuren. How I remember Evert Beth. Synthese, 179(2): 207–210, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9777-4;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9777-4.pdf.

Parikh:2011:BDI

[188] Rohit Parikh. Beth definability, interpolation and language splitting. Synthese, 179(2):211–221, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9778-3;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9778-3.pdf.

Degremont:2011:DWC

[189] Cédric Dégremont and Jonathan Zvesper. Dynamics we can believe in: a view from the Amsterdam School on the centenary of Evert Willem Beth. Synthese, 179(2):223–238, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9779-2;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9779-
2.pdf.

Schulz:2011:IYW

[190] Katrin Schulz. “If you’d wiggled A, then B would’ve changed”. Synthese, 179(2):239–251, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9780-9;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9780-9.pdf.

Gabbay:2011:RIT

[191] D. M. Gabbay. Reactive intuitionistic tableaux. Synthese, 179(2): 253–269, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9781-8;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9781-8.pdf.

Dieks:2011:WBP

[192] Dennis Dieks. E. W. Beth as a philosopher of physics. Synthese, 179 (2):271–284, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9782-7;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9782-7.pdf.

Baltag:2011:QLD

[193] Alexandru Baltag and Sonja Smets. Quantum logic as a dynamic logic. Synthese, 179(2):285–306, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9783-6;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9783-
6.pdf.

Schurz:2011:SCB

[194] Gerhard Schurz. Structural correspondence between theories and convergence to truth. Synthese, 179(2):307–320, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9784-5;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9784-
5.pdf.

Dizadji-Bahmani:2011:CRB

[195] F. Dizadji-Bahmani, R. Frigg, and S. Hartmann. Confirmation and reduction: a Bayesian account. Synthese, 179(2):321–338, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

010-9775-6;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9775-6.pdf.

Leitgeb:2011:LGP

[196] Hannes Leitgeb. Logic in general philosophy of science: old things and new things. Synthese, 179(2):339–350, March 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9776-5;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9776-
5.pdf.

Dodd:2011:QMT

[197] Dylan Dodd. Quasi-miracles, typicality, and counterfactuals. Synthese, 179(3):351–360, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9656-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9656-z.pdf.

McArthur:2011:DTC

[198] Daniel James McArthur. Discovery, theory change and structural realism. Synthese, 179(3):361–376, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9672-z;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9672-
z.pdf.

Walton:2011:DRI

[199] Douglas Walton. Defeasible reasoning and informal fallacies. Synthese, 179(3):377–407, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9657-y;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9657-y.pdf.

Kelp:2011:DVE

[200] Christoph Kelp. In defence of virtue epistemology. Synthese, 179(3): 409–433, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9681-y;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9681-y.pdf.

Landry:2011:HSA

[201] Elaine Landry. How to be a structuralist all the way down. Synthese, 179(3):435–454, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9691-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9691-9.pdf.

Nanay:2011:RR

[202] Bence Nanay. Replication without replicators. Synthese, 179(3):455–477, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9702-x;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9702-x.pdf.

Littlejohn:2011:EAA

[203] Clayton Mitchell Littlejohn. Evidence and armchair access. Synthese, 179(3):479–500, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9703-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9703-9.pdf.

Gaifman:2011:EVT

[204] Haim Gaifman. Erratum to: Vagueness, tolerance and contextual logic. Synthese, 179(3):501–502, April 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9746-y;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9746-y.pdf.

Frigg:2011:P

[205] Roman Frigg, Stephan Hartmann, and Cyrille Imbert. Preface. Synthese, 180(1):1–2, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9562-4;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9562-4.pdf.

Psillos:2011:LAR

[206] Stathis Psillos. Living with the abstract: realism and models. Synthese, 180(1):3–17, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9563-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9563-3.pdf.

Pincock:2011:MR

[207] Christopher Pincock. Modeling reality. Synthese, 180(1):19–32, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

009-9564-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9564-2.pdf.

Bokulich:2011:HSM

[208] Alisa Bokulich. How scientific models can explain. Synthese, 180 (1):33–45, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9565-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9565-1.pdf.

Maki:2011:MLT

[209] Uskali Mäki. Models and the locus of their truth. Synthese, 180(1):47–63, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9566-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9566-0.pdf.

Sprenger:2011:SPM

[210] Jan Sprenger. Science without (parametric) models: the case of bootstrap resampling. Synthese, 180(1):65–76, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9567-z;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9567-
z.pdf.

Frigg:2011:PSH

[211] Roman Frigg and Julian Reiss. The philosophy of simulation: hot new issues or same old stew? Synthese, 180(1):77, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9577-x;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9577-x.pdf.

Schurz:2011:EIS

[212] Gerhard Schurz and Ioannis Votsis. Editorial introduction to scientific realism quo vadis? Theories, structures, underdetermination and reference. Synthese, 180(2):79–85, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9610-0;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9610-
0.pdf.

Ladyman:2011:SRV

[213] James Ladyman. Structural realism versus standard scientific realism: the case of phlogiston and dephlogisticated air. Synthese, 180(2):87–101, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9607-8;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9607-8.pdf.

Schurz:2011:SCI

[214] Gerhard Schurz. Structural correspondence, indirect reference, and partial truth: phlogiston theory and Newtonian mechanics. Synthese, 180 (2):103–120, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9608-7;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9608-7.pdf.

Votsis:2011:SIP

[215] Ioannis Votsis. Saving the intuitions: polylithic reference. Synthese, 180(2):121–137, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9601-1;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9601-1.pdf.

Fahrbach:2011:HGS

[216] Ludwig Fahrbach. How the growth of science ends theory change. Synthese, 180(2):139–155, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9602-0;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9602-0.pdf.

Worrall:2011:URE

[217] John Worrall. Underdetermination, realism and empirical equivalence. Synthese, 180(2):157–172, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9599-4;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9599-4.pdf.

Hoyningen-Huene:2011:RMA

[218] Paul Hoyningen-Huene. Reconsidering the miracle argument on the supposition of transient underdetermination. Synthese, 180(2):173–187, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

009-9600-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9600-2.pdf.

Carrier:2011:UET

[219] Martin Carrier. Underdetermination as an epistemological test tube: expounding hidden values of the scientific community. Synthese, 180 (2):189–204, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9597-6;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9597-6.pdf.

French:2011:MUW

[220] Steven French. Metaphysical underdetermination: why worry? Synthese, 180(2):205–221, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9598-5;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9598-5.pdf.

Muller:2011:WW

[221] F. A. Muller. Withering away, weakly. Synthese, 180(2):223–233, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9609-6;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9609-6.pdf.

Lyre:2011:SUP

[222] Holger Lyre. Is structural underdetermination possible? Synthese, 180 (2):235–247, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9603-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9603-z.pdf.

Friedman:2011:CTT

[223] Michael Friedman. Carnap on theoretical terms: structuralism without metaphysics. Synthese, 180(2):249–263, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9604-y;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9604-
y.pdf.

Leitgeb:2011:NLC

[224] Hannes Leitgeb. New life for Carnap’s Aufbau? Synthese, 180(2): 265–299, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9605-x;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9605-x.pdf.

Psillos:2011:CRF

[225] Stathis Psillos. Choosing the realist framework. Synthese, 180(2): 301–316, May 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9606-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9606-9.pdf.

Bobzien:2011:DTH

[226] Susanne Bobzien. In defense of true higher-order vagueness. Synthese, 180(3):317–335, June 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9704-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9704-8.pdf.

Roberts:2011:HGD

[227] Bryan W. Roberts. How Galileo dropped the ball and Fermat picked it up. Synthese, 180(3):337–356, June 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9705-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9705-
7.pdf.

Bayer:2011:RAD

[228] Benjamin Bayer. A role for abstractionism in a direct realist foundationalism. Synthese, 180(3):357–389, June 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9707-5;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9707-
5.pdf.

vanderSchaar:2011:CAF

[229] Maria van der Schaar. The cognitive act and the first-person perspective: an epistemology for constructive type theory. Synthese, 180(3):391–417, June 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9708-4;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9708-4.pdf.

Campos:2011:DBP

[230] Daniel G. Campos. On the distinction between Peirce’s abduction and Lipton’s inference to the best explanation. Synthese, 180(3):419–442, June 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9709-3;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9709-3.pdf.

Mandik:2011:SN

[231] Pete Mandik. Supervenience and neuroscience. Synthese, 180(3):443–463, June 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9712-8;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9712-8.pdf.

Irzik:2011:ISS

[232] Gürol Irzik and Elliott Sober. Introduction to the Synthese special issue on Hans Reichenbach, Istanbul, and experience and prediction. Synthese, 181(1):1–2, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9596-7;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9596-7.pdf.

Sober:2011:RCU

[233] Elliott Sober. Reichenbach’s cubical universe and the problem of the external world. Synthese, 181(1):3–21, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9593-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9593-
x.pdf.

Psillos:2011:RAS

[234] Stathis Psillos. On Reichenbach’s argument for scientific realism. Synthese, 181(1):23–40, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9594-9;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9594-9.pdf.

Padovani:2011:RRP

[235] Flavia Padovani. Relativizing the relativized a priori: Reichenbach’s axioms of coordination divided. Synthese, 181(1):41–62, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9590-0;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9590-0.pdf.

Berkovski:2011:RWA

[236] Sandy Berkovski. Reichenbach and Weyl on apriority and mathematical applicability. Synthese, 181(1):63–77, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9591-z;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9591-
z.pdf.

Bagce:2011:RRP

[237] Samet Bagce. Reichenbach on the relative a priori and the context of discovery/justification distinction. Synthese, 181(1):79–93, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9588-7;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9588-7.pdf.

Galavotti:2011:HRI

[238] Maria Carla Galavotti. On Hans Reichenbach’s inductivism. Synthese, 181(1):95–111, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9589-6;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9589-6.pdf.

Peijnenburg:2011:GLR

[239] Jeanne Peijnenburg and David Atkinson. Grounds and limits: Reichenbach and foundationalist epistemology. Synthese, 181(1):113–124, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9586-9;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9586-9.pdf.

Eberhardt:2011:RSP

[240] Frederick Eberhardt. Reliability via synthetic a priori: Reichenbach’s doctoral thesis on probability. Synthese, 181(1):125–136, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9587-8;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9587-8.pdf.

Stadler:2011:REP

[241] Friedrich Stadler. The road to experience and prediction from within: Hans Reichenbach’s scientific correspondence from Berlin to Istanbul. Synthese, 181(1):137–155, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9595-8;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9595-8.pdf.

Irzik:2011:HRI

[242] Gürol Irzik. Hans Reichenbach in Istanbul. Synthese, 181(1):157–180, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9592-y;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9592-y.pdf.

Moffett:2011:AMS

[243] Marc Moffett and Greg Ray. The 37th annual meeting of the society for exact philosophy. Synthese, 181(2):181–184, July 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9795-2;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9795-
2.pdf.

Bandyoapdhyay:2011:LSP

[244] Prasanta S. Bandyoapdhyay, Davin Nelson, Mark Greenwood, Gordon Brittan, and Jesse Berwald. The logic of Simpson’s paradox. Synthese, 181(2):185–208, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9797-0;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9797-0.pdf.

Hiddleston:2011:RMM

[245] Eric Hiddleston. Reductionism and the micro–macro mirroring thesis. Synthese, 181(2):209–226, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9798-z;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9798-z.pdf.

Hitchcock:2011:TCC

[246] Christopher Hitchcock. Trumping and contrastive causation. Synthese, 181(2):227–240, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9799-y;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9799-y.pdf.

Hoffmann:2011:TKP

[247] Glen Hoffmann. Two kinds of a priori infallibility. Synthese, 181 (2):241–253, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9800-9;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9800-9.pdf.

Johns:2011:SOD

[248] Richard Johns. Self-organisation in dynamical systems: a limiting result. Synthese, 181(2):255–275, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9801-8;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9801-8.pdf.

Marti:2011:GTR

[249] Genoveva Martí and José Martínez-Fernández. General terms, rigidity and the trivialization problem. Synthese, 181(2):277–293, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9802-7;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9802-7.pdf.

Schotch:2011:WT

[250] Peter K. Schotch and Gillman Payette. Worlds and times. Synthese, 181(2):295–315, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9803-6;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9803-6.pdf.

Pollock:2011:RDA

[251] John L. Pollock. Reasoning defeasibly about probabilities. Synthese, 181(2):317–352, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9804-5;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9804-5.pdf.

Sher:2011:LMW

[252] Gila Sher. Is logic in the mind or in the world? Synthese, 181(2):353–365, July 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9796-1;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9796-1.pdf.

Peterson:2011:QBR

[253] Daniel Peterson. Qeauty and the books: a response to Lewis’s quantum Sleeping Beauty problem. Synthese, 181(3):367–374, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9715-5;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9715-5.pdf.

Ward:2011:KWW

[254] Dave Ward, Tom Roberts, and Andy Clark. Knowing what we can do: actions, intentions, and the construction of phenomenal experience. Synthese, 181(3):375–394, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9714-6;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9714-
6.pdf.

Dougherty:2011:CAC

[255] Trent Dougherty and Patrick Rysiew. Clarity about concessive knowledge attributions: reply to Dodd. Synthese, 181(3):395–403, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9713-7;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9713-7.pdf.

Boschetti:2011:CEC

[256] Fabio Boschetti. Causality, emergence, computation and unreasonable expectations. Synthese, 181(3):405–412, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9720-8;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9720-
8.pdf.

Schulte:2011:TTT

[257] Peter Schulte. Truthmakers: a tale of two explanatory projects. Synthese, 181(3):413–431, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9716-4;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9716-4.pdf.

Jantzen:2011:NTE

[258] Benjamin C. Jantzen. No two entities without identity. Synthese, 181 (3):433–450, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9717-3;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9717-3.pdf.

Wrenn:2011:PSN

[259] Chase Wrenn. Practical success and the nature of truth. Synthese, 181 (3):451–470, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9733-3;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9733-3.pdf.

Coffman:2011:TCA

[260] E. J. Coffman. Two claims about epistemic propriety. Synthese, 181 (3):471–488, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9742-2;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9742-2.pdf.

Raftopoulos:2011:AFR

[261] Athanasios Raftopoulos. Ambiguous figures and representationalism. Synthese, 181(3):489–514, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9743-1;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9743-1.pdf.

Vaesen:2011:KCE

[262] Krist Vaesen. Knowledge without credit, exhibit 4: extended cognition. Synthese, 181(3):515–529, August 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9744-0;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9744-0.pdf.

Machamer:2011:PDT

[263] Peter Machamer. Phenomena, data and theories: a special issue of Synthese. Synthese, 182(1):1–5, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9622-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9622-
9.pdf.

Bogen:2011:SPS

[264] Jim Bogen. ‘Saving the phenomena’ and saving the phenomena. Synthese, 182(1):7–22, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9619-4;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9619-4.pdf.

Apel:2011:MER

[265] Jochen Apel. On the meaning and the epistemological relevance of the notion of a scientific phenomenon. Synthese, 182(1):23–38, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9620-y;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9620-y.pdf.

Schindler:2011:BWD

[266] Samuel Schindler. Bogen and Woodward’s data-phenomena distinction, forms of theory-ladenness, and the reliability of data. Synthese, 182 (1):39–55, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9615-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9615-8.pdf.

Feest:2011:WES

[267] Uljana Feest. What exactly is stabilized when phenomena are stabilized? Synthese, 182(1):57–71, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9616-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9616-
7.pdf.

McAllister:2011:WDP

[268] James W. McAllister. What do patterns in empirical data tell us about the structure of the world? Synthese, 182(1):73–87, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9613-x;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9613-x.pdf.

Votsis:2011:DMT

[269] Ioannis Votsis. Data meet theory: up close and inferentially personal. Synthese, 182(1):89–100, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9614-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9614-
9.pdf.

Massimi:2011:DPK

[270] Michela Massimi. From data to phenomena: a Kantian stance. Synthese, 182(1):101–116, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9611-z;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9611-z.pdf.

Tal:2011:DPB

[271] Eran Tal. From data to phenomena and back again: computersimulated signatures. Synthese, 182(1):117–129, September 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9612-y;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9612-
y.pdf.

Lowe:2011:DPC

[272] Benedikt Löwe and Thomas Müller. Data and phenomena in conceptual modelling. Synthese, 182(1):131–148, September 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9621-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9621-
x.pdf.

Falkenburg:2011:WPP

[273] Brigitte Falkenburg. What are the phenomena of physics? Synthese, 182(1):149–163, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9617-6;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9617-6.pdf.

Woodward:2011:DPR

[274] James F. Woodward. Data and phenomena: a restatement and defense. Synthese, 182(1):165–179, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9618-5;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9618-
5.pdf.

Allen:2011:SSI

[275] Colin Allen and Tony Beavers. Synthese special issue: representing philosophy. Synthese, 182(2):181–183, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9664-z;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9664-

z.pdf. See erratum [305].

Grenon:2011:FOP

[276] Pierre Grenon and Barry Smith. Foundations of an ontology of philosophy. Synthese, 182(2):185–204, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9658-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9658-
x.pdf.

Buckner:2011:EOT

[277] Cameron Buckner, Mathias Niepert, and Colin Allen. From encyclopedia to ontology: toward dynamic representation of the discipline of philosophy. Synthese, 182(2):205–233, September 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9659-9;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9659-
9.pdf.

Pasin:2011:ORA

[278] Michele Pasin and Enrico Motta. Ontological requirements for annotation and navigation of philosophical resources. Synthese, 182(2):235– 267, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9660-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9660-3.pdf.

Menzel:2011:KRW

[279] Christopher Menzel. Knowledge representation, the world wide Web, and the evolution of logic. Synthese, 182(2):269–295, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9661-2;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9661-2.pdf.

Morrow:2011:NM

[280] David R. Morrow and Chris Alen Sula. Naturalized metaphilosophy. Synthese, 182(2):297–313, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9662-1;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9662-
1.pdf.

Beavers:2011:NEI

[281] Anthony F. Beavers. Noesis and the encyclopedic Internet vision. Synthese, 182(2):315–333, September 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-009-9663-0;   http://link.springer.com/
content/pdf/10.1007/s11229-009-9663-0.pdf.

Zhang:2011:IDC

[282] Jiji Zhang and Peter Spirtes. Intervention, determinism, and the causal minimality condition. Synthese, 182(3):335–347, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9751-1;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9751-1.pdf.

Walton:2011:DSS

[283] Douglas Walton. A dialogue system specification for explanation. Synthese, 182(3):349–374, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9745-z;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9745-z.pdf.

Tiehen:2011:DMC

[284] Justin T. Tiehen. Disproportional mental causation. Synthese, 182 (3):375–391, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9749-8;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9749-8.pdf.

Bradley:2011:SLN

[285] D. J. Bradley. Self-location is no problem for conditionalization. Synthese, 182(3):393–411, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9748-9;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9748-9.pdf.

Lyon:2011:DPN

[286] Aidan Lyon. Deterministic probability: neither chance nor credence. Synthese, 182(3):413–432, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9750-2;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9750-2.pdf.

Johnson:2011:PFD

[287] Daniel M. Johnson. Proper function and defeating experiences. Synthese, 182(3):433–447, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9752-0;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9752-0.pdf.

Mueller:2011:DKM

[288] Axel Mueller. Does Kantian mental content externalism help metaphysical realists? Synthese, 182(3):449–473, October 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9753-z;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9753-
z.pdf.

Staley:2011:IEA

[289] Kent Staley and Aaron Cobb. Internalist and externalist aspects of justification in scientific inquiry. Synthese, 182(3):475–492, October 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9754-y;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9754-y.pdf.

Betti:2011:AMO

[290] Arianna Betti, Willem R. de Jong, and Marije Martijn. The axiomatic method, the order of concepts and the hierarchy of sciences: an introduction. Synthese, 183(1):1–5, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9671-0;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9671-
0.pdf.

vandenBerg:2011:KCP

[291] Hein van den Berg. Kant’s conception of proper science. Synthese, 183 (1):7–26, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9665-y;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9665-y.pdf.

Ziv:2011:BKI

[292] Anita Konzelmann Ziv. Bolzanian knowing: infallibility, virtue and foundational truth. Synthese, 183(1):27–45, November 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9666-x;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9666-
x.pdf.

Schlimm:2011:CRA

[293] Dirk Schlimm. On the creative role of axiomatics. The discovery of lattices by Schröder, Dedekind, Birkhoff, and others. Synthese, 183(1): 47–68, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9667-9;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9667-9.pdf.

Hintikka:2011:WAM

[294] Jaakko Hintikka. What is the axiomatic method? Synthese, 183(1): 69–85, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9668-8;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9668-8.pdf.

Muller:2011:RRS

[295] F. A. Muller. Reflections on the revolution at Stanford. Synthese, 183 (1):87–114, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9669-7;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9669-7.pdf. See response [296].

Suppes:2011:FDS

[296] Patrick Suppes. Future development of scientific structures closer to experiments: Response to F. A. Muller. Synthese, 183(1):115–126, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9670-1;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9670-1.pdf. See [295].

Alspector-Kelly:2011:WSD

[297] Marc Alspector-Kelly. Why safety doesn’t save closure. Synthese, 183(2): 127–142, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9755-x;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9755-x.pdf.

Russell:2011:ICS

[298] Gillian Russell. Indexicals, context-sensitivity and the failure of implication. Synthese, 183(2):143–160, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9756-9;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9756-
9.pdf.

Muldoon:2011:RIM

[299] Ryan Muldoon and Michael Weisberg. Robustness and idealization in models of cognitive labor. Synthese, 183(2):161–174, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9757-8;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9757-8.pdf.

Avnur:2011:OPN

[300] Yuval Avnur. An old problem for the new rationalism. Synthese, 183(2): 175–185, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9759-6;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9759-6.pdf.

vanderSchaar:2011:AGT

[301] Maria van der Schaar. Assertion and grounding: a theory of assertion for constructive type theory. Synthese, 183(2):187–210, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9758-7;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9758-7.pdf.

Cobreros:2011:PVP

[302] Pablo Cobreros. Paraconsistent vagueness: a positive argument. Synthese, 183(2):211–227, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9760-0;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9760-0.pdf.

Schroer:2011:CDP

[303] Robert Schroer. Can determinable properties earn their keep? Synthese, 183(2):229–247, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9767-6;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9767-6.pdf.

Paternotte:2011:BRA

[304] Cedric Paternotte. Being realistic about common knowledge: a Lewisian approach. Synthese, 183(2):249–276, November 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9770-y;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9770-
y.pdf.

Allen:2011:ESS

[305] Colin Allen and Tony Beavers. Erratum to: Synthese special issue: representing philosophy. Synthese, 183(2):277, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9690-x;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9690-x.pdf. See [275].

Weaver:2011:EWC

[306] Christopher Gregory Weaver. Erratum to: What could be caused must actually be caused. Synthese, 183(2):279, November 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9852-x;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9852-
x.pdf.

Masuzawa:2011:EII

[307] Takuya Masuzawa and Koji Hasebe. Erratum to: Iterative information update and stability of strategies. Synthese, 183(2):281, November 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9858-4;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9858-4.pdf.

Piccinini:2011:IPN

[308] Gualtiero Piccinini and Carl Craver. Integrating psychology and neuroscience: functional analyses as mechanism sketches. Synthese, 183(3): 283–311, December 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9898-4;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9898-4.pdf.

Weiskopf:2011:MMP

[309] Daniel A. Weiskopf. Models and mechanisms in psychological explanation. Synthese, 183(3):313–338, December 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9958-9;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9958-
9.pdf.

Kaplan:2011:EDC

[310] David Michael Kaplan. Explanation and description in computational neuroscience. Synthese, 183(3):339–373, December 2011. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9970-0;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9970-
0.pdf.

Couch:2011:MCR

[311] Mark B. Couch. Mechanisms and constitutive relevance. Synthese, 183 (3):375–388, December 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9882-z;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9882-z.pdf. See erratum [442].

Waskan:2011:MEL

[312] Jonathan Waskan. Mechanistic explanation at the limit. Synthese, 183 (3):389–408, December 2011. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9869-1;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9869-1.pdf.

Woodward:2011:MR

[313] James Woodward. Mechanisms revisited. Synthese, 183(3):409–427, December 2011. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.

1007/s11229-011-9870-3;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9870-3.pdf.

Peijnenburg:2012:ISI

[314] Jeanne Peijnenburg, Branden Fitelson, and Igor Douven. Introduction to the special issue: Probability, confirmation and fallacies. Synthese, 184(1):1, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9700-z;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9700-z.pdf.

Tentori:2012:HCF

[315] Katya Tentori and Vincenzo Crupi. How the conjunction fallacy is tied to probabilistic confirmation: Some remarks on Schupbach (2009). Synthese, 184(1):3–12, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9701-y;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9701-
y.pdf.

Schupbach:2012:CFT

[316] Jonah N. Schupbach. Is the conjunction fallacy tied to probabilistic confirmation? Synthese, 184(1):13–27, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9698-2;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9698-
2.pdf.

Shogenji:2012:DEJ

[317] Tomoji Shogenji. The degree of epistemic justification and the conjunction fallacy. Synthese, 184(1):29–48, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9699-1;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9699-

1.pdf. See commentary [318] and defense [417].

Atkinson:2012:CJC

[318] David Atkinson. Confirmation and justification. A commentary on Shogenji’s measure. Synthese, 184(1):49–61, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9696-4;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9696-

4.pdf. See [317].

Kuipers:2012:RPL

[319] Theo A. F. Kuipers. A realist partner for Linda: confirming a theoretical hypothesis more than its observational sub-hypothesis. Synthese, 184 (1):63–71, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-009-9697-3;   http://link.springer.com/content/
pdf/10.1007/s11229-009-9697-3.pdf.

Hartmann:2012:WBC

[320] Stephan Hartmann and Wouter Meijs. Walter the banker: the conjunction fallacy reconsidered. Synthese, 184(1):73–87, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9694-6;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9694-6.pdf.

Festa:2012:EOH

[321] Roberto Festa. “For unto every one that hath shall be given”. Matthew properties for incremental confirmation. Synthese, 184(1):89–100, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
009-9695-5;   http://link.springer.com/content/pdf/10.1007/
s11229-009-9695-5.pdf.

Peijnenburg:2012:CCP

[322] Jeanne Peijnenburg. A case of confusing probability and confirmation. Synthese, 184(1):101–107, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9692-8;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9692-
8.pdf.

Douven:2012:QPC

[323] Igor Douven and Jos Uffink. Quantum probabilities and the conjunction principle. Synthese, 184(1):109–114, January 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9693-7;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9693-
7.pdf.

DeBrabanter:2012:I

[324] Philippe De Brabanter and Mikhail Kissine. Introduction. Synthese, 184 (2):115–120, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9727-1;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9727-1.pdf.

Predelli:2012:IIR

[325] Stefano Predelli. Indexicality, intensionality, and relativist postsemantics. Synthese, 184(2):121–136, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9728-0;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9728-
0.pdf.

Stojanovic:2012:DS

[326] Isidora Stojanovic. Domain-sensitivity. Synthese, 184(2):137–155, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9729-z;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9729-z.pdf.

Moltmann:2012:TKF

[327] Friederike Moltmann. Two kinds of first-person-oriented content. Synthese, 184(2):157–177, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9730-6;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9730-6.pdf.

Corazza:2012:SMV

[328] Eros Corazza and Jérôme Dokic. Situated minimalism versus free enrichment. Synthese, 184(2):179–198, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9731-5;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9731-
5.pdf.

Kissine:2012:CCE

[329] Mikhail Kissine. From contexts to circumstances of evaluation: is the trade-off always innocuous? Synthese, 184(2):199–216, January 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9732-4;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9732-4.pdf.

Moretti:2012:WOC

[330] Luca Moretti. Wright, Okasha and Chandler on transmission failure. Synthese, 184(3):217–234, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9771-x;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9771-x.pdf.

Patton:2012:ETB

[331] Lydia Patton. Experiment and theory building. Synthese, 184(3): 235–246, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9772-9;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9772-9.pdf.

Turri:2012:KJT

[332] John Turri. Is knowledge justified true belief? Synthese, 184(3): 247–259, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9773-8;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9773-8.pdf.

Slutsky:2012:CDU

[333] David Slutsky. Confusion and dependence in uses of history. Synthese, 184(3):261–286, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9785-4;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9785-4.pdf.

Moon:2012:WDE

[334] Andrew Moon. Warrant does entail truth. Synthese, 184(3):287– 297, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9815-2;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9815-2.pdf.

Weaver:2012:WCC

[335] Christopher Gregory Weaver. What could be caused must actually be caused. Synthese, 184(3):299–317, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9814-3;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9814-
3.pdf.

Burgess:2012:NPC

[336] Simon Burgess. Newcomb’s problem and its conditional evidence: a common cause of confusion. Synthese, 184(3):319–339, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9816-1;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9816-1.pdf.

Wayne:2012:ESL

[337] Andrew Wayne. Emergence and singular limits. Synthese, 184(3): 341–356, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9817-0;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9817-0.pdf.

Chase:2012:LQR

[338] James Kennedy Chase. The logic of Quinean revisability. Synthese, 184 (3):357–373, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9819-y;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9819-y.pdf.

Effingham:2012:UCM

[339] N. Effingham. An unwelcome consequence of the multiverse thesis. Synthese, 184(3):375–386, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9820-5;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9820-5.pdf.

Gustafsson:2012:CSA

[340] Johan E. Gustafsson and Martin Peterson. A computer simulation of the argument from disagreement. Synthese, 184(3):387–405, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9822-3;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9822-3.pdf.

Jones:2012:DCC

[341] Todd Jones. Do customs compete with conditioning? Turf battles and division of labor in social explanation. Synthese, 184(3):407–430, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

010-9794-3;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9794-3.pdf.

Floridi:2012:SIN

[342] Luciano Floridi. Semantic information and the network theory of account. Synthese, 184(3):431–454, February 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9821-4;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9821-
4.pdf.

Bedau:2012:IPP

[343] Mark A. Bedau. Introduction to philosophical problems about life. Synthese, 185(1):1–3, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9872-1;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9872-1.pdf.

Pennock:2012:NBD

[344] Robert T. Pennock. Negotiating boundaries in the definition of life: Wittgensteinian and Darwinian insights on resolving conceptual border conflicts. Synthese, 185(1):5–20, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9873-0;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9873-
0.pdf.

Ruiz-Mirazo:2012:AEM

[345] Kepa Ruiz-Mirazo and Alvaro Moreno. Autonomy in evolution: from minimal to complex life. Synthese, 185(1):21–52, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9874-z;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9874-z.pdf.

Escobar:2012:AD

[346] Jorge M. Escobar. Autopoiesis and Darwinism. Synthese, 185(1): 53–72, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9875-y;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9875-y.pdf.

Bedau:2012:FAD

[347] Mark A. Bedau. A functional account of degrees of minimal chemical life. Synthese, 185(1):73–88, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9876-x;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9876-
x.pdf.

Sandler:2012:AVR

[348] Ronald Sandler. Is artefactualness a value-relevant property of living things? Synthese, 185(1):89–102, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9877-9;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9877-
9.pdf.

Shields:2012:DL

[349] Christopher Shields. The dialectic of life. Synthese, 185(1):103– 124, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9878-8;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9878-8.pdf.

Cleland:2012:LD

[350] Carol E. Cleland. Life without definitions. Synthese, 185(1):125– 144, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9879-7;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9879-7.pdf.

Machery:2012:WSW

[351] Edouard Machery. Why I stopped worrying about the definition of life … and why you should as well. Synthese, 185(1):145–164, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9880-1;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9880-1.pdf.

Bersini:2012:PSP

[352] Hugues Bersini, Pasquale Stano, Pier Luigi Luisi, and Mark A. Bedau. Philosophical and scientific perspectives on emergence. Synthese, 185 (2):165–169, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9718-2;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9718-2.pdf.

Mitchell:2012:ELF

[353] Sandra D. Mitchell. Emergence: logical, functional and dynamical. Synthese, 185(2):171–186, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9719-1;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9719-1.pdf.

Boschetti:2012:CEC

[354] Fabio Boschetti. Causality, emergence, computation and unreasonable expectations. Synthese, 185(2):187–194, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0077-z;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0077-
z.pdf.

Huneman:2012:DPO

[355] Philippe Huneman. Determinism, predictability and open-ended evolution: lessons from computational emergence. Synthese, 185(2): 195–214, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9721-7;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9721-7.pdf.

Bich:2012:CEL

[356] Leonardo Bich. Complex emergence and the living organization: an epistemological framework for biology. Synthese, 185(2):215–232, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9722-6;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9722-6.pdf.

Bitbol:2012:DCF

[357] Michel Bitbol. Downward causation without foundations. Synthese, 185 (2):233–255, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9723-5;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9723-5.pdf.

Bersini:2012:EPB

[358] Hugues Bersini. Emergent phenomena belong only to biology. Synthese, 185(2):257–272, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9724-4;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9724-4.pdf.

Damiano:2012:CEL

[359] Luisa Damiano. Co-emergences in life and science: a double proposal for biological emergentism. Synthese, 185(2):273–294, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9725-3;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9725-3.pdf.

Corning:2012:REE

[360] Peter A. Corning. The re-emergence of emergence, and the causal role of synergy in emergent evolution. Synthese, 185(2):295–317, March 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9726-2;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9726-2.pdf.

Ballantyne:2012:LI

[361] Nathan Ballantyne. Luck and interests. Synthese, 185(3):319–334, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9747-x;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9747-x.pdf.

Stenlund:2012:DSF

[362] Sören Stenlund. Different senses of finitude: An inquiry into Hilbert’s finitism. Synthese, 185(3):335–363, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9823-2;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9823-
2.pdf.

Hendrickson:2012:CRP

[363] Noel Hendrickson. Counterfactual reasoning and the problem of selecting antecedent scenarios. Synthese, 185(3):365–386, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9824-1;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9824-1.pdf.

Novaes:2012:RLH

[364] Catarina Dutilh Novaes. Reassessing logical hylomorphism and the demarcation of logical constants. Synthese, 185(3):387–410, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9825-0;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9825-0.pdf.

Glass:2012:IBE

[365] David H. Glass. Inference to the best explanation: does it track truth? Synthese, 185(3):411–427, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9829-9;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9829-
9.pdf.

Gal:2012:NDP

[366] Ofer Gal and Raz Chen-Morris. Nature’s drawing: problems and resolutions in the mathematization of motion. Synthese, 185(3):429–466, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9978-5;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9978-5.pdf.

Schuster:2012:PMS

[367] John A. Schuster. Physico-mathematics and the search for causes in Descartes’ optics — 1619–1637. Synthese, 185(3):467–499, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9979-4;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9979-4.pdf.

Dupre:2012:KOH

[368] Sven Dupré. Kepler’s optics without hypotheses. Synthese, 185(3): 501–525, April 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9977-6;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9977-6.pdf.

Mumma:2012:DMH

[369] John Mumma and Marco Panza. Diagrams in mathematics: history and philosophy. Synthese, 186(1):1–5, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9988-3;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9988-
3.pdf.

Saito:2012:TDT

[370] Ken Saito. Traditions of the diagram, tradition of the text: A case study. Synthese, 186(1):7–20, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0073-3;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0073-3.pdf.

DeYoung:2012:MDM

[371] Gregg De Young. Mathematical diagrams from manuscript to print: examples from the Arabic Euclidean transmission. Synthese, 186(1):21–54, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0070-6;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0070-6.pdf.

Panza:2012:TRD

[372] Marco Panza. The twofold role of diagrams in Euclid’s plane geometry. Synthese, 186(1):55–102, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0074-2;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0074-2.pdf.

Mumma:2012:CGR

[373] John Mumma. Constructive geometrical reasoning and diagrams. Synthese, 186(1):103–119, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9981-x;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9981-x.pdf.

Coliva:2012:HDR

[374] Annalisa Coliva. Human diagrammatic reasoning and seeing-as. Synthese, 186(1):121–148, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9982-9;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9982-9.pdf.

Shin:2012:FID

[375] Sun-Joo Shin. The forgotten individual: diagrammatic reasoning in mathematics. Synthese, 186(1):149–168, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0075-1;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0075-
1.pdf.

DePierris:2012:HSG

[376] Graciela De Pierris. Hume on space, geometry, and diagrammatic reasoning. Synthese, 186(1):169–189, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0071-5;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0071-
5.pdf.

Dunlop:2012:MFM

[377] Katherine Dunlop. The mathematical form of measurement and the argument for Proposition I in Newton’s Principia. Synthese, 186(1):191–229, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9983-8;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9983-8.pdf.

Friedman:2012:KGS

[378] Michael Friedman. Kant on geometry and spatial intuition. Synthese, 186(1):231–255, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0066-2;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0066-2.pdf.

Tournes:2012:DTD

[379] Dominique Tournès. Diagrams in the theory of differential equations (eighteenth to nineteenth centuries). Synthese, 186(1):257–288, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0069-z;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0069-z.pdf.

Macbeth:2012:DRF

[380] Danielle Macbeth. Diagrammatic reasoning in Frege’s begriffsschrift. Synthese, 186(1):289–314, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0068-0;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0068-0.pdf.

Smadja:2012:LAD

[381] Ivahn Smadja. Local axioms in disguise: Hilbert on Minkowski diagrams. Synthese, 186(1):315–370, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9984-7;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9984-7.pdf.

Feferman:2012:LRI

[382] Solomon Feferman. And so on …: reasoning with infinite diagrams. Synthese, 186(1):371–386, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9985-6;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9985-6.pdf.

Halimi:2012:DS

[383] Brice Halimi. Diagrams as sketches. Synthese, 186(1):387–409, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9986-5;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9986-5.pdf.

SanGines:2012:SLD

[384] Aránzazu San Ginés. Seeing the language: a diagrammatic approach to natural discourse. Synthese, 186(1):411–439, May 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9987-4;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9987-
4.pdf.

Smadja:2012:ELA

[385] Ivahn Smadja. Erratum to: Local axioms in disguise: Hilbert on Minkowski diagrams. Synthese, 186(1):441–442, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0039-x;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0039-
x.pdf.

Wheeler:2012:I

[386] Gregory Wheeler. Introduction. Synthese, 186(2):443–446, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9993-6;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9993-6.pdf. See erratum [443].

Harper:2012:BC

[387] William Harper, Sheldon J. Chow, and Gemma Murray. Bayesian chance. Synthese, 186(2):447–474, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9994-5;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9994-5.pdf.

Howson:2012:MUI

[388] Colin Howson. Modelling uncertain inference. Synthese, 186(2):475–492, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9995-4;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9995-4.pdf.

Levi:2012:DC

[389] Isaac Levi. Deductive closure. Synthese, 186(2):493–499, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9996-3;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9996-3.pdf.

Teng:2012:WAF

[390] Choh Man Teng. When adjunction fails. Synthese, 186(2):501–510, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0002-x;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0002-x.pdf.

Makinson:2012:LQB

[391] David Makinson. Logical questions behind the lottery and preface paradoxes: lossy rules for uncertain inference. Synthese, 186(2):511–529, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9997-2;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9997-2.pdf.

Lin:2012:GLS

[392] Hanti Lin and Kevin T. Kelly. A geo-logical solution to the lottery paradox, with applications to conditional logic. Synthese, 186(2):531–575, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9998-1;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9998-1.pdf.

Cozman:2012:SPD

[393] Fabio G. Cozman. Sets of probability distributions, independence, and convexity. Synthese, 186(2):577–600, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9999-0;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9999-
0.pdf.

Helzner:2012:RE

[394] Jeffrey Helzner. On the representation of error. Synthese, 186(2):601–613, May 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0001-y;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0001-y.pdf.

vanBenthem:2012:NLP

[395] Johan van Benthem and Sonja Smets. New logical perspectives on physics. Synthese, 186(3):615–617, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-011-9911-y;   http://link.springer.
com/content/pdf/10.1007/s11229-011-9911-y.pdf.

Aiello:2012:LPS

[396] Marco Aiello, Guram Bezhanishvili, Isabelle Bloch, and Valentin Goranko. Logic for physical space. Synthese, 186(3):619–632, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9913-9;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9913-9.pdf.

Andreka:2012:LRS

[397] Hajnal Andréka, Judit X. Madarász, István Németi, and Gergely Székely. A logic road from special relativity to general relativity. Synthese, 186 (3):633–649, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9914-8;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9914-8.pdf.

Coecke:2012:PCQ

[398] Bob Coecke and Robert W. Spekkens. Picturing classical and quantum Bayesian inference. Synthese, 186(3):651–696, June 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9917-5;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9917-
5.pdf.

Abramsky:2012:BTM

[399] Samson Abramsky. Big toy models. Synthese, 186(3):697–718, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9912-x;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9912-x.pdf.

Heunen:2012:BOA

[400] Chris Heunen, Nicolaas P. Landsman, and Bas Spitters. Bohrification of operator algebras and quantum logic. Synthese, 186(3):719–752, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9918-4;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9918-4.pdf.

Baltag:2012:DTQ

[401] Alexandru Baltag and Sonja Smets. The dynamic turn in quantum logic. Synthese, 186(3):753–773, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9915-7;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9915-7.pdf.

vanBenthem:2012:LET

[402] Johan van Benthem. The logic of empirical theories revisited. Synthese, 186(3):775–792, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9916-6;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9916-6.pdf.

Costa:2012:MFD

[403] Horacio Arló Costa and Jeffrey Helzner. More foundations of the decision sciences: introduction. Synthese, 187(1):1–10, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0021-7;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0021-7.pdf.

Gilboa:2012:RBW

[404] Itzhak Gilboa, Andrew Postlewaite, and David Schmeidler. Rationality of belief or: why savage’s axioms are neither necessary nor sufficient for rationality. Synthese, 187(1):11–31, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0034-2;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0034-2.pdf.

Hajek:2012:RIP

[405] Alan Hájek and Michael Smithson. Rationality and indeterminate probabilities. Synthese, 187(1):33–48, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0033-3;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0033-
3.pdf.

Xiao:2012:WDC

[406] Erte Xiao and Cristina Bicchieri. Words or deeds? Choosing what to know about others. Synthese, 187(1):49–63, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0026-2;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0026-2.pdf.

McClennen:2012:RC

[407] Edward McClennen. Rational cooperation. Synthese, 187(1):65–93, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0032-4;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0032-4.pdf.

Spohn:2012:RYD

[408] Wolfgang Spohn. Reversing 30 years of discussion: why causal decision theorists should one-box. Synthese, 187(1):95–122, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0023-5;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0023-5.pdf.

Joyce:2012:RIC

[409] James M. Joyce. Regret and instability in causal decision theory. Synthese, 187(1):123–145, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0022-6;   http://link.springer.com/
content/pdf/10.1007/s11229-011-0022-6.pdf.

Gaifman:2012:DUM

[410] Haim Gaifman and Anubav Vasudevan. Deceptive updating and minimal information methods. Synthese, 187(1):147–178, July 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0028-0;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0028-
0.pdf.

List:2012:TJA

[411] Christian List. The theory of judgment aggregation: an introductory review. Synthese, 187(1):179–207, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0025-3;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0025-
3.pdf.

Hartmann:2012:JAP

[412] Stephan Hartmann and Jan Sprenger. Judgment aggregation and the problem of tracking the truth. Synthese, 187(1):209–221, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0031-5;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0031-5.pdf.

Weirich:2012:CA

[413] Paul Weirich. Collective acts. Synthese, 187(1):223–241, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0027-1;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0027-1.pdf.

Gigerenzer:2012:HFC

[414] Gerd Gigerenzer and Thomas Sturm. How (far) can rationality be naturalized? Synthese, 187(1):243–268, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0030-6;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0030-
6.pdf.

Hertwig:2012:PRD

[415] Ralph Hertwig. The psychology and rationality of decisions from experience. Synthese, 187(1):269–292, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-011-0024-4;   http://link.springer.
com/content/pdf/10.1007/s11229-011-0024-4.pdf.

Rustichini:2012:DME

[416] Aldo Rustichini. Decision making and equilibria. Synthese, 187(1): 293–304, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0029-z;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0029-z.pdf.

Schubert:2012:CRR

[417] Stefan Schubert. Coherence reasoning and reliability: a defense of the Shogenji measure. Synthese, 187(2):305–319, July 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9856-6;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9856-

6.pdf. See [317].

Shieber:2012:PDI

[418] Joseph Shieber. A partial defense of intuition on naturalist grounds. Synthese, 187(2):321–341, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9864-6;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9864-6.pdf.

Abrams:2012:MP

[419] Marshall Abrams. Mechanistic probability. Synthese, 187(2):343–375, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9830-3;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9830-3.pdf.

Anderson:2012:KC

[420] David James Anderson. Knowledge and conviction. Synthese, 187 (2):377–392, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9831-2;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9831-2.pdf.

Cooper:2012:ODB

[421] Gordon Cooper and Stephen M. Humphry. The ontological distinction between units and entities. Synthese, 187(2):393–401, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9832-1;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9832-1.pdf.

Mormann:2012:MSC

[422] Thomas Mormann. On the mereological structure of complex states of affairs. Synthese, 187(2):403–418, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9828-x;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9828-
x.pdf.

Tillman:2012:RJI

[423] Chris Tillman. Reconciling justificatory internalism and content externalism. Synthese, 187(2):419–440, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-010-9827-y;   http://link.springer.
com/content/pdf/10.1007/s11229-010-9827-y.pdf.

Placek:2012:IMN

[424] Tomasz Placek and Nuel Belnap. Indeterminism is a modal notion: branching spacetimes and Earman’s pruning. Synthese, 187(2):441–469, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9846-8;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9846-8.pdf.

vanRooij:2012:IUH

[425] Iris van Rooij, Cory D. Wright, and Todd Wareham. Intractability and the use of heuristics in psychological explanations. Synthese, 187 (2):471–487, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9847-7;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9847-7.pdf.

Busch:2012:CNI

[426] Jacob Busch. Can the new indispensability argument be saved from Euclidean rescues? Synthese, 187(2):489–508, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9848-6;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9848-
6.pdf.

Booth:2012:ATC

[427] Anthony Robert Booth. All things considered duties to believe. Synthese, 187(2):509–517, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9857-5;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9857-5.pdf.

Corsentino:2012:PP

[428] Anthony Corsentino. Predicates in perspective. Synthese, 187(2): 519–545, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9859-3;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9859-3.pdf.

Myrvold:2012:EVV

[429] Wayne C. Myrvold. Epistemic values and the value of learning. Synthese, 187(2):547–568, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9860-x;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9860-x.pdf.

Raatikainen:2012:RII

[430] Panu Raatikainen. Ramsification and inductive inference. Synthese, 187(2):569–577, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9861-9;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9861-9.pdf.

Goforth:2012:ECA

[431] David Goforth and David Robinson. Effective choice in all the symmetric 2 × 2 games. Synthese, 187(2):579–605, July 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9862-8;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9862-
8.pdf.

Schubert:2012:CCR

[432] Stefan Schubert. Is coherence conducive to reliability? Synthese, 187 (2):607–621, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9865-5;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9865-5.pdf.

deBoer:2012:TDS

[433] Mathijs de Boer, Dov M. Gabbay, Xavier Parent, and Marija Slavkovic. Two dimensional standard Deontic logic [including a detailed analysis of the 1985 Jones-Pörn deontic logic system]. Synthese, 187(2):623–660, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9866-4;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9866-4.pdf.

DeClercq:2012:SPG

[434] Rafael De Clercq. On some putative graph-theoretic counterexamples to the Principle of the Identity of Indiscernibles. Synthese, 187(2):661–672, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9867-3;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9867-3.pdf.

Fairweather:2012:DQV

[435] Abrol Fairweather. Duhem-Quine virtue epistemology. Synthese, 187 (2):673–692, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9868-2;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9868-2.pdf.

Dambock:2012:TSR

[436] Christian Damböck. Theory structuralism in a rigid framework. Synthese, 187(2):693–713, July 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0009-3;   http://link.springer.com/
content/pdf/10.1007/s11229-011-0009-3.pdf.

Larvor:2012:HTA

[437] Brendan Larvor. How to think about informal proofs. Synthese, 187(2):715–730, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0007-5;   https://link.springer.
com/content/pdf/10.1007/s11229-011-0007-5.pdf.

Urbaniak:2012:PTE

[438] Rafal Urbaniak. “Platonic” thought experiments: how on earth? Synthese, 187(2):731–752, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0008-4;   https://link.springer.
com/content/pdf/10.1007/s11229-011-0008-4.pdf.

Decker:2012:DEA

[439] Jason Decker. Disagreement, evidence, and agnosticism. Synthese, 187(2):753–783, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0010-x;   https://link.springer.
com/content/pdf/10.1007/s11229-011-0010-x.pdf.

Dragulinescu:2012:SMM

[440] Stefan Dragulinescu. On ‘stabilising’ medical mechanisms, truth-makers and epistemic causality: a critique to Williamson and Russo’s approach. Synthese, 187(2):785–800, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0011-9;   https://link.springer.
com/content/pdf/10.1007/s11229-011-0011-9.pdf.

Arenhart:2012:MEN

[441] Jonas R. Becker Arenhart. Many entities, no identity. Synthese, 187(2):801–812, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0013-7;   https://link.springer.
com/content/pdf/10.1007/s11229-011-0013-7.pdf.

Couch:2012:EMC

[442] Mark B. Couch. Erratum to: Mechanisms and constitutive relevance. Synthese, 187(2):813, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0080-4;   https://link.springer.
com/content/pdf/10.1007/s11229-012-0080-4.pdf. See [311].

Wheeler:2012:EI

[443] Gregory Wheeler. Erratum to: Introduction. Synthese, 187(2):815, July 2012. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-012-0139-2;   https://link.springer.com/content/pdf/
10.1007/s11229-012-0139-2.pdf. See [386].

Lindstrom:2012:IPL

[444] Sten Lindström, Erik Palmgren, and Dag Westerståhl. Introduction: The philosophy of logical consequence and inference. Synthese, 187(3): 817–820, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0180-1;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0180-1.pdf.

Boyer:2012:BPT

[445] Julien Boyer and Gabriel Sandu. Between proof and truth. Synthese, 187 (3):821–832, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9903-y;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9903-y.pdf. See erratum [455].

Eklund:2012:MTL

[446] Matti Eklund. Multitude, tolerance and language-transcendence. Synthese, 187(3):833–847, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9904-x;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9904-x.pdf.

Hakli:2012:DDT

[447] Raul Hakli and Sara Negri. Does the deduction theorem fail for modal logic? Synthese, 187(3):849–867, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9905-9;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9905-
9.pdf.

Pagin:2012:AIC

[448] Peter Pagin. Assertion, inference, and consequence. Synthese, 187(3): 869–885, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9906-8;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9906-8.pdf.

Prawitz:2012:ESV

[449] Dag Prawitz. The epistemic significance of valid inference. Synthese, 187 (3):887–898, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9907-7;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9907-7.pdf.

Read:2012:MTC

[450] Stephen Read. The medieval theory of consequence. Synthese, 187 (3):899–912, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9908-6;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9908-6.pdf.

Sandqvist:2012:AIM

[451] Tor Sandqvist. Acceptance, inference, and the multiple-conclusion sequent. Synthese, 187(3):913–924, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-011-9909-5;   http://link.springer.
com/content/pdf/10.1007/s11229-011-9909-5.pdf.

Schroeder-Heister:2012:CHC

[452] Peter Schroeder-Heister. The categorical and the hypothetical: a critique of some fundamental assumptions of standard semantics. Synthese, 187 (3):925–942, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9910-z;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9910-z.pdf.

Sundholm:2012:IVC

[453] Göran Sundholm. “Inference versus consequence” revisited: inference, consequence, conditional, implication. Synthese, 187(3):943–956, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9901-0;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9901-0.pdf.

Westerstaahl:2012:CCB

[454] Dag Westerståhl. From constants to consequence, and back. Synthese, 187(3):957–971, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9902-z;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9902-z.pdf.

Boyer:2012:EBP

[455] Julien Boyer and Gabriel Sandu. Erratum to: Between proof and truth. Synthese, 187(3):973–974, August 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0043-1;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0043-

1.pdf. See [445].

Malpass:2012:SSI

[456] Alex Malpass and Chris Gifford. Synthese special issue introduction. Synthese, 188(1):1–3, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0123-x;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0123-x.pdf.

Belnap:2012:NDB

[457] Nuel Belnap. Newtonian determinism to branching space-times indeterminism in two moves. Synthese, 188(1):5–21, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0063-5;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0063-5.pdf.

Placek:2012:IBH

[458] Tomasz Placek. On individuals in branching histories. Synthese, 188 (1):23–39, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0082-2;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0082-2.pdf.

Muller:2012:BLP

[459] Thomas Müller. Branching in the landscape of possibilities. Synthese, 188 (1):41–65, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0059-6;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0059-6.pdf.

Wilson:2012:EQM

[460] Alastair Wilson. Everettian quantum mechanics without branching time. Synthese, 188(1):67–84, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0048-9;

 http://link.springer.com/content/pdf/10.1007/s11229-011-0048-
9.pdf.

Farr:2012:TEB

[461] Matt Farr. On A- and B-theoretic elements of branching spacetimes. Synthese, 188(1):85–116, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0046-y;   http://link.springer.com/
content/pdf/10.1007/s11229-011-0046-y.pdf.

Malpass:2012:FTR

[462] Alex Malpass and Jacek Wawer. A future for the thin red line. Synthese, 188(1):117–142, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0064-4;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0064-4.pdf.

deAlmeida:2012:GE

[463] Claudio de Almeida and Stephen Hetherington. Guest editorial. Synthese, 188(2):143, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9929-1;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9929-1.pdf.

Steup:2012:BCI

[464] Matthias Steup. Belief control and intentionality. Synthese, 188(2): 145–163, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9919-3;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9919-3.pdf.

Brueckner:2012:EP

[465] Anthony Brueckner and Alex Bundy. On “epistemic permissiveness”. Synthese, 188(2):165–177, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-011-9921-9;   http://link.springer.
com/content/pdf/10.1007/s11229-011-9921-9.pdf.

Adams:2012:TCC

[466] Fred Adams, John A. Barker, and Julia Figurelli. Towards closure on closure. Synthese, 188(2):179–196, September 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9922-8;

 http://link.springer.com/content/pdf/10.1007/s11229-011-9922-
8.pdf.

deAlmeida:2012:ECS

[467] Claudio de Almeida. Epistemic closure, skepticism and defeasibility. Synthese, 188(2):197–215, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9923-7;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9923-
7.pdf.

Hetherington:2012:GIG

[468] Stephen Hetherington. The gettier-illusion: Gettier-partialism and infallibilism. Synthese, 188(2):217–230, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9924-6;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9924-
6.pdf.

Williams:2012:MPB

[469] John N. Williams. Moore-paradoxical belief, conscious belief and the epistemic Ramsey test. Synthese, 188(2):231–246, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9925-5;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9925-5.pdf.

Adler:2012:CFP

[470] Jonathan E. Adler. Contextualism and fallibility: pragmatic encroachment, possibility, and strength of epistemic position. Synthese, 188(2): 247–272, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9926-4;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9926-4.pdf.

Reed:2012:KDC

[471] Baron Reed. Knowledge, doubt, and circularity. Synthese, 188(2):273– 287, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9927-3;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9927-3.pdf. See reply [473].

Battaly:2012:SRK

[472] Heather Battaly. Sosa’s reflective knowledge: How damaging is epistemic circularity? Synthese, 188(2):289–308, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9928-2;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9928-2.pdf. See reply [473].

Sosa:2012:RKR

[473] Ernest Sosa. On reflective knowledge: replies to Battaly and Reed. Synthese, 188(2):309–321, September 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9920-x;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9920-x.pdf. See [472, 471].

Ohrstrom:2012:P

[474] Peter Øhrstrøm, Per F. V. Hasle, and Ulrik Sandborg-Petersen. Preface. Synthese, 188(3):323–324, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9934-4;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9934-4.pdf.

Ohrstrom:2012:LA

[475] Peter Øhrstrøm and Per F. V. Hasle. From a logical angle. Synthese, 188 (3):325–330, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9941-5;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9941-5.pdf.

Hasle:2012:PPP

[476] Per F. V. Hasle. The problem of predestination: as a prelude to A. N. Prior’s tense logic. Synthese, 188(3):331–347, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9942-4;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9942-4.pdf.

Uckelman:2012:APM

[477] Sara L. Uckelman. Arthur Prior and medieval logic. Synthese, 188 (3):349–366, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9943-3;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9943-3.pdf.

Ploug:2012:BTI

[478] Thomas Ploug and Peter Øhrstrøm. Branching time, indeterminism and tense logic. Synthese, 188(3):367–379, October 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9944-2;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9944-
2.pdf.

Prior:2012:FUE

[479] A. N. Prior. Faith, unbelief and evil: a fragment of a dialogue. Synthese, 188(3):381–397, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9945-1;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9945-1.pdf.

Jakobsen:2012:IFU

[480] David Jakobsen. An introduction to ‘faith, unbelief and evil’. Synthese, 188(3):399–409, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9946-0;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9946-0.pdf.

Prior:2012:PPL

[481] A. N. Prior. The paradox of the prisoner in logical form. Synthese, 188 (3):411–416, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9947-z;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9947-z.pdf.

Ohrstrom:2012:TK

[482] Peter Øhrstrøm, Lasse Burri Gram-Hansen, and Ulrik Sandborg-Petersen. Time and knowledge. Synthese, 188(3):417–422, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9948-y;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9948-y.pdf.

Prior:2012:LOO

[483] A. N. Prior. The logic of obligation and the obligations of the logician. Synthese, 188(3):423–448, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9935-3;

 http://link.springer.com/content/pdf/10.1007/s11229-011-9935-

3.pdf. See discussion [484].

Ohrstrom:2012:PDH

[484] Peter Øhrstrøm, Jörg Zeller, and Ulrik Sandborg-Petersen. Prior’s defence of Hintikka’s theorem. A discussion of Prior’s ‘The logic of obligation and the obligations of the logician’. Synthese, 188(3):449–454, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9936-2;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9936-2.pdf. See [483].

Prior:2012:FFP

[485] A. N. Prior. The fable of the four preachers. Synthese, 188(3): 455–457, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9937-1;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9937-1.pdf.

Ujvari:2012:PFL

[486] Márta Ujvári. Prior’s Fable and the limits of de re possibility. Synthese, 188(3):459–467, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9938-0;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9938-0.pdf.

Muller:2012:LPS

[487] Thomas Müller and Niko Strobach. A letter on the present state of affairs. Synthese, 188(3):469–485, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9939-z;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9939-
z.pdf.

Uckelman:2012:PIJ

[488] Sara L. Uckelman. Prior on an insolubilium of Jean Buridan. Synthese, 188(3):487–498, October 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9940-6;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9940-6.pdf.

Briggs:2012:PSS

[489] Rachael Briggs and Mark Jago. Propositions and same-saying: introduction. Synthese, 189(1):1–10, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0091-1;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0091-
1.pdf.

Briggs:2012:TN

[490] Rachael Briggs. Truthmaking without necessitation. Synthese, 189(1): 11–28, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0093-z;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0093-z.pdf.

Fine:2012:DPW

[491] Kit Fine. A difficulty for the possible worlds analysis of counterfactuals. Synthese, 189(1):29–57, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0094-y;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0094-
y.pdf.

Jago:2012:CW

[492] Mark Jago. Constructing worlds. Synthese, 189(1):59–74, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0095-x;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0095-x.pdf.

Rabern:2012:AIA

[493] Brian Rabern. Against the identification of assertoric content with compositional value. Synthese, 189(1):75–96, November 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0096-9;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0096-
9.pdf.

Ripley:2012:SCT

[494] David Ripley. Structures and circumstances: two ways to fine-grain propositions. Synthese, 189(1):97–118, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0100-4;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0100-
4.pdf.

Schaffer:2012:NP

[495] Jonathan Schaffer. Necessitarian propositions. Synthese, 189(1):119– 162, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0097-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0097-8.pdf.

Schellenberg:2012:SFS

[496] Susanna Schellenberg. Sameness of Fregean sense. Synthese, 189(1): 163–175, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0098-7;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0098-7.pdf.

Schroeter:2012:BOW

[497] Laura Schroeter. Bootstrapping our way to samesaying. Synthese, 189(1): 177–197, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0099-6;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0099-6.pdf.

Weber:2012:EPM

[498] Clas Weber. Eternalism and propositional multitasking: in defence of the operator argument. Synthese, 189(1):199–219, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0092-0;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0092-0.pdf.

Coliva:2012:I

[499] Annalisa Coliva, Sebastiano Moruzzi, and Giorgio Volpe. Introduction. Synthese, 189(2):221–234, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0174-z;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0174-
z.pdf.

Coliva:2012:VFW

[500] Annalisa Coliva. Varieties of failure (of warrant transmission: what else?!). Synthese, 189(2):235–254, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0006-6;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0006-
6.pdf.

Pritchard:2012:WGO

[501] Duncan Pritchard. Wittgenstein and the groundlessness of our believing. Synthese, 189(2):255–272, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0057-8;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0057-
8.pdf.

Wedgwood:2012:JI

[502] Ralph Wedgwood. Justified inference. Synthese, 189(2):273–295, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0012-8;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0012-8.pdf.

Avnur:2012:MFE

[503] Yuval Avnur. Mere faith and entitlement. Synthese, 189(2):297– 315, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0053-z;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0053-z.pdf.

Volpe:2012:CYB

[504] Giorgio Volpe. Cornerstones: You’d better believe them. Synthese, 189(2):317–336, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0051-1;   http://link.springer.com/
content/pdf/10.1007/s11229-011-0051-1.pdf.

Dodd:2012:ESA

[505] Dylan Dodd. Evidentialism and skeptical arguments. Synthese, 189(2): 337–352, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0067-1;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0067-1.pdf.

Millar:2012:SPK

[506] Alan Millar. Scepticism, perceptual knowledge, and doxastic responsibility. Synthese, 189(2):353–372, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0005-7;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0005-
7.pdf.

Gerken:2012:DJS

[507] Mikkel Gerken. Discursive justification and skepticism. Synthese, 189(2): 373–394, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0076-0;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0076-0.pdf.

Kallestrup:2012:BRG

[508] Jesper Kallestrup. Bootstrap and rollback: generalizing epistemic circularity. Synthese, 189(2):395–413, November 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9990-9;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9990-
9.pdf.

Andersen:2012:CRM

[509] Holly Andersen. The case for regularity in mechanistic causal explanation. Synthese, 189(3):415–432, December 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9965-x;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9965-
x.pdf.

Towl:2012:LCK

[510] Brandon Towl. Laws and constrained kinds: a lesson from motor neuroscience. Synthese, 189(3):433–450, December 2012. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9950-4;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9950-
4.pdf.

Garson:2012:FSC

[511] Justin Garson. Function, selection, and construction in the brain. Synthese, 189(3):451–481, December 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0122-y;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0122-y.pdf.

Herschbach:2012:MVS

[512] Mitchell Herschbach. Mirroring versus simulation: on the representational function of simulation. Synthese, 189(3):483–513, December 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9969-6;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9969-6.pdf.

Spaulding:2012:MNE

[513] Shannon Spaulding. Mirror neurons are not evidence for the simulation theory. Synthese, 189(3):515–534, December 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0086-y;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0086-
y.pdf.

Jennings:2012:SA

[514] Carolyn Dicey Jennings. The subject of attention. Synthese, 189(3): 535–554, December 2012. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0164-1;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0164-1.pdf.

Douven:2013:FEP

[515] Igor Douven. The formal epistemology project. Synthese, 190(1): 1–2, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0178-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0178-8.pdf.

DeCruz:2013:MSE

[516] Helen De Cruz and Johan De Smedt. Mathematical symbols as epistemic actions. Synthese, 190(1):3–19, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-010-9837-9;
 http://link.springer.com/content/pdf/10.1007/s11229-010-9837-
9.pdf.

Krzyzanowska:2013:BAR

[517] Karolina Krzyzanowska. Belief ascription and the Ramsey test. Synthese, 190(1):21–36, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0160-5;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0160-5.pdf.

Wenmackers:2013:FIL

[518] Sylvia Wenmackers and Leon Horsten. Fair infinite lotteries. Synthese, 190(1):37–61, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9836-x;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9836-x.pdf.

Fischer:2013:SRR

[519] Martin Fischer. Some remarks on restricting the knowability principle. Synthese, 190(1):63–88, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9833-0;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9833-0.pdf.

Heylen:2013:MEA

[520] Jan Heylen. Modal-epistemic arithmetic and the problem of quantifying in. Synthese, 190(1):89–111, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0154-3;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0154-
3.pdf.

Sequoiah-Grayson:2013:ECC

[521] Sebastian Sequoiah-Grayson. Epistemic closure and commutative, nonassociative residuated structures. Synthese, 190(1):113–128, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
010-9834-z;   http://link.springer.com/content/pdf/10.1007/
s11229-010-9834-z.pdf.

Chandler:2013:CCS

[522] Jake Chandler. Contrastive confirmation: some competing accounts. Synthese, 190(1):129–138, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-010-9845-9;   http://link.springer.com/
content/pdf/10.1007/s11229-010-9845-9.pdf.

Dietz:2013:CC

[523] Richard Dietz. Comparative concepts. Synthese, 190(1):139–170, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0152-5;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0152-5.pdf.

Douven:2013:PB

[524] Igor Douven and Christoph Kelp. Proper bootstrapping. Synthese, 190 (1):171–185, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0115-x;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0115-x.pdf.

Vorms:2013:P

[525] Marion Vorms and Christopher Pincock. Preface. Synthese, 190(2): 187–188, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0140-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0140-9.pdf.

Teller:2013:CMP

[526] Paul Teller. The concept of measurement-precision. Synthese, 190(2): 189–202, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0141-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0141-8.pdf.

Epstein:2013:PTH

[527] Brian Epstein and Patrick Forber. The perils of tweaking: how to use macrodata to set parameters in complex simulation models. Synthese, 190(2):203–218, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0142-7;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0142-7.pdf.

Bolinska:2013:ERI

[528] Agnes Bolinska. Epistemic representation, informativeness and the aim of faithful representation. Synthese, 190(2):219–234, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0143-6;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0143-6.pdf.

Purves:2013:FTF

[529] Gordon Michael Purves. Finding truth in fictions: identifying nonfictions in imaginary cracks. Synthese, 190(2):235–251, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0144-5;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0144-5.pdf.

Gelfert:2013:SMB

[530] Axel Gelfert. Strategies of model-building in condensed matter physics: trade-offs as a demarcation criterion between physics and biology? Synthese, 190(2):253–272, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0145-4;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0145-
4.pdf.

Gildenhuys:2013:CPG

[531] Peter Gildenhuys. Classical population genetics and the semantic approach to scientific theories. Synthese, 190(2):273–291, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0146-3;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0146-3.pdf.

Vorms:2013:MDT

[532] Marion Vorms. Models of data and theoretical hypotheses: a case-study in classical genetics. Synthese, 190(2):293–319, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0147-2;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0147-2.pdf.

Svetlova:2013:ICC

[533] Ekaterina Svetlova. De-idealization by commentary: the case of financial valuation models. Synthese, 190(2):321–337, January 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0148-1;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0148-1.pdf.

Whitcomb:2013:OWU

[534] Dennis Whitcomb. One wage of unknowability. Synthese, 190(3): 339–352, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0017-3;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0017-3.pdf.

Mork:2013:UCS

[535] Jonas Clausen Mork. Uncertainty, credal sets and second order probability. Synthese, 190(3):353–378, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0042-2;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0042-
2.pdf.

Diez:2013:GTE

[536] José Díez, Kareem Khalifa, and Bert Leuridan. General theories of explanation: buyer beware. Synthese, 190(3):379–396, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0020-8;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0020-8.pdf.

McEvoy:2013:EMC

[537] Mark McEvoy. Experimental mathematics, computers and the a priori. Synthese, 190(3):397–412, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0035-1;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0035-
1.pdf.

Hindriks:2013:LPS

[538] Frank Hindriks. The location problem in social ontology. Synthese, 190 (3):413–437, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0036-0;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0036-0.pdf.

Laraudogoitia:2013:ZFI

[539] Jon Pérez Laraudogoitia. Zeno and flow of information. Synthese, 190 (3):439–447, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0037-z;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0037-z.pdf.

Barros:2013:NCC

[540] D. Benjamin Barros. Negative causation in causal and mechanistic explanation. Synthese, 190(3):449–469, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0040-4;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0040-
4.pdf.

Matsubara:2013:RUS

[541] Keizo Matsubara. Realism, underdetermination and string theory dualities. Synthese, 190(3):471–489, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0041-3;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0041-
3.pdf.

Shaw:2013:SBR

[542] James R. Shaw. De se belief and rational choice. Synthese, 190(3): 491–508, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0044-0;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0044-0.pdf.

Hoeltje:2013:EI

[543] Miguel Hoeltje, Benjamin Schnieder, and Alex Steinberg. Explanation by induction? Synthese, 190(3):509–524, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0045-z;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0045-z.pdf.

Culina:2013:LPC

[544] Boris Culina. Logic of paradoxes in classical set theories. Synthese, 190 (3):525–547, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0047-x;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0047-x.pdf.

Lee:2013:SRC

[545] Chunghyoung Lee. The staccato roller coaster: a simple physical model of the staccato run. Synthese, 190(3):549–562, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0049-8;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0049-8.pdf.

Rotondo:2013:UCD

[546] Andrew Rotondo. Undermining, circularity, and disagreement. Synthese, 190(3):563–584, February 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0050-2;   http://link.springer.com/
content/pdf/10.1007/s11229-011-0050-2.pdf.

Tummolini:2013:CTA

[547] Luca Tummolini, Giulia Andrighetto, Cristiano Castelfranchi, and Rosaria Conte. A convention or (tacit) agreement betwixt us: on reliance and its normative consequences. Synthese, 190(4):585–618, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0194-8;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0194-8.pdf.

Bigaj:2013:HEC

[548] Tomasz Bigaj. How to evaluate counterfactuals in the quantum world. Synthese, 190(4):619–637, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0195-7;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0195-
7.pdf.

Brogaard:2013:RC

[549] Berit Brogaard and Joe Salerno. Remarks on counterpossibles. Synthese, 190(4):639–660, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0196-6;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0196-6.pdf. See critique [2085].

Cantwell:2013:CCD

[550] John Cantwell. Conditionals in causal decision theory. Synthese, 190 (4):661–679, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0197-5;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0197-5.pdf.

Djordjevic:2013:SC

[551] Vladan Djordjevic. Similarity and cotenability. Synthese, 190(4): 681–691, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0198-4;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0198-4.pdf.

Meyer:2013:MPC

[552] Ulrich Meyer. Modal property comprehension. Synthese, 190(4): 693–707, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0199-3;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0199-3.pdf.

Tennant:2013:PCP

[553] Neil Tennant. Parts, classes and parts of classes: an anti-realist reading of Lewisian mereology. Synthese, 190(4):709–742, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0200-1;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0200-1.pdf.

Torza:2013:HLK

[554] Alessandro Torza. How to Lewis a Kripke–Hintikka. Synthese, 190 (4):743–779, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0201-0;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0201-0.pdf.

Helzner:2013:EE

[555] Jeffrey Helzner. Epistemology and economics. Synthese, 190(5): 781–786, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0185-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0185-9.pdf.

Dietrich:2013:RPB

[556] Franz Dietrich and Christian List. Reasons for (prior) belief in Bayesian epistemology. Synthese, 190(5):787–808, March 2013. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0224-6;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0224-
6.pdf.

Eriksson:2013:IPB

[557] Lina Eriksson and Wlodek Rabinowicz. The interference problem for the betting interpretation of degrees of belief. Synthese, 190(5): 809–830, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0187-7;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0187-7.pdf.

Arlo-Costa:2013:FFH

[558] Horacio Arló-Costa and Arthur Paul Pedersen. Fast and frugal heuristics: rationality and the limits of naturalism. Synthese, 190(5):831–850, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0188-6;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0188-6.pdf.

Hill:2013:AE

[559] Brian Hill. Awareness and equilibrium. Synthese, 190(5):851–869, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0189-5;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0189-5.pdf.

Sillari:2013:RFC

[560] Giacomo Sillari. Rule-following as coordination: a game-theoretic approach. Synthese, 190(5):871–890, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-012-0190-z;   http://link.springer.
com/content/pdf/10.1007/s11229-012-0190-z.pdf.

Roy:2013:SAI

[561] Olivier Roy and Eric Pacuit. Substantive assumptions in interaction: a logical perspective. Synthese, 190(5):891–908, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0191-y;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0191-y.pdf.

Bicchieri:2013:SSB

[562] Cristina Bicchieri and Hugo Mercier. Self-serving biases and public justifications in trust games. Synthese, 190(5):909–922, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0192-x;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0192-x.pdf.

Skyrms:2013:CTS

[563] Brian Skyrms. The core theory of subjunctive conditionals. Synthese, 190 (5):923–928, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0193-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0193-9.pdf.

Helzner:2013:RTT

[564] Jeffrey Helzner. Rationalizing two-tiered choice functions through conditional choice. Synthese, 190(6):929–951, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0056-9;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0056-
9.pdf.

Asmus:2013:VRS

[565] C. M. Asmus. Vagueness and revision sequences. Synthese, 190(6): 953–974, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0052-0;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0052-0.pdf.

Mackonis:2013:IBE

[566] Adolfas Mackonis. Inference to the best explanation, coherence and other explanatory virtues. Synthese, 190(6):975–995, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0054-y;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0054-y.pdf.

Wilkenfeld:2013:URM

[567] Daniel A. Wilkenfeld. Understanding as representation manipulability. Synthese, 190(6):997–1016, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-0055-x;   http://link.springer.com/
content/pdf/10.1007/s11229-011-0055-x.pdf.

Glynn:2013:CFP

[568] Luke Glynn. Causal foundationalism, physical causation, and difference-making. Synthese, 190(6):1017–1037, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0058-7;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0058-
7.pdf.

Gardenfors:2013:TCD

[569] Peter Gärdenfors and Frank Zenker. Theory change as dimensional change: conceptual spaces applied to the dynamics of empirical theories. Synthese, 190(6):1039–1058, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-0060-0;
 http://link.springer.com/content/pdf/10.1007/s11229-011-0060-
0.pdf.

McMyler:2013:ESA

[570] Benjamin McMyler. The epistemic significance of address. Synthese, 190 (6):1059–1078, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9871-2;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9871-2.pdf.

Lam:2013:CPE

[571] Barry Lam. Calibrated probabilities and the epistemology of disagreement. Synthese, 190(6):1079–1098, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9881-0;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9881-
0.pdf.

Kim:2013:WN

[572] Joongol Kim. What are numbers? Synthese, 190(6):1099–1112, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9883-y;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9883-y.pdf. See comments [1986].

Wright:2013:DKI

[573] Stephen Wright. Does Klein’s infinitism offer a response to Agrippa’s trilemma? Synthese, 190(6):1113–1130, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9884-x;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9884-
x.pdf.

Vermaas:2013:DSA

[574] Pieter E. Vermaas, Massimiliano Carrara, Stefano Borgo, and Pawel Garbacz. The design stance and its artefacts. Synthese, 190(6):1131–1152, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9885-9;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9885-9.pdf.

Khalifa:2013:UEO

[575] Kareem Khalifa. Is understanding explanatory or objectual? Synthese, 190(6):1153–1171, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9886-8;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9886-8.pdf.

Petersen:2013:UE

[576] Steve Petersen. Utilitarian epistemology. Synthese, 190(6):1173–1184, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9887-7;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9887-7.pdf.

Ahlstrom-Vij:2013:EI

[577] Kristoffer Ahlstrom-Vij, Klemens Kappel, and Nikolaj Jang Lee Linding Pedersen. The epistemology of inclusiveness. Synthese, 190(7): 1185–1188, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0265-5;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0265-5.pdf.

Goldberg:2013:IFA

[578] Sanford C. Goldberg. Inclusiveness in the face of anticipated disagreement. Synthese, 190(7):1189–1207, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0102-2;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0102-
2.pdf.

Landemore:2013:DCD

[579] Hélène Landemore. Deliberation, cognitive diversity, and democratic inclusiveness: an epistemic argument for the random selection of representatives. Synthese, 190(7):1209–1231, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0062-6;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0062-
6.pdf.

Thompson:2013:GMG

[580] Christopher Thompson. A general model of a group search procedure, applied to epistemic democracy. Synthese, 190(7):1233–1252, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0108-9;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0108-9.pdf.

Peter:2013:PEV

[581] Fabienne Peter. The procedural epistemic value of deliberation. Synthese, 190(7):1253–1266, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0119-6;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0119-6.pdf.

Trout:2013:DSE

[582] J. D. Trout. Democracy and scientific expertise: illusions of political and epistemic inclusion. Synthese, 190(7):1267–1291, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0226-4;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0226-4.pdf.

Miller:2013:WCK

[583] Boaz Miller. When is consensus knowledge based? Distinguishing shared knowledge from mere agreement. Synthese, 190(7):1293–1316, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0225-5;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0225-5.pdf.

Fricker:2013:EJC

[584] Miranda Fricker. Epistemic justice as a condition of political freedom? Synthese, 190(7):1317–1332, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0227-3;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0227-
3.pdf.

Hartmann:2013:EFE

[585] Stephan Hartmann, Chiara Lisciandra, and Edouard Machery. Editorial: Formal epistemology meets experimental philosophy. Synthese, 190 (8):1333–1335, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0269-1;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0269-1.pdf.

Colyvan:2013:INM

[586] Mark Colyvan. Idealisations in normative models. Synthese, 190(8): 1337–1350, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0166-z;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0166-z.pdf.

Devetag:2013:YBP

[587] Giovanna Devetag, Hykel Hosni, and Giacomo Sillari. You better play 7: mutual versus common knowledge of advice in a weaklink experiment. Synthese, 190(8):1351–1381, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0177-9;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0177-
9.pdf.

Overton:2013:ESD

[588] James A. Overton. “Explain” in scientific discourse. Synthese, 190 (8):1383–1405, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0109-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0109-8.pdf.

Unterhuber:2013:NTP

[589] Matthias Unterhuber and Gerhard Schurz. The new tweety puzzle: arguments against monistic Bayesian approaches in epistemology and cognitive science. Synthese, 190(8):1407–1435, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0159-y;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0159-y.pdf.

Vallinder:2013:DCS

[590] Aron Vallinder and Erik J. Olsson. Do computer simulations support the argument from disagreement? Synthese, 190(8):1437–1454, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0107-x;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0107-x.pdf.

Wagner:2013:CP

[591] Carl G. Wagner. The corroboration paradox. Synthese, 190(8): 1455–1469, May 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0106-y;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0106-y.pdf.

Northcott:2013:VCA

[592] Robert Northcott. Verisimilitude: a causal approach. Synthese, 190 (9):1471–1488, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9895-7;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9895-7.pdf.

Pust:2013:SBE

[593] Joel Pust. Sleeping Beauty, evidential support and indexical knowledge: reply to Horgan. Synthese, 190(9):1489–1501, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9888-6;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9888-6.pdf.

Bertrand:2013:PES

[594] Michael Bertrand. Proper environment and the SEP account of biological function. Synthese, 190(9):1503–1517, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9889-5;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9889-
5.pdf.

Pynn:2013:BET

[595] Geoff Pynn. The Bayesian explanation of transmission failure. Synthese, 190(9):1519–1531, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9890-z;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9890-z.pdf.

Zamora-Bonilla:2013:WGT

[596] Jesús Zamora-Bonilla. Why are good theories good? reflections on epistemic values, confirmation, and formal epistemology. Synthese, 190 (9):1533–1553, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9893-9;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9893-9.pdf.

Spanos:2013:FIP

[597] Aris Spanos. A frequentist interpretation of probability for modelbased inductive inference. Synthese, 190(9):1555–1585, June 2013. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9892-x;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9892-
x.pdf.

Thurow:2013:DVB

[598] Joshua C. Thurow. The defeater version of Benacerraf’s problem for a priori knowledge. Synthese, 190(9):1587–1603, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9894-8;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9894-8.pdf.

Wen:2013:SGC

[599] Xuefeng Wen and Shier Ju. Semantic games with chance moves revisited: from IF logic to partial logic. Synthese, 190(9):1605–1620, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9897-5;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9897-5.pdf.

Bain:2013:CTS

[600] Jonathan Bain. Category-theoretic structure and radical ontic structural realism. Synthese, 190(9):1621–1635, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9896-6;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9896-
6.pdf.

Wolpert:2013:LNP

[601] David H. Wolpert and Gregory Benford. The lesson of Newcomb’s paradox. Synthese, 190(9):1637–1646, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9899-3;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9899-
3.pdf.

Oddie:2013:CCL

[602] Graham Oddie. The content, consequence and likeness approaches to verisimilitude: compatibility, trivialization, and underdetermination. Synthese, 190(9):1647–1687, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9930-8;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9930-8.pdf.

French:2013:EPM

[603] Rohan French. Expressive power, mood, and actuality. Synthese, 190 (9):1689–1699, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9933-5;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9933-5.pdf.

Lockard:2013:EI

[604] Matthew Lockard. Epistemic instrumentalism. Synthese, 190(9): 1701–1718, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9932-6;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9932-6.pdf.

Wray:2013:STR

[605] K. Brad Wray. Success and truth in the realism/anti-realism debate. Synthese, 190(9):1719–1729, June 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9931-7;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9931-7.pdf.

Pagin:2013:PLM

[606] Peter Pagin, Robert van Rooij, and Jonas Akerman. Philosophy of language and mind. Synthese, 190(10):1731–1733, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0258-4;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0258-4.pdf.

French:2013:PES

[607] Craig French. Perceptual experience and seeing that p. Synthese, 190 (10):1735–1751, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0259-3;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0259-3.pdf.

Fritz:2013:LET

[608] Peter Fritz. A logic for epistemic two-dimensional semantics. Synthese, 190(10):1753–1770, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0260-x;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0260-x.pdf.

Hansen:2013:SIC

[609] Nat Hansen. A slugfest of intuitions: contextualism and experimental design. Synthese, 190(10):1771–1792, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0261-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0261-
9.pdf.

Guillot:2013:LSS

[610] Marie Guillot. The limits of selflessness: semantic relativism and the epistemology of de se thoughts. Synthese, 190(10):1793–1816, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0262-8;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0262-8.pdf.

Pacherie:2013:IJA

[611] Elisabeth Pacherie. Intentional joint agency: shared intention lite. Synthese, 190(10):1817–1839, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0263-7;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0263-7.pdf.

Recanati:2013:PCD

[612] François Recanati. Perceptual concepts: in defence of the indexical model. Synthese, 190(10):1841–1855, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0264-6;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0264-
6.pdf.

Hoffmann:2013:PI

[613] Michael H. G. Hoffmann, Jan C. Schmidt, and Nancy J. Nersessian. Philosophy of and as interdisciplinarity. Synthese, 190(11):1857–1864, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0214-8;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0214-8.pdf.

Holbrook:2013:WIC

[614] J. Britt Holbrook. What is interdisciplinary communication? Reflections on the very idea of disciplinary integration. Synthese, 190(11):1865–1879, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0179-7;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0179-7.pdf.

Andersen:2013:EDI

[615] Hanne Andersen and Susann Wagenknecht. Epistemic dependence in interdisciplinary groups. Synthese, 190(11):1881–1898, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0172-1;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0172-1.pdf.

Fuller:2013:DIP

[616] Steve Fuller. Deviant interdisciplinarity as philosophical practice: prolegomena to deep intellectual history. Synthese, 190(11):1899–1916, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0208-6;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0208-6.pdf.

Frodeman:2013:PD

[617] Robert Frodeman. Philosophy dedisciplined. Synthese, 190(11):1917– 1936, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0181-0;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0181-0.pdf.

ORourke:2013:PIC

[618] Michael O’Rourke and Stephen J. Crowley. Philosophical intervention and cross-disciplinary science: the story of the toolbox project. Synthese, 190(11):1937–1954, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0175-y;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0175-
y.pdf.

Tuana:2013:EPP

[619] Nancy Tuana. Embedding philosophers in the practices of science: bringing humanities to the sciences. Synthese, 190(11):1955–1973, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0171-2;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0171-2.pdf.

Hoffmann:2013:EPI

[620] Michael H. G. Hoffmann, Jan C. Schmidt, and Nancy J. Nersessian. Erratum to: Philosophy of and as interdisciplinarity. Synthese, 190(11):1975, July 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0266-4;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0266-4.pdf.

Marusic:2013:SKG

[621] Berislav Marusić. The self-knowledge gambit. Synthese, 190(12): 1977–1999, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9949-x;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9949-x.pdf.

Bartels:2013:WMP

[622] Andreas Bartels. Why metrical properties are not powers. Synthese, 190(12):2001–2013, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9951-3;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9951-3.pdf.

Alexander:2013:AVF

[623] Samuel Alexander. An axiomatic version of Fitch’s paradox. Synthese, 190(12):2015–2020, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9954-0;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9954-0.pdf.

Brown:2013:IEH

[624] Jessica Brown. Intuitions, evidence and hopefulness. Synthese, 190(12): 2021–2046, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9952-2;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9952-2.pdf.

Torrengo:2013:GPP

[625] Giuliano Torrengo. The grounding problem and presentist explanations. Synthese, 190(12):2047–2063, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9955-z;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9955-
z.pdf.

Tallant:2013:OPP

[626] Jonathan Tallant. Optimus prime: paraphrasing prime number talk. Synthese, 190(12):2065–2083, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9959-8;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9959-
8.pdf.

Rast:2013:CDR

[627] Erich H. Rast. On contextual domain restriction in categorial grammar. Synthese, 190(12):2085–2115, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9960-2;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9960-
2.pdf.

Pitts:2013:ICR

[628] J. Brian Pitts. Irrelevant conjunction and the ratio measure or historical skepticism. Synthese, 190(12):2117–2139, August 2013. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9961-1;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9961-
1.pdf.

Marquis:2013:MFF

[629] Jean-Pierre Marquis. Mathematical forms and forms of mathematics: leaving the shores of extensional mathematics. Synthese, 190(12): 2141–2164, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9962-0;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9962-0.pdf.

Warglien:2013:SCS

[630] Massimo Warglien and Peter Gärdenfors. Semantics, conceptual spaces, and the meeting of minds. Synthese, 190(12):2165–2193, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9963-z;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9963-z.pdf.

Seselja:2013:AAE

[631] Dunja Seselja and Christian Straßer. Abstract argumentation and explanation applied to scientific debates. Synthese, 190(12):2195–2217, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9964-y;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9964-y.pdf.

Nottelmann:2013:DCE

[632] Nikolaj Nottelmann. The deontological conception of epistemic justification: a reassessment. Synthese, 190(12):2219–2241, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9967-8;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9967-8.pdf.

Werndl:2013:CBD

[633] Charlotte Werndl. On choosing between deterministic and indeterministic models: underdetermination and indirect evidence. Synthese, 190(12):

2243–2265, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9966-9;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9966-9.pdf.

Nagatsu:2013:LUT

[634] Michiru Nagatsu. The limits of unification for theory appraisal: a case of economics and psychology. Synthese, 190(12):2267–2289, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9971-z;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9971-z.pdf.

Charlow:2013:WWK

[635] Nate Charlow. What we know and what to do. Synthese, 190(12): 2291–2323, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9974-9;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9974-9.pdf.

Douskos:2013:LAI

[636] Christos Douskos. The linguistic argument for intellectualism. Synthese, 190(12):2325–2340, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9972-y;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9972-y.pdf.

Montminy:2013:RCC

[637] Martin Montminy. The role of context in contextualism. Synthese, 190(12):2341–2366, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-011-9975-8;   http://link.springer.com/
content/pdf/10.1007/s11229-011-9975-8.pdf.

Grim:2013:HSF

[638] Patrick Grim, Robert Rosenberger, Adam Rosenfeld, Brian Anderson, and Robb E. Eason. How simulations fail. Synthese, 190(12): 2367–2390, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9976-7;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9976-7.pdf.

Stegenga:2013:ITA

[639] Jacob Stegenga. An impossibility theorem for amalgamating evidence. Synthese, 190(12):2391–2411, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9973-x;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9973-
x.pdf.

Baron:2013:TIA

[640] Sam Baron. A truthmaker indispensability argument. Synthese, 190(12): 2413–2427, August 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-9989-2;   http://link.springer.com/content/
pdf/10.1007/s11229-011-9989-2.pdf.

Michaelian:2013:IEC

[641] Kourken Michaelian. The information effect: constructive memory, testimony, and epistemic luck. Synthese, 190(12):2429–2456, August 2013. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-011-9992-7;   https://link.springer.com/content/pdf/
10.1007/s11229-011-9992-7.pdf.

Williams:2013:CPS

[642] John N. Williams. The completeness of the pragmatic solution to Moore’s paradox in belief: a reply to Chan. Synthese, 190(12):2457– 2476, August 2013. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-011-9991-8;   https://link.springer.com/content/pdf/
10.1007/s11229-011-9991-8.pdf. See [41].

Moretti:2013:ETI

[643] L. Moretti and N. J. L. L. Pedersen. Epistemic transmission and interaction. Synthese, 190(13):2477–2479, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0289-x;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0289-
x.pdf.

Moretti:2013:WWT

[644] Luca Moretti and Tommaso Piazza. When warrant transmits and when it doesn’t: towards a general framework. Synthese, 190(13):2481– 2503, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-011-0018-2;   http://link.springer.com/content/
pdf/10.1007/s11229-011-0018-2.pdf.

Bjerring:2013:IWL

[645] Jens Christian Bjerring. Impossible worlds and logical omniscience: an impossibility result. Synthese, 190(13):2505–2524, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0038-y;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0038-y.pdf.

Shogenji:2013:CCT

[646] Tomoji Shogenji. Coherence of the contents and the transmission of probabilistic support. Synthese, 190(13):2525–2545, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-0003-9;   http://link.springer.com/content/pdf/10.1007/
s11229-011-0003-9.pdf.

DeLanghe:2013:PDU

[647] Rogier De Langhe. Peer disagreement under multiple epistemic systems. Synthese, 190(13):2547–2556, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0149-0;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0149-
0.pdf.

Olsson:2013:NAC

[648] Erik J. Olsson and Aron Vallinder. Norms of assertion and communication in social networks. Synthese, 190(13):2557–2571, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0313-1;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0313-1.pdf.

Pedersen:2013:EAB

[649] Nikolaj Jang Lee Linding Pedersen and Jesper Kallestrup. The epistemology of absence-based inference. Synthese, 190(13):2573–2593, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0255-7;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0255-7.pdf.

Aloni:2013:KW

[650] Maria Aloni, Paul Égré, and Tikitu de Jager. Knowing whether A or B. Synthese, 190(14):2595–2621, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-009-9646-1;
 http://link.springer.com/content/pdf/10.1007/s11229-009-9646-
1.pdf.

Grune-Yanoff:2013:PCC

[651] Till Grüne-Yanoff. Preference change and conservatism: comparing the Bayesian and the AGM models of preference revision. Synthese, 190(14): 2623–2641, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-010-9863-7;   http://link.springer.com/content/
pdf/10.1007/s11229-010-9863-7.pdf.

Wedgwood:2013:GSN

[652] Ralph Wedgwood. Gandalf’s solution to the Newcomb problem. Synthese, 190(14):2643–2675, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-011-9900-1;
 http://link.springer.com/content/pdf/10.1007/s11229-011-9900-
1.pdf.

Maffezioli:2013:CFK

[653] Paolo Maffezioli, Alberto Naibo, and Sara Negri. The Church– Fitch knowability paradox in the light of structural proof theory. Synthese, 190(14):2677–2716, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0061-7;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0061-
7.pdf.

Moon:2013:REK

[654] Andrew Moon. Remembering entails knowing. Synthese, 190(14):2717– 2729, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0065-3;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0065-3.pdf.

Sharon:2013:ECU

[655] Assaf Sharon and Levi Spectre. Epistemic closure under deductive inference: what is it and can we afford it? Synthese, 190(14):2731– 2748, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0078-y;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0078-y.pdf.

Weaver:2013:CFP

[656] Christopher Gregory Weaver. A Church–Fitch proof for the universality of causation. Synthese, 190(14):2749–2772, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0079-x;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0079-x.pdf.

Sullivan:2013:MPC

[657] Arthur Sullivan. Multiple propositions, contextual variability, and the semantics/pragmatics interface. Synthese, 190(14):2773–2800, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0084-0;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0084-0.pdf.

Dukich:2013:TTE

[658] John M. Dukich. Two types of empirical adequacy: a partial structures approach. Synthese, 190(14):2801–2820, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0085-z;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0085-z.pdf.

Rowbottom:2013:EEC

[659] Darrell P. Rowbottom. Empirical evidence claims are a priori. Synthese, 190(14):2821–2834, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0087-x;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0087-x.pdf.

Hetherington:2013:CKA

[660] Stephen Hetherington. Concessive knowledge-attributions: fallibilism and gradualism. Synthese, 190(14):2835–2851, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0088-9;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0088-9.pdf.

Mazzola:2013:CDE

[661] Claudio Mazzola. Correlations, deviations and expectations: the extended principle of the common cause. Synthese, 190(14):2853– 2866, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0089-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0089-8.pdf.

Scorzato:2013:RSS

[662] Luigi Scorzato. On the role of simplicity in science. Synthese, 190(14): 2867–2895, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0101-3;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0101-3.pdf.

Peregrin:2013:CLF

[663] Jaroslav Peregrin and Vladimír Svoboda. Criteria for logical formalization. Synthese, 190(14):2897–2924, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0104-0;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0104-
0.pdf.

Laraudogoitia:2013:ND

[664] Jon Pérez Laraudogoitia. On Norton’s dome. Synthese, 190(14):2925– 2941, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0105-z;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0105-z.pdf.

Evans:2013:PBR

[665] Ian Evans. The problem of the basing relation. Synthese, 190(14):2943– 2957, September 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0111-1;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0111-1.pdf.

Tallant:2013:IP

[666] Jonathan Tallant. Intuitions in physics. Synthese, 190(15):2959– 2980, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0113-z;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0113-z.pdf.

Beardman:2013:NFD

[667] Stephanie Beardman. A non-factualist defense of the reflection principle. Synthese, 190(15):2981–2999, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0114-y;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0114-
y.pdf.

Brasoveanu:2013:GQF

[668] Adrian Brasoveanu. The grammar of quantification and the fine structure of interpretation contexts. Synthese, 190(15):3001–3051, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0118-7;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0118-7.pdf.

Jackson:2013:CAE

[669] Magdalena Balcerak Jackson. Conceptual analysis and epistemic progress. Synthese, 190(15):3053–3074, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0120-0;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0120-
0.pdf.

Jonsson:2013:SMJ

[670] Martin L. Jönsson and Elias Assarsson. Shogenji’s measure of justification and the inverse conjunction fallacy. Synthese, 190(15):3075– 3085, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0125-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0125-8.pdf.

Northcott:2013:DE

[671] Robert Northcott. Degree of explanation. Synthese, 190(15):3087– 3105, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0124-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0124-9.pdf.

Guala:2013:NLC

[672] Francesco Guala. The normativity of Lewis conventions. Synthese, 190(15):3107–3122, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0131-x;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0131-x.pdf.

Cariani:2013:AR

[673] Fabrizio Cariani. Aggregating with reason(s). Synthese, 190(15):3123– 3147, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0133-8;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0133-8.pdf.

Martini:2013:PAB

[674] Carlo Martini. A puzzle about belief updating. Synthese, 190(15): 3149–3160, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0132-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0132-9.pdf.

Rieger:2013:CMP

[675] Adam Rieger. Conditionals are material: the positive arguments. Synthese, 190(15):3161–3174, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0134-7;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0134-7.pdf.

Azzouni:2013:INL

[676] Jody Azzouni. Inconsistency in natural languages. Synthese, 190(15): 3175–3184, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0136-5;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0136-5.pdf.

Budzynska:2013:CES

[677] Katarzyna Budzynska. Circularity in ethotic structures. Synthese, 190(15):3185–3207, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0135-6;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0135-6.pdf.

Mizrahi:2013:PIB

[678] Moti Mizrahi. The pessimistic induction: a bad argument gone too far. Synthese, 190(15):3209–3226, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0138-3;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0138-
3.pdf.

Alegre:2013:PNS

[679] María Caamaño Alegre. Pragmatic norms in science: making them explicit. Synthese, 190(15):3227–3246, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0150-7;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0150-
7.pdf.

Spector:2013:MES

[680] David Spector. Margin for error semantics and signal perception. Synthese, 190(15):3247–3263, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0155-2;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0155-2.pdf.

Hermens:2013:SQM

[681] Ronnie Hermens. Speakable in quantum mechanics. Synthese, 190(15): 3265–3286, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0158-z;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0158-z.pdf.

Baave:2013:FD

[682] Arvid Båve. Formulating deflationism. Synthese, 190(15):3287–3305, October 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0163-2;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0163-2.pdf.

Cevolani:2013:VBC

[683] Gustavo Cevolani, Roberto Festa, and Theo A. F. Kuipers. Verisimilitude and belief change for nomic conjunctive theories. Synthese, 190(16):

3307–3324, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0165-0;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0165-0.pdf.

Hansson:2013:MPC

[684] Sven Ove Hansson. Maximal and perimaximal contraction. Synthese, 190(16):3325–3348, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0167-y;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0167-y.pdf.

Artemov:2013:DKS

[685] Sergei Artemov and Tudor Protopopescu. Discovering knowability: a semantic analysis. Synthese, 190(16):3349–3376, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0168-x;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0168-x.pdf.

Schwarz:2013:VMT

[686] Wolfgang Schwarz. Variations on a Montagovian theme. Synthese, 190(16):3377–3395, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0173-0;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0173-0.pdf.

Kralemann:2013:MIM

[687] Björn Kralemann and Claas Lattmann. Models as icons: modeling models in the semiotic framework of Peirce’s theory of signs. Synthese, 190(16):3397–3420, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0176-x;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0176-x.pdf.

Chen:2013:AIV

[688] Min Chen and Luciano Floridi. An analysis of information visualisation. Synthese, 190(16):3421–3438, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0183-y;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0183-
y.pdf.

Halpin:2013:BAA

[689] John Halpin. Briggs on antirealist accounts of scientific law. Synthese, 190(16):3439–3449, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0202-z;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0202-z.pdf.

ElSkaf:2013:UES

[690] Rawad El Skaf and Cyrille Imbert. Unfolding in the empirical sciences: experiments, thought experiments and computer simulations. Synthese, 190(16):3451–3474, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0203-y;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0203-
y.pdf.

Arenhart:2013:WI

[691] Jonas Rafael Becker Arenhart. Wither away individuals. Synthese, 190(16):3475–3494, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0204-x;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0204-x.pdf.

Lavelle:2013:CEM

[692] Jane Suilin Lavelle, George Botterill, and Suzanne Lock. Contrastive explanation and the many absences problem. Synthese, 190(16):3495– 3510, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0205-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0205-9.pdf.

Park:2013:SBU

[693] Ilho Park. Simultaneous belief updates via successive Jeffrey conditionalization. Synthese, 190(16):3511–3533, November 2013. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0207-7;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0207-
7.pdf.

Staffel:2013:CTR

[694] Julia Staffel. Can there be reasoning with degrees of belief? Synthese, 190(16):3535–3551, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0209-5;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0209-5.pdf.

Betz:2013:JIB

[695] Gregor Betz. Justifying inference to the best explanation as a practical meta-syllogism on dialectical structures. Synthese, 190(16):3553– 3578, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0210-z;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0210-z.pdf.

Corner:2013:NTA

[696] Adam Corner and Ulrike Hahn. Normative theories of argumentation: are some norms better than others? Synthese, 190(16):3579– 3610, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0211-y;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0211-y.pdf.

Isaac:2013:MR

[697] Alistair M. C. Isaac. Modeling without representation. Synthese, 190(16): 3611–3623, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0213-9;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0213-9.pdf.

Tanona:2013:DCC

[698] Scott Tanona. Decoherence and the Copenhagen cut. Synthese, 190(16): 3625–3649, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0216-6;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0216-6.pdf.

Peels:2013:DDR

[699] Rik Peels. Does doxastic responsibility entail the ability to believe otherwise? Synthese, 190(17):3651–3669, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0217-5;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0217-
5.pdf.

daCosta:2013:MOP

[700] Newton da Costa, Olimpia Lombardi, and Mariano Lastiri. A modal ontology of properties for quantum mechanics. Synthese, 190(17):3671– 3693, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0218-4;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0218-4.pdf.

Berry:2013:DRM

[701] Sharon Berry. Default reasonableness and the mathoids. Synthese, 190(17):3695–3713, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0219-3;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0219-3.pdf.

Carter:2013:KVC

[702] J. Adam Carter, Benjamin Jarvis, and Katherine Rubin. Knowledge and the value of cognitive ability. Synthese, 190(17):3715–3729, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0220-x;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0220-x.pdf.

Tulodziecki:2013:UMP

[703] Dana Tulodziecki. Underdetermination, methodological practices, and realism. Synthese, 190(17):3731–3750, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0221-9;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0221-
9.pdf.

Becker:2013:WRD

[704] Kelly Becker. Why reliabilism does not permit easy knowledge. Synthese, 190(17):3751–3775, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0222-8;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0222-8.pdf.

Pelling:2013:AS

[705] Charlie Pelling. Assertion and safety. Synthese, 190(17):3777–3796, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.

1007/s11229-012-0223-7;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0223-7.pdf.

Bowen:2013:DEP

[706] Adam J. Bowen. Dissolving an epistemological puzzle of time perception. Synthese, 190(17):3797–3817, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0228-2;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0228-
2.pdf.

Giovanelli:2013:TCP

[707] Marco Giovanelli. Talking at cross-purposes: how Einstein and the logical empiricists never agreed on what they were disagreeing about. Synthese, 190(17):3819–3863, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0229-1;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0229-
1.pdf.

Mantel:2013:ARA

[708] Susanne Mantel. Acting for reasons, apt action, and knowledge. Synthese, 190(17):3865–3888, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0230-8;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0230-
8.pdf.

Cowling:2013:IP

[709] Sam Cowling. Ideological parsimony. Synthese, 190(17):3889–3908, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0231-7;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0231-7.pdf.

Ramsey:2013:D

[710] Grant Ramsey. Driftability. Synthese, 190(17):3909–3928, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0232-6;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0232-6.pdf.

Baril:2013:PEA

[711] Anne Baril. Pragmatic encroachment in accounts of epistemic excellence. Synthese, 190(17):3929–3952, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0234-4;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0234-
4.pdf.

Johnson:2013:BTO

[712] Daniel M. Johnson. B-theory old and new: on ontological commitment. Synthese, 190(17):3953–3970, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0239-z;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0239-
z.pdf.

Roth:2013:FPS

[713] Martin Roth. Folk psychology as science. Synthese, 190(17):3971– 3982, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0240-6;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0240-6.pdf.

Carter:2013:HMO

[714] Jessica Carter. Handling mathematical objects: representations and context. Synthese, 190(17):3983–3999, November 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0241-5;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0241-
5.pdf.

Holm:2013:NZP

[715] Ruurik Holm. Non-zero probabilities for universal generalizations. Synthese, 190(18):4001–4007, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-013-0244-x;   http://link.springer.
com/content/pdf/10.1007/s11229-013-0244-x.pdf.

Lewis:2013:DAS

[716] Peter J. Lewis. The Doomsday argument and the simulation argument. Synthese, 190(18):4009–4022, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0245-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0245-
9.pdf.

Baskent:2013:STP

[717] Can Baskent. Some topological properties of paraconsistent models. Synthese, 190(18):4023–4040, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-013-0246-8;   http://link.springer.
com/content/pdf/10.1007/s11229-013-0246-8.pdf.

Brun:2013:IEA

[718] Georg Brun and Hans Rott. Interpreting enthymematic arguments using belief revision. Synthese, 190(18):4041–4063, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0248-6;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0248-6.pdf.

Stapleford:2013:IED

[719] Scott Stapleford. Imperfect epistemic duties and the justificational fecundity of evidence. Synthese, 190(18):4065–4075, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0249-5;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0249-5.pdf.

Beebe:2013:WWR

[720] James R. Beebe. Weakness of will, reasonability, and compulsion. Synthese, 190(18):4077–4093, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0250-z;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0250-
z.pdf.

Centrone:2013:NMD

[721] Stefania Centrone. Notes on Mally’s deontic logic and the collapse of Seinsollen and Sein. Synthese, 190(18):4095–4116, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0251-y;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0251-y.pdf.

Levin:2013:AME

[722] Janet Levin. Armchair methodology and epistemological naturalism. Synthese, 190(18):4117–4136, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0253-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0253-
9.pdf.

Schindler:2013:KMH

[723] Samuel Schindler. The Kuhnian mode of HPS. Synthese, 190(18):4137– 4154, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0252-x;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0252-x.pdf.

Arciszewski:2013:RDL

[724] Michal Arciszewski. Reducing the Dauer larva: molecular models of biological phenomena in Caenorhabditis elegans research. Synthese, 190(18): 4155–4179, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0254-8;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0254-8.pdf.

Otero:2013:PRN

[725] Manuel Pérez Otero. Purposes of reasoning and (a new vindication of) Moore’s proof of an external world. Synthese, 190(18):4181–4200, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0256-6;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0256-6.pdf.

Carter:2013:ECE

[726] J. Adam Carter. Extended cognition and epistemic luck. Synthese, 190(18):4201–4214, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0267-3;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0267-3.pdf.

Tohme:2013:AEC

[727] Fernando Tohmé and Ricardo Crespo. Abduction in economics: a conceptual framework and its model. Synthese, 190(18):4215–4237, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0268-2;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0268-2.pdf.

Ross:2013:EPP

[728] Peter W. Ross and Dale Turner. Existence problems in philosophy and science. Synthese, 190(18):4239–4259, December 2013. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0270-8;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0270-
8.pdf.

Percival:2013:BPW

[729] Philip Percival. Branching of possible worlds. Synthese, 190(18):4261– 4291, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0271-7;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0271-7.pdf.

Blome-Tillmann:2013:KI

[730] Michael Blome-Tillmann. Knowledge and implicatures. Synthese, 190 (18):4293–4319, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0274-4;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0274-4.pdf.

Wray:2013:PIE

[731] K. Brad Wray. The pessimistic induction and the exponential growth of science reassessed. Synthese, 190(18):4321–4330, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0276-2;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0276-2.pdf.

Deutsch:2013:CT

[732] David Deutsch. Constructor theory. Synthese, 190(18):4331–4359, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0279-z;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0279-z.pdf.

Briesen:2013:RBE

[733] Jochen Briesen. Reliabilism, bootstrapping, and epistemic circularity. Synthese, 190(18):4361–4372, December 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0272-6;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0272-
6.pdf.

Imbert:2014:ISS

[734] Cyrille Imbert, Ryan Muldoon, Jan Sprenger, and Kevin Zollman. Introduction, SI of Synthese “The collective dimension of science”. Synthese, 191(1):1–2, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0349-2;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0349-2.pdf.

Martini:2014:ESV

[735] Carlo Martini. Experts in science: a view from the trenches. Synthese, 191(1):3–15, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0321-1;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0321-1.pdf.

Boyer:2014:BHW

[736] Thomas Boyer. Is a bird in the hand worth two in the bush? Or, whether scientists should publish intermediate results. Synthese, 191 (1):17–35, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0242-4;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0242-4.pdf.

deRidder:2014:EDC

[737] Jeroen de Ridder. Epistemic dependence and collective scientific knowledge. Synthese, 191(1):37–53, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0283-3;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0283-
3.pdf.

Mayo-Wilson:2014:RTN

[738] Conor Mayo-Wilson. Reliability of testimonial norms in scientific communities. Synthese, 191(1):55–78, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0320-2;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0320-
2.pdf.

Green:2014:EDC

[739] Adam Green. Evaluating distributed cognition. Synthese, 191(1): 79–95, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0305-1;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0305-1.pdf.

Bonilla:2014:NCA

[740] Jesús Zamora Bonilla. The nature of co-authorship: a note on recognition sharing and scientific argumentation. Synthese, 191(1):97–108, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0238-0;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0238-0.pdf.

Vaesen:2014:MTS

[741] Krist Vaesen and Wybo Houkes. Modelling the truth of scientific beliefs with cultural evolutionary theory. Synthese, 191(1):109–125, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0257-5;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0257-5.pdf.

Chirimuuta:2014:MMC

[742] M. Chirimuuta. Minimal models and canonical neural computations: the distinctness of computational explanation in neuroscience. Synthese, 191 (2):127–153, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0369-y;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0369-y.pdf.

DeBrigard:2014:MRR

[743] Felipe De Brigard. Is memory for remembering? Recollection as a form of episodic hypothetical thinking. Synthese, 191(2):155–185, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0247-7;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0247-7.pdf.

Koralus:2014:ACS

[744] Philipp Koralus. Attention, consciousness, and the semantics of questions. Synthese, 191(2):187–211, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0382-1;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0382-
1.pdf.

Morgan:2014:RGM

[745] Alex Morgan. Representations gone mental. Synthese, 191(2):213– 244, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0328-7;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0328-7.pdf.

Schlosser:2014:NSF

[746] Markus E. Schlosser. The neuroscientific study of free will: A diagnosis of the controversy. Synthese, 191(2):245–262, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0312-2;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0312-2.pdf.

Sebastian:2014:DEW

[747] Miguel Ángel Sebastián. Dreams: an empirical way to settle the discussion between cognitive and non-cognitive theories of consciousness. Synthese, 191(2):263–285, January 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0385-y;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0385-
y.pdf.

Secchi:2014:MTA

[748] Luigi Secchi. The main two arguments for probabilism are flawed. Synthese, 191(3):287–295, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0286-0;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0286-0.pdf.

DeBruin:2014:DPF

[749] L. C. De Bruin and A. Newen. The developmental paradox of false belief understanding: a dual-system solution. Synthese, 191(3):297–320, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0127-6;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0127-6.pdf.

Hollebrandse:2014:CFS

[750] Bart Hollebrandse, Angeliek van Hout, and Petra Hendriks. Children’s first and second-order false-belief reasoning in a verbal and a low-verbal task. Synthese, 191(3):321–333, February 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0169-9;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0169-
9.pdf.

vanderVaart:2014:TMA

[751] Elske van der Vaart and Charlotte K. Hemelrijk. ‘Theory of mind’ in animals: ways to make progress. Synthese, 191(3):335–354, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0170-3;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0170-3.pdf.

Raijmakers:2014:CSU

[752] Maartje E. J. Raijmakers, Dorothy J. Mandell, Sara E. van Es, and Marian Counihan. Children’s strategy use when playing strategic games. Synthese, 191(3):355–370, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-012-0212-x;
 http://link.springer.com/content/pdf/10.1007/s11229-012-0212-
x.pdf.

Degremont:2014:ETB

[753] Cédric Dégremont, Lena Kurzen, and Jakub Szymanik. Exploring the tractability border in epistemic tasks. Synthese, 191(3):371–408, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
012-0215-7;   http://link.springer.com/content/pdf/10.1007/
s11229-012-0215-7.pdf.

Demey:2014:ADP

[754] Lorenz Demey. Agreeing to disagree in probabilistic dynamic epistemic logic. Synthese, 191(3):409–438, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0280-6;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0280-
6.pdf.

Kumar:2014:KNK

[755] Victor Kumar. ‘Knowledge’ as a natural kind term. Synthese, 191(3): 439–457, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0281-5;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0281-5.pdf.

Baron:2014:OME

[756] Sam Baron. Optimisation and mathematical explanation: doing the Lévy walk. Synthese, 191(3):459–479, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0284-2;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0284-
2.pdf.

Berto:2014:WED

[757] Francesco Berto and Jacopo Tagliabue. The world is either digital or analogue. Synthese, 191(3):481–497, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0285-1;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0285-
1.pdf.

Fontaine:2014:TSA

[758] Matthieu Fontaine and Shahid Rahman. Towards a semantics for the artifactual theory of fiction and beyond. Synthese, 191(3):499–516, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0287-z;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0287-z.pdf.

DAgostino:2014:VIY

[759] Fred D’Agostino. Verballed? Incommensurability 50 years on. Synthese, 191(3):517–538, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0288-y;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0288-y.pdf.

Nathan:2014:DNK

[760] Marco J. Nathan and Andrea Borghini. Development and natural kinds. Synthese, 191(3):539–556, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0290-4;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0290-
4.pdf.

Tang:2014:ISN

[761] Hao Tang. “It is not a something, but not a nothing either!” — McDowell on Wittgenstein. Synthese, 191(3):557–567, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0291-3;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0291-3.pdf.

Fischer:2014:PIH

[762] Eugen Fischer. Philosophical intuitions, heuristics, and metaphors. Synthese, 191(3):569–606, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0292-2;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0292-2.pdf.

McKenna:2014:NS

[763] Robin McKenna. Normative scorekeeping. Synthese, 191(3):607– 625, February 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0293-1;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0293-1.pdf.

Peijnenburg:2014:IRD

[764] Jeanne Peijnenburg and Sylvia Wenmackers. Infinite regress in decision theory, philosophy of science, and formal epistemology. Synthese, 191 (4):627–628, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0397-2;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0397-2.pdf.

Bartha:2014:SSP

[765] Paul Bartha, John Barker, and Alan Hájek. Satan, Saint Peter and Saint Petersburg. Synthese, 191(4):629–660, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0379-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0379-
9.pdf.

Lin:2014:RPD

[766] Hanti Lin. On the regress problem of deciding how to decide. Synthese, 191(4):661–670, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0398-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0398-1.pdf.

Norton:2014:MDP

[767] John D. Norton. A material dissolution of the problem of induction. Synthese, 191(4):671–690, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0356-3;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0356-3.pdf.

Bewersdorf:2014:IPJ

[768] Benjamin Bewersdorf. Infinitism and probabilistic justification. Synthese, 191(4):691–699, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0367-0;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0367-0.pdf.

Herzberg:2014:DIC

[769] Frederik Herzberg. The dialectics of infinitism and coherentism: inferential justification versus holism and coherence. Synthese, 191(4): 701–723, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0273-5;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0273-5.pdf.

Harbecke:2014:RSC

[770] Jens Harbecke. The role of supervenience and constitution in neuroscientific research. Synthese, 191(5):725–743, March 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0308-y;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0308-
y.pdf.

vanDitmarsch:2014:DL

[771] Hans van Ditmarsch. Dynamics of lying. Synthese, 191(5):745–777, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0275-3;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0275-3.pdf. See correction [1948].

Brown:2014:CR

[772] Campbell Brown. The composition of reasons. Synthese, 191(5): 779–800, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0299-8;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0299-8.pdf.

Hudson:2014:SPA

[773] Robert Hudson. Saving Pritchard’s anti-luck virtue epistemology: the case of temp. Synthese, 191(5):801–815, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0294-0;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0294-
0.pdf.

Michael:2014:MSE

[774] John Michael, Wayne Christensen, and Søren Overgaard. Mindreading as social expertise. Synthese, 191(5):817–840, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0295-z;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0295-z.pdf.

Rami:2014:UAP

[775] Dolf Rami. On the unification argument for the predicate view on proper names. Synthese, 191(5):841–862, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0296-y;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0296-
y.pdf.

Dyrkolbotn:2014:PDL

[776] Sjur Dyrkolbotn and Michal Walicki. Propositional discourse logic. Synthese, 191(5):863–899, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0297-x;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0297-x.pdf.

Shiver:2014:MBT

[777] Anthony Shiver. Mereological bundle theory and the identity of indiscernibles. Synthese, 191(5):901–913, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0298-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0298-
9.pdf.

Calosi:2014:QMP

[778] Claudio Calosi. Quantum mechanics and priority monism. Synthese, 191 (5):915–928, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0300-6;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0300-6.pdf.

Fiebich:2014:MEF

[779] Anika Fiebich. Mindreading with ease? Fluency and belief reasoning in 4- to 5-year-olds. Synthese, 191(5):929–944, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0301-5;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0301-5.pdf.

Rohwer:2014:LUK

[780] Yasha Rohwer. Lucky understanding without knowledge. Synthese, 191 (5):945–959, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0322-0;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0322-0.pdf.

Johnsen:2014:RQE

[781] Bredo C. Johnsen. Reclaiming Quine’s epistemology. Synthese, 191 (5):961–988, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0302-4;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0302-4.pdf.

Howson:2014:FAA

[782] Colin Howson. Finite additivity, another lottery paradox and conditionalisation. Synthese, 191(5):989–1012, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0303-3;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0303-
3.pdf.

Waskan:2014:EAP

[783] Jonathan Waskan, Ian Harmon, Zachary Horne, Joseph Spino, and John Clevenger. Explanatory anti-psychologism overturned by lay and scientific case classifications. Synthese, 191(5):1013–1035, March 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0304-2;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0304-2.pdf.

Floridi:2014:ICS

[784] Luciano Floridi. Information closure and the sceptical objection. Synthese, 191(6):1037–1050, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0306-0;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0306-0.pdf.

Pruss:2014:ITS

[785] Alexander R. Pruss. Infinitesimals are too small for countably infinite fair lotteries. Synthese, 191(6):1051–1057, April 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0307-z;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0307-
z.pdf.

Fischer:2014:WID

[786] Robert William Fischer. Why it doesn’t matter whether the virtues are truth-conducive. Synthese, 191(6):1059–1073, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0309-x;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0309-x.pdf.

Duddy:2014:RPT

[787] Conal Duddy. Reconciling probability theory and coherentism. Synthese, 191(6):1075–1084, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0310-4;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0310-4.pdf.

Diekemper:2014:EP

[788] Joseph Diekemper. The existence of the past. Synthese, 191(6): 1085–1104, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0311-3;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0311-3.pdf.

deSa:2014:LVL

[789] Dan López de Sa. Lewis vs Lewis on the problem of the many. Synthese, 191(6):1105–1117, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0314-0;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0314-0.pdf.

Hannon:2014:FVK

[790] Michael Hannon. Fallibilism and the value of knowledge. Synthese, 191 (6):1119–1146, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0315-z;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0315-z.pdf.

Tugby:2014:CDE

[791] Matthew Tugby. Categoricalism, dispositionalism, and the epistemology of properties. Synthese, 191(6):1147–1162, April 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0316-y;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0316-
y.pdf.

Currie:2014:NMP

[792] Adrian Mitchell Currie. Narratives, mechanisms and progress in historical science. Synthese, 191(6):1163–1183, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0317-x;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0317-
x.pdf.

Rossi:2014:STI

[793] Mauro Rossi. Simulation theory and interpersonal utility comparisons reconsidered. Synthese, 191(6):1185–1210, April 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0318-9;

 http://link.springer.com/content/pdf/10.1007/s11229-013-0318-
9.pdf.

Rowbottom:2014:AS

[794] Darrell P. Rowbottom. Aimless science. Synthese, 191(6):1211–1221, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0319-8;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0319-8.pdf.

Ranalli:2014:LPP

[795] Chris Ranalli. Luck, propositional perception, and the entailment thesis. Synthese, 191(6):1223–1247, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0323-z;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0323-
z.pdf.

Galliani:2014:TSD

[796] P. Galliani. Transition semantics: the dynamics of dependence logic. Synthese, 191(6):1249–1276, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0327-8;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0327-8.pdf.

Rescorla:2014:TCI

[797] Michael Rescorla. A theory of computational implementation. Synthese, 191(6):1277–1307, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0324-y;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0324-y.pdf.

Caso:2014:ART

[798] Ramiro Caso. Assertion and relative truth. Synthese, 191(6):1309–1325, April 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0329-6;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0329-6.pdf.

Wang:2014:CSC

[799] Linton Wang and Wei-Fen Ma. Comparative syllogism and counterfactual knowledge. Synthese, 191(6):1327–1348, April 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0330-0;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0330-
0.pdf.

Baumgartner:2014:EIR

[800] Michael Baumgartner. Exhibiting interpretational and representational validity. Synthese, 191(7):1349–1373, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0331-z;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0331-
z.pdf.

Keefe:2014:WLP

[801] Rosanna Keefe. What logical pluralism cannot be. Synthese, 191(7): 1375–1390, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0333-x;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0333-x.pdf.

Ballantyne:2014:DLP

[802] Nathan Ballantyne. Does luck have a place in epistemology? Synthese, 191(7):1391–1407, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0334-9;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0334-9.pdf.

Lukits:2014:PME

[803] Stefan Lukits. The principle of maximum entropy and a problem in probability kinematics. Synthese, 191(7):1409–1431, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0335-8;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0335-8.pdf.

Tang:2014:IPB

[804] Weng Hong Tang. Intentionality and partial belief. Synthese, 191(7): 1433–1450, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0336-7;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0336-7.pdf.

Fillion:2014:EAM

[805] Nicolas Fillion and Robert M. Corless. On the epistemological analysis of modeling and computational error in the mathematical sciences. Synthese, 191(7):1451–1467, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0339-4;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0339-
4.pdf.

Fresco:2014:IIP

[806] Nir Fresco and Marty J. Wolf. The instructional information processing account of digital computation. Synthese, 191(7):1469–1492, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0338-5;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0338-5.pdf.

Kraal:2014:ARE

[807] Anders Kraal. The aim of Russell’s early logicism: a reinterpretation. Synthese, 191(7):1493–1510, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0342-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0342-
9.pdf.

Tallant:2014:AMN

[808] Jonathan Tallant. Against mereological nihilism. Synthese, 191(7): 1511–1527, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0343-8;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0343-8.pdf.

Kappel:2014:EEA

[809] Klemens Kappel and Emil F. L. Moeller. Epistemic expressivism and the argument from motivation. Synthese, 191(7):1529–1547, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0347-4;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0347-4.pdf.

Ghijsen:2014:PDE

[810] Harmen Ghijsen. Phenomenalist dogmatist experientialism and the distinctiveness problem. Synthese, 191(7):1549–1566, May 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0348-3;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0348-
3.pdf.

Burgess:2014:DNI

[811] John P. Burgess. On a derivation of the necessity of identity. Synthese, 191(7):1567–1585, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0351-8;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0351-8.pdf.

Frise:2014:SFF

[812] Matthew Frise. Speaking freely: on free will and the epistemology of testimony. Synthese, 191(7):1587–1603, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0350-9;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0350-
9.pdf.

Urbaniak:2014:PQM

[813] Rafal Urbaniak. Plural quantifiers: a modal interpretation. Synthese, 191 (7):1605–1626, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0354-5;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0354-5.pdf.

Fuhrmann:2014:KPK

[814] André Fuhrmann. Knowability as potential knowledge. Synthese, 191 (7):1627–1648, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0340-y;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0340-y.pdf.

Fuhrmann:2014:EKP

[815] André Fuhrmann. Erratum to: Knowability as potential knowledge. Synthese, 191(7):1649, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0359-0;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0359-0.pdf.

Clarke:2014:MMC

[816] Brendan Clarke, Bert Leuridan, and Jon Williamson. Modelling mechanisms with causal cycles. Synthese, 191(8):1651–1681, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0360-7;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0360-7.pdf.

Carter:2014:NMA

[817] J. Adam Carter and Emma C. Gordon. A new maneuver against the epistemic relativist. Synthese, 191(8):1683–1695, May 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0357-2;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0357-
2.pdf.

Camilleri:2014:TCE

[818] Kristian Camilleri. Toward a constructivist epistemology of thought experiments in science. Synthese, 191(8):1697–1716, May 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0358-1;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0358-
1.pdf.

Lutz:2014:PPE

[819] Matt Lutz. The pragmatics of pragmatic encroachment. Synthese, 191 (8):1717–1740, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0361-6;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0361-6.pdf.

Schindler:2014:EFR

[820] Samuel Schindler. Explanatory fictions-for real? Synthese, 191(8): 1741–1755, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0362-5;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0362-5.pdf.

Kremer:2014:IFI

[821] Philip Kremer. Indeterminacy of fair infinite lotteries. Synthese, 191 (8):1757–1760, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0364-3;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0364-3.pdf.

Werner:2014:SSD

[822] Preston J. Werner. Seemings: still dispositions to believe. Synthese, 191 (8):1761–1774, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0363-4;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0363-4.pdf.

Gijsbers:2014:HAC

[823] Victor Gijsbers and Leon de Bruin. How agency can solve interventionism’s problem of circularity. Synthese, 191(8):1775–1791, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0366-1;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0366-1.pdf.

Roberts:2014:PR

[824] Pendaran Roberts. Parsing the rainbow. Synthese, 191(8):1793–1811, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0368-z;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0368-z.pdf.

Smit:2014:DIA

[825] J. P. Smit, Filip Buekens, and Stan du Plessis. Developing the incentivized action view of institutional reality. Synthese, 191(8):1813–1830, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0370-5;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0370-5.pdf.

Rueger:2014:IPR

[826] Alexander Rueger. Idealized and perspectival representations: some reasons for making a distinction. Synthese, 191(8):1831–1845, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0371-4;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0371-4.pdf.

Rizza:2014:ATT

[827] Davide Rizza. Arrow’s theorem and theory choice. Synthese, 191(8): 1847–1856, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0372-3;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0372-3.pdf.

Benton:2014:IPP

[828] Matthew A. Benton and John Turri. Iffy predictions and proper expectations. Synthese, 191(8):1857–1866, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0377-y;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0377-
y.pdf.

Booth:2014:SRM

[829] Anthony Robert Booth. On some recent moves in defence of doxastic compatibilism. Synthese, 191(8):1867–1880, May 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0378-x;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0378-
x.pdf.

Masterton:2014:WDF

[830] George Masterton. What to do with a forecast? Synthese, 191(8): 1881–1907, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0384-z;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0384-z.pdf.

Giordani:2014:FII

[831] Alessandro Giordani. On the factivity of implicit intersubjective knowledge. Synthese, 191(8):1909–1923, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0381-2;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0381-
2.pdf.

Hausman:2014:SGC

[832] Daniel M. Hausman, Reuben Stern, and Naftali Weinberger. Systems without a graphical causal representation. Synthese, 191(8):1925–1930, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0380-3;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0380-3.pdf.

Palermos:2014:KCI

[833] Spyridon Orestis Palermos. Knowledge and cognitive integration. Synthese, 191(8):1931–1951, May 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0383-0;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0383-0.pdf.

Pedersen:2014:RT

[834] Nikolaj Jang Lee Linding Pedersen, Kristoffer Ahlström-Vij, and Klemens Kappel. Rational trust. Synthese, 191(9):1953–1955, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0451-0;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0451-0.pdf.

Frost-Arnold:2014:CAR

[835] Karen Frost-Arnold. The cognitive attitude of rational trust. Synthese, 191(9):1957–1974, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-012-0151-6;   http://link.springer.com/
content/pdf/10.1007/s11229-012-0151-6.pdf.

Faulkner:2014:PRT

[836] Paul Faulkner. The practical rationality of trust. Synthese, 191(9): 1975–1989, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0103-1;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0103-1.pdf.

Vallinder:2014:TVO

[837] Aron Vallinder and Erik J. Olsson. Trust and the value of overconfidence: a Bayesian perspective on social network communication. Synthese, 191 (9):1991–2007, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0375-0;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0375-0.pdf.

Kappel:2014:BT

[838] Klemens Kappel. Believing on trust. Synthese, 191(9):2009–2028, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0376-z;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0376-z.pdf.

Hawley:2014:PPT

[839] Katherine Hawley. Partiality and prejudice in trusting. Synthese, 191 (9):2029–2045, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0129-4;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0129-4.pdf.

Hawley:2014:EPP

[840] Katherine Hawley. Erratum to: Partiality and prejudice in trusting. Synthese, 191(9):2047, June 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0453-y;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0453-y.pdf.

Lloyd:2014:OCM

[841] Elisabeth A. Lloyd and Vanessa J. Schweizer. Objectivity and a comparison of methodological scenario approaches for climate change research. Synthese, 191(10):2049–2088, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0353-6;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0353-
6.pdf.

Markus:2014:IAC

[842] Keith A. Markus. An incremental approach to causal inference in the behavioral sciences. Synthese, 191(10):2089–2113, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0386-x;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0386-x.pdf.

Giraud:2014:CFS

[843] Thibaut Giraud. Constructing formal semantics from an ontological perspective. The case of second-order logics. Synthese, 191(10):2115–2145, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0387-9;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0387-9.pdf.

Kramer:2014:ICT

[844] Stephan Krämer. Implicit commitment in theory choice. Synthese, 191 (10):2147–2165, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0388-8;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0388-8.pdf.

Huber:2014:NFC

[845] Franz Huber. New foundations for counterfactuals. Synthese, 191(10): 2167–2193, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0391-0;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0391-0.pdf.

Dambock:2014:KNS

[846] Christian Damböck. Kuhn’s notion of scientific progress: “reduction” between incommensurable theories in a rigid structuralist framework. Synthese, 191(10):2195–2213, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0392-z;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0392-
z.pdf.

Walter:2014:WEF

[847] Sven Walter. Willusionism, epiphenomenalism, and the feeling of conscious will. Synthese, 191(10):2215–2238, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0393-y;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0393-
y.pdf.

Welch:2014:PC

[848] John R. Welch. Plausibilistic coherence. Synthese, 191(10):2239–2253, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0395-9;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0395-9.pdf.

Staples:2014:CRE

[849] Mark Staples. Critical rationalism and engineering: ontology. Synthese, 191(10):2255–2279, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0396-3;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0396-3.pdf.

Loeb:2014:TTT

[850] Iris Loeb. Towards transfinite type theory: rereading Tarski’s Wahrheitsbegriff. Synthese, 191(10):2281–2299, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0399-0;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0399-
0.pdf.

Dallmann:2014:NAC

[851] Justin M. Dallmann. A normatively adequate credal reductivism. Synthese, 191(10):2301–2313, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0402-9;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0402-9.pdf.

Boucher:2014:WPS

[852] Sandy C. Boucher. What is a philosophical stance? Paradigms, policies and perspectives. Synthese, 191(10):2315–2332, July 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0400-y;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0400-
y.pdf.

Pexton:2014:HDA

[853] Mark Pexton. How dimensional analysis can explain. Synthese, 191 (10):2333–2351, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0401-x;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0401-x.pdf.

Zenker:2014:EIS

[854] Frank Zenker and Carlo Proietti. Editors’ introduction: social dynamics and collective rationality. Synthese, 191(11):2353–2358, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

014-0430-5;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0430-5.pdf.

Brossel:2014:HRD

[855] Peter Brössel and Anna-Maria A. Eder. How to resolve doxastic disagreement. Synthese, 191(11):2359–2381, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0431-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0431-
4.pdf.

Cevolani:2014:TAB

[856] Gustavo Cevolani. Truth approximation, belief merging, and peer disagreement. Synthese, 191(11):2383–2401, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0486-2;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0486-
2.pdf.

Liu:2014:LDB

[857] Fenrong Liu, Jeremy Seligman, and Patrick Girard. Logical dynamics of belief change in the community. Synthese, 191(11):2403–2431, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0432-3;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0432-3.pdf.

Masterton:2014:TVC

[858] George Masterton. Topological variability of collectives and its import for social epistemology. Synthese, 191(11):2433–2443, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0433-2;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0433-2.pdf.

Bjerring:2014:RPI

[859] Jens Christian Bjerring, Jens Ulrik Hansen, and Nikolaj Jang Lee Linding Pedersen. On the rationality of pluralistic ignorance. Synthese, 191 (11):2445–2470, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0434-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0434-1.pdf.

Rendsvig:2014:PIB

[860] Rasmus K. Rendsvig. Pluralistic ignorance in the bystander effect: informational dynamics of unresponsive witnesses in situations calling for intervention. Synthese, 191(11):2471–2498, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0435-0;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0435-0.pdf.

DeLanghe:2014:SII

[861] Rogier De Langhe. To specialize or to innovate? An internalist account of pluralistic ignorance in economics. Synthese, 191(11):2499–2511, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0436-z;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0436-z.pdf.

Robson:2014:SEA

[862] Jon Robson. A social epistemology of aesthetics: belief polarization, echo chambers and aesthetic judgement. Synthese, 191(11):2513–2528, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0437-y;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0437-y.pdf.

Kenyon:2014:FPD

[863] Tim Kenyon. False polarization: debiasing as applied social epistemology. Synthese, 191(11):2529–2547, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0438-x;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0438-
x.pdf.

Baumgaertner:2014:YNM

[864] Bert Baumgaertner. Yes, no, maybe so: a veritistic approach to echo chambers using a trichotomous belief model. Synthese, 191(11): 2549–2569, July 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0439-9;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0439-9.pdf.

Schafer:2014:DPE

[865] Karl Schafer. Doxastic planning and epistemic internalism. Synthese, 191(12):2571–2591, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0412-7;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0412-7.pdf.

Keren:2014:TBP

[866] Arnon Keren. Trust and belief: a preemptive reasons account. Synthese, 191(12):2593–2615, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0416-3;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0416-3.pdf.

Kraft:2014:TAA

[867] Tim Kraft. Transmission arguments against knowledge closure are still fallacious. Synthese, 191(12):2617–2632, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0403-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0403-
8.pdf.

Iacona:2014:OTR

[868] Andrea Iacona. Ockhamism without thin red lines. Synthese, 191(12): 2633–2652, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0405-6;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0405-6.pdf.

Sustar:2014:FDB

[869] Predrag Sustar and Zdenka Brzović. The function debate: between “cheap tricks” and evolutionary neutrality. Synthese, 191(12):2653– 2671, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0407-4;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0407-4.pdf.

Kummerfeld:2014:MCR

[870] Erich Kummerfeld and David Danks. Model change and reliability in scientific inference. Synthese, 191(12):2673–2693, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

014-0408-3;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0408-3.pdf.

Barrett:2014:FAM

[871] David Barrett. Functional analysis and mechanistic explanation. Synthese, 191(12):2695–2714, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0410-9;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0410-9.pdf.

Klein:2014:WTI

[872] Ralf-Thomas Klein. Where there are internal defeaters, there are “confirmers”. Synthese, 191(12):2715–2728, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0415-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0415-
4.pdf.

Fridland:2014:TLC

[873] Ellen Fridland. They’ve lost control: reflections on skill. Synthese, 191(12):2729–2750, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0411-8;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0411-8.pdf.

Intemann:2014:TLS

[874] Kristen Intemann and Inmaculada de Melo-Martín. Are there limits to scientists’ obligations to seek and engage dissenters? Synthese, 191(12): 2751–2765, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0414-5;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0414-5.pdf.

Besong:2014:MID

[875] Brian Besong. Moral intuitionism and disagreement. Synthese, 191(12): 2767–2789, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0420-7;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0420-7.pdf.

Johnsen:2014:HGR

[876] Bredo Johnsen. Hume, Goodman and radical inductive skepticism. Synthese, 191(12):2791–2813, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0417-2;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0417-2.pdf.

Loeb:2014:UMT

[877] Iris Loeb. Uniting model theory and the universalist tradition of logic: Carnap’s early axiomatics. Synthese, 191(12):2815–2833, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0425-2;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0425-2.pdf.

Tebben:2014:DDC

[878] Nicholas Tebben. Deontology and doxastic control. Synthese, 191(12): 2835–2847, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0423-4;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0423-4.pdf.

May:2014:VCF

[879] Joshua May. On the very concept of free will. Synthese, 191(12): 2849–2866, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0426-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0426-1.pdf.

Baron:2014:CTW

[880] Sam Baron and Kristie Miller. Causation in a timeless world. Synthese, 191(12):2867–2886, August 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0427-0;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0427-0.pdf.

Bueno:2014:SI

[881] Otávio Bueno and Peter Vickers. Is science inconsistent? Synthese, 191(13):2887–2889, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0463-9;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0463-9.pdf.

Vickers:2014:TFI

[882] Peter Vickers. Theory flexibility and inconsistency in science. Synthese, 191(13):2891–2906, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0464-8;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0464-8.pdf.

Verelst:2014:NVL

[883] Karin Verelst. Newton versus Leibniz: intransparency versus inconsistency. Synthese, 191(13):2907–2940, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0465-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0465-
7.pdf.

Saatsi:2014:ISR

[884] Juha Saatsi. Inconsistency and scientific realism. Synthese, 191(13): 2941–2955, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0466-6;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0466-6.pdf.

Pincock:2014:HAI

[885] Christopher Pincock. How to avoid inconsistent idealizations. Synthese, 191(13):2957–2972, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0467-5;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0467-5.pdf.

Batterman:2014:IPC

[886] Robert W. Batterman. The inconsistency of Physics (with a capital “P”). Synthese, 191(13):2973–2992, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0468-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0468-
4.pdf.

Azzouni:2014:NCS

[887] Jody Azzouni. A new characterization of scientific theories. Synthese, 191(13):2993–3008, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0469-3;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0469-3.pdf.

Davey:2014:CGS

[888] Kevin Davey. Can good science be logically inconsistent? Synthese, 191(13):3009–3026, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0470-x;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0470-x.pdf.

Frisch:2014:MSR

[889] Mathias Frisch. Models and scientific representations or: who is afraid of inconsistency? Synthese, 191(13):3027–3040, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0471-9;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0471-9.pdf.

daCosta:2014:PIQ

[890] Newton C. A. da Costa and Décio Krause. Physics, inconsistency, and quasi-truth. Synthese, 191(13):3041–3055, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0472-8;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0472-8.pdf.

Benham:2014:CPI

[891] Richard Benham, Chris Mortensen, and Graham Priest. Chunk and permeate III: the Dirac delta function. Synthese, 191(13):3057–3062, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0473-7;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0473-7.pdf.

vanBendegem:2014:IMM

[892] Jean Paul van Bendegem. Inconsistency in mathematics and the mathematics of inconsistency. Synthese, 191(13):3063–3078, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0474-6;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0474-6.pdf.

Brown:2014:SS

[893] M. Bryson Brown. The shape of science. Synthese, 191(13):3079– 3109, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0475-5;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0475-5.pdf.

Seselja:2014:EJC

[894] Dunja Seselja and Christian Straßer. Epistemic justification in the context of pursuit: a coherentist approach. Synthese, 191(13):3111– 3141, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0476-4;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0476-4.pdf.

McAllister:2014:MDE

[895] James W. McAllister. Methodological dilemmas and emotion in science. Synthese, 191(13):3143–3158, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0477-3;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0477-
3.pdf.

Singer:2014:SBS

[896] Daniel Jeremy Singer. Sleeping beauty should be imprecise. Synthese, 191(14):3159–3172, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0429-y;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0429-y.pdf.

Tang:2014:WDI

[897] Hao Tang. Wittgenstein and the dualism of the inner and the outer. Synthese, 191(14):3173–3194, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0441-2;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0441-
2.pdf.

Lutz:2014:GEA

[898] Sebastian Lutz. Generalizing empirical adequacy i: multiplicity and approximation. Synthese, 191(14):3195–3225, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0440-3;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0440-3.pdf.

Perrine:2014:DNR

[899] Timothy Perrine. In defense of non-reductionism in the epistemology of testimony. Synthese, 191(14):3227–3237, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0443-0;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0443-0.pdf.

Dimmock:2014:KCP

[900] Paul Dimmock and Torfinn Thomesen Huvenes. Knowledge, conservatism, and pragmatics. Synthese, 191(14):3239–3269, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0442-1;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0442-1.pdf.

Hicks:2014:NDS

[901] Daniel J. Hicks. A new direction for science and values. Synthese, 191(14):3271–3295, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0447-9;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0447-9.pdf.

Puddifoot:2014:DER

[902] Katherine Puddifoot. A defence of epistemic responsibility: why laziness and ignorance are bad after all. Synthese, 191(14):3297–3309, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0445-y;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0445-y.pdf.

Tucker:2014:WIJ

[903] Chris Tucker. On what inferentially justifies what: the vices of reliabilism and proper functionalism. Synthese, 191(14):3311–3328, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0446-x;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0446-x.pdf.

Unnsteinsson:2014:CSS

[904] Elmar Geir Unnsteinsson. Compositionality and sandbag semantics. Synthese, 191(14):3329–3350, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0449-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0449-
7.pdf.

Yap:2014:IEL

[905] Audrey Yap. Idealization, epistemic logic, and epistemology. Synthese, 191(14):3351–3366, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0448-8;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0448-8.pdf.

Wilkenfeld:2014:FEN

[906] Daniel A. Wilkenfeld. Functional explaining: a new approach to the philosophy of explanation. Synthese, 191(14):3367–3391, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0452-z;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0452-z.pdf.

Smid:2014:TOO

[907] Jeroen Smid. Tarski’s one and only concept of truth. Synthese, 191(14): 3393–3406, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0450-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0450-1.pdf.

Kane:2014:RJA

[908] Jeffrey Kane and Pavel Naumov. The Ryoan-ji axiom for common knowledge on hypergraphs. Synthese, 191(14):3407–3426, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0455-9;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0455-9.pdf.

Corkum:2014:PTD

[909] Phil Corkum. Presentism, truthmakers and distributional properties. Synthese, 191(14):3427–3446, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0456-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0456-
8.pdf.

Long:2014:IIS

[910] Bruce Raymond Long. Information is intrinsically semantic but alethically neutral. Synthese, 191(14):3447–3467, September 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0457-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0457-
7.pdf.

Kummerfeld:2014:EMC

[911] Erich Kummerfeld and David Danks. Erratum to: Model change and methodological virtues in scientific inference. Synthese, 191(14):3469– 3472, September 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0454-x;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0454-x.pdf.

Zardini:2014:CCI

[912] Elia Zardini. Context and consequence. An intercontextual substructural logic. Synthese, 191(15):3473–3500, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0490-6;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0490-
6.pdf.

King:2014:PIV

[913] Nathan L. King. Perseverance as an intellectual virtue. Synthese, 191(15):3501–3523, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0418-1;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0418-1.pdf.

Pruss:2014:RPC

[914] Alexander R. Pruss. Regular probability comparisons imply the Banach–Tarski Paradox. Synthese, 191(15):3525–3540, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0458-6;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0458-6.pdf.

Dinges:2014:ECC

[915] Alexander Dinges. Epistemic contextualism can be stated properly. Synthese, 191(15):3541–3556, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0459-5;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0459-5.pdf.

Akiba:2014:DID

[916] Ken Akiba. A defense of indeterminate distinctness. Synthese, 191(15): 3557–3573, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0462-x;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0462-x.pdf.

Marcus:2014:HPI

[917] Russell Marcus. The holistic presumptions of the indispensability argument. Synthese, 191(15):3575–3594, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0481-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0481-
7.pdf.

Barberousse:2014:AWC

[918] Anouk Barberousse and Marion Vorms. About the warrants of computerbased empirical knowledge. Synthese, 191(15):3595–3620, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0482-6;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0482-6.pdf.

Hedberg:2014:ESI

[919] Trevor Hedberg. Epistemic supererogation and its implications. Synthese, 191(15):3621–3637, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0483-5;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0483-5.pdf.

Degenaar:2014:RHR

[920] Jan Degenaar and Erik Myin. Representation-hunger reconsidered. Synthese, 191(15):3639–3648, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0484-4;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0484-4.pdf.

Skiles:2014:TDT

[921] Alexander Skiles. Is there a dilemma for the truthmaker nonmaximalist? Synthese, 191(15):3649–3659, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0485-3;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0485-
3.pdf.

Schippers:2014:CSA

[922] Michael Schippers. Coherence, striking agreement, and reliability. Synthese, 191(15):3661–3684, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0488-0;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0488-0.pdf.

Macdonald:2014:MME

[923] Cynthia Macdonald. In my ‘mind’s eye’: introspectionism, detectivism, and the basis of authoritative self-knowledge. Synthese, 191(15):3685– 3710, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0487-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0487-1.pdf.

James:2014:HRT

[924] Steven P. James. Hallucinating real things. Synthese, 191(15):3711– 3732, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0492-4;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0492-4.pdf.

Beirlaen:2014:CLA

[925] Mathieu Beirlaen and Atocha Aliseda. A conditional logic for abduction. Synthese, 191(15):3733–3758, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0496-0;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0496-
0.pdf.

Kulp:2014:PTM

[926] Christopher B. Kulp. The pre-theoreticality of moral intuitions. Synthese, 191(15):3759–3778, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0497-z;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0497-z.pdf.

King:2014:EPI

[927] Nathan L. King. Erratum to: Perseverance as an intellectual virtue. Synthese, 191(15):3779–3801, October 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0511-5;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0511-
5.pdf.

Boehm:2014:HDC

[928] Miren Boehm. Hume’s definitions of ‘cause’: Without idealizations, within the bounds of science. Synthese, 191(16):3803–3819, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0499-x;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0499-x.pdf.

Schippers:2014:PMC

[929] Michael Schippers. Probabilistic measures of coherence: from adequacy constraints towards pluralism. Synthese, 191(16):3821–3845, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0501-7;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0501-7.pdf.

Palmer:2014:DFC

[930] David Palmer. Deterministic Frankfurt cases. Synthese, 191(16):3847– 3864, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0500-8;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0500-8.pdf.

Talbot:2014:WNE

[931] Brian Talbot. Why so negative? Evidence aggregation and armchair philosophy. Synthese, 191(16):3865–3896, November 2014. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0509-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0509-
z.pdf.

Friederich:2014:PPA

[932] Simon Friederich, Robert Harlander, and Koray Karaca. Philosophical perspectives on ad hoc hypotheses and the Higgs mechanism. Synthese, 191(16):3897–3917, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0504-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0504-
4.pdf.

Tiehen:2014:PAA

[933] Justin Tiehen. A psychofunctionalist argument against nonconceptualism. Synthese, 191(16):3919–3934, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0505-3;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0505-
3.pdf.

Doucet:2014:NPW

[934] Mathieu Doucet and John Turri. Non-psychological weakness of will: self-control, stereotypes, and consequences. Synthese, 191(16):3935– 3954, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0508-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0508-0.pdf.

Park:2014:CMC

[935] Ilho Park. Confirmation measures and collaborative belief updating. Synthese, 191(16):3955–3975, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0507-1;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0507-
1.pdf.

DiBello:2014:ECA

[936] Marcello Di Bello. Epistemic closure, assumptions and topics of inquiry. Synthese, 191(16):3977–4002, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0522-2;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0522-
2.pdf.

Cowie:2014:DIA

[937] Christopher Cowie. In defence of instrumentalism about epistemic normativity. Synthese, 191(16):4003–4017, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0510-6;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0510-
6.pdf.

Stefansson:2014:DBC

[938] H. Orri Stefánsson. Desires, beliefs and conditional desirability. Synthese, 191(16):4019–4035, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0512-4;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0512-4.pdf.

Bertolotti:2014:EAG

[939] Tommaso Bertolotti and Lorenzo Magnani. An epistemological analysis of gossip and gossip-based knowledge. Synthese, 191(17):4037– 4067, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0514-2;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0514-2.pdf.

Anacona:2014:BAS

[940] Maribel Anacona, Luis Carlos Arboleda, and F. Javier Pérez-Fernández. On Bourbaki’s axiomatic system for set theory. Synthese, 191(17):4069– 4098, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0515-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0515-1.pdf.

Dennis:2014:CIK

[941] Peter Dennis. Criteria for indefeasible knowledge: John Mcdowell and ‘epistemological disjunctivism’. Synthese, 191(17):4099–4113, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0516-0;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0516-0.pdf.

McEvoy:2014:CTR

[942] Mark McEvoy. Causal tracking reliabilism and the Gettier prob-

lem.  Synthese, 191(17):4115–4130, November 2014.  CODEN

SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0523-1;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0523-
1.pdf.

Dawson:2014:WPA

[943] Ryan Dawson. Wittgenstein on pure and applied mathematics. Synthese, 191(17):4131–4148, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0520-4;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0520-4.pdf.

Meynell:2014:IIN

[944] Letitia Meynell. Imagination and insight: a new account of the content of thought experiments. Synthese, 191(17):4149–4168, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0519-x;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0519-x.pdf.

Smith:2014:EUC

[945] Ryan Smith. Explanation, understanding, and control. Synthese, 191 (17):4169–4200, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0521-3;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0521-3.pdf.

Pantsar:2014:EFA

[946] Markus Pantsar. An empirically feasible approach to the epistemology of arithmetic. Synthese, 191(17):4201–4229, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0526-y;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0526-y.pdf.

Buehler:2014:IUC

[947] Denis Buehler. Incomplete understanding of complex numbers: Girolamo Cardano: a case study in the acquisition of mathematical concepts. Synthese, 191(17):4231–4252, November 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0527-x;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0527-
x.pdf.

Boos:2014:RIF

[948] William Boos. Reflective inquiry and “the fate of reason”. Synthese, 191(18):4253–4314, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0533-z;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0533-z.pdf.

Ahmed:2014:CDT

[949] Arif Ahmed and Adam Caulton. Causal decision theory and EPR correlations. Synthese, 191(18):4315–4352, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0536-9;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0536-
9.pdf.

Kim:2014:LGI

[950] Brian Kim. The locality and globality of instrumental rationality: the normative significance of preference reversals. Synthese, 191(18):4353– 4376, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0529-8;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0529-8.pdf.

Catren:2014:NII

[951] Gabriel Catren and Julien Page. On the notions of indiscernibility and indeterminacy in the light of the Galois–Grothendieck theory. Synthese, 191(18):4377–4408, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0528-9;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0528-
9.pdf.

Muldoon:2014:WTD

[952] Ryan Muldoon, Chiara Lisciandra, and Stephan Hartmann. Why are there descriptive norms? Because we looked for them. Synthese, 191(18): 4409–4429, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0534-y;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0534-y.pdf.

Belleri:2014:YCS

[953] Delia Belleri. You can say what you think: vindicating the effability of our thoughts. Synthese, 191(18):4431–4450, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0537-8;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0537-8.pdf.

Brouwer:2014:PR

[954] Thomas N. P. A. Brouwer. A paradox of rejection. Synthese, 191(18): 4451–4464, December 2014. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0541-z;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0541-z.pdf.

Davis:2015:FDC

[955] John K. Davis. Faultless disagreement, cognitive command, and epistemic peers. Synthese, 192(1):1–24, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0543-x;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0543-
x.pdf.

Schmitt:2015:FTR

[956] Margaret Schmitt. Freedom and (theoretical) reason. Synthese, 192 (1):25–41, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0547-6;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0547-6.pdf.

Gervais:2015:MNM

[957] Raoul Gervais. Mechanistic and non-mechanistic varieties of dynamical models in cognitive science: explanatory power, understanding, and the ‘mere description’ worry. Synthese, 192(1):43–66, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0548-5;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0548-5.pdf.

Mulligan:2015:DPT

[958] Thomas Mulligan. Disagreement, peerhood, and three paradoxes of conciliationism. Synthese, 192(1):67–78, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0553-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0553-
8.pdf.

John:2015:IRC

[959] Stephen John. Inductive risk and the contexts of communication. Synthese, 192(1):79–96, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0554-7;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0554-7.pdf.

Rose:2015:PTF

[960] David Rose. Persistence through function preservation. Synthese, 192 (1):97–146, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0555-6;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0555-6.pdf.

Asay:2015:MDM

[961] Jamin Asay and S. Seth Bordner. A modest defense of manifestationalism. Synthese, 192(1):147–161, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0556-5;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0556-
5.pdf.

Fulda:2015:MFD

[962] Fermín Fulda. A mechanistic framework for Darwinism or why Fodor’s objection fails. Synthese, 192(1):163–183, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0557-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0557-
4.pdf.

Perrine:2015:UTT

[963] Timothy Perrine. Undermining truthmaker theory. Synthese, 192(1): 185–200, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0558-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0558-3.pdf.

Napoletano:2015:CWS

[964] Toby Napoletano. Compositionality as weak supervenience. Synthese, 192(1):201–220, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0562-7;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0562-7.pdf.

Friederich:2015:RTL

[965] Simon Friederich. Re-thinking local causality. Synthese, 192(1):221– 240, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0563-6;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0563-6.pdf.

Bassett:2015:CBT

[966] Robert Bassett. A critique of benchmark theory. Synthese, 192(1): 241–267, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0566-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0566-3.pdf.

Lewis:2015:GSM

[967] Cory Travers Lewis and Christopher Belanger. The generality of scientific models: a measure theoretic approach. Synthese, 192(1):269–285, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0567-2;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0567-2.pdf.

DeBianchi:2015:EDS

[968] Silvia De Bianchi and J. D. Wells. Explanation and the dimensionality of space. Synthese, 192(1):287–303, January 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0568-1;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0568-
1.pdf.

Earnshaw:2015:GSC

[969] Eugene Earnshaw. Group selection and contextual analysis. Synthese, 192(1):305–316, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0569-0;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0569-0.pdf.

Griffith:2015:HNT

[970] Aaron M. Griffith. How negative truths are made true. Synthese, 192 (1):317–335, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0570-7;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0570-7.pdf.

Staples:2015:CRE

[971] Mark Staples. Critical rationalism and engineering: methodology. Synthese, 192(1):337–362, January 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0571-6;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0571-6.pdf.

McKenna:2015:ECD

[972] Robin McKenna. Epistemic contextualism defended. Synthese, 192(2): 363–383, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0572-5;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0572-5.pdf.

Turri:2015:KNA

[973] John Turri. Knowledge and the norm of assertion: a simple test. Synthese, 192(2):385–392, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0573-4;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0573-4.pdf.

Rasmussen:2015:BTD

[974] Joshua Rasmussen. Building thoughts from dust: a Cantorian puzzle. Synthese, 192(2):393–404, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0575-2;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0575-2.pdf.

Nuffer:2015:WDM

[975] Gerhard Nuffer. What difference might and may make. Synthese, 192 (2):405–429, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0576-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0576-1.pdf.

Rini:2015:HTP

[976] Regina A. Rini. How not to test for philosophical expertise. Synthese, 192 (2):431–452, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0579-y;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0579-y.pdf.

Brown:2015:CAC

[977] Joshua D. K. Brown. Chemical atomism: a case study in confirmation and ontology. Synthese, 192(2):453–485, February 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0581-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0581-
4.pdf.

Graber:2015:CTW

[978] Abraham Graber. Creating truths by winning arguments: the problem of methodological artifacts in philosophy. Synthese, 192(2):487–503, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0580-5;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0580-5.pdf.

daCosta:2015:FFS

[979] Newton C. A. da Costa and Federico Holik. A formal framework for the study of the notion of undefined particle number in quantum mechanics. Synthese, 192(2):505–523, February 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0583-2;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0583-
2.pdf.

Jespersen:2015:I

[980] Bjørn Jespersen and Marie Duzí. Introduction. Synthese, 192(3): 525–534, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0665-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0665-9.pdf.

Pollard:2015:AHS

[981] Carl Pollard. Agnostic hyperintensional semantics. Synthese, 192(3): 535–562, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0373-2;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0373-2.pdf.

Fox:2015:TTL

[982] Chris Fox and Shalom Lappin. Type-theoretic logic with an operational account of intensionality. Synthese, 192(3):563–584, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0390-1;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0390-1.pdf.

Jago:2015:HP

[983] Mark Jago. Hyperintensional propositions. Synthese, 192(3):585– 601, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0461-y;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0461-y.pdf.

Wieckowski:2015:CBR

[984] Bartosz Wieckowski. Constructive belief reports. Synthese, 192(3): 603–633, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0540-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0540-0.pdf.

Duzi:2015:TQH

[985] Marie Duzí and Bjørn Jespersen. Transparent quantification into hyperintensional objectual attitudes. Synthese, 192(3):635–677, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0578-z;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0578-z.pdf.

Asay:2015:EL

[986] Jamin Asay. Epistemicism and the liar. Synthese, 192(3):679–699, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0596-x;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0596-x.pdf.

Verdejo:2015:SCA

[987] Víctor M. Verdejo. The systematicity challenge to anti-representational dynamicism. Synthese, 192(3):701–722, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0597-9;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0597-
9.pdf.

Krempel:2015:WKC

[988] Raquel Krempel. Wittgenstein on knowledge: a critique. Synthese, 192 (3):723–734, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0593-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0593-0.pdf.

Rescorla:2015:SER

[989] Michael Rescorla. Some epistemological ramifications of the Borel– Kolmogorov paradox. Synthese, 192(3):735–767, March 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0586-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0586-
z.pdf.

Hannon:2015:UCK

[990] Michael Hannon. The universal core of knowledge. Synthese, 192(3): 769–786, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0587-y;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0587-y.pdf.

Jager:2015:LME

[991] Christoph Jäger and Eva Bänninger-Huber. Looking into meta-emotions. Synthese, 192(3):787–811, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0588-x;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0588-x.pdf.

Klausen:2015:GKR

[992] Søren Harnow Klausen. Group knowledge: a real-world approach. Synthese, 192(3):813–839, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0589-9;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0589-9.pdf.

Kroedel:2015:SAD

[993] Thomas Kroedel. A simple argument for downward causation. Synthese, 192(3):841–858, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0600-5;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0600-5.pdf.

Rancourt:2015:ERS

[994] Benjamin T. Rancourt. Epistemic relativism and semantic blindness. Synthese, 192(3):859–876, March 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0611-2;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0611-2.pdf.

Paseau:2015:MIR

[995] A. C. Paseau. A measure of inferential-role preservation. Synthese, 192(3):??, March 2015. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s11229-015-0705-5.pdf. See erratum [1397].

Pietarinen:2015:EP

[996] Ahti-Veikko Pietarinen. Editorial preface. Synthese, 192(4):877–879, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0656-2;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0656-2.pdf.

Pietarinen:2015:TPE

[997] Ahti-Veikko Pietarinen. Two papers on existential graphs by Charles Peirce. Synthese, 192(4):881–922, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0498-y;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0498-
y.pdf.

Gangle:2015:SID

[998] Rocco Gangle and Gianluca Caterina. The sheet of indication: a diagrammatic semantics for Peirce’s EG-alpha. Synthese, 192(4):923–940, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0495-1;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0495-1.pdf.

Pietarinen:2015:EBQ

[999] Ahti-Veikko Pietarinen. Exploring the beta quadrant. Synthese, 192 (4):941–970, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0677-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0677-5.pdf.

He:2015:TIT

[1000] Chuansheng He. E-type interpretation without E-type pronoun: how Peirce’s Graphs capture the uniqueness implication of donkey pronouns in discourse anaphora. Synthese, 192(4):971–990, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0325-x;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0325-x.pdf.

Hilpinen:2015:CSR

[1001] Risto Hilpinen. Conception, sense, and reference in Peircean semiotics. Synthese, 192(4):991–1018, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0326-9;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0326-9.pdf.

Stjernfelt:2015:D

[1002] Frederik Stjernfelt. Dicisigns. Synthese, 192(4):1019–1054, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0406-5;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0406-5.pdf.

Moore:2015:PTT

[1003] Matthew E. Moore. Peirce’s topical theory of continuity. Synthese, 192 (4):1055–1071, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0337-6;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0337-6.pdf.

Pietarinen:2015:PDT

[1004] Ahti-Veikko Pietarinen and Frederik Stjernfelt. Peirce and diagrams: two contributors to an actual discussion review each other. Synthese, 192 (4):1073–1088, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0658-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0658-8.pdf.

Stanev:2015:ESR

[1005] Roger Stanev. Early stopping of RCTs: two potential issues for error statistics. Synthese, 192(4):1089–1116, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0602-3;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0602-
3.pdf.

Kulicki:2015:CPE

[1006] Piotr Kulicki and Robert Trypuz. Completely and partially executable sequences of actions in deontic context. Synthese, 192(4):1117–1138, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0604-1;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0604-1.pdf.

Evans:2015:RNE

[1007] Peter W. Evans. Retrocausality at no extra cost. Synthese, 192(4): 1139–1155, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0605-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0605-0.pdf.

Zinke:2015:ERV

[1008] Alexandra Zinke. On exhibiting representational validity. Synthese, 192 (4):1157–1171, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0607-y;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0607-y.pdf.

Krabbe:2015:TNA

[1009] Erik C. W. Krabbe and Jan Albert van Laar. That’s no argument! the dialectic of non-argumentation. Synthese, 192(4):1173–1197, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0609-9;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0609-9.pdf.

Floridi:2015:MS

[1010] Luciano Floridi, Nir Fresco, and Giuseppe Primiero. On malfunctioning software. Synthese, 192(4):1199–1220, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0610-3;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0610-
3.pdf.

Turri:2015:SAS

[1011] John Turri. Selfless assertions: some empirical evidence. Synthese, 192 (4):1221–1233, April 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0621-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0621-0.pdf.

Spolaore:2015:AFT

[1012] Giuseppe Spolaore. Agency and fictional truth: a formal study on fiction-making. Synthese, 192(5):1235–1265, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0613-0;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0613-
0.pdf.

Weber:2015:IB

[1013] Zach Weber and A. J. Cotnoir. Inconsistent boundaries. Synthese, 192 (5):1267–1294, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0614-z;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0614-z.pdf.

Brenner:2015:MNS

[1014] Andrew Brenner. Mereological nihilism and the special arrangement question. Synthese, 192(5):1295–1314, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0619-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0619-
7.pdf.

Reeder:2015:ZAI

[1015] Patrick Reeder. Zeno’s arrow and the infinitesimal calculus. Synthese, 192 (5):1315–1335, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0620-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0620-1.pdf.

Parent:2015:EKW

[1016] T. Parent. Externalism and “knowing what” one thinks. Synthese, 192 (5):1337–1350, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0624-x;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0624-x.pdf.

Warren:2015:CCC

[1017] Jared Warren. Conventionalism, consistency, and consistency sentences. Synthese, 192(5):1351–1371, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0626-8;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0626-8.pdf.

McQueen:2015:MAP

[1018] Kelvin J. McQueen. Mass additivity and a priori entailment. Synthese, 192(5):1373–1392, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0627-7;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0627-7.pdf.

Felka:2015:PNS

[1019] Katharina Felka. On the presuppositions of number sentences. Synthese, 192(5):1393–1412, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0629-5;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0629-5.pdf.

Steinkruger:2015:AAS

[1020] Philipp Steinkrüger. Aristotle’s assertoric syllogistic and modern relevance logic. Synthese, 192(5):1413–1444, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0631-y;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0631-
y.pdf.

Buskell:2015:HSO

[1021] Andrew Buskell. How to be skilful: opportunistic robustness and normative sensitivity. Synthese, 192(5):1445–1466, May 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0634-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0634-
8.pdf.

Staffel:2015:MOI

[1022] Julia Staffel. Measuring the overall incoherence of credence functions. Synthese, 192(5):1467–1493, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0640-x;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0640-x.pdf.

Kyle:2015:NOI

[1023] Brent G. Kyle. The new and old ignorance puzzles: How badly do we need closure? Synthese, 192(5):1495–1525, May 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0642-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0642-
8.pdf.

Benetreau-Dupin:2015:BWK

[1024] Yann Benétreau-Dupin. The Bayesian who knew too much. Synthese, 192 (5):1527–1542, May 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0647-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0647-3.pdf.

Toader:2015:OUN

[1025] Iulian D. Toader. Objectivity and understanding: a new reading of Carnap’s Aufbau. Synthese, 192(5):1543–1557, May 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0648-2;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0648-
2.pdf.

Dawid:2015:MWD

[1026] Richard Dawid and Karim P. Y. Thébault. Many worlds: decoherent or incoherent? Synthese, 192(5):1559–1580, May 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0650-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0650-
8.pdf.

Hamami:2015:LQ

[1027] Yacin Hamami and Floris Roelofsen. Logics of questions. Synthese, 192 (6):1581–1584, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0736-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0736-y.pdf.

Wisniewski:2015:IEL

[1028] Andrzej Wiśniewski and Dorota Leszczyńska-Jasion. Inferential erotetic logic meets inquisitive semantics. Synthese, 192(6):1585–1608, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0355-4;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0355-4.pdf.

Hamami:2015:IMI

[1029] Yacin Hamami. The interrogative model of inquiry meets dynamic epistemic logics. Synthese, 192(6):1609–1642, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0460-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0460-
z.pdf.

Ciardelli:2015:IDE

[1030] Ivano A. Ciardelli and Floris Roelofsen. Inquisitive dynamic epistemic logic. Synthese, 192(6):1643–1687, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0404-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0404-
7.pdf.

Ciardelli:2015:SLD

[1031] Ivano Ciardelli, Jeroen Groenendijk, and Floris Roelofsen. On the semantics and logic of declaratives and interrogatives. Synthese, 192(6): 1689–1728, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0352-7;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0352-7.pdf.

Spector:2015:USE

[1032] Benjamin Spector and Paul Egré. A uniform semantics for embedded interrogatives: an answer, not necessarily the answer. Synthese, 192 (6):1729–1784, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0722-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0722-4.pdf.

Huisman:2015:RFA

[1033] Leendert Huisman. Reflecting on finite additivity. Synthese, 192(6): 1785–1797, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0652-6;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0652-6.pdf.

Donaldson:2015:PM

[1034] Thomas Donaldson. Platitudes in mathematics. Synthese, 192(6): 1799–1820, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0653-5;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0653-5.pdf.

Naumov:2015:LC

[1035] Pavel Naumov and Jia Tao. Logic of confidence. Synthese, 192(6): 1821–1838, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0655-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0655-3.pdf.

Radulescu:2015:LI

[1036] Alexandru Radulescu. The logic of indexicals. Synthese, 192(6): 1839–1860, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0659-7;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0659-7.pdf.

Stapleford:2015:EVA

[1037] Scott Stapleford. Epistemic versus all things considered requirements. Synthese, 192(6):1861–1881, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0660-1;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0660-1.pdf.

Dohrn:2015:EAH

[1038] Daniel Dohrn. Egan and agents: How evidential decision theory can deal with Egan’s dilemma. Synthese, 192(6):1883–1908, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0661-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0661-0.pdf.

Steward:2015:YSC

[1039] Stephen Steward. Ya shouldn’ta couldn’ta wouldn’ta. Synthese, 192 (6):1909–1921, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0663-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0663-y.pdf.

Sinigaglia:2015:PAR

[1040] Corrado Sinigaglia and Stephen A. Butterfill. On a puzzle about relations between thought, experience and the motoric. Synthese, 192 (6):1923–1936, June 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0672-x;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0672-x.pdf.

Szekely:2015:LRT

[1041] Gergely Székely. Logic and relativity theory. Synthese, 192(7):1937–1938, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0622-z;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0622-z.pdf.

Beziau:2015:RUL

[1042] Jean-Yves Beziau. The relativity and universality of logic. Synthese, 192 (7):1939–1954, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0419-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0419-0.pdf.

Cooper:2015:MDC

[1043] S. Barry Cooper. The machine as data: a computational view of emergence and definability. Synthese, 192(7):1955–1988, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0803-4;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0803-4.pdf.

Wuthrich:2015:QIT

[1044] Christian Wüthrich. A quantum-information-theoretic complement to a general-relativistic implementation of a beyond-Turing computer. Synthese, 192(7):1989–2008, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0502-6;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0502-6.pdf.

Benda:2015:AFR

[1045] Thomas Benda. An axiomatic foundation of relativistic spacetime. Synthese, 192(7):2009–2024, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0345-6;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0345-6.pdf.

Ferenczi:2015:PDS

[1046] Miklós Ferenczi. Probabilities defined on standard and non-standard cylindric set algebras. Synthese, 192(7):2025–2033, July 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0444-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0444-
z.pdf.

Friend:2015:ESH

[1047] Michèle Friend. On the epistemological significance of the Hungarian Project. Synthese, 192(7):2035–2051, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0608-x;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0608-
x.pdf.

Gomori:2015:FSS

[1048] Márton Gömöri and László E. Szabó. Formal statement of the special principle of relativity. Synthese, 192(7):2053–2076, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
013-0374-1;   http://link.springer.com/content/pdf/10.1007/
s11229-013-0374-1.pdf.

Govindarajalulu:2015:PVP

[1049] Naveen Sundar Govindarajalulu, Selmer Bringsjord, and Joshua Taylor. Proof verification and proof discovery for relativity. Synthese, 192 (7):2077–2094, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0424-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0424-3.pdf.

Jenei:2015:CCG

[1050] Sándor Jenei and Franco Montagna. A classification of certain grouplike FL_{e}-chains. Synthese, 192(7):2095–2121, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0409-2;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0409-

2.pdf. See erratum [1160].

Manzano:2015:VH

[1051] María Manzano and Enrique Alonso. Visions of Henkin. Synthese, 192 (7):2123–2138, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-013-0389-7;   http://link.springer.com/content/
pdf/10.1007/s11229-013-0389-7.pdf. See note [1422].

Marcus:2015:SSE

[1052] Solomon Marcus. Starting from the scenario Euclid–Bolyai–Einstein. Synthese, 192(7):2139–2149, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0346-5;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0346-5.pdf.

Mikulas:2015:ETR

[1053] Szabolcs Mikulás. The equational theories of representable residuated semigroups. Synthese, 192(7):2151–2158, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0513-3;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0513-
3.pdf.

Mikulas:2015:CET

[1054] Szabolcs Mikulás, Ildikó Sain, and András Simon. Complexity of equational theory of relational algebras with standard projection elements. Synthese, 192(7):2159–2182, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0689-1;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0689-
1.pdf.

Molnar:2015:ARD

[1055] Attila Molnár and Gergely Székely. Axiomatizing relativistic dynamics using formal thought experiments. Synthese, 192(7):2183–2222, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0545-8;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0545-8.pdf.

Sayed-Ahmed:2015:NEA

[1056] Tarek Sayed-Ahmed. Neat embeddings as adjoint situations. Synthese, 192(7):2223–2259, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-013-0344-7;   http://link.springer.com/
content/pdf/10.1007/s11229-013-0344-7.pdf.

Stannett:2015:MOS

[1057] Mike Stannett. Motion and observation in a single-particle universe. Synthese, 192(7):2261–2271, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0489-z;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0489-z.pdf.

Svarny:2015:DBE

[1058] Petr Svarný. Does branching explain flow of time or the other way around? Synthese, 192(7):2273–2292, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0638-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0638-
4.pdf.

Valyi:2015:ASF

[1059] Sándor Vályi. On the axiomatizability of some first-order spatiotemporal theories. Synthese, 192(7):2293–2309, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0365-2;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0365-
2.pdf.

Yuan:2015:REB

[1060] Yongfeng Yuan and Shier Ju. Rational evaluation in belief revision. Synthese, 192(7):2311–2336, July 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0802-5;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0802-5.pdf.

Sher:2015:MJH

[1061] Gila Sher. In memoriam: Jaakko Hintikka. Synthese, 192(8):2337– 2338, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0882-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0882-2.pdf.

Larvor:2015:HPI

[1062] Brendan P. Larvor, Benedikt Löwe, and Dirk Schlimm. History and philosophy of infinity. Synthese, 192(8):2339–2344, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0857-3;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0857-3.pdf.

Nawar:2015:AF

[1063] Tamer Nawar. Aristotelian finitism. Synthese, 192(8):2345–2360, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0827-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0827-9.pdf.

Uckelman:2015:LCS

[1064] Sara L. Uckelman. The logic of categorematic and syncategorematic infinity. Synthese, 192(8):2361–2377, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0670-z;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0670-
z.pdf.

Cortese:2015:IBM

[1065] João Figueiredo Nobre Cortese. Infinity between mathematics and apologetics: Pascal’s notion of infinite distance. Synthese, 192(8): 2379–2393, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0599-7;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0599-7.pdf.

DeBianchi:2015:WSG

[1066] Silvia De Bianchi. When series go in indefinitum, ad infinitum and in infinitum: concepts of infinity in Kant’s antinomy of pure reason. Synthese, 192(8):2395–2412, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0813-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0813-
2.pdf.

Incurvati:2015:CF

[1067] Luca Incurvati. On the concept of finitism. Synthese, 192(8):2413– 2436, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0639-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0639-3.pdf.

Rin:2015:TRC

[1068] Benjamin Rin. Transfinite recursion and computation in the iterative conception of set. Synthese, 192(8):2437–2462, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0560-9;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0560-9.pdf.

Antos:2015:MCS

[1069] Carolin Antos, Sy-David Friedman, Radek Honzik, and Claudio Ternullo. Multiverse conceptions in set theory. Synthese, 192(8):2463–2488, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0819-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0819-9.pdf.

Pantsar:2015:SHI

[1070] Markus Pantsar. In search of ℵ_{0}: how infinity can be created. Synthese, 192(8):2489–2511, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0775-4;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0775-4.pdf.

Koren:2015:HCV

[1071] Ladislav Koren. Hinge commitments vis-à-vis the transmission problem. Synthese, 192(8):2513–2534, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0664-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0664-
x.pdf.

Nissan-Rozen:2015:TRD

[1072] Ittay Nissan-Rozen. A triviality result for the “desire by necessity” thesis. Synthese, 192(8):2535–2556, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0666-8;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0666-
8.pdf.

Roberts:2015:NED

[1073] Robert C. Roberts and Ryan West. Natural epistemic defects and corrective virtues. Synthese, 192(8):2557–2576, August 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0669-5;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0669-
5.pdf.

Logan:2015:SSC

[1074] Shay Allen Logan. The semantics of social constructivism. Synthese, 192 (8):2577–2598, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0674-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0674-8.pdf.

Sebens:2015:KCE

[1075] Charles T. Sebens. Killer collapse: empirically probing the philosophically unsatisfactory region of GRW. Synthese, 192(8):2599–2615, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0680-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0680-x.pdf.

Worsnip:2015:NSW

[1076] Alex Worsnip. Narrow-scoping for wide-scopers. Synthese, 192(8): 2617–2646, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0681-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0681-9.pdf.

Stuchlik:2015:EDE

[1077] Joshua Stuchlik. Epistemological disjunctivism and easy knowledge. Synthese, 192(8):2647–2665, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0683-7;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0683-7.pdf.

Demircioglu:2015:GPE

[1078] Erhan Demircioğlu. The given in perceptual experience. Synthese, 192 (8):2667–2693, August 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0730-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0730-4.pdf.

Jackson:2015:ISI

[1079] Magdalena Balcerak Jackson. Introduction to the special issue “the roles of experience in a priori knowledge”. Synthese, 192(9):2695–2699, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0838-6;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0838-6.pdf.

Casullo:2015:FCP

[1080] Albert Casullo. Four challenges to the a priori–a posteriori distinction. Synthese, 192(9):2701–2724, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0341-x;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0341-
x.pdf.

Jenkins:2015:TCP

[1081] C. S. I. Jenkins and Masashi Kasaki. The traditional conception of the a priori. Synthese, 192(9):2725–2746, October 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-013-0394-x;
 http://link.springer.com/content/pdf/10.1007/s11229-013-0394-
x.pdf.

Grundmann:2015:HRS

[1082] Thomas Grundmann. How reliabilism saves the apriori/aposteriori distinction. Synthese, 192(9):2747–2768, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0422-5;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0422-
5.pdf.

Smithies:2015:IRL

[1083] Declan Smithies. Ideal rationality and logical omniscience. Synthese, 192(9):2769–2793, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0735-z;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0735-z.pdf.

Eder:2015:FRS

[1084] Günther Eder and Esther Ramharter. Formal reconstructions of St. Anselm’s ontological argument. Synthese, 192(9):2795–2825, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0682-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0682-8.pdf.

Praem:2015:PTE

[1085] Sara Kier Praëm and Asbjørn Steglich-Petersen. Philosophical thought experiments as heuristics for theory discovery. Synthese, 192(9):2827– 2842, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0684-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0684-6.pdf.

Ardourel:2015:DSP

[1086] Vincent Ardourel. A discrete solution for the paradox of Achilles and the tortoise. Synthese, 192(9):2843–2861, October 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0688-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0688-
2.pdf.

Wong:2015:NAN

[1087] Wai hung Wong and Zanja Yudell. A normative account of the need for explanation. Synthese, 192(9):2863–2885, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0690-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0690-8.pdf.

Conitzer:2015:DBA

[1088] Vincent Conitzer. A Dutch book against sleeping beauties who are evidential decision theorists. Synthese, 192(9):2887–2899, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0691-7;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0691-7.pdf.

Kraemer:2015:NPI

[1089] Daniel M. Kraemer. Natural probabilistic information. Synthese, 192(9): 2901–2919, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0692-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0692-6.pdf.

DesAutels:2015:TPI

[1090] Lane DesAutels. Toward a propensity interpretation of stochastic mechanism for the life sciences. Synthese, 192(9):2921–2953, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0694-4;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0694-4.pdf.

Palermos:2015:AEV

[1091] Spyridon Orestis Palermos. Active externalism, virtue reliabilism and scientific knowledge. Synthese, 192(9):2955–2986, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0695-3;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0695-3.pdf.

Taylor:2015:JEI

[1092] Samuel A. Taylor. Is justification easy or impossible? Getting acquainted with a middle road. Synthese, 192(9):2987–3009, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0697-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0697-1.pdf.

Schwarz:2015:LMU

[1093] Wolfgang Schwarz. Lost memories and useless coins: revisiting the absentminded driver. Synthese, 192(9):3011–3036, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0699-z;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0699-z.pdf.

Baker:2015:WTU

[1094] Derek Baker. Why transparency undermines economy. Synthese, 192(9): 3037–3050, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0700-x;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0700-x.pdf.

Griffith:2015:EHN

[1095] Aaron M. Griffith. Erratum to: How negative truths are made true. Synthese, 192(9):3051, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0623-y;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0623-y.pdf.

Hausman:2015:ESG

[1096] Daniel M. Hausman, Reuben Stern, and Naftali Weinberger. Erratum to: Systems without a graphical causal representation. Synthese, 192 (9):3053, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0686-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0686-4.pdf.

Sole:2015:IST

[1097] Albert Solé and Carl Hoefer. Introduction: space-time and the wave function. Synthese, 192(10):3055–3070, October 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0826-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0826-
x.pdf.

Barrett:2015:PWM

[1098] Jeffrey A. Barrett. Pure wave mechanics and the very idea of empirical adequacy. Synthese, 192(10):3071–3104, October 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0698-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0698-
0.pdf.

Ney:2015:FPO

[1099] Alyssa Ney. Fundamental physical ontologies and the constraint of empirical coherence: a defense of wave function realism. Synthese, 192(10): 3105–3124, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0633-9;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0633-9.pdf.

Norsen:2015:CWF

[1100] Travis Norsen, Damiano Marian, and Xavier Oriols. Can the wave function in configuration space be replaced by single-particle wave functions in physical space? Synthese, 192(10):3125–3151, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

014-0577-0;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0577-0.pdf.

Callender:2015:OWO

[1101] Craig Callender. One world, one beable. Synthese, 192(10):3153– 3177, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0582-3;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0582-3.pdf.

Dorato:2015:LNR

[1102] Mauro Dorato. Laws of nature and the reality of the wave function. Synthese, 192(10):3179–3201, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0696-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0696-
2.pdf.

Suarez:2015:BD

[1103] Mauricio Suárez. Bohmian dispositions. Synthese, 192(10):3203– 3228, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0741-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0741-1.pdf.

Egg:2015:POQ

[1104] Matthias Egg and Michael Esfeld. Primitive ontology and quantum state in the GRW matter density theory. Synthese, 192(10):3229–3245, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0590-3;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0590-3.pdf.

Myrvold:2015:WW

[1105] Wayne C. Myrvold. What is a wavefunction? Synthese, 192(10): 3247–3274, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0635-7;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0635-7.pdf.

vanStrien:2015:CNM

[1106] Marij van Strien. Continuity in nature and in mathematics: Boltzmann and Poincaré. Synthese, 192(10):3275–3295, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0701-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0701-9.pdf.

Henschen:2015:CPC

[1107] Tobias Henschen. Ceteris paribus conditions and the interventionist account of causality. Synthese, 192(10):3297–3311, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0703-7;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0703-7.pdf.

Grunbaum:2015:FAH

[1108] Thor Grünbaum. The feeling of agency hypothesis: a critique. Synthese, 192(10):3313–3337, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0704-6;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0704-6.pdf.

Navarro:2015:NAB

[1109] Jesús Navarro. No achievement beyond intention. Synthese, 192(10): 3339–3369, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0708-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0708-2.pdf.

Serban:2015:SLM

[1110] Maria Serban. The scope and limits of a mechanistic view of computational explanation. Synthese, 192(10):3371–3396, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0709-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0709-1.pdf.

Knowles:2015:GWH

[1111] Robert Knowles and David Liggins. Good weasel hunting. Synthese, 192(10):3397–3412, October 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0711-7;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0711-7.pdf.

Jantzen:2015:OM

[1112] Benjamin C. Jantzen, Deborah G. Mayo, and Lydia Patton. Ontology & methodology. Synthese, 192(11):3413–3423, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0972-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0972-1.pdf.

Ruetsche:2015:SGL

[1113] Laura Ruetsche. The Shaky Game +25, or: on locavoracity. Synthese, 192(11):3425–3442, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0551-x;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0551-x.pdf.

Patton:2015:MRM

[1114] Lydia Patton. Methodological realism and modal resourcefulness: out of the web and into the mine. Synthese, 192(11):3443–3462, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0917-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0917-8.pdf.

Shech:2015:SMG

[1115] Elay Shech. Scientific misrepresentation and guides to ontology: the need for representational code and contents. Synthese, 192(11):3463– 3485, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0506-2;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0506-2.pdf.

Clatterbuck:2015:DBW

[1116] Hayley Clatterbuck. Drift beyond Wright–Fisher. Synthese, 192(11): 3487–3507, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0598-8;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0598-8.pdf.

Hoover:2015:OSS

[1117] Kevin D. Hoover. The ontological status of shocks and trends in macroeconomics. Synthese, 192(11):3509–3532, November 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0503-5;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0503-
5.pdf.

Spanos:2015:ESM

[1118] Aris Spanos and Deborah G. Mayo. Error statistical modeling and inference: Where methodology meets ontology. Synthese, 192(11):3533– 3555, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0744-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0744-y.pdf.

Angner:2015:NSV

[1119] Erik Angner. “To navigate safely in the vast sea of empirical facts”. Synthese, 192(11):3557–3575, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0552-9;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0552-
9.pdf.

Woodward:2015:MOI

[1120] James Woodward. Methodology, ontology, and interventionism. Synthese, 192(11):3577–3599, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-014-0479-1;   http://link.springer.
com/content/pdf/10.1007/s11229-014-0479-1.pdf.

Danks:2015:GDS

[1121] David Danks. Goal-dependence in (scientific) ontology. Synthese, 192 (11):3601–3616, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0649-1;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0649-1.pdf.

Jantzen:2015:PSN

[1122] Benjamin C. Jantzen. Projection, symmetry, and natural kinds. Synthese, 192(11):3617–3646, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-014-0637-5;   http://link.springer.
com/content/pdf/10.1007/s11229-014-0637-5.pdf.

LeMorvan:2015:IKN

[1123] Pierre Le Morvan. On the ignorance, knowledge, and nature of propositions. Synthese, 192(11):3647–3662, November 2015. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0712-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0712-
6.pdf.

Evers:2015:SEN

[1124] Daan Evers. Street on evolution and the normativity of epistemic reasons. Synthese, 192(11):3663–3676, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0714-4;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0714-
4.pdf.

Borges:2015:SD

[1125] Rodrigo Borges. On synchronic dogmatism. Synthese, 192(11):3677– 3693, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0715-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0715-3.pdf.

Petkov:2015:EUC

[1126] Stefan Petkov. Explanatory unification and conceptualization. Synthese, 192(11):3695–3717, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0716-2;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0716-2.pdf.

Ferguson:2015:TPS

[1127] Thomas Macaulay Ferguson. Two paradoxes of semantic information. Synthese, 192(11):3719–3730, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0717-1;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0717-
1.pdf.

Romero:2015:WTI

[1128] Felipe Romero. Why there isn’t inter-level causation in mechanisms. Synthese, 192(11):3731–3755, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0718-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0718-
0.pdf.

Clark:2015:WEM

[1129] Andy Clark. What ‘extended me’ knows. Synthese, 192(11):3757– 3775, November 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0719-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0719-z.pdf.

Eronen:2015:UTM

[1130] Markus Eronen and Raphael van Riel. Understanding through modeling: the explanatory power of inadequate representation. Synthese, 192(12): 3777–3780, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0973-0;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0973-0.pdf.

deRegt:2015:SUT

[1131] Henk W. de Regt. Scientific understanding: truth or dare? Synthese, 192(12):3781–3797, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0538-7;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0538-7.pdf.

Kelp:2015:UP

[1132] Christoph Kelp. Understanding phenomena. Synthese, 192(12):3799– 3816, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0616-x;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0616-x.pdf.

Kuorikoski:2015:ERS

[1133] Jaakko Kuorikoski and Petri Ylikoski. External representations and scientific understanding. Synthese, 192(12):3817–3837, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0591-2;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0591-2.pdf.

vanRiel:2015:CMB

[1134] Raphael van Riel. The content of model-based information. Synthese, 192(12):3839–3858, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0728-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0728-y.pdf.

Toon:2015:WU

[1135] Adam Toon. Where is the understanding? Synthese, 192(12):3859– 3875, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0702-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0702-8.pdf.

Brossel:2015:RES

[1136] Peter Brössel. On the role of explanatory and systematic power in scientific reasoning. Synthese, 192(12):3877–3913, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0870-6;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0870-6.pdf.

Buckner:2015:FKS

[1137] Cameron Buckner. Functional kinds: a skeptical look. Synthese, 192(12): 3915–3942, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0606-z;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0606-z.pdf.

Irvine:2015:MRN

[1138] Elizabeth Irvine. Models, robustness, and non-causal explanation: a foray into cognitive science and biology. Synthese, 192(12):3943– 3959, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0524-0;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0524-0.pdf.

Eronen:2015:RR

[1139] Markus I. Eronen. Robustness and reality. Synthese, 192(12):3961– 3977, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0801-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0801-6.pdf.

Frigg:2015:AFA

[1140] Roman Frigg, Leonard A. Smith, and David A. Stainforth. An assessment of the foundational assumptions in high-resolution climate projections: the case of UKCP09. Synthese, 192(12):3979–4008, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0739-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0739-8.pdf.

Turri:2015:EFN

[1141] John Turri. Evidence of factive norms of belief and decision. Synthese, 192(12):4009–4030, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0727-z;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0727-z.pdf.

Nicolai:2015:DTO

[1142] Carlo Nicolai. Deflationary truth and the ontology of expressions. Synthese, 192(12):4031–4055, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0729-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0729-
x.pdf.

Moeller:2015:CKC

[1143] Emil Frederik Lundbjerg Moeller. Consuming knowledge claims across contexts. Synthese, 192(12):4057–4070, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0732-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0732-
2.pdf.

Grossi:2015:SAL

[1144] Davide Grossi and Fernando R. Velázquez-Quesada. Syntactic awareness in logical dynamics. Synthese, 192(12):4071–4105, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0733-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0733-1.pdf.

Conitzer:2015:CRC

[1145] Vincent Conitzer. Can rational choice guide us to correct de se beliefs? Synthese, 192(12):4107–4119, December 2015. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0737-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0737-
x.pdf.

Girard:2016:PD

[1146] Patrick Girard and Koji Tanaka. Paraconsistent dynamics. Synthese, 193(1):1–14, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0740-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0740-2.pdf.

Tulenheimo:2016:TFC

[1147] Tero Tulenheimo. The two faces of compatibility with justified beliefs. Synthese, 193(1):15–30, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0742-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0742-
0.pdf.

Giovannini:2016:BGB

[1148] Eduardo N. Giovannini. Bridging the gap between analytic and synthetic geometry: Hilbert’s axiomatic approach. Synthese, 193(1):31–70, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0743-z;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0743-z.pdf.

Miller:2016:AJE

[1149] Seumas Miller. Assertions, joint epistemic actions and social practices. Synthese, 193(1):71–94, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0745-x;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0745-x.pdf.

Giberman:2016:MPN

[1150] Daniel Giberman. Moving parts: a new indexical treatment of context-shifting predication. Synthese, 193(1):95–124, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0747-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0747-8.pdf.

Engelhardt:2016:WWT

[1151] Jeff Engelhardt. What we talk about when we talk about content externalism. Synthese, 193(1):125–143, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0748-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0748-
7.pdf.

Kornmesser:2016:FBA

[1152] Stephan Kornmesser. A frame-based approach for theoretical concepts. Synthese, 193(1):145–166, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0750-0;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0750-0.pdf.

Horvath:2016:CAN

[1153] Joachim Horvath. Conceptual analysis and natural kinds: the case of knowledge. Synthese, 193(1):167–184, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0751-z;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0751-
z.pdf.

Liebesman:2016:DVU

[1154] David Liebesman. Does vagueness underlie the mass/count distinction? Synthese, 193(1):185–203, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0752-y;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0752-
y.pdf.

Christensen:2016:MH

[1155] Ryan Christensen. McGee on Horwich. Synthese, 193(1):205–218, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0753-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0753-x.pdf.

Stewart:2016:CCV

[1156] Rush T. Stewart. Conditional choice with a vacuous second tier. Synthese, 193(1):219–243, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0754-9;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0754-9.pdf.

Ismael:2016:HDC

[1157] Jenann Ismael. How do causes depend on us? The many faces of perspectivalism. Synthese, 193(1):245–267, January 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0757-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0757-
6.pdf.

Hartl:2016:MSY

[1158] Peter Hartl. Modal scepticism, Yablo-style conceivability, and analogical reasoning. Synthese, 193(1):269–291, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0759-4;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0759-
4.pdf.

Wright:2016:TKJ

[1159] Stephen Wright. The transmission of knowledge and justification. Synthese, 193(1):293–311, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0760-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0760-y.pdf.

Jenei:2016:ECC

[1160] Sándor Jenei and Franco Montagna. Erratum to: A classification of certain group-like FL_{e}-chains. Synthese, 193(1):313, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0778-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0778-1.pdf. See [1050].

Clark:2016:EWE

[1161] Andy Clark. Erratum to: What ‘extended me’ knows. Synthese, 193 (1):315, January 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0788-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0788-z.pdf.

Molinini:2016:IEO

[1162] Daniele Molinini, Fabrice Pataut, and Andrea Sereni. Indispensability and explanation: an overview and introduction. Synthese, 193(2): 317–332, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0998-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0998-4.pdf.

Baker:2016:PIB

[1163] Alan Baker. Parsimony and inference to the best mathematical explanation. Synthese, 193(2):333–350, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0723-3;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0723-
3.pdf.

Pataut:2016:CPI

[1164] Fabrice Pataut. Comments on “parsimony and inference to the best mathematical explanation”. Synthese, 193(2):351–363, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0706-4;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0706-4.pdf.

Baron:2016:EDI

[1165] Sam Baron. The explanatory dispensability of idealizations. Synthese, 193(2):365–386, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0517-z;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0517-z.pdf.

DeBianchi:2016:WER

[1166] Silvia De Bianchi. Which explanatory role for mathematics in scientific models? Reply to “the explanatory dispensability of idealizations”. Synthese, 193(2):387–401, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0795-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0795-
0.pdf.

Molinini:2016:EEE

[1167] Daniele Molinini. Evidence, explanation and enhanced indispensability. Synthese, 193(2):403–422, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0494-2;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0494-2.pdf. See reply [1168].

Sereni:2016:EEM

[1168] Andrea Sereni. Equivalent explanations and mathematical realism. Reply to “evidence, explanation, and enhanced indispensability”. Synthese, 193 (2):423–434, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0491-5;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0491-5.pdf. See [1167].

Busch:2016:SSR

[1169] Jacob Busch and Joe Morrison. Should scientific realists be Platonists? Synthese, 193(2):435–449, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0676-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0676-
6.pdf.

Hunt:2016:IPC

[1170] Josh Hunt. Indispensability and the problem of compatible explanations. Synthese, 193(2):451–467, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0667-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0667-
7.pdf.

Panza:2016:VIA

[1171] Marco Panza and Andrea Sereni. The varieties of indispensability arguments. Synthese, 193(2):469–516, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0977-9;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0977-
9.pdf.

Galinon:2016:NIR

[1172] Henri Galinon. Naturalizing indispensability: a rejoinder to ‘the varieties of indispensability arguments’. Synthese, 193(2):517–530, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0978-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0978-8.pdf.

Liggins:2016:GIA

[1173] David Liggins. Grounding and the indispensability argument. Synthese, 193(2):531–548, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0478-2;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0478-2.pdf. See reply [1174].

Plebani:2016:NCG

[1174] Matteo Plebani. Nominalistic content, grounding, and covering generalizations: Reply to ‘Grounding and the indispensability argument’. Synthese, 193(2):549–558, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0428-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0428-

z.pdf. See [1173].

Gladziejewski:2016:PCR

[1175] Pawel Gladziejewski. Predictive coding and representationalism. Synthese, 193(2):559–582, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0762-9;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0762-9.pdf.

Janssen-Lauret:2016:CIO

[1176] Frédérique Janssen-Lauret. Committing to an individual: ontological commitment, reference and epistemology. Synthese, 193(2):583– 604, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0763-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0763-8.pdf.

Taylor:2016:QMF

[1177] David E. Taylor. Quine on matters of fact. Synthese, 193(2):605– 636, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0765-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0765-6.pdf.

Nathan:2016:MMB

[1178] Marco J. Nathan and Guillermo Del Pinal. Mapping the mind: bridge laws and the psycho-neural interface. Synthese, 193(2):637–657, February 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0769-2;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0769-2.pdf.

Aagotnes:2016:ISI

[1179] Thomas Ågotnes, Giacomo Bonanno, and Wiebe van der Hoek. Introduction to the special issue. Synthese, 193(3):659–662, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1017-0;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1017-0.pdf.

Apt:2016:PSN

[1180] Krzysztof R. Apt, Evangelos Markakis, and Sunil Simon. Paradoxes in social networks with multiple products. Synthese, 193(3):663–687, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0864-4;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0864-4.pdf.

Conitzer:2016:SMS

[1181] Vincent Conitzer. On Stackelberg mixed strategies. Synthese, 193(3): 689–703, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0927-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0927-6.pdf.

LiCalzi:2016:BCC

[1182] Marco LiCalzi and Nadia Maagli. Bargaining over a common categorisation. Synthese, 193(3):705–723, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0790-5;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0790-
5.pdf.

Lorini:2016:MLI

[1183] Emiliano Lorini. A minimal logic for interactive epistemology. Synthese, 193(3):725–755, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0960-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0960-5.pdf.

Tsakas:2016:CBE

[1184] Elias Tsakas. Correlated-belief equilibrium. Synthese, 193(3):757– 779, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0791-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0791-4.pdf.

Bradfield:2016:POB

[1185] Julian Bradfield, Julian Gutierrez, and Michael Wooldridge. Partialorder Boolean games: informational independence in a logic-based model of strategic interaction. Synthese, 193(3):781–811, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0991-y;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0991-y.pdf.

Renne:2016:LTE

[1186] Bryan Renne, Joshua Sack, and Audrey Yap. Logics of temporalepistemic actions. Synthese, 193(3):813–849, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0773-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0773-
6.pdf.

Katz:2016:ART

[1187] Matthew Katz. Analog representations and their users. Synthese, 193 (3):851–871, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0774-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0774-5.pdf.

Fagan:2016:SCS

[1188] Melinda Bonnie Fagan. Stem cells and systems models: clashing views of explanation. Synthese, 193(3):873–907, March 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0776-3;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0776-
3.pdf.

Schwartzkopff:2016:STR

[1189] Robert Schwartzkopff. Singular terms revisited. Synthese, 193(3): 909–936, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0777-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0777-2.pdf.

Hartimo:2016:SRH

[1190] Mirja Hartimo and Mitsuhiro Okada. Syntactic reduction in Husserl’s early phenomenology of arithmetic. Synthese, 193(3):937–969, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0779-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0779-0.pdf.

Jackson:2016:RTF

[1191] Alexander Jackson. From relative truth to finean non-factualism. Synthese, 193(3):971–989, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0787-0;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0787-0.pdf.

Miller:2016:WHA

[1192] Boaz Miller. What is Hacking’s argument for entity realism? Synthese, 193(3):991–1006, March 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0789-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0789-y.pdf.

Gebharter:2016:ISI

[1193] Alexander Gebharter and Gerhard Schurz. Introduction to the special issue “Causation, probability, and truth — the philosophy of Clark Glymour”. Synthese, 193(4):1007–1010, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-1007-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-1007-
7.pdf.

Zhang:2016:TFF

[1194] Jiji Zhang and Peter Spirtes. The three faces of faithfulness. Synthese, 193(4):1011–1027, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0673-9;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0673-9.pdf.

Eberhardt:2016:GGC

[1195] Frederick Eberhardt. Green and grue causal variables. Synthese, 193 (4):1029–1046, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0832-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0832-z.pdf.

Woodward:2016:PVC

[1196] James Woodward. The problem of variable choice. Synthese, 193(4): 1047–1072, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0810-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0810-5.pdf.

Schurz:2016:CTC

[1197] Gerhard Schurz and Alexander Gebharter. Causality as a theoretical concept: explanatory warrant and empirical content of the theory of causal nets. Synthese, 193(4):1073–1103, April 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0630-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0630-
z.pdf.

Schurz:2016:ECT

[1198] Gerhard Schurz and Alexander Gebharter. Erratum to: Causality as a theoretical concept: explanatory warrant and empirical content of the theory of causal nets. Synthese, 193(4):1105–1106, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1019-y;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1019-y.pdf.

Hagmayer:2016:CBN

[1199] York Hagmayer. Causal Bayes nets as psychological theories of causal reasoning: evidence from psychological research. Synthese, 193(4): 1107–1126, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0734-0;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0734-0.pdf.

Nager:2016:CPE

[1200] Paul M. Näger. The causal problem of entanglement. Synthese, 193 (4):1127–1155, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0668-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0668-6.pdf.

Hitchcock:2016:CID

[1201] Christopher Hitchcock. Conditioning, intervening, and decision. Synthese, 193(4):1157–1176, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0710-8;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0710-8.pdf.

Hoffmann-Kolss:2016:BPC

[1202] Vera Hoffmann-Kolss. Of brains and planets: on a causal criterion for mind-brain identities. Synthese, 193(4):1177–1189, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0671-y;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0671-y.pdf.

Kelly:2016:RRR

[1203] Kevin T. Kelly, Konstantin Genin, and Hanti Lin. Realism, rhetoric, and reliability. Synthese, 193(4):1191–1223, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0993-9;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0993-
9.pdf.

Wenmackers:2016:NTA

[1204] Sylvia Wenmackers and Jan-Willem Romeijn. New theory about old evidence. Synthese, 193(4):1225–1250, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0632-x;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0632-
x.pdf.

Glymour:2016:CGR

[1205] Clark Glymour. Clark Glymour’s responses to the contributions to the Synthese special issue “Causation, probability, and truth: the philosophy of Clark Glymour”. Synthese, 193(4):1251–1285, April 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1021-4;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1021-4.pdf.

Bechtel:2016:INR

[1206] William Bechtel. Investigating neural representations: the tale of place cells. Synthese, 193(5):1287–1321, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0480-8;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0480-
8.pdf.

Calvo:2016:PPN

[1207] Paco Calvo. The philosophy of plant neurobiology: a manifesto. Synthese, 193(5):1323–1343, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1040-1;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1040-1.pdf.

Cheng:2016:WEM

[1208] Sen Cheng and Markus Werning. What is episodic memory if it is a natural kind? Synthese, 193(5):1345–1385, May 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0628-6;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0628-
6.pdf.

Hochstein:2016:OMM

[1209] Eric Hochstein. One mechanism, many models: a distributed theory of mechanistic explanation. Synthese, 193(5):1387–1407, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0844-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0844-8.pdf.

Koralus:2016:CVC

[1210] Philipp Koralus. Can visual cognitive neuroscience learn anything from the philosophy of language? Ambiguity and the topology of neural network models of multistable perception. Synthese, 193(5):1409–1432, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0518-y;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0518-y.pdf.

Malanowski:2016:EMU

[1211] Sarah Malanowski. Is episodic memory uniquely human? Evaluating the episodic-like memory research program. Synthese, 193(5):1433–1455, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0966-z;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0966-z.pdf.

Milkowski:2016:ECI

[1212] Marcin Milkowski. Explanatory completeness and idealization in large brain simulations: a mechanistic perspective. Synthese, 193(5):1457– 1478, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0731-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0731-3.pdf.

Olin:2016:BPS

[1213] Lauren Olin. Burge on perception and sensation. Synthese, 193(5): 1479–1508, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0531-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0531-1.pdf.

Boone:2016:CNR

[1214] Worth Boone and Gualtiero Piccinini. The cognitive neuroscience revolution. Synthese, 193(5):1509–1534, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-015-0783-4;   http://link.springer.
com/content/pdf/10.1007/s11229-015-0783-4.pdf.

dePasquale:2016:CMR

[1215] Jean-Frédéric de Pasquale and Pierre Poirier. Convolution and modal representations in Thagard and Stewart’s neural theory of creativity: a critical analysis. Synthese, 193(5):1535–1560, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0934-7;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0934-7.pdf.

Robins:2016:OMF

[1216] Sarah K. Robins. Optogenetics and the mechanism of false memory. Synthese, 193(5):1561–1583, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1045-9;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1045-9.pdf.

Stinson:2016:MPR

[1217] Catherine Stinson. Mechanisms in psychology: ripping nature at its seams. Synthese, 193(5):1585–1614, May 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0871-5;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0871-
5.pdf.

Whittington:2016:LKV

[1218] Lee John Whittington. Luck, knowledge and value. Synthese, 193(6): 1615–1633, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0794-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0794-1.pdf.

Sylvan:2016:ID

[1219] Kurt Sylvan. The illusion of discretion. Synthese, 193(6):1635–1665, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0796-z;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0796-z.pdf.

Woods:2016:ADC

[1220] Jack Woods. Assertion, denial, content, and (logical) form. Synthese, 193 (6):1667–1680, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0797-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0797-y.pdf.

Kim:2016:DIB

[1221] Namjoong Kim. A dilemma for the imprecise Bayesian. Synthese, 193 (6):1681–1702, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0800-7;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0800-7.pdf.

Vranas:2016:NFI

[1222] Peter B. M. Vranas. New foundations for imperative logic III: A general definition of argument validity. Synthese, 193(6):1703–1753, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0805-2;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0805-2.pdf.

McClelland:2016:CSR

[1223] Tom McClelland. Can self-representationalism explain away the apparent irreducibility of consciousness? Synthese, 193(6):1755–1776, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0806-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0806-1.pdf.

Rochefort-Maranda:2016:CIV

[1224] Guillaume Rochefort-Maranda. On the correct interpretation of p values and the importance of random variables. Synthese, 193(6):1777–1793, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0807-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0807-0.pdf.

Ohtani:2016:WCP

[1225] Hiroshi Ohtani. Wittgenstein on context and philosophical pictures. Synthese, 193(6):1795–1816, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0809-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0809-y.pdf.

Mix:2016:NEA

[1226] Lucas Mix. Nested explanation in Aristotle and Mayr. Synthese, 193 (6):1817–1832, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0811-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0811-4.pdf.

Hahn:2016:NFA

[1227] Ulrike Hahn and Jos Hornikx. A normative framework for argument quality: argumentation schemes with a Bayesian foundation. Synthese, 193 (6):1833–1873, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0815-0;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0815-0.pdf.

Meadows:2016:SS

[1228] Toby Meadows. Sets and supersets. Synthese, 193(6):1875–1907, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

015-0818-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0818-x.pdf.

Kroedel:2016:GMC

[1229] Thomas Kroedel and Moritz Schulz. Grounding mental causation. Synthese, 193(6):1909–1923, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0820-3;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0820-3.pdf.

daSilva:2016:HHC

[1230] Jairo Jose da Silva. Husserl and Hilbert on completeness, still. Synthese, 193(6):1925–1947, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0821-2;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0821-2.pdf.

Leonard:2016:WMH

[1231] Matt Leonard. What is mereological harmony? Synthese, 193(6): 1949–1965, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0822-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0822-1.pdf.

Stefansson:2016:DC

[1232] H. Orri Stefánsson. Desirability of conditionals. Synthese, 193(6): 1967–1981, June 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0823-0;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0823-0.pdf.

Lombardi:2016:WSI

[1233] Olimpia Lombardi, Federico Holik, and Leonardo Vanni. What is Shannon information? Synthese, 193(7):1983–2012, July 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0824-z;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0824-
z.pdf.

DAlfonso:2016:BMA

[1234] Simon D’Alfonso. Belief merging with the aim of truthlikeness. Synthese, 193(7):2013–2034, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0825-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0825-y.pdf.

Tsai:2016:EEA

[1235] Cheng hung Tsai. Ethical expertise and the articulacy requirement. Synthese, 193(7):2035–2052, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0828-8;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0828-8.pdf.

Olen:2016:WSE

[1236] Peter Olen and Stephen P. Turner. Was Sellars an error theorist? Synthese, 193(7):2053–2075, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0829-7;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0829-7.pdf.

Todd:2016:BMF

[1237] Patrick Todd. On behalf of a mutable future. Synthese, 193(7): 2077–2095, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0830-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0830-1.pdf.

Galeazzi:2016:ELM

[1238] Paolo Galeazzi and Emiliano Lorini. Epistemic logic meets epistemic game theory: a comparison between multi-agent Kripke models and type spaces. Synthese, 193(7):2097–2127, July 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0834-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0834-
x.pdf.

Arruda:2016:CSA

[1239] Caroline T. Arruda and Daniel J. Povinelli. Chimps as secret agents. Synthese, 193(7):2129–2158, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0835-9;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0835-9.pdf.

Elster:2016:TBT

[1240] Jon Elster. Tool-box or toy-box? Hard obscurantism in economic modeling. Synthese, 193(7):2159–2184, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0836-8;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0836-
8.pdf.

Russo:2016:KDW

[1241] Andrew Russo. Kim’s dilemma: why mental causation is not productive. Synthese, 193(7):2185–2203, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0837-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0837-
7.pdf.

Endicott:2016:FSP

[1242] Ronald Endicott. Functionalism, superduperfunctionalism, and physicalism: lessons from supervenience. Synthese, 193(7):2205–2235, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0839-5;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0839-5.pdf.

Baron:2016:MF

[1243] Sam Baron. Metaphysics as fairness. Synthese, 193(7):2237–2259, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0842-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0842-x.pdf.

Barrios:2016:SE

[1244] Edison Barrios. Simple is not easy. Synthese, 193(7):2261–2305, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0843-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0843-9.pdf.

Howson:2016:DII

[1245] Colin Howson. Does information inform confirmation? Synthese, 193 (7):2307–2321, July 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0918-7;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0918-7.pdf.

Carter:2016:BC

[1246] J. Adam Carter, Benjamin W. Jarvis, and Katherine Rubin. Belief without credence. Synthese, 193(8):2323–2351, August 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0846-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0846-
6.pdf.

De:2016:ICT

[1247] Michael De. Intrinsicality and counterpart theory. Synthese, 193(8): 2353–2365, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0847-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0847-5.pdf.

Balaguer:2016:CAX

[1248] Mark Balaguer. Conceptual analysis and x-phi. Synthese, 193(8): 2367–2388, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0848-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0848-4.pdf.

Weatherall:2016:FBY

[1249] James Owen Weatherall. Fiber bundles, Yang–Mills theory, and general relativity. Synthese, 193(8):2389–2425, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0849-3;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0849-
3.pdf.

Lee:2016:EK

[1250] Christian Ryan Lee. Excluded knowledge. Synthese, 193(8):2427– 2452, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0858-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0858-2.pdf.

Chase:2016:VV

[1251] James Kennedy Chase. Voting and vagueness. Synthese, 193(8): 2453–2468, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0859-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0859-1.pdf.

Varga:2016:IEC

[1252] Somogy Varga. Interaction and extended cognition. Synthese, 193(8): 2469–2496, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0861-7;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0861-7.pdf.

Keven:2016:ENM

[1253] Nazim Keven. Events, narratives and memory. Synthese, 193(8): 2497–2517, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0862-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0862-6.pdf.

Hlobil:2016:SNU

[1254] Ulf Hlobil. Social norms and unthinkable options. Synthese, 193(8): 2519–2537, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0863-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0863-5.pdf.

Sanz:2016:DVJ

[1255] Wagner de Campos Sanz and Hermógenes Oliveira. On Dummett’s verificationist justification procedure. Synthese, 193(8):2539–2559, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0865-3;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0865-3.pdf.

Champagne:2016:BPO

[1256] Marc Champagne. Brandom, Peirce, and the overlooked friction of contrapiction. Synthese, 193(8):2561–2576, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0866-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0866-
2.pdf.

Dinges:2016:SPI

[1257] Alexander Dinges. Skeptical pragmatic invariantism: good, but not good enough. Synthese, 193(8):2577–2593, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0867-1;

 http://link.springer.com/content/pdf/10.1007/s11229-015-0867-
1.pdf.

Wallach:2016:NCT

[1258] Efraim Wallach. Niche construction theory as an explanatory framework for human phenomena. Synthese, 193(8):2595–2618, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0868-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0868-0.pdf.

Wilkins:2016:WPN

[1259] Shane Maxwell Wilkins. Why paraphrase nihilism fails. Synthese, 193(8): 2619–2632, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0869-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0869-z.pdf.

Beebe:2016:DBP

[1260] James R. Beebe. Do bad people know more? Interactions between attributions of knowledge and blame. Synthese, 193(8):2633–2657, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0872-4;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0872-4.pdf.

Madva:2016:WIA

[1261] Alex Madva. Why implicit attitudes are (probably) not beliefs. Synthese, 193(8):2659–2684, August 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0874-2;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0874-2.pdf.

Feintzeig:2016:UIC

[1262] Benjamin Feintzeig. Unitary inequivalence in classical systems. Synthese, 193(9):2685–2705, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0875-1;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0875-1.pdf.

Hammond:2016:DJA

[1263] Paul Hammond. Distinguishing joint actions from collective actions. Synthese, 193(9):2707–2720, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0876-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0876-
0.pdf.

Boyd:2016:PEE

[1264] Kenneth Boyd. Pragmatic encroachment and epistemically responsible action. Synthese, 193(9):2721–2745, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0878-y;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0878-
y.pdf.

Oesterheld:2016:FPU

[1265] Caspar Oesterheld. Formalizing preference utilitarianism in physical world models. Synthese, 193(9):2747–2759, September 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0883-1;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0883-
1.pdf.

King:2016:SAM

[1266] Martin King. On structural accounts of model-explanations. Synthese, 193(9):2761–2778, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0885-z;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0885-z.pdf.

Roberts:2016:RAP

[1267] Pendaran Roberts and Kelly Ann Schmidtke. Relationalism about perceptible properties and the principle of charity. Synthese, 193(9):2779– 2803, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0886-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0886-y.pdf.

Koscholke:2016:ARO

[1268] Jakob Koscholke and Michael Schippers. Against relative overlap measures of coherence. Synthese, 193(9):2805–2814, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0887-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0887-x.pdf.

Brownstein:2016:DBI

[1269] Michael Brownstein and Eliot Michaelson. Doing without believing: Intellectualism, knowledge-how, and belief-attribution. Synthese, 193(9): 2815–2836, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0888-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0888-9.pdf.

Christias:2016:CSA

[1270] Dionysis Christias. Can Sellars’ argument for scientific realism be used against his own scientia mensura principle? Synthese, 193(9):2837– 2863, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0890-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0890-2.pdf.

Robson:2016:TS

[1271] David Robson. Topological supervenience. Synthese, 193(9):2865– 2897, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0891-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0891-1.pdf.

DelPinal:2016:PCC

[1272] Guillermo Del Pinal. Prototypes as compositional components of concepts. Synthese, 193(9):2899–2927, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0892-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0892-
0.pdf.

Harmon:2016:EAI

[1273] Ian Harmon and Zachary Horne. Evidence for anti-intellectualism about know-how from a sentence recognition task. Synthese, 193(9):2929– 2947, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0894-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0894-y.pdf.

Corns:2016:PES

[1274] Jennifer Corns. Pain eliminativism: scientific and traditional. Synthese, 193(9):2949–2971, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0897-8;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0897-8.pdf.

Barnett:2016:CMP

[1275] Zach Barnett and Han Li. Conciliationism and merely possible disagreement. Synthese, 193(9):2973–2985, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0898-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0898-
7.pdf.

Boyce:2016:DPF

[1276] Kenny Boyce and Andrew Moon. In defense of proper functionalism: cognitive science takes on Swampman. Synthese, 193(9):2987–3001, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0899-6;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0899-6.pdf.

Sagoff:2016:TGC

[1277] Mark Sagoff. Are there general causal forces in ecology? Synthese, 193(9):3003–3024, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0907-x;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0907-x.pdf.

Quinn:2016:PIB

[1278] Aleta Quinn. Phylogenetic inference to the best explanation and the bad lot argument. Synthese, 193(9):3025–3039, September 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0908-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0908-9.pdf.

Simion:2016:AKE

[1279] Mona Simion. Assertion: knowledge is enough. Synthese, 193(10): 3041–3056, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0914-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0914-y.pdf.

Kuipers:2016:MPG

[1280] Theo A. F. Kuipers. Models, postulates, and generalized nomic truth approximation. Synthese, 193(10):3057–3077, October 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0916-9;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0916-
9.pdf.

DiPaolo:2016:IAE

[1281] Joshua DiPaolo and Robert Mark Simpson. Indoctrination anxiety and the etiology of belief. Synthese, 193(10):3079–3098, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0919-6;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0919-6.pdf.

Talbott:2016:NPP

[1282] William J. Talbott. A non-probabilist principle of higher-order reasoning. Synthese, 193(10):3099–3145, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0922-y;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0922-
y.pdf.

Poggiolesi:2016:DNC

[1283] Francesca Poggiolesi. On defining the notion of complete and immediate formal grounding. Synthese, 193(10):3147–3167, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0923-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0923-x.pdf.

Saebo:2016:HQM

[1284] Kjell Johan Sæbø. “How” questions and the manner–method distinction. Synthese, 193(10):3169–3194, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0924-9;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0924-
9.pdf.

Hofer-Szabo:2016:GDB

[1285] Gábor Hofer-Szabó and Péter Vecsernyés. A generalized definition of Bell’s local causality. Synthese, 193(10):3195–3207, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0925-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0925-8.pdf.

Jantzen:2016:DLW

[1286] Benjamin C. Jantzen. Discovery without a ‘logic’ would be a miracle. Synthese, 193(10):3209–3238, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0926-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0926-
7.pdf.

Blutner:2016:QCB

[1287] Reinhard Blutner and Peter beim Graben. Quantum cognition and bounded rationality. Synthese, 193(10):3239–3291, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0928-5;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0928-5.pdf.

Gaertner:2016:ECT

[1288] Wulf Gaertner and Nicolas Wüthrich. Evaluating competing theories via a common language of qualitative verdicts. Synthese, 193(10): 3293–3309, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0929-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0929-4.pdf.

Baetu:2016:IME

[1289] Tudor M. Baetu. From interventions to mechanistic explanations. Synthese, 193(10):3311–3327, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0930-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0930-y.pdf.

Giraud:2016:MM

[1290] Thibaut Giraud. On modal Meinongianism. Synthese, 193(10):3329– 3346, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0933-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0933-8.pdf.

Michael:2016:MTA

[1291] Michaelis Michael. On a “most telling” argument for paraconsistent logic. Synthese, 193(10):3347–3362, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0935-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0935-
6.pdf.

Samoilova:2016:TIU

[1292] Kateryna Samoilova. Transparency and introspective unification. Synthese, 193(10):3363–3381, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0936-5;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0936-5.pdf.

Schippers:2016:CAC

[1293] Michael Schippers. Competing accounts of contrastive coherence. Synthese, 193(10):3383–3395, October 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0937-4;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0937-4.pdf.

Albretsen:2016:P

[1294] Jørgen Albretsen, Per Hasle, and Peter Øhrstrøm. Preface. Synthese, 193(11):3397–3399, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1240-8;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1240-8.pdf.

Hasle:2016:PPS

[1295] Per Hasle and Peter Øhrstrøm. Prior’s paradigm for the study of time and its methodological motivation. Synthese, 193(11):3401–3416, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1161-6;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1161-6.pdf.

Strobach:2016:AYM

[1296] Niko Strobach. An angry young man. Synthese, 193(11):3417–3427, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0901-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0901-3.pdf.

Jakobsen:2016:ICL

[1297] David Jakobsen. An illusion close to life. Synthese, 193(11):3429– 3439, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1076-2;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1076-2.pdf.

Faroldi:2016:ECN

[1298] Federico L. G. Faroldi. Ethical copula, negation, and responsibility judgments. Synthese, 193(11):3441–3448, November 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0911-1;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0911-
1.pdf.

Rini:2016:LLB

[1299] Adriane Rini. The logic of logic and the basis of ethics. Synthese, 193(11): 3449–3457, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0946-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0946-3.pdf.

Markoska-Cubrinovska:2016:PWC

[1300] Aneta Markoska-Cubrinovska. Possible worlds in “the craft of formal logic”. Synthese, 193(11):3459–3471, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0912-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0912-
0.pdf.

Rybarikova:2016:PAL

[1301] Zuzana Rybaríková. Prior on Aristotle’s logical squares. Synthese, 193(11):3473–3482, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0900-4;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0900-4.pdf.

Maruyama:2016:PTN

[1302] Yoshihiro Maruyama. Prior’s tonk, notions of logic, and levels of inconsistency: vindicating the pluralistic unity of science in the light of categorical logical positivism. Synthese, 193(11):3483–3495, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0932-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0932-9.pdf.

Slater:2016:PI

[1303] Hartley Slater. Prior’s individuals. Synthese, 193(11):3497–3506, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0881-3;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0881-3.pdf.

Copeland:2016:PTS

[1304] B. Jack Copeland. Prior, translational semantics, and the Barcan formula. Synthese, 193(11):3507–3519, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0955-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0955-
2.pdf.

Felappi:2016:WFA

[1305] Giulia Felappi. Why fuss about these quirks of the vernacular? Propositional attitude sentences in Prior’s Nachlass. Synthese, 193(11):3521– 3534, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0903-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0903-1.pdf.

VanCleve:2016:OOP

[1306] James Van Cleve. Objectivity without objects: a Priorian program. Synthese, 193(11):3535–3549, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0915-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0915-
x.pdf.

Levine:2016:PBB

[1307] James Levine. Prior, Berkeley, and the Barcan formula. Synthese, 193(11):3551–3565, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0909-8;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0909-8.pdf.

Pleitz:2016:SPP

[1308] Martin Pleitz. Solving Prior’s problem with a Priorean tool. Synthese, 193(11):3567–3577, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0931-x;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0931-x.pdf.

Doyle:2016:SEF

[1309] James Doyle. ‘Spurious egocentricity’ and the first person. Synthese, 193(11):3579–3589, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0948-1;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0948-1.pdf.

LaVine:2016:PTG

[1310] Matt La Vine. Prior’s thank-goodness argument reconsidered. Synthese, 193(11):3591–3606, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0904-0;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0904-0.pdf.

Cresswell:2016:PSM

[1311] M. J. Cresswell. Prior on the semantics of modal and tense logic. Synthese, 193(11):3607–3623, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0949-0;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0949-
0.pdf.

Fernando:2016:PTS

[1312] Tim Fernando. Prior and temporal sequences for natural language. Synthese, 193(11):3625–3637, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0902-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0902-
2.pdf.

Akama:2016:PPA

[1313] Seiki Akama, Tetsuya Murai, and Yasuo Kudo. Partial and paraconsistent approaches to future contingents in tense logic. Synthese, 193(11): 3639–3649, November 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0905-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0905-z.pdf.

Nelson:2016:PPE

[1314] Michael Nelson. Prior and possibly not existing. Synthese, 193 (11):3651–3663, November 2016. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0906-y;   https://link.springer.
com/content/pdf/10.1007/s11229-015-0906-y.pdf.

Blackburn:2016:APN

[1315] Patrick Blackburn and Klaus Frovin Jørgensen. Arthur Prior and ‘now’. Synthese, 193(11):3665–3676, November 2016. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0921-z;
 https://link.springer.com/content/pdf/10.1007/s11229-015-0921-
z.pdf.

Blackburn:2016:RPH

[1316] Patrick Blackburn and Klaus Frovin Jørgensen. Reichenbach, Prior and hybrid tense logic. Synthese, 193(11):3677–3689, November 2016. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-015-0920-0;   https://link.springer.com/content/pdf/
10.1007/s11229-015-0920-0.pdf.

Meyer:2016:TPF

[1317] Ulrich Meyer. Tense, propositions, and facts. Synthese, 193 (11):3691–3699, November 2016. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0910-2;   https://link.springer.
com/content/pdf/10.1007/s11229-015-0910-2.pdf.

Park:2016:WAC

[1318] Woosuk Park. Where have all the Californian tense-logicians gone? Synthese, 193(11):3701–3712, November 2016. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0896-9;
 https://link.springer.com/content/pdf/10.1007/s11229-015-0896-
9.pdf.

Tulenheimo:2016:WTS

[1319] Tero Tulenheimo. Worlds, times and selves revisited. Synthese, 193 (11):3713–3725, November 2016. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0895-x;   https://link.springer.
com/content/pdf/10.1007/s11229-015-0895-x.pdf.

Beni:2016:SRA

[1320] Majid Davoody Beni. Structural realist account of the self. Synthese, 193(12):3727–3740, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1098-9;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1098-9.pdf.

Burnston:2016:CNL

[1321] Daniel C. Burnston. Computational neuroscience and localized neural function. Synthese, 193(12):3741–3762, December 2016. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1099-8;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1099-
8.pdf.

Dacey:2016:RAP

[1322] Mike Dacey. Rethinking associations in psychology. Synthese, 193(12): 3763–3786, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1167-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1167-0.pdf.

Ferretti:2016:PAP

[1323] Gabriele Ferretti. Pictures, action properties and motor related effects. Synthese, 193(12):3787–3817, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1097-x;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1097-
x.pdf.

Hochstetter:2016:ABA

[1324] Gregor Hochstetter. Attention in bodily awareness. Synthese, 193(12): 3819–3842, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1141-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1141-x.pdf.

Levy:2016:UNF

[1325] Arnon Levy. The unity of neuroscience: a flat view. Synthese, 193(12): 3843–3863, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1256-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1256-0.pdf.

Malinowska:2016:CNC

[1326] Joanna K. Malinowska. Cultural neuroscience and the category of race: the case of the other-race effect. Synthese, 193(12):3865–3887, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1108-y;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1108-y.pdf.

Plebe:2016:NPC

[1327] Alessio Plebe and Marco Mazzone. Neural plasticity and concepts ontogeny. Synthese, 193(12):3889–3929, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1131-z;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1131-
z.pdf.

Rusanen:2016:CE

[1328] Anna-Mari Rusanen and Otto Lappi. On computational explanations. Synthese, 193(12):3931–3949, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1101-5;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1101-
5.pdf.

Zednik:2016:BRE

[1329] Carlos Zednik and Frank Jäkel. Bayesian reverse-engineering considered as a research strategy for cognitive science. Synthese, 193(12): 3951–3985, December 2016. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1180-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1180-3.pdf.

Schiemer:2017:CLR

[1330] Georg Schiemer. Carnap on logic and rationality. Synthese, 194(1): 1–14, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1252-4;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1252-4.pdf.

Loeb:2017:RUL

[1331] Iris Loeb. The role of universal language in the early work of Carnap and Tarski. Synthese, 194(1):15–31, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0601-4;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0601-
4.pdf.

Schiemer:2017:CEM

[1332] Georg Schiemer, Richard Zach, and Erich Reck. Carnap’s early metatheory: scope and limits. Synthese, 194(1):33–65, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0877-z;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0877-z.pdf.

Awodey:2017:CIL

[1333] Steve Awodey. Carnap and the invariance of logical truth. Synthese, 194 (1):67–78, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0781-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0781-6.pdf.

Creath:2017:LA

[1334] Richard Creath. The logical and the analytic. Synthese, 194(1): 79–96, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0685-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0685-5.pdf.

Wagner:2017:CTS

[1335] Pierre Wagner. Carnapian and Tarskian semantics. Synthese, 194(1): 97–119, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0853-7;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0853-7.pdf.

Olen:2017:FSR

[1336] Peter Olen. A forgotten strand of reception history: understanding pure semantics. Synthese, 194(1):121–141, January 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0678-4;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0678-

4.pdf. See [1396].

Steinberger:2017:FCN

[1337] Florian Steinberger. Frege and Carnap on the normativity of logic. Synthese, 194(1):143–162, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0880-4;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0880-4.pdf.

Carus:2017:CR

[1338] A. W. Carus. Carnapian rationality. Synthese, 194(1):163–184, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0574-3;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0574-3.pdf.

Carnap:2017:VC

[1339] Rudolf Carnap. Value concepts (1958). Synthese, 194(1):185–194, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0793-2;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0793-2.pdf.

Novaes:2017:CEF

[1340] Catarina Dutilh Novaes and Erich Reck. Carnapian explication, formalisms as cognitive tools, and the paradox of adequate formalization. Synthese, 194(1):195–215, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0816-z;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0816-
z.pdf.

Lutz:2017:CES

[1341] Sebastian Lutz. Carnap on empirical significance. Synthese, 194(1): 217–252, January 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0561-8;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0561-8.pdf.

Dawid:2017:PLD

[1342] Richard Dawid. A philosophical look at the discovery the Higgs boson. Synthese, 194(2):253–257, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1246-2;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1246-2.pdf.

Franklin:2017:MPP

[1343] Allan Franklin. The missing piece of the puzzle: the discovery of the Higgs boson. Synthese, 194(2):259–274, February 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0550-y;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0550-
y.pdf.

Beauchemin:2017:AMA

[1344] Pierre-Hugues Beauchemin. Autopsy of measurements with the AT-LAS detector at the LHC. Synthese, 194(2):275–312, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0944-5;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0944-5.pdf.

Perovic:2017:ERA

[1345] Slobodan Perovic. Experimenter’s regress argument, empiricism, and the calibration of the large hadron collider. Synthese, 194(2):313–332, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0749-6;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0749-6.pdf.

Karaca:2017:CSE

[1346] Koray Karaca. A case study in experimental exploration: exploratory data selection at the large hadron collider. Synthese, 194(2):333– 354, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1206-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1206-x.pdf.

Staley:2017:PWF

[1347] Kent W. Staley. Pragmatic warrant for frequentist statistical practice: the case of high energy physics. Synthese, 194(2):355–376, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1111-3;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1111-3.pdf.

Dawid:2017:BPD

[1348] Richard Dawid. Bayesian perspectives on the discovery of the Higgs particle. Synthese, 194(2):377–394, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0943-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0943-
6.pdf.

Cousins:2017:JLP

[1349] Robert D. Cousins. The Jeffreys–Lindley paradox and discovery criteria in high energy physics. Synthese, 194(2):395–432, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0525-z;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0525-z.pdf. See erratum [2077].

Stoltzner:2017:VEH

[1350] Michael Stöltzner. The variety of explanations in the Higgs sector. Synthese, 194(2):433–460, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1112-2;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1112-2.pdf.

Wuthrich:2017:HDD

[1351] Adrian Wüthrich. The Higgs discovery as a diagnostic causal inference. Synthese, 194(2):461–476, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0941-8;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0941-
8.pdf.

Wells:2017:HNS

[1352] James D. Wells. Higgs naturalness and the scalar boson proliferation instability problem. Synthese, 194(2):477–490, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0618-8;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0618-8.pdf.

McCoy:2017:PGR

[1353] C. D. McCoy. Prediction in general relativity. Synthese, 194(2):491– 509, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0954-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0954-3.pdf.

Levy:2017:ESF

[1354] Neil Levy. Embodied savoir-faire: knowledge-how requires motor representations. Synthese, 194(2):511–530, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0956-1;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0956-
1.pdf.

Dodd:2017:HKB

[1355] Jordan Dodd. Hope, knowledge, and blindspots. Synthese, 194(2): 531–543, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0957-0;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0957-0.pdf.

Laraudogoitia:2017:SIC

[1356] Jon Pérez Laraudogoitia. A simple and interesting classical mechanical supertask. Synthese, 194(2):545–570, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0958-z;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0958-
z.pdf.

Hamlin:2017:TTU

[1357] Colin Hamlin. Towards a theory of universes: structure theory and the mathematical universe hypothesis. Synthese, 194(2):571–591, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0959-y;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0959-y.pdf.

Young:2017:KHA

[1358] Garry Young. Knowledge how, ability, and the type-token distinction. Synthese, 194(2):593–607, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0961-4;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0961-
4.pdf.

Baron:2017:FF

[1359] Sam Baron. Feel the flow. Synthese, 194(2):609–630, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0964-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0964-1.pdf.

Locke:2017:INL

[1360] Dustin Locke. Implicature and non-local pragmatic encroachment. Synthese, 194(2):631–654, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0965-0;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0965-0.pdf.

Gangopadhyay:2017:FSC

[1361] Nivedita Gangopadhyay. The future of social cognition: paradigms, concepts and experiments. Synthese, 194(3):655–672, March 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1162-5;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1162-
5.pdf.

Carruthers:2017:MAE

[1362] Peter Carruthers. Mindreading in adults: evaluating two-systems views. Synthese, 194(3):673–688, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0792-3;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0792-3.pdf.

Rakoczy:2017:DDD

[1363] Hannes Rakoczy. In defense of a developmental dogma: children acquire propositional attitude folk psychology around age 4. Synthese, 194 (3):689–707, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0860-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0860-8.pdf.

Smith:2017:WE

[1364] Joel Smith. What is empathy for? Synthese, 194(3):709–722, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0771-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0771-8.pdf.

Cassam:2017:WAK

[1365] Quassim Cassam. What asymmetry? Knowledge of self, knowledge of others, and the inferentialist challenge. Synthese, 194(3):723–741, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0772-7;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0772-7.pdf.

Overgaard:2017:UT

[1366] Søren Overgaard. The unobservability thesis. Synthese, 194(3):743– 760, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0804-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0804-3.pdf.

Newen:2017:DLC

[1367] Albert Newen. Defending the liberal-content view of perceptual experience: direct social perception of emotions and person impressions. Synthese, 194(3):761–785, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1030-3;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1030-3.pdf.

Varga:2017:CMP

[1368] Somogy Varga. The case for mind perception. Synthese, 194(3): 787–807, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0994-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0994-8.pdf.

Nichols:2017:EMA

[1369] Shaun Nichols. The essence of mentalistic agents. Synthese, 194(3): 809–825, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0893-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0893-z.pdf.

Hutto:2017:BSC

[1370] Daniel D. Hutto. Basic social cognition without mindreading: minding minds without attributing contents. Synthese, 194(3):827–846, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0831-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0831-0.pdf.

Karakostas:2017:CSQ

[1371] Vassilios Karakostas and Elias Zafiris. Contextual semantics in quantum mechanics from a categorical point of view. Synthese, 194(3): 847–886, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0970-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0970-3.pdf.

Fultot:2017:MAI

[1372] Martin Flament Fultot. Modulation: an alternative to instructions and forces. Synthese, 194(3):887–916, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0976-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0976-
x.pdf.

Skovgaard-Olsen:2017:PLO

[1373] Niels Skovgaard-Olsen. The problem of logical omniscience, the preface paradox, and doxastic commitments. Synthese, 194(3):917–939, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0979-7;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0979-7.pdf.

Pitcovski:2017:GBP

[1374] Eli Pitcovski. Getting the big picture. Synthese, 194(3):941–962, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0981-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0981-0.pdf.

Huang:2017:PRC

[1375] Minyao Huang. A plea for radical contextualism. Synthese, 194(3): 963–988, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0982-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0982-z.pdf.

Willard-Kyle:2017:DGM

[1376] Christopher Willard-Kyle. Do great minds really think alike? Synthese, 194(3):989–1026, March 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0984-x;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0984-x.pdf.

Holm:2017:TO

[1377] Sune Holm and John Basl. Teleological organisation. Synthese, 194 (4):1027–1029, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1340-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1340-0.pdf.

Odenbaugh:2017:NEM

[1378] Jay Odenbaugh. Nothing in ethics makes sense except in the light of evolution? Natural goodness, normativity, and naturalism. Synthese, 194 (4):1031–1055, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0675-7;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0675-7.pdf.

Basl:2017:TTI

[1379] John Basl. A trilemma for teleological individualism. Synthese, 194 (4):1057–1074, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1316-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1316-0.pdf.

Holm:2017:TB

[1380] Sune Holm. Teleology and biocentrism. Synthese, 194(4):1075–1087, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1300-0;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1300-0.pdf.

Mossio:2017:WMB

[1381] Matteo Mossio and Leonardo Bich. What makes biological organisation teleological? Synthese, 194(4):1089–1114, April 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0594-z;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0594-
z.pdf.

Dussault:2017:PEP

[1382] Antoine C. Dussault and Frédéric Bouchard. A persistence enhancing propensity account of ecological function to explain ecosystem evolution. Synthese, 194(4):1115–1145, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1065-5;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1065-
5.pdf.

Neander:2017:FAS

[1383] Karen Neander. Functional analysis and the species design. Synthese, 194 (4):1147–1168, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0940-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0940-9.pdf.

Leahy:2017:TAE

[1384] Brian Leahy and Maximilian Huber. Two arguments for the etiological theory over the modal theory of biological function. Synthese, 194 (4):1169–1187, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0544-9;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0544-9.pdf.

Horne:2017:OEU

[1385] Zachary Horne and Jonathan Livengood. Ordering effects, updating effects, and the specter of global skepticism. Synthese, 194(4):1189–1218, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0985-9;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0985-9.pdf.

Stein:2017:TTM

[1386] Jordan Stein. Tharp’s theorems of metaphysics and the notion of necessary truth. Synthese, 194(4):1219–1231, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0987-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0987-
7.pdf.

Kelp:2017:LJ

[1387] Christoph Kelp. Lotteries and justification. Synthese, 194(4):1233–1244, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0989-5;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0989-5.pdf.

Cabrera:2017:CTB

[1388] Frank Cabrera. Can there be a Bayesian explanationism? On the prospects of a productive partnership. Synthese, 194(4):1245–1272, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0990-z;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0990-z.pdf.

Wilkenfeld:2017:MU

[1389] Daniel A. Wilkenfeld. MUDdy understanding. Synthese, 194(4): 1273–1293, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0992-x;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0992-x.pdf.

Henderson:2017:NMA

[1390] Leah Henderson. The no miracles argument and the base rate fallacy. Synthese, 194(4):1295–1302, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0995-7;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0995-7.pdf.

Koscholke:2017:PCM

[1391] Jakob Koscholke and Marc Jekel. Probabilistic coherence measures: a psychological study of coherence assessment. Synthese, 194(4): 1303–1322, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0996-6;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0996-6.pdf.

Arenhart:2017:RVQ

[1392] Jonas Rafael Becker Arenhart. The received view on quantum nonindividuality: formal and metaphysical analysis. Synthese, 194(4): 1323–1347, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0997-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0997-5.pdf.

Skiba:2017:FIP

[1393] Lukas Skiba. Fictionalism and the incompleteness problem. Synthese, 194 (4):1349–1362, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-1000-1;   http://link.springer.com/content/
pdf/10.1007/s11229-015-1000-1.pdf.

Howick:2017:RPD

[1394] Jeremy Howick. The relativity of ‘placebos’: defending a modified version of Grünbaum’s definition. Synthese, 194(4):1363–1396, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-1001-0;   http://link.springer.com/content/pdf/10.1007/
s11229-015-1001-0.pdf.

Raleigh:2017:AID

[1395] Thomas Raleigh. Against an inferentialist dogma. Synthese, 194(4): 1397–1421, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-1002-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-1002-z.pdf.

Olen:2017:EFS

[1396] Peter Olen. Erratum to: A forgotten strand of reception history: understanding pure semantics. Synthese, 194(4):1423, April 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0725-1;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0725-1.pdf. See [1336].

Paseau:2017:EMI

[1397] A. C. Paseau. Erratum to: A measure of inferential-role preservation. Synthese, 194(4):1425, April 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0738-9;   https://link.springer.
com/content/pdf/10.1007/s11229-015-0738-9.pdf. See [995].

Dutant:2017:I

[1398] Julien Dutant, Davide Fassio, and Anne Meylan. Introduction. Synthese, 194(5):1427–1431, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1352-9;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1352-9.pdf.

Olsson:2017:EVR

[1399] Erik J. Olsson. Engel vs. Rorty on truth. Synthese, 194(5):1433–1450, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0968-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0968-x.pdf.

McHugh:2017:EDC

[1400] Conor McHugh. Engel on doxastic correctness. Synthese, 194(5): 1451–1462, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0767-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0767-4.pdf.

Carter:2017:AUE

[1401] J. Adam Carter. Assertion, uniqueness and epistemic hypocrisy. Synthese, 194(5):1463–1476, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0766-5;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0766-5.pdf.

Pritchard:2017:EPE

[1402] Duncan Pritchard. Engel on pragmatic encroachment and epistemic value. Synthese, 194(5):1477–1486, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0755-8;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0755-
8.pdf.

Kelp:2017:CK

[1403] Christoph Kelp and Mona Simion. Commodious knowledge. Synthese, 194(5):1487–1502, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0938-3;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0938-3.pdf.

Keith:2017:CSS

[1404] Lehrer Keith. Common sense and skepticism: a lecture. Synthese, 194 (5):1503–1516, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0770-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0770-9.pdf.

Egre:2017:KRT

[1405] Paul Egré. Knowledge as de re true belief? Synthese, 194(5):1517–1529, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1115-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1115-z.pdf.

Dokic:2017:KPA

[1406] Jérôme Dokic. Knowledge, perception, and the art of camouflage. Synthese, 194(5):1531–1539, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0758-5;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0758-5.pdf.

Douven:2017:HAO

[1407] Igor Douven. How to account for the oddness of missing-link conditionals. Synthese, 194(5):1541–1554, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0756-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0756-
7.pdf.

Engel:2017:RNB

[1408] Pascal Engel. Replies: On norms of belief and knowledge. Synthese, 194 (5):1555–1564, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1358-3;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1358-3.pdf.

Huber:2017:WFR

[1409] Franz Huber. Why follow the royal rule? Synthese, 194(5):1565–1590, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-1004-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-1004-x.pdf.

Newhard:2017:PTI

[1410] Jay Newhard. Plain truth and the incoherence of alethic functionalism. Synthese, 194(5):1591–1611, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-1006-8;   http://link.springer.com/
content/pdf/10.1007/s11229-015-1006-8.pdf.

Kwong:2017:OMC

[1411] Jack M. C. Kwong. Is open-mindedness conducive to truth? Synthese, 194(5):1613–1626, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-1008-6;   http://link.springer.com/
content/pdf/10.1007/s11229-015-1008-6.pdf.

Dumsday:2017:MES

[1412] Travis Dumsday. MaxCon extended simples and the dispositionalist ontology of laws. Synthese, 194(5):1627–1641, May 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-1009-5;
 http://link.springer.com/content/pdf/10.1007/s11229-015-1009-
5.pdf.

Warren:2017:EVN

[1413] Jared Warren. Epistemology versus non-causal realism. Synthese, 194 (5):1643–1662, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-1010-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-1010-z.pdf.

Machuca:2017:NPR

[1414] Diego E. Machuca. A neo-Pyrrhonian response to the disagreeing about disagreement argument. Synthese, 194(5):1663–1680, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-1012-x;   http://link.springer.com/content/pdf/10.1007/
s11229-015-1012-x.pdf.

Starzak:2017:IJG

[1415] Tobias Starzak. Interpretations without justification: a general argument against Morgan’s canon. Synthese, 194(5):1681–1701, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1013-4;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1013-4.pdf.

Raz:2017:SHE

[1416] Tim Räz. The silent hexagon: explaining comb structures. Synthese, 194 (5):1703–1724, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1014-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1014-3.pdf.

Roberts:2017:ERC

[1417] Pendaran Roberts. An ecumenical response to color contrast cases. Synthese, 194(5):1725–1742, May 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1016-1;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1016-1.pdf.

Fang:2017:HMO

[1418] Wei Fang. Holistic modeling: an objection to Weisberg’s weighted feature-matching account. Synthese, 194(5):1743–1764, May 2017. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1018-z;
 https://link.springer.com/content/pdf/10.1007/s11229-016-1018-
z.pdf.

Alford-Duguid:2017:EPH

[1419] Dominic Alford-Duguid and Michael Arsenault. On the explanatory power of hallucination. Synthese, 194(5):1765–1785, May 2017. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1020-5;
 https://link.springer.com/content/pdf/10.1007/s11229-016-1020-
5.pdf.

Paternotte:2017:VVS

[1420] Cedric Paternotte and Milena Ivanova. Virtues and vices in scientific practice. Synthese, 194(5):1787–1807, May 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1023-2;
 https://link.springer.com/content/pdf/10.1007/s11229-016-1023-
2.pdf.

Pezet:2017:FP

[1421] Robert E. Pezet. A foundation for presentism. Synthese, 194(5):1809– 1837, May 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-016-1024-1;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1024-1.pdf.

Manzano:2017:NVH

[1422] María Manzano and Enrique Alonso. A note on Visions of Henkin. Synthese, 194(6):1839–1840, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1022-3;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1022-3.pdf. See [1051].

Clapp:2017:DP

[1423] Lenny Clapp. On denying presuppositions. Synthese, 194(6):1841–1900, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1026-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1026-z.pdf.

Gao:2017:RAK

[1424] Jie Gao. Rational action without knowledge (and vice versa). Synthese, 194(6):1901–1917, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1027-y;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1027-y.pdf.

McKeon:2017:SIB

[1425] Matthew W. McKeon. Statements of inference and begging the question. Synthese, 194(6):1919–1943, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1028-x;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1028-x.pdf.

Loew:2017:CPF

[1426] Christian Loew. Causation, physics, and fit. Synthese, 194(6):1945–1965, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

016-1029-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1029-9.pdf.

Pence:2017:GDF

[1427] Charles H. Pence. Is genetic drift a force? Synthese, 194(6):1967–1988, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1031-2;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1031-2.pdf.

Hennig:2017:MP

[1428] Boris Hennig. The man without properties. Synthese, 194(6):1989–2006, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1033-0;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1033-0.pdf.

Valaris:2017:WRM

[1429] Markos Valaris. What reasoning might be. Synthese, 194(6):2007–2024, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1034-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1034-z.pdf.

Thorn:2017:PMS

[1430] Paul D. Thorn. On the preference for more specific reference classes. Synthese, 194(6):2025–2051, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1035-y;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1035-y.pdf.

Tugby:2017:PR

[1431] Matthew Tugby. The problem of retention. Synthese, 194(6):2053–2075, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1036-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1036-x.pdf.

Greenwood:2017:SSB

[1432] John D. Greenwood. Solitary social belief. Synthese, 194(6):2077–2099, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

016-1037-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1037-9.pdf.

Lennertz:2017:PCN

[1433] Benjamin Lennertz. Probabilistic consistency norms and quantificational credences. Synthese, 194(6):2101–2119, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1039-7;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1039-
7.pdf.

Yuan:2017:RMR

[1434] Yongfeng Yuan. Rational metabolic revision based on core beliefs. Synthese, 194(6):2121–2146, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1042-z;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1042-z.pdf.

Broncano-Berrocal:2017:REV

[1435] Fernando Broncano-Berrocal. A robust enough virtue epistemology. Synthese, 194(6):2147–2174, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1043-y;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1043-y.pdf.

Carter:2017:MAL

[1436] J. Adam Carter and Martin Peterson. The modal account of luck revisited. Synthese, 194(6):2175–2184, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1047-7;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1047-
7.pdf.

Rohrs:2017:SPC

[1437] Benjamin Rohrs. Supervaluational propositional content. Synthese, 194 (6):2185–2201, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1051-y;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1051-y.pdf.

Lauwers:2017:ILL

[1438] Luc Lauwers. Infinite lotteries, large and small sets. Synthese, 194 (6):2203–2209, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1052-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1052-x.pdf.

Clarke:2017:HDL

[1439] Christopher Clarke. How to define levels of explanation and evaluate their indispensability. Synthese, 194(6):2211–2231, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1053-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1053-9.pdf.

Buijsman:2017:ARM

[1440] Stefan Buijsman. Accessibility of reformulated mathematical content. Synthese, 194(6):2233–2250, June 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1054-8;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1054-8.pdf.

Huttemann:2017:SCE

[1441] Andreas Hüttemann. Structuralism, causation and explanation. Synthese, 194(7):2251–2253, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1465-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1465-1.pdf.

Saatsi:2017:SC

[1442] Juha Saatsi. Structuralism with and without causation. Synthese, 194 (7):2255–2271, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0595-y;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0595-y.pdf.

Chakravartty:2017:PCM

[1443] Anjan Chakravartty. Particles, causation, and the metaphysics of structure. Synthese, 194(7):2273–2289, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0913-z;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0913-
z.pdf.

Reutlinger:2017:CFR

[1444] Alexander Reutlinger. Are causal facts really explanatorily emergent? Ladyman and Ross on higher-level causal facts and renormalization group explanation. Synthese, 194(7):2291–2305, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0530-2;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0530-2.pdf.

Dorato:2017:DVS

[1445] Mauro Dorato. Dynamical versus structural explanations in scientific revolutions. Synthese, 194(7):2307–2327, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0546-7;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0546-
7.pdf.

Esfeld:2017:HAQ

[1446] Michael Esfeld. How to account for quantum non-locality: ontic structural realism and the primitive ontology of quantum physics. Synthese, 194(7):2329–2344, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0549-4;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0549-4.pdf.

Trogdon:2017:RP

[1447] Kelly Trogdon. Revelation and physicalism. Synthese, 194(7):2345–2366, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1055-7;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1055-7.pdf.

Smid:2017:IMT

[1448] Jeroen Smid. ‘Identity’ as a mereological term. Synthese, 194(7): 2367–2385, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1056-6;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1056-6.pdf.

Oderberg:2017:FRP

[1449] David S. Oderberg. Finality revived: powers and intentionality. Synthese, 194(7):2387–2425, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1057-5;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1057-5.pdf.

Dellsen:2017:RAR

[1450] Finnur Dellsén. Realism and the absence of rivals. Synthese, 194(7): 2427–2446, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1059-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1059-3.pdf.

Andersen:2017:BEC

[1451] Mikkel Birkegaard Andersen, Thomas Bolander, Hans van Ditmarsch, and Martin Holm Jensen. Bisimulation and expressivity for conditional belief, degrees of belief, and safe belief. Synthese, 194(7):2447–2487, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1060-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1060-x.pdf.

Zhang:2017:PDP

[1452] Xiaoxing Zhang. The paradox of the diffusiveness of power. Synthese, 194 (7):2489–2500, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1061-9;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1061-9.pdf.

Sullivan-Bissett:2017:MD

[1453] Ema Sullivan-Bissett. Malfunction defended. Synthese, 194(7):2501– 2522, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1062-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1062-8.pdf.

Chierchia:2017:CBM

[1454] Gennaro Chierchia. Clouds and blood. More on vagueness and the mass/count distinction. Synthese, 194(7):2523–2538, July 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1063-7;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1063-
7.pdf.

Austin:2017:AEE

[1455] Christopher J. Austin. Aristotelian essentialism: essence in the age of evolution. Synthese, 194(7):2539–2556, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1066-4;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1066-
4.pdf.

Morganti:2017:MNM

[1456] Matteo Morganti and Tuomas E. Tahko. Moderately naturalistic metaphysics. Synthese, 194(7):2557–2580, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1068-2;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1068-
2.pdf.

Ivanova:2017:PAS

[1457] Milena Ivanova. Poincaré’s aesthetics of science. Synthese, 194(7): 2581–2594, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1069-1;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1069-1.pdf.

Gyenis:2017:CUC

[1458] Z. Gyenis, G. Hofer-Szabó, and M. Rédei. Conditioning using conditional expectations: the Borel–Kolmogorov paradox. Synthese, 194(7): 2595–2630, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1070-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1070-8.pdf.

Andow:2017:DNP

[1459] James Andow. Do non-philosophers think epistemic consequentialism is counterintuitive? Synthese, 194(7):2631–2643, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1071-7;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1071-7.pdf.

Merlo:2017:MRV

[1460] Giovanni Merlo. Multiple reference and vague objects. Synthese, 194 (7):2645–2666, July 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1075-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1075-3.pdf.

Kruse:2017:ISI

[1461] Andrea Kruse and Heinrich Wansing. Introduction to the special issue “doxastic agency and epistemic responsibility”. Synthese, 194(8): 2667–2671, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1153-6;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1153-6.pdf.

Steup:2017:BI

[1462] Matthias Steup. Believing intentionally. Synthese, 194(8):2673–2694, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0780-7;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0780-7.pdf.

Rott:2017:NDV

[1463] Hans Rott. Negative doxastic voluntarism and the concept of belief. Synthese, 194(8):2695–2720, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1032-1;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1032-1.pdf.

Nottelmann:2017:ADV

[1464] Nikolaj Nottelmann. Against a descriptive vindication of doxastic voluntarism. Synthese, 194(8):2721–2744, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0768-3;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0768-
3.pdf.

McHugh:2017:AC

[1465] Conor McHugh. Attitudinal control. Synthese, 194(8):2745–2762, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0643-7;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0643-7.pdf.

Flowerree:2017:ABI

[1466] A. K. Flowerree. Agency of belief and intention. Synthese, 194(8): 2763–2784, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1138-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1138-5.pdf.

Wagner:2017:AFW

[1467] Verena Wagner. On the analogy of free will and free belief. Synthese, 194 (8):2785–2810, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0851-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0851-9.pdf.

Kruse:2017:WDR

[1468] Andrea Kruse. Why doxastic responsibility is not based on direct doxastic control. Synthese, 194(8):2811–2842, August 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0951-6;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0951-
6.pdf.

Wansing:2017:RLI

[1469] Heinrich Wansing. Remarks on the logic of imagination. A step towards understanding doxastic control through imagination. Synthese, 194(8): 2843–2861, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0945-4;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0945-4.pdf.

Goldberg:2017:SK

[1470] Sanford C. Goldberg. Should have known. Synthese, 194(8):2863– 2894, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0662-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0662-z.pdf.

Peels:2017:RBE

[1471] Rik Peels. Responsible belief and epistemic justification. Synthese, 194 (8):2895–2915, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1038-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1038-8.pdf.

Kramer:2017:HCI

[1472] Stephan Krämer. A hyperintensional criterion of irrelevance. Synthese, 194(8):2917–2930, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1078-0;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1078-0.pdf.

McDonnell:2017:WPP

[1473] Jane McDonnell. Wigner’s puzzle and the Pythagorean heuristic. Synthese, 194(8):2931–2948, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1080-6;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1080-6.pdf.

Phillips:2017:IVO

[1474] Ben Phillips. Inscrutability and visual objects. Synthese, 194(8): 2949–2971, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1083-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1083-3.pdf.

Fleisher:2017:VD

[1475] Will Fleisher. Virtuous distinctions. Synthese, 194(8):2973–3003, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1084-2;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1084-2.pdf.

Gaukroger:2017:WBC

[1476] Cressida Gaukroger. Why broad content can’t influence behaviour. Synthese, 194(8):3005–3020, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1085-1;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1085-1.pdf.

Atkins:2017:RAS

[1477] Philip Atkins. A Russellian account of suspended judgment. Synthese, 194(8):3021–3046, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1089-x;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1089-x.pdf.

Warmbrod:2017:TCT

[1478] Ken Warmbrod. Time, change and time without change. Synthese, 194 (8):3047–3067, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1090-4;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1090-4.pdf.

Appley:2017:TNO

[1479] Bryan C. Appley and Gregory Stoutenburg. Two new objections to explanationism. Synthese, 194(8):3069–3084, August 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1093-1;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1093-
1.pdf.

Baker:2017:RPE

[1480] Jordan Baker. Rejecting Pereboom’s empirical objection to agentcausation. Synthese, 194(8):3085–3100, August 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1094-0;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1094-
0.pdf.

Kim:2017:HES

[1481] Brian Kim and Anubav Vasudevan. How to expect a surprising exam. Synthese, 194(8):3101–3133, August 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1096-y;
 https://link.springer.com/content/pdf/10.1007/s11229-016-1096-
y.pdf.

Heersmink:2017:DSP

[1482] Richard Heersmink. Distributed selves: personal identity and extended memory systems. Synthese, 194(8):3135–3151, August 2017. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1102-4;
 https://link.springer.com/content/pdf/10.1007/s11229-016-1102-
4.pdf.

Hughes:2017:EIC

[1483] Michael Hughes. Epistemic inconsistency and categorical coherence: a study of probabilistic measures of coherence. Synthese, 194(8):3153– 3185, August 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-016-1105-1;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1105-1.pdf.

Psillos:2017:SRQ

[1484] Stathis Psillos and Emma Ruttkamp-Bloem. Scientific realism: quo vadis? Introduction: new thinking about scientific realism. Synthese, 194(9):3187–3201, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1493-x;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1493-x.pdf.

Lyons:2017:ESH

[1485] Timothy D. Lyons. Epistemic selectivity, historical threats, and the non-epistemic tenets of scientific realism. Synthese, 194(9):3203–3219, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1103-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1103-3.pdf.

Vickers:2017:USR

[1486] Peter Vickers. Understanding the selective realist defence against the PMI. Synthese, 194(9):3221–3232, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1082-4;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1082-
4.pdf.

Saatsi:2017:RRR

[1487] Juha Saatsi. Replacing recipe realism. Synthese, 194(9):3233–3244, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0962-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0962-3.pdf.

Leconte:2017:PSP

[1488] Gauvain Leconte. Predictive success, partial truth and Duhemian realism. Synthese, 194(9):3245–3265, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1305-8;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1305-
8.pdf.

Alai:2017:RHO

[1489] Mario Alai. Resisting the historical objections to realism: Is Doppelt’s a viable solution? Synthese, 194(9):3267–3290, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1087-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1087-z.pdf.

Niiniluoto:2017:ORA

[1490] Ilkka Niiniluoto. Optimistic realism about scientific progress. Synthese, 194(9):3291–3309, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0974-z;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0974-z.pdf.

French:2017:SRR

[1491] Steven French. (structural) realism and its representational vehicles. Synthese, 194(9):3311–3326, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0879-x;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0879-
x.pdf.

Forbes:2017:PEA

[1492] Curtis Forbes. A pragmatic, existentialist approach to the scientific realism debate. Synthese, 194(9):3327–3346, September 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1015-2;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1015-
2.pdf.

Spurrett:2017:PEH

[1493] David Spurrett. Physicalism as an empirical hypothesis. Synthese, 194(9):3347–3360, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0986-8;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0986-8.pdf.

Berenstain:2017:AMP

[1494] Nora Berenstain. The applicability of mathematics to physical modality. Synthese, 194(9):3361–3377, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1067-3;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1067-
3.pdf.

Chakravartty:2017:RNT

[1495] Anjan Chakravartty. Reflections on new thinking about scientific realism. Synthese, 194(9):3379–3392, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1514-9;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1514-
9.pdf.

Baccelli:2017:DBR

[1496] Jean Baccelli. Do bets reveal beliefs? Synthese, 194(9):3393–3419, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0939-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0939-2.pdf.

Peet:2017:EIU

[1497] Andrew Peet. Epistemic injustice in utterance interpretation. Synthese, 194(9):3421–3443, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0942-7;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0942-7.pdf.

Dawid:2017:IR

[1498] Philip Dawid. On individual risk. Synthese, 194(9):3445–3474, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0953-4;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0953-4.pdf.

Heilmann:2017:HF

[1499] Conrad Heilmann and Stefan Wintein. How to be fairer. Synthese, 194(9):3475–3499, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0967-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0967-y.pdf.

Weber:2017:UAR

[1500] Erik Weber and Merel Lefevere. Unification, the answer to resemblance questions. Synthese, 194(9):3501–3521, September 2017. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0969-9;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0969-
9.pdf.

Miyazono:2017:DFE

[1501] Kengo Miyazono. Does functionalism entail extended mind? Synthese, 194(9):3523–3541, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0971-2;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0971-2.pdf.

Fernandez:2017:SMP

[1502] Anthony Vincent Fernandez. The subject matter of phenomenological research: existentials, modes, and prejudices. Synthese, 194(9):3543– 3562, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1106-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1106-0.pdf.

Barrett:2017:QCM

[1503] Thomas William Barrett and Hans Halvorson. Quine’s conjecture on many-sorted logic. Synthese, 194(9):3563–3582, September 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1107-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1107-z.pdf.

Tsementzis:2017:UFS

[1504] Dimitris Tsementzis. Univalent foundations as structuralist foundations. Synthese, 194(9):3583–3617, September 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1109-x;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1109-
x.pdf;   https://link.springer.com/content/pdf/10.1007/s11229-
016-1109-x.pdf.

Glazer:2017:LAS

[1505] Trip Glazer. Looking angry and sounding sad: The perceptual analysis of emotional expression. Synthese, 194(9):3619–3643, September 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/

s11229-016-1113-1;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1113-1.pdf.

Burnston:2017:CPC

[1506] Daniel C. Burnston. Cognitive penetration and the cognition–perception interface. Synthese, 194(9):3645–3668, September 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1116-y;
 https://link.springer.com/content/pdf/10.1007/s11229-016-1116-

y.pdf. See correction [1989].

Munton:2017:FFF

[1507] Jessie Munton. Frege, fiction and force. Synthese, 194(9):3669– 3692, September 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1117-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1117-x.pdf;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1117-x.pdf.

Dietz:2017:VPI

[1508] Richard Dietz. Vagueness and probability: introduction. Synthese, 194(10):3693–3698, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1347-6;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1347-6.pdf.

Fine:2017:PV

[1509] Kit Fine. The possibility of vagueness. Synthese, 194(10):3699–3725, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0625-9.

Pagin:2017:THO

[1510] Peter Pagin. Tolerance and higher-order vagueness. Synthese, 194(10): 3727–3760, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0798-x.

Smith:2017:UAT

[1511] Nicholas J. J. Smith. Undead argument: the truth-functionality objection to fuzzy theories of vagueness. Synthese, 194(10):3761–3787, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0651-7.

Keefe:2017:DBE

[1512] Rosanna Keefe. Degrees of belief, expected and actual. Synthese, 194(10): 3789–3800, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1049-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1049-5.pdf.

Lassiter:2017:AVB

[1513] Daniel Lassiter and Noah D. Goodman. Adjectival vagueness in a Bayesian model of interpretation. Synthese, 194(10):3801–3836, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0786-1.

Egre:2017:VJP

[1514] Paul Égré. Vague judgment: a probabilistic account. Synthese, 194(10): 3837–3865, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1092-2.

Douven:2017:VST

[1515] Igor Douven. Verities, the sorites, and Theseus’ ship. Synthese, 194(10): 3867–3878, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0724-2.

Ichinose:2017:NPM

[1516] Masaki Ichinose. Normativity, probability, and meta-vagueness. Synthese, 194(10):3879–3900, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0950-7.

Rabinowicz:2017:VP

[1517] Wlodek Rabinowicz. From values to probabilities. Synthese, 194(10): 3901–3929, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0693-5.

Lyon:2017:VC

[1518] Aidan Lyon. Vague credence. Synthese, 194(10):3931–3954, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0782-5.

Cresto:2017:LTU

[1519] Eleonora Cresto. Lost in translation: unknowable propositions in probabilistic frameworks. Synthese, 194(10):3955–3977, October 2017. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0884-0.

Fitelson:2017:DI

[1520] Branden Fitelson and Alan Hájek. Declarations of independence. Synthese, 194(10):3979–3995, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0559-2.

Smith:2017:IPB

[1521] Martin Smith. Intuitionistc probability and the Bayesian objection to dogmatism. Synthese, 194(10):3997–4009, October 2017. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1120-2;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1120-
2.pdf.

Hernandez-Conde:2017:CAC

[1522] José V. Hernández-Conde. A case against convexity in conceptual spaces. Synthese, 194(10):4011–4037, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1123-z.

Prettyman:2017:PCI

[1523] Adrienne Prettyman. Perceptual content is indexed to attention. Synthese, 194(10):4039–4054, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1125-x.

Mundici:2017:CFC

[1524] Daniele Mundici. Coherence of de Finetti coherence. Synthese, 194(10): 4055–4063, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1126-9.

Boult:2017:ENJ

[1525] Cameron Boult. Epistemic normativity and the justification-excuse distinction. Synthese, 194(10):4065–4081, October 2017. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-016-1127-8.

Martini:2017:HVT

[1526] Francesco Martini. Hearsay viewed through the lens of trust, reputation and coherence. Synthese, 194(10):4083–4099, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/s11229-016-1128-7.

Desmond:2017:SBE

[1527] Hugh Desmond. Symmetry breaking and the emergence of pathdependence. Synthese, 194(10):4101–4131, October 2017. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-016-1130-0.

Stern:2017:IDT

[1528] Reuben Stern. Interventionist decision theory. Synthese, 194(10): 4133–4153, October 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1133-x.

Koons:2017:MRC

[1529] Jeremy Randel Koons. A myth resurgent: classical foundationalism and the new Sellarsian critique. Synthese, 194(10):4155–4169, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1134-9.

Cowling:2017:CAH

[1530] Sam Cowling. Conceivability arguments for haecceitism. Synthese, 194(10):4171–4190, October 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1136-7.

Buckner:2017:WCA

[1531] Cameron Buckner and Ellen Fridland. What is cognition? Angsty monism, permissive pluralism(s), and the future of cognitive science. Synthese, 194(11):4191–4195, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1505-x;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1505-
x.pdf.

Ramsey:2017:MCR

[1532] William Ramsey. Must cognition be representational? Synthese, 194(11): 4197–4214, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0644-6.

Rowlands:2017:AAR

[1533] Mark Rowlands. Arguing about representation. Synthese, 194(11): 4215–4232, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0646-4.

Allen:2017:DC

[1534] Colin Allen. On (not) defining cognition. Synthese, 194(11):4233– 4249, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1454-4.

Newen:2017:WCP

[1535] Albert Newen. What are cognitive processes? An example-based approach. Synthese, 194(11):4251–4268, November 2017. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-015-0812-3.

Aizawa:2017:CB

[1536] Ken Aizawa. Cognition and behavior. Synthese, 194(11):4269– 4288, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0645-5.

Figdor:2017:PDP

[1537] Carrie Figdor. On the proper domain of psychological predicates. Synthese, 194(11):4289–4310, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-014-0603-2.

Greif:2017:WEE

[1538] Hajo Greif. What is the extension of the extended mind? Synthese, 194(11):4311–4336, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0799-9;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0799-9.pdf.

Fridland:2017:AM

[1539] Ellen Fridland. Automatically minded. Synthese, 194(11):4337– 4363, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0617-9.

Michael:2017:PUW

[1540] John Michael. Putting unicepts to work: a teleosemantic perspective on the infant mindreading puzzle. Synthese, 194(11):4365–4388, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0850-x.

Pittard:2017:DRR

[1541] John Pittard. Disagreement, reliability, and resilience. Synthese, 194(11): 4389–4409, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1137-6.

Lahtinen:2017:TUF

[1542] Valtteri Lahtinen and Antti Stenvall. Towards a unified framework for decomposability of processes. Synthese, 194(11):4411–4427, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1139-4.

Sudbery:2017:LFQ

[1543] Anthony Sudbery. The logic of the future in quantum theory. Synthese, 194(11):4429–4453, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1142-9;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1142-9.pdf.

Hufendiek:2017:ANE

[1544] Rebekka Hufendiek. Affordances and the normativity of emotions. Synthese, 194(11):4455–4476, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1144-7.

Wieland:2017:RSI

[1545] Jan Willem Wieland. Responsibility for strategic ignorance. Synthese, 194(11):4477–4497, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1145-6;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1145-6.pdf.

Heesen:2017:ASC

[1546] Remco Heesen. Academic superstars: competent or lucky? Synthese, 194(11):4499–4518, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1146-5;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1146-5.pdf.

Poyhonen:2017:VCD

[1547] Samuli Pöyhönen. Value of cognitive diversity in science. Synthese, 194(11):4519–4540, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1147-4.

Gundersen:2017:LRC

[1548] Eline Busck Gundersen. Lewis’s revised conditional analysis revisited. Synthese, 194(11):4541–4558, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1151-8.

Westra:2017:SMP

[1549] Evan Westra. Spontaneous mindreading: a problem for the two-systems account. Synthese, 194(11):4559–4581, November 2017. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-016-1159-0.

Immerman:2017:QCS

[1550] Daniel Immerman. Question closure to solve the surprise test. Synthese, 194(11):4583–4596, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1160-7.

Dodd:2017:BC

[1551] Dylan Dodd. Belief and certainty. Synthese, 194(11):4597–4621, November 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1163-4.

Kleinschmidt:2017:RFD

[1552] Shieva Kleinschmidt. Refining four-dimensionalism. Synthese, 194(11): 4623–4640, November 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1164-3.

Colombo:2017:IOS

[1553] Matteo Colombo, Raoul Gervais, and Jan Sprenger. Introduction: objectivity in science. Synthese, 194(12):4641–4642, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1582-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1582-x.pdf.

Bialek:2017:IRB

[1554] Max Bialek. Interest relativism in the best system analysis of laws. Synthese, 194(12):4643–4655, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1203-0;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1203-
0.pdf.

Eigi:2017:DMS

[1555] Jaana Eigi. Different motivations, similar proposals: objectivity in scientific community and democratic science policy. Synthese, 194(12): 4657–4669, December 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1077-1.

Koskinen:2017:WEC

[1556] Inkeri Koskinen. Where is the epistemic community? On democratisation of science and social accounts of objectivity. Synthese, 194(12): 4671–4686, December 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1173-2.

Kusch:2017:ERS

[1557] Martin Kusch. Epistemic relativism, scepticism, pluralism. Synthese, 194(12):4687–4703, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1041-0;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1041-0.pdf.

Ludwig:2017:OLK

[1558] David Ludwig. The objectivity of local knowledge. lessons from ethnobiology. Synthese, 194(12):4705–4720, December 2017. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1210-1;

 http://link.springer.com/content/pdf/10.1007/s11229-016-1210-
1.pdf.

Truijens:2017:DNS

[1559] Femke L. Truijens. Do the numbers speak for themselves? A critical analysis of procedural objectivity in psychotherapeutic efficacy research. Synthese, 194(12):4721–4740, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1188-8.

Weber:2017:HOB

[1560] Marcel Weber. How objective are biological functions? Synthese, 194(12):4741–4755, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1483-z.

Bertolotti:2017:TCC

[1561] Tommaso Bertolotti and Lorenzo Magnani. Theoretical considerations on cognitive niche construction. Synthese, 194(12):4757–4779, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1165-2.

Knowles:2017:GEF

[1562] Jonathan Knowles. Global expressivism and the flight from metaphysics. Synthese, 194(12):4781–4797, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1166-1.

Nissan-Rozen:2017:NMG

[1563] Ittay Nissan-Rozen. Newcomb meets Gettier. Synthese, 194(12):4799– 4814, December 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1169-y.

Statham:2017:MCR

[1564] Georgie Statham. The manipulation of chemical reactions: probing the limits of interventionism. Synthese, 194(12):4815–4838, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1170-5;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1170-5.pdf.

Islami:2017:MMH

[1565] Arezoo Islami. A match not made in heaven: on the applicability of mathematics in physics. Synthese, 194(12):4839–4861, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1171-4.

Ellerman:2017:QMS

[1566] David Ellerman. Quantum mechanics over sets: a pedagogical model with non-commutative finite probability theory as its quantum probability calculus. Synthese, 194(12):4863–4896, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/s11229-016-1175-0.

Drake:2017:DPP

[1567] J. Drake. Doxastic permissiveness and the promise of truth. Synthese, 194(12):4897–4912, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1176-z.

Cronin:2017:RTD

[1568] Patrick Cronin. Regularity theories disconfirmed: a revamped argument and a wager. Synthese, 194(12):4913–4933, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/s11229-016-1177-y.

Fisher:2017:CCI

[1569] Tyrus Fisher. Causal counterfactuals are not interventionist counterfactuals. Synthese, 194(12):4935–4957, December 2017. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-016-1183-0.

Gaultier:2017:SPK

[1570] Benoit Gaultier. Skills, procedural knowledge, and knowledge-how. Synthese, 194(12):4959–4981, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1184-z.

Fisher:2017:CDC

[1571] Tyrus Fisher. Counterlegal dependence and causation’s arrows: causal models for backtrackers and counterlegals. Synthese, 194(12):4983– 5003, December 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1189-7.

Kosterec:2017:NT

[1572] Milos Kosterec. On the number of types. Synthese, 194(12):5005– 5021, December 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1190-1.

Whittle:2017:SRP

[1573] Bruno Whittle. Self-referential propositions. Synthese, 194(12):5023– 5037, December 2017. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1191-0. See correction [1947].

Fahrbach:2017:SRE

[1574] Ludwig Fahrbach. Scientific revolutions and the explosion of scientific evidence. Synthese, 194(12):5039–5072, December 2017. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-016-1193-y.

Raerinne:2017:EEB

[1575] Jani Raerinne. Explanations of exceptions in biology: corrective asymmetry versus autonomy. Synthese, 194(12):5073–5092, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1195-9.

Jung:2017:SDP

[1576] Jaemin Jung. Steadfastness, deference, and permissive rationality. Synthese, 194(12):5093–5112, December 2017. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1197-7.

Kostic:2018:MTE

[1577] Daniel Kostić. Mechanistic and topological explanations: an introduction. Synthese, 195(1):1–10, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-016-1257-z;   http://link.springer.
com/content/pdf/10.1007/s11229-016-1257-z.pdf.

Rosenberg:2018:MMI

[1578] Alex Rosenberg. Making mechanism interesting. Synthese, 195(1): 11–33, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0713-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0713-5.pdf.

Harinen:2018:MMC

[1579] Totte Harinen. Mutual manipulability and causal inbetweenness. Synthese, 195(1):35–54, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0564-5;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0564-5.pdf.

Rathkopf:2018:NRC

[1580] Charles Rathkopf. Network representation and complex systems. Synthese, 195(1):55–78, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0726-0;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0726-0.pdf.

Kostic:2018:TR

[1581] Daniel Kostić. The topological realization. Synthese, 195(1):79–98, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1248-0;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1248-0.pdf.

Felline:2018:MMS

[1582] Laura Felline. Mechanisms meet structural explanation. Synthese, 195 (1):99–114, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0746-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0746-9.pdf.

Huneman:2018:DPE

[1583] Philippe Huneman. Diversifying the picture of explanations in biological sciences: ways of combining topology with mechanisms. Synthese, 195 (1):115–146, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0808-z;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0808-z.pdf.

Darrason:2018:MTE

[1584] Marie Darrason. Mechanistic and topological explanations in medicine: the case of medical genetics and network medicine. Synthese, 195(1): 147–173, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0983-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0983-y.pdf.

Lewandowsky:2018:AWM

[1585] Stephan Lewandowsky, John Cook, and Elisabeth Lloyd. The ‘Alice in Wonderland’ mechanics of the rejection of (climate) science: simulating coherence by conspiracism. Synthese, 195(1):175–196, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1198-6;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1198-6.pdf.

Goodman:2018:SCB

[1586] Rachel Goodman. On the supposed connection between proper names and singular thought. Synthese, 195(1):197–223, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1202-1;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1202-1.pdf.

Howson:2018:RPC

[1587] Colin Howson. Repelling a Prussian charge with a solution to a paradox of Dubins. Synthese, 195(1):225–233, January 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1205-y;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1205-
y.pdf.

Murray:2018:RFO

[1588] Samuel Murray. Reference fiction, and omission. Synthese, 195(1): 235–257, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1211-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1211-0.pdf.

Bales:2018:RRC

[1589] Adam Bales. Richness and rationality: causal decision theory and the WAR argument. Synthese, 195(1):259–267, January 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1214-x;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1214-
x.pdf.

Mousavian:2018:COR

[1590] Seyed N. Mousavian and Mohammad Saleh Zarepour. Concept originalism, reference-shift and belief reports. Synthese, 195(1):269–285, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1217-7;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1217-7.pdf.

Lockhart:2018:WWT

[1591] Thomas Lockhart. Why warrant transmits across epistemological disjunctivist Moorean-style arguments. Synthese, 195(1):287–319, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1218-6;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1218-6.pdf.

Ciardelli:2018:QIT

[1592] Ivano Ciardelli. Questions as information types. Synthese, 195(1): 321–365, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1221-y;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1221-y.pdf. See correction [2054].

Beirlaen:2018:LDD

[1593] Mathieu Beirlaen, Bert Leuridan, and Frederik Van De Putte. A logic for the discovery of deterministic causal regularities. Synthese, 195(1): 367–399, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1222-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1222-x.pdf.

Wallach:2018:BRP

[1594] Efraim Wallach. Bayesian representation of a prolonged archaeological debate. Synthese, 195(1):401–431, January 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1224-8;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1224-
8.pdf.

Thomason:2018:LNP

[1595] Richmond H. Thomason. The Little Nell Problem: reasonable and resolute maintenance of agent intentions. Synthese, 195(1):433–440, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1229-3;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1229-3.pdf.

Baskent:2018:YPE

[1596] Can Baskent. A Yabloesque paradox in epistemic game theory. Synthese, 195(1):441–464, January 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1231-9;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1231-9.pdf.

Floridi:2018:WMK

[1597] Luciano Floridi. What a maker’s knowledge could be. Synthese, 195 (1):465–481, January 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1232-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1232-8.pdf.

Jacot:2018:GSI

[1598] Justine Jacot and Philip Pärnamets. GIRL special issue introduction. Synthese, 195(2):483–490, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1656-9;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1656-
9.pdf.

vanRooij:2018:RAI

[1599] Iris van Rooij, Cory D. Wright, Johan Kwisthout, and Todd Wareham. Rational analysis, intractability, and the prospects of ‘as if’-explanations. Synthese, 195(2):491–510, February 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0532-0;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0532-
0.pdf.

Cownden:2018:ILA

[1600] Daniel Cownden, Kimmo Eriksson, and Pontus Strimling. The implications of learning across perceptually and strategically distinct situations. Synthese, 195(2):511–528, February 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0641-9;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0641-
9.pdf.

Rich:2018:CAE

[1601] Patricia Rich. Comparing the axiomatic and ecological approaches to rationality: fundamental agreement theorems in SCOP. Synthese, 195 (2):529–547, February 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0584-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0584-1.pdf.

Gardenfors:2018:LCL

[1602] Peter Gärdenfors. Levels of communication and lexical semantics. Synthese, 195(2):549–569, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0493-3;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0493-3.pdf.

Spychalska:2018:LFM

[1603] Maria Spychalska. At least not false, at most possible: between truth and assertibility of superlative quantifiers. Synthese, 195(2):571–602, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0615-y;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0615-y.pdf.

Bruner:2018:DLL

[1604] Justin Bruner, Cailin O’Connor, Hannah Rubin, and Simon M. Huttegger. David Lewis in the lab: experimental results on the emergence of meaning. Synthese, 195(2):603–621, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0535-x;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0535-x.pdf.

Barrett:2018:EAC

[1605] Jeffrey A. Barrett. The evolution, appropriation, and composition of rules. Synthese, 195(2):623–636, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-014-0421-6;
 http://link.springer.com/content/pdf/10.1007/s11229-014-0421-
6.pdf.

Teng:2018:PFS

[1606] Lu Teng. Is phenomenal force sufficient for immediate perceptual justification? Synthese, 195(2):637–656, February 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1233-7;

 http://link.springer.com/content/pdf/10.1007/s11229-016-1233-
7.pdf.

Brenner:2018:SSC

[1607] Andrew Brenner. Science and the special composition question. Synthese, 195(2):657–678, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1234-6;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1234-6.pdf.

Borgoni:2018:BSK

[1608] Cristina Borgoni. Basic self-knowledge and transparency. Synthese, 195 (2):679–696, February 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1235-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1235-5.pdf.

MacLeod:2018:WMI

[1609] Miles MacLeod. What makes interdisciplinarity difficult? Some consequences of domain specificity in interdisciplinary practice. Synthese, 195 (2):697–720, February 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1236-4;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1236-4.pdf.

Sandgren:2018:WWW

[1610] Alexander Sandgren. Which witch is which? Exotic objects and intentional identity. Synthese, 195(2):721–739, February 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1237-3;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1237-
3.pdf.

Watson:2018:CSS

[1611] David Watson and Luciano Floridi. Crowdsourced science: sociotechnical epistemology in the e-research paradigm. Synthese, 195(2): 741–764, February 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1238-2;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1238-2.pdf.

Sansom:2018:AUT

[1612] Roger Sansom and Jannai Shields. Asymmetry in the unificationist theory of causal explanation. Synthese, 195(2):765–783, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1241-7;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1241-7.pdf.

Kramer:2018:TTG

[1613] Stephan Krämer. Towards a theory of ground-theoretic content. Synthese, 195(2):785–814, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1242-6;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1242-6.pdf.

Longworth:2018:UWW

[1614] Guy Longworth. Understanding what was said. Synthese, 195(2): 815–834, February 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1243-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1243-5.pdf.

Beckers:2018:PAD

[1615] Sander Beckers and Joost Vennekens. A principled approach to defining actual causation. Synthese, 195(2):835–862, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1247-1;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1247-1.pdf.

Isaac:2018:TPE

[1616] Manuel Gustavo Isaac. Toward a phenomenological epistemology of mathematical logic. Synthese, 195(2):863–874, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1249-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1249-z.pdf.

Sandoz:2018:AME

[1617] Raphaël Sandoz. Applying mathematics to empirical sciences: flashback to a puzzling disciplinary interaction. Synthese, 195(2):875–898, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1251-5;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1251-5.pdf.

Chase:2018:FCK

[1618] James Chase and Penelope Rush. Factivity, consistency and knowability. Synthese, 195(2):899–918, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1253-3;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1253-3.pdf.

Khudairi:2018:GCM

[1619] Hasen Khudairi. Grounding, conceivability, and the mind-body problem. Synthese, 195(2):919–926, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1254-2;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1254-

2.pdf. See correction [2083].

Ulatowski:2018:MAT

[1620] Joseph Ulatowski and Cory Wright. Minimalism about truth: special issue introduction. Synthese, 195(3):927–933, March 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1376-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1376-
1.pdf.

Howat:2018:CAP

[1621] Andrew W. Howat. Constituting assertion: a pragmatist critique of Horwich’s ‘Truth’. Synthese, 195(3):935–954, March 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1196-8;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1196-
8.pdf.

Kijania-Placek:2018:CMA

[1622] Katarzyna Kijania-Placek. Can minimalism about truth embrace polysemy? Synthese, 195(3):955–985, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1228-4;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1228-
4.pdf.

Wright:2018:TEM

[1623] Cory Wright. Truth, explanation, minimalism. Synthese, 195(3): 987–1009, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1129-6;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1129-6.pdf.

Simmons:2018:TQM

[1624] Keith Simmons. Three questions for minimalism. Synthese, 195(3): 1011–1034, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1135-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1135-8.pdf.

Marques:2018:I

[1625] Teresa Marques. This is not an instance of (e). Synthese, 195(3): 1035–1063, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1293-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1293-8.pdf.

Gupta:2018:IPG

[1626] Anil Gupta and Shawn Standefer. Intersubstitutivity principles and the generalization function of truth. Synthese, 195(3):1065–1075, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1318-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1318-y.pdf.

Cieslinski:2018:MGP

[1627] Cezary Cieśliński. Minimalism and the generalisation problem: on Horwich’s second solution. Synthese, 195(3):1077–1101, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1227-5;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1227-5.pdf.

Ferrari:2018:VMT

[1628] Filippo Ferrari. The value of minimalist truth. Synthese, 195(3): 1103–1125, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1207-9;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1207-9.pdf.

Horwich:2018:TNC

[1629] Paul Horwich. Is truth a normative concept? Synthese, 195(3): 1127–1138, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1208-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1208-8.pdf.

Campbell:2018:DDA

[1630] Douglas I. Campbell. Doxastic desire and attitudinal monism. Synthese, 195(3):1139–1161, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1255-1;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1255-1.pdf.

Tebben:2018:BIV

[1631] Nicholas Tebben. Belief isn’t voluntary, but commitment is. Synthese, 195(3):1163–1179, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1258-y;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1258-y.pdf.

DiBella:2018:QPN

[1632] Nicholas DiBella. The qualitative paradox of non-conglomerability. Synthese, 195(3):1181–1210, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1261-3;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1261-3.pdf.

McDonnell:2018:TPC

[1633] Neil McDonnell. Transitivity and proportionality in causation. Synthese, 195(3):1211–1229, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1263-1;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1263-1.pdf.

Poggiolesi:2018:CLN

[1634] Francesca Poggiolesi. On constructing a logic for the notion of complete and immediate formal grounding. Synthese, 195(3):1231–1254, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1265-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1265-z.pdf.

Leuschner:2018:IAT

[1635] Anna Leuschner. Is it appropriate to ‘target’ inappropriate dissent? On the normative consequences of climate skepticism. Synthese, 195 (3):1255–1271, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1267-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1267-x.pdf.

Hartmann:2018:VDT

[1636] Stephan Hartmann and Soroush Rafiee Rad. Voting, deliberation and truth. Synthese, 195(3):1273–1293, March 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1268-9;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1268-
9.pdf.

Christias:2018:PCM

[1637] Dionysis Christias. On the proper construal of the manifest-scientific image distinction: Brandom contra Sellars. Synthese, 195(3):1295–1320, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1271-1;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1271-1.pdf.

Bolos:2018:SVE

[1638] Anthony Bolos and James Henry Collin. A sensitive virtue epistemology. Synthese, 195(3):1321–1335, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1273-z;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1273-
z.pdf.

Lange:2018:TSE

[1639] Marc Lange. Transitivity, self-explanation, and the explanatory circularity argument against Humean accounts of natural law. Synthese, 195 (3):1337–1353, March 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1274-y;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1274-y.pdf.

Wright:2018:RLD

[1640] Jonathan Wright. Rigged lotteries: a diachronic problem for reducing belief to credence. Synthese, 195(3):1355–1373, March 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1275-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1275-x.pdf.

McFarland:2018:ISS

[1641] Andrew McFarland. Introduction for Synthese special issue causation in the metaphysics of science: natural kinds. Synthese, 195(4):1375–1378, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1631-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1631-5.pdf.

Khalidi:2018:NKN

[1642] Muhammad Ali Khalidi. Natural kinds as nodes in causal networks. Synthese, 195(4):1379–1396, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0841-y;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0841-y.pdf.

Bird:2018:MNK

[1643] Alexander Bird. The metaphysics of natural kinds. Synthese, 195(4): 1397–1426, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0833-y;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0833-y.pdf.

Magnus:2018:TON

[1644] P. D. Magnus. Taxonomy, ontology, and natural kinds. Synthese, 195 (4):1427–1439, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0785-2;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0785-2.pdf.

McFarland:2018:CPI

[1645] Andrew McFarland. Causal powers and isomeric chemical kinds. Synthese, 195(4):1441–1457, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1044-x;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1044-x.pdf.

Longy:2018:DWN

[1646] Françoise Longy. Do we need two notions of natural kind to account for the history of ”jade”? Synthese, 195(4):1459–1486, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1213-y;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1213-y.pdf.

Slater:2018:AEK

[1647] Matthew H. Slater. Anchoring in ecosystemic kinds. Synthese, 195 (4):1487–1508, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1302-y;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1302-y.pdf.

Hartimo:2018:HCD

[1648] Mirja Hartimo. Husserl on completeness, definitely. Synthese, 195 (4):1509–1527, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1278-7;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1278-7.pdf.

Strollo:2018:SNV

[1649] Andrea Strollo. A simple notion of validity for alethic pluralism. Synthese, 195(4):1529–1546, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1280-0;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1280-0.pdf.

Peet:2018:EUT

[1650] Andrew Peet. Etiology, understanding, and testimonial belief. Synthese, 195(4):1547–1567, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1281-z;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1281-z.pdf.

Delhotel:2018:RFC

[1651] Jean-Michel Delhôtel. Relativistic frameworks and the case for (or against) incommensurability. Synthese, 195(4):1569–1585, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1283-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1283-x.pdf.

Willemsen:2018:OEN

[1652] Pascale Willemsen. Omissions and expectations: a new approach to the things we failed to do. Synthese, 195(4):1587–1614, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1284-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1284-9.pdf.

Wallbridge:2018:PCL

[1653] Kevin Wallbridge. The peculiar case of Lehrer’s lawyer. Synthese, 195 (4):1615–1630, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1285-8;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1285-8.pdf.

Roche:2018:TPB

[1654] William Roche. Is there a place in Bayesian confirmation theory for the Reverse Matthew Effect? Synthese, 195(4):1631–1648, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1286-7;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1286-7.pdf.

Urbaniak:2018:CLC

[1655] Rafal Urbaniak and Bert Leuridan. Challenging Lewis’s challenge to the best system account of lawhood. Synthese, 195(4):1649–1666, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1287-6;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1287-6.pdf.

Lewerentz:2018:MIC

[1656] Lukas Lewerentz and Benjamin Marschall. Metasemantics, intentions and circularity. Synthese, 195(4):1667–1679, April 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1290-y;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1290-
y.pdf.

Girard:2018:PCP

[1657] Patrick Girard and Marcus A. Triplett. Prioritised ceteris paribus logic for counterfactual reasoning. Synthese, 195(4):1681–1703, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1296-5;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1296-5.pdf.

Ganson:2018:SML

[1658] Todd Ganson. Sensory malfunctions, limitations, and trade-offs. Synthese, 195(4):1705–1713, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1298-3;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1298-3.pdf.

Godden:2018:PAA

[1659] David Godden and Frank Zenker. A probabilistic analysis of argument cogency. Synthese, 195(4):1715–1740, April 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1299-2;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1299-
2.pdf.

Dolev:2018:OKU

[1660] Yuval Dolev. Is ontology the key to understanding tense? Synthese, 195 (4):1741–1749, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1303-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1303-x.pdf.

Green:2018:NAS

[1661] Sara Green, Maria Serban, Raphael Scholl, Nicholaos Jones, Ingo Brigandt, and William Bechtel. Network analyses in systems biology: new strategies for dealing with biological complexity. Synthese, 195(4): 1751–1777, April 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1307-6;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1307-6.pdf.

White:2018:MND

[1662] Ben White. Metaphysical necessity dualism. Synthese, 195(4):1779–1798, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1308-5;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1308-5.pdf.

Trueman:2018:PVP

[1663] Robert Trueman. The Prenective View of propositional content. Synthese, 195(4):1799–1825, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1309-4;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1309-4.pdf.

Moon:2018:NDN

[1664] Andrew Moon. The nature of doubt and a new puzzle about belief, doubt, and confidence. Synthese, 195(4):1827–1848, April 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1310-y;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1310-y.pdf.

Arstila:2018:WMU

[1665] Valtteri Arstila. What makes unique hues unique? Synthese, 195 (5):1849–1872, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1313-3;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1313-3.pdf.

Holtzman:2018:NCS

[1666] Geoffrey S. Holtzman. A neuropsychological challenge to the sentimentalism/rationalism distinction. Synthese, 195(5):1873–1889, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1344-9;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1344-9.pdf.

Villalobos:2018:EAC

[1667] Mario Villalobos and Joe Dewhurst. Enactive autonomy in computational systems. Synthese, 195(5):1891–1908, May 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1386-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1386-
z.pdf.

Rescorla:2018:IAP

[1668] Michael Rescorla. An interventionist approach to psychological explanation. Synthese, 195(5):1909–1940, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1553-2;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1553-
2.pdf.

Williams:2018:SIR

[1669] Daniel Williams and Lincoln Colling. From symbols to icons: the return of resemblance in the cognitive neuroscience revolution. Synthese, 195(5):1941–1967, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1578-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1578-6.pdf.

Buehler:2018:CES

[1670] Denis Buehler. The central executive system. Synthese, 195(5): 1969–1991, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1589-3;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1589-3.pdf.

Haraldsen:2018:RT

[1671] Fredrik Haraldsen. Rigidity and triviality. Synthese, 195(5):1993–1999, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1311-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1311-x.pdf.

Ovenden:2018:GCA

[1672] Chris Ovenden. Guidance control and the anti-akrasia chip. Synthese, 195(5):2001–2019, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1312-4;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1312-4.pdf.

Levy:2018:DNQ

[1673] Yair Levy. Does the normative question about rationality rest on a mistake? Synthese, 195(5):2021–2038, May 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1314-2;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1314-
2.pdf.

Ohtani:2018:PPA

[1674] Hiroshi Ohtani. Philosophical pictures about mathematics: Wittgenstein and contradiction. Synthese, 195(5):2039–2063, May 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1317-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1317-
z.pdf.

Genot:2018:SI

[1675] Emmanuel J. Genot. Strategies of inquiry. Synthese, 195(5):2065–2088, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1319-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1319-x.pdf.

Schlosser:2018:ECT

[1676] Markus E. Schlosser. Embodied cognition and temporally extended agency. Synthese, 195(5):2089–2112, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1320-4;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1320-
4.pdf.

Knappik:2018:BFP

[1677] Franz Knappik. Bayes and the first person: consciousness of thoughts, inner speech and probabilistic inference. Synthese, 195(5):2113–2140, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1321-3;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1321-3.pdf.

Normand:2018:CIH

[1678] Scott Normand. Criteria of identity and the hermeneutic goal of ante rem structuralism. Synthese, 195(5):2141–2153, May 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1325-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1325-
z.pdf.

Lam:2018:ITL

[1679] Derek Lam. Is imagination too liberal for modal epistemology? Synthese, 195(5):2155–2174, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1329-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1329-8.pdf.

Pupa:2018:ACR

[1680] Francesco Pupa. The argument from convention revisited. Synthese, 195(5):2175–2204, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1330-2;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1330-2.pdf.

Rettler:2018:DDB

[1681] Lindsay Rettler. In defense of doxastic blame. Synthese, 195(5): 2205–2226, May 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1332-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1332-0.pdf.

Bright:2018:DBD

[1682] Liam Kofi Bright. Du Bois’ democratic defence of the value free ideal. Synthese, 195(5):2227–2245, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1333-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1333-
z.pdf.

Bradley:2018:CEC

[1683] Darren Bradley. Carnap’s epistemological critique of metaphysics. Synthese, 195(5):2247–2265, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1335-x;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1335-x.pdf.

Politi:2018:SRS

[1684] Vincenzo Politi. Scientific revolutions, specialization and the discovery of the structure of DNA: toward a new picture of the development of the sciences. Synthese, 195(5):2267–2293, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1339-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1339-6.pdf. See erratum [1809].

Malinsky:2018:IS

[1685] Daniel Malinsky. Intervening on structure. Synthese, 195(5):2295–2312, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1341-z;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1341-z.pdf.

Demircioglu:2018:EIC

[1686] Erhan Demircioglu. Epistemic infinitism and the conditional character of inferential justification. Synthese, 195(5):2313–2334, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1529-2;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1529-2.pdf.

Fisher:2018:LAM

[1687] A. R. J. Fisher. On Lewis against magic: a study of method in metaphysics. Synthese, 195(5):2335–2353, May 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0679-3;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0679-

3.pdf. See correction [2079].

Kirchhoff:2018:PBE

[1688] Michael Kirchhoff. Predictive brains and embodied, enactive cognition: an introduction to the special issue. Synthese, 195(6):2355–2366, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1534-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1534-5.pdf.

Orlandi:2018:PPS

[1689] Nico Orlandi. Predictive perceptual systems. Synthese, 195(6):2367– 2386, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1373-4;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1373-4.pdf.

Kiefer:2018:CMH

[1690] Alex Kiefer and Jakob Hohwy. Content and misrepresentation in hierarchical generative models. Synthese, 195(6):2387–2415, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1435-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1435-7.pdf.

Bruineberg:2018:ABS

[1691] Jelle Bruineberg, Julian Kiverstein, and Erik Rietveld. The anticipating brain is not a scientist: the free-energy principle from an ecological–enactive perspective. Synthese, 195(6):2417–2444, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1239-1;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1239-1.pdf.

Hutto:2018:GPP

[1692] Daniel D. Hutto. Getting into predictive processing’s great guessing game: Bootstrap heaven or hell? Synthese, 195(6):2445–2458, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1385-0;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1385-0.pdf.

Allen:2018:CAT

[1693] Micah Allen and Karl J. Friston. From cognitivism to autopoiesis: towards a computational framework for the embodied mind. Synthese, 195 (6):2459–2482, June 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1288-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1288-5.pdf.

Fabry:2018:BBE

[1694] Regina E. Fabry. Betwixt and between: the enculturated predictive processing approach to cognition. Synthese, 195(6):2483–2518, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1334-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1334-y.pdf.

Kirchhoff:2018:AFE

[1695] Michael D. Kirchhoff. Autopoiesis, free energy, and the life-mind continuity thesis. Synthese, 195(6):2519–2540, June 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1100-6;

 http://link.springer.com/content/pdf/10.1007/s11229-016-1100-
6.pdf.

Klein:2018:WDP

[1696] Colin Klein. What do predictive coders want? Synthese, 195(6): 2541–2557, June 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1250-6;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1250-6.pdf.

Miller:2018:HEP

[1697] Mark Miller and Andy Clark. Happily entangled: prediction, emotion, and the embodied mind. Synthese, 195(6):2559–2575, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1399-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1399-7.pdf. See correction [2078].

Windt:2018:PBD

[1698] Jennifer M. Windt. Predictive brains, dreaming selves, sleeping bodies: how the analysis of dream movement can inform a theory of selfand world-simulation in dreams. Synthese, 195(6):2577–2625, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1525-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1525-6.pdf.

Gallagher:2018:AIE

[1699] Shaun Gallagher and Micah Allen. Active inference, enactivism and the hermeneutics of social cognition. Synthese, 195(6):2627–2648, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1269-8;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1269-8.pdf.

Richard:2018:LMD

[1700] Sebastien Richard. Leśniewski on metalogic and definitions. Synthese, 195(6):2649–2676, June 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1343-x;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1343-x.pdf.

Textor:2018:FRC

[1701] Mark Textor. Frege’s recognition criterion for thoughts and its problems. Synthese, 195(6):2677–2696, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1345-8;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1345-8.pdf.

Berto:2018:CPS

[1702] Francesco Berto and Tom Schoonen. Conceivability and possibility: some dilemmas for humeans. Synthese, 195(6):2697–2715, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1346-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1346-7.pdf.

Listwa:2018:FSP

[1703] Daniel Listwa. The Faulty Signal Problem: counterfactual asymmetries in causal decision theory and rational deliberation. Synthese, 195 (6):2717–2739, June 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1348-5;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1348-5.pdf.

Bochner:2018:STC

[1704] Gregory Bochner. Singular truth-conditions without singular propositions. Synthese, 195(6):2741–2760, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1354-7;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1354-7.pdf.

Keas:2018:STV

[1705] Michael N. Keas. Systematizing the theoretical virtues. Synthese, 195 (6):2761–2793, June 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1355-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1355-6.pdf.

Rice:2018:IMH

[1706] Collin Rice. Idealized models, holistic distortions, and universality. Synthese, 195(6):2795–2819, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1357-4;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1357-4.pdf.

Dinges:2018:KII

[1707] Alexander Dinges. Knowledge, intuition and implicature. Synthese, 195 (6):2821–2843, June 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1359-2;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1359-2.pdf.

Moretti:2018:DCE

[1708] Luca Moretti and Tommaso Piazza. Defeaters in current epistemology: introduction to the special issue. Synthese, 195(7):2845–2854, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1551-4;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1551-4.pdf.

Bagnoli:2018:DPK

[1709] Carla Bagnoli. Defeaters and practical knowledge. Synthese, 195(7): 2855–2875, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1095-z;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1095-z.pdf.

Carter:2018:MED

[1710] J. Adam Carter. Meta-epistemic defeat. Synthese, 195(7):2877–2896, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1187-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1187-9.pdf.

Casullo:2018:PSD

[1711] Albert Casullo. Pollock and Sturgeon on defeaters. Synthese, 195 (7):2897–2906, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1073-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1073-5.pdf.

Steup:2018:DDJ

[1712] Matthias Steup. Destructive defeat and justificational force: the dialectic of dogmatism, conservatism, and meta-evidentialism. Synthese, 195(7):2907–2933, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1182-1;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1182-1.pdf.

Pedersen:2018:NRA

[1713] Nikolaj Jang Lee Linding Pedersen. Non-rational action in the face of disagreement: an argument against (strong) non-conformism. Synthese, 195(7):2935–2966, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1086-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1086-0.pdf. See correction [2082].

Brogaard:2018:DHM

[1714] Berit Brogaard. In defense of hearing meanings. Synthese, 195(7): 2967–2983, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1178-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1178-x.pdf.

Gluer:2018:DL

[1715] Kathrin Glüer. Defeating looks. Synthese, 195(7):2985–3012, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1186-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1186-x.pdf.

Graham:2018:FRA

[1716] Peter J. Graham. Formulating reductionism about testimonial warrant and the challenge from childhood testimony. Synthese, 195(7):3013–3033, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1140-y;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1140-y.pdf.

McCain:2018:EAP

[1717] Kevin McCain. Explanationist aid for phenomenal conservatism. Synthese, 195(7):3035–3050, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1064-6;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1064-6.pdf.

McGrath:2018:DPE

[1718] Matthew McGrath. Defeating pragmatic encroachment? Synthese, 195 (7):3051–3064, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1264-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1264-0.pdf.

Pritchard:2018:ALV

[1719] Duncan Pritchard. Anti-luck virtue epistemology and epistemic defeat. Synthese, 195(7):3065–3077, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1074-4;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1074-4.pdf.

McAllister:2018:SSG

[1720] Blake McAllister. Seemings as sui generis. Synthese, 195(7):3079–3096, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1360-9;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1360-9.pdf.

Cooper:2018:TDT

[1721] Andrew Cooper. Two directions for teleology: naturalism and idealism. Synthese, 195(7):3097–3119, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1364-5;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1364-5.pdf.

Dellsen:2018:DCU

[1722] Finnur Dellsén. Deductive cogency, understanding, and acceptance. Synthese, 195(7):3121–3141, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1365-4;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1365-4.pdf.

Topey:2018:QHA

[1723] Brett Topey. Quinean holism, analyticity, and diachronic rational norms. Synthese, 195(7):3143–3171, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1366-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1366-
3.pdf.

Radulescu:2018:DBI

[1724] Alexandru Radulescu. The difference between indexicals and demonstratives. Synthese, 195(7):3173–3196, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1367-2;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1367-2.pdf.

Fuller:2018:WCD

[1725] Jonathan Fuller. What are chronic diseases? Synthese, 195(7):3197–3220, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1368-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1368-1.pdf.

Kogelmann:2018:WWC

[1726] Brian Kogelmann. What we choose, what we prefer. Synthese, 195 (7):3221–3240, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1369-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1369-0.pdf.

Svoboda:2018:LTD

[1727] Vladimír Svoboda. A Lewisian taxonomy for deontic logic. Synthese, 195(7):3241–3266, July 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1370-7;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1370-7.pdf.

Roberts:2018:LIA

[1728] Pendaran Roberts, James Andow, and Kelly Ann Schmidtke. Lay intuitions about epistemic normativity. Synthese, 195(7):3267–3287, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1371-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1371-6.pdf.

Ben-Yami:2018:ISS

[1729] Hanoch Ben-Yami, Robyn Carston, and Markus Werning. Introduction to the 2nd Synthese Special Issue: trends in philosophy of language and mind. Synthese, 195(8):3289–3291, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1174-1;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1174-1.pdf.

Alvarez:2018:RAA

[1730] Maria Alvarez. Reasons for action, acting for reasons, and rationality. Synthese, 195(8):3293–3310, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-1005-9;
 http://link.springer.com/content/pdf/10.1007/s11229-015-1005-
9.pdf.

Garcia-Carpintero:2018:STI

[1731] Manuel García-Carpintero. De se thoughts and immunity to error through misidentification. Synthese, 195(8):3311–3333, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0817-y;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0817-y.pdf.

Marques:2018:R

[1732] Teresa Marques. Retractions. Synthese, 195(8):3335–3359, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0852-8;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0852-8.pdf.

Nida-Rumelin:2018:EPF

[1733] Martine Nida-Rümelin. The experience property frame work: a misleading paradigm. Synthese, 195(8):3361–3387, August 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1121-1;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1121-
1.pdf.

Toribio:2018:VER

[1734] Josefa Toribio. Visual experience: rich but impenetrable. Synthese, 195 (8):3389–3406, August 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0889-8;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0889-8.pdf.

Kannisto:2018:KFE

[1735] Toni Kannisto. Kant and Frege on existence. Synthese, 195(8):3407–3432, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1372-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1372-5.pdf.

Poznic:2018:TVT

[1736] Michael Poznic. Thin versus thick accounts of scientific representation. Synthese, 195(8):3433–3451, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1374-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1374-
3.pdf.

Sullivan-Bissett:2018:EDT

[1737] Ema Sullivan-Bissett. Explaining doxastic transparency: aim, norm, or function? Synthese, 195(8):3453–3476, August 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1377-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1377-
0.pdf.

Mollo:2018:FIM

[1738] Dimitri Coelho Mollo. Functional individuation, mechanistic implementation: the proper way of seeing the mechanistic view of concrete computation. Synthese, 195(8):3477–3497, August 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1380-5;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1380-
5.pdf.

Lee:2018:MFC

[1739] Poong Shil Lee. Mental files, concepts, and bodies of information. Synthese, 195(8):3499–3518, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1381-4;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1381-4.pdf.

Collins:2018:RA

[1740] John Collins. The redundancy of the act. Synthese, 195(8):3519–3545, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1382-3;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1382-3.pdf.

Akagi:2018:RPC

[1741] Mikio Akagi. Rethinking the problem of cognition. Synthese, 195(8): 3547–3570, August 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1383-2;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1383-2.pdf.

Kopec:2018:PAE

[1742] Matthew Kopec. A pluralistic account of epistemic rationality. Synthese, 195(8):3571–3596, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1388-x;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1388-x.pdf.

Habgood-Coote:2018:KHS

[1743] Joshua Habgood-Coote. Knowing-how, showing, and epistemic norms. Synthese, 195(8):3597–3620, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1389-9;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1389-9.pdf.

Ma:2018:GGC

[1744] Minghui Ma and Ahti-Veikko Pietarinen. Gamma graph calculi for modal logics. Synthese, 195(8):3621–3650, August 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1390-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1390-
3.pdf.

Wen:2018:JAN

[1745] Xuefeng Wen. Judgment aggregation in nonmonotonic logic. Synthese, 195(8):3651–3683, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1391-2;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1391-2.pdf.

Schipper:2018:ANT

[1746] Arthur Schipper. Aboutness and negative truths: a modest strategy for truthmaker theorists. Synthese, 195(8):3685–3722, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1396-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1396-x.pdf.

Clark:2018:GMC

[1747] Michael J. Clark and Nathan Wildman. Grounding, mental causation, and overdetermination. Synthese, 195(8):3723–3733, August 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1402-3;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1402-3.pdf.

Bewersdorf:2018:ISI

[1748] Benjamin Bewersdorf and Jeanne Peijnenburg. Introduction to the special issue on epistemic justification. Synthese, 195(9):3735, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1440-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1440-x.pdf.

Cresto:2018:KAR

[1749] Eleonora Cresto. Knowledge attribution revisited: a deflationary account. Synthese, 195(9):3737–3753, September 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1282-y;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1282-
y.pdf.

Dutant:2018:JDI

[1750] Julien Dutant and Clayton Littlejohn. Just do it? When to do what you judge you ought to do. Synthese, 195(9):3755–3772, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1220-z;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1220-z.pdf.

Booth:2018:AID

[1751] Anthony Robert Booth. Advice for infallibilists: DIVORCE and RETREAT! Synthese, 195(9):3773–3789, September 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1421-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1421-
0.pdf.

Sharadin:2018:EIR

[1752] Nathaniel Sharadin. Epistemic instrumentalism and the reason to believe in accord with the evidence. Synthese, 195(9):3791–3809, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1245-3;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1245-3.pdf.

Bondy:2018:PEL

[1753] Patrick Bondy and Duncan Pritchard. Propositional epistemic luck, epistemic risk, and epistemic justification. Synthese, 195(9):3811– 3820, September 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1262-2;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1262-2.pdf.

deGrefte:2018:EJE

[1754] Job de Grefte. Epistemic justification and epistemic luck. Synthese, 195(9):3821–3836, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1306-7;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1306-7.pdf.

Spohn:2018:EJS

[1755] Wolfgang Spohn. Epistemic justification: its subjective and its objective ways. Synthese, 195(9):3837–3856, September 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1393-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1393-
0.pdf.

Smith:2018:LEJ

[1756] Martin Smith. The logic of epistemic justification. Synthese, 195(9): 3857–3875, September 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1422-z;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1422-z.pdf.

Schurz:2018:OJN

[1757] Gerhard Schurz. Optimality justifications: new foundations for foundation-oriented epistemology. Synthese, 195(9):3877–3897, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1363-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1363-6.pdf.

Roche:2018:FIR

[1758] William Roche. Foundationalism with infinite regresses of probabilistic support. Synthese, 195(9):3899–3917, September 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1289-4;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1289-
4.pdf.

Bich:2018:DLP

[1759] Leonardo Bich and Sara Green. Is defining life pointless? Operational definitions at the frontiers of biology. Synthese, 195(9):3919– 3946, September 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1397-9;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1397-9.pdf.

Palmira:2018:TPT

[1760] Michele Palmira. Towards a pluralist theory of singular thought. Synthese, 195(9):3947–3974, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1401-4;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1401-4.pdf.

Balaguer:2018:WDA

[1761] Mark Balaguer. Why the debate about composition is factually empty (or why there’s no fact of the matter whether anything exists). Synthese, 195(9):3975–4008, September 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1403-2;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1403-
2.pdf.

Spaulding:2018:DYS

[1762] Shannon Spaulding. Do you see what I see? How social differences influence mindreading. Synthese, 195(9):4009–4030, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1404-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1404-1.pdf.

Fine:2018:II

[1763] Kit Fine. Ignorance of ignorance. Synthese, 195(9):4031–4045, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1406-z;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1406-z.pdf.

Laraudogoitia:2018:RID

[1764] Jon Pérez Laraudogoitia. Rigidity, instability and dimensionality. Synthese, 195(9):4047–4062, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1407-y;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1407-y.pdf.

Dawid:2018:NMA

[1765] Richard Dawid and Stephan Hartmann. The no miracles argument without the base rate fallacy. Synthese, 195(9):4063–4079, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1408-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1408-x.pdf.

Tzinman:2018:RDP

[1766] Rina Tzinman. Is Romeo dead? On the persistence of organisms. Synthese, 195(9):4081–4105, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1409-9;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1409-
9.pdf.

Bohn:2018:NAW

[1767] Einar Duenger Bohn. Normativity all the way down: from normative realism to pannormism. Synthese, 195(9):4107–4124, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1410-3;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1410-3.pdf.

Fumagalli:2018:WAS

[1768] Roberto Fumagalli. Who is afraid of scientific imperialism? Synthese, 195(9):4125–4146, September 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1411-2;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1411-2.pdf.

Hannon:2018:IRJ

[1769] Michael Hannon. Intuitions, reflective judgments, and experimental philosophy. Synthese, 195(9):4147–4168, September 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1412-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1412-
1.pdf.

Smart:2018:MI

[1770] Paul R. Smart. Mandevillian intelligence. Synthese, 195(9):4169– 4200, September 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1414-z;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1414-z.pdf.

vanderHoek:2018:LRI

[1771] Wiebe van der Hoek, Wesley H. Holiday, and Wen fang Wang. Logic, rationality and interaction (LORI-5): introduction to the special issue. Synthese, 195(10):4201–4204, October 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1640-4;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1640-
4.pdf.

Gaifman:2018:SMR

[1772] Haim Gaifman and Yang Liu. A simpler and more realistic subjective decision theory. Synthese, 195(10):4205–4241, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1594-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1594-6.pdf.

Georgatos:2018:GM

[1773] Konstantinos Georgatos. Geodesic merging. Synthese, 195(10):4243– 4264, October 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1432-x;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1432-x.pdf.

Ghosh:2018:SST

[1774] Sujata Ghosh and Rineke Verbrugge. Studying strategies and types of players: experiments, logics and cognitive models. Synthese, 195(10):

4265–4307, October 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1338-7;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1338-7.pdf.

Hawke:2018:IDE

[1775] Peter Hawke and Shane Steinert-Threlkeld. Informational dynamics of epistemic possibility modals. Synthese, 195(10):4309–4342, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1216-8;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1216-8.pdf.

Klein:2018:KBN

[1776] Dominik Klein, Olivier Roy, and Norbert Gratzl. Knowledge, belief, normality, and introspection. Synthese, 195(10):4343–4372, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1353-8;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1353-8.pdf.

Kocurek:2018:EPF

[1777] Alexander W. Kocurek. On the expressive power of first-order modal logic with two-dimensional operators. Synthese, 195(10):4373–4417, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1172-3;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1172-3.pdf.

Wang:2018:LGD

[1778] Yanjing Wang. A logic of goal-directed knowing how. Synthese, 195(10): 4419–4439, October 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1272-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1272-0.pdf.

Wang:2018:TVS

[1779] Wen fang Wang. Three-valued semantic pluralism: a defense of a threevalued solution to the sorites paradox. Synthese, 195(10):4441–4476, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1517-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1517-6.pdf.

Jhun:2018:MCC

[1780] Jennifer Jhun, Patricia Palacios, and James Owen Weatherall. Market crashes as critical phenomena? explanation, idealization, and universality in econophysics. Synthese, 195(10):4477–4505, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1415-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1415-y.pdf.

Salas:2018:WPB

[1781] Javier González de Prado Salas. Whose purposes? Biological teleology and intentionality. Synthese, 195(10):4507–4524, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1416-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1416-x.pdf.

Yates:2018:IFI

[1782] David Yates. Inverse functionalism and the individuation of powers. Synthese, 195(10):4525–4550, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1417-9;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1417-9.pdf.

Madison:2018:JE

[1783] B. J. C. Madison. On justifications and excuses. Synthese, 195(10): 4551–4562, October 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1418-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1418-8.pdf.

Runyan:2018:ACL

[1784] Jason D. Runyan. Agent-causal libertarianism, statistical neural laws and wild coincidences. Synthese, 195(10):4563–4580, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1419-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1419-7.pdf.

Aagotnes:2018:TL

[1785] Thomas Ågotnes, Hans van Ditmarsch, and Yanjing Wang. True lies. Synthese, 195(10):4581–4615, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1423-y;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1423-y.pdf.

Buckareff:2018:IJS

[1786] Andrei A. Buckareff. I’m just sitting around doing nothing: on exercising intentional agency in omitting to act. Synthese, 195(10):4617–4635, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1424-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1424-x.pdf.

Peressini:2018:TNI

[1787] Anthony F. Peressini. There is nothing it is like to see red: holism and subjective experience. Synthese, 195(10):4637–4666, October 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1425-9;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1425-9.pdf.

Rauzy:2018:CEI

[1788] Jean-Baptiste Rauzy, Stefano Cossara, and Xiaoxing Zhang. Cartesian epistemology: an introduction. Synthese, 195(11):4667–4669, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1705-z;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1705-z.pdf.

Fumerton:2018:CEI

[1789] Richard Fumerton. Cartesian epistemology and infallible justification. Synthese, 195(11):4671–4681, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1456-2;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1456-
2.pdf.

Hetherington:2018:RPK

[1790] Stephen Hetherington. The redundancy problem: From knowledgeinfallibilism to knowledge-minimalism. Synthese, 195(11):4683–4702, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1091-3;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1091-3.pdf.

Guillon:2018:EDA

[1791] Jean Baptiste Guillon. The Evil Demon argument as based on closure plus meta-coherence. Synthese, 195(11):4703–4731, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1215-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1215-9.pdf.

Pritchard:2018:CRS

[1792] Duncan Pritchard. Contextualism and radical scepticism. Synthese, 195(11):4733–4750, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1122-0;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1122-0.pdf.

Rabouin:2018:LIE

[1793] David Rabouin. Logic of imagination. Echoes of Cartesian epistemology in contemporary philosophy of mathematics and beyond. Synthese, 195(11):4751–4783, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1562-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1562-1.pdf.

Saint-Germier:2018:CIC

[1794] Pierre Saint-Germier. Conceivability, inconceivability and Cartesian modal epistemology. Synthese, 195(11):4785–4816, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1194-x;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1194-x.pdf.

Colombo:2018:BCS

[1795] Matteo Colombo. Bayesian cognitive science, predictive brains, and the nativism debate. Synthese, 195(11):4817–4838, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1427-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1427-7.pdf.

Shech:2018:IES

[1796] Elay Shech. Idealizations, essential self-adjointness, and minimal model explanation in the Aharonov–Bohm effect. Synthese, 195(11):4839– 4863, November 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1428-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1428-6.pdf.

Majeed:2018:WCP

[1797] Raamy Majeed. Why the Canberra plan won’t help you do serious metaphysics. Synthese, 195(11):4865–4882, November 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1430-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1430-
z.pdf.

Roush:2018:ESE

[1798] Sherrilyn Roush. The epistemic superiority of experiment to simulation. Synthese, 195(11):4883–4906, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1431-y;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1431-
y.pdf.

Polzler:2018:CES

[1799] Thomas Pölzler. Can the empirical sciences contribute to the moral realism/anti-realism debate? Synthese, 195(11):4907–4930, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1434-8;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1434-8.pdf.

Hausmann:2018:CAU

[1800] Marco Hausmann. The consequence argument ungrounded. Synthese, 195(11):4931–4950, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1436-6;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1436-6.pdf.

Clarke:2018:ABC

[1801] Roger Clarke. Assertion, belief, and context. Synthese, 195(11):4951– 4977, November 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1437-5;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1437-5.pdf.

McLear:2018:MAA

[1802] Colin McLear. Motion and the affection argument. Synthese, 195(11): 4979–4995, November 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1438-4;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1438-4.pdf.

Morvarid:2018:FEL

[1803] Hashem Morvarid. Finean essence, local necessity, and pure logical properties. Synthese, 195(11):4997–5005, November 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1441-9;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1441-
9.pdf.

Butler:2018:ENS

[1804] Jesse M. Butler. An entirely non-self-referential Yabloesque paradox. Synthese, 195(11):5007–5019, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1443-7;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1443-
7.pdf.

Nutting:2018:ORS

[1805] Eileen S. Nutting. Ontological realism and sentential form. Synthese, 195(11):5021–5036, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1446-4;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1446-4.pdf.

Richmond:2018:TTH

[1806] Alasdair Richmond. Time travel, hyperspace and Cheshire Cats. Synthese, 195(11):5037–5058, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1448-2;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1448-2.pdf.

DAlessandro:2018:AST

[1807] William D’Alessandro. Arithmetic, set theory, reduction and explanation. Synthese, 195(11):5059–5089, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1450-8;

 http://link.springer.com/content/pdf/10.1007/s11229-017-1450-
8.pdf.

Longenecker:2018:NCP

[1808] Michael Tze-Sung Longenecker. Non-concrete parts of material objects. Synthese, 195(11):5091–5111, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1451-7;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1451-
7.pdf.

Politi:2018:ESR

[1809] Vincenzo Politi. Erratum to: Scientific revolutions, specialization and the discovery of the structure of DNA: toward a new picture of the development of the sciences. Synthese, 195(11):5113, November 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1394-z;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1394-z.pdf. See [1684].

Downey:2018:PPR

[1810] Adrian Downey. Predictive processing and the representation wars: a victory for the eliminativist (via fictionalism). Synthese, 195(12): 5115–5139, December 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1442-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1442-8.pdf.

Lyre:2018:SDM

[1811] Holger Lyre. Structures, dynamics and mechanisms in neuroscience: an integrative account. Synthese, 195(12):5141–5158, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1616-4;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1616-4.pdf.

Pernu:2018:MCN

[1812] Tuomas K. Pernu. Mental causation via neuroprosthetics? A critical analysis. Synthese, 195(12):5159–5174, December 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1713-z;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1713-
z.pdf.

Haas:2018:ESP

[1813] Julia Haas. An empirical solution to the puzzle of weakness of will. Synthese, 195(12):5175–5195, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1712-0;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1712-
0.pdf.

Heinzelmann:2018:DD

[1814] Nora Heinzelmann. Deontology defended. Synthese, 195(12):5197– 5216, December 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1762-3;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1762-3.pdf.

Gottlieb:2018:CLM

[1815] Joseph Gottlieb. Consciousness and the limits of memory. Synthese, 195(12):5217–5243, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-018-1793-9;   http://link.springer.com/
content/pdf/10.1007/s11229-018-1793-9.pdf.

Paz:2018:ECA

[1816] Abel Wajnerman Paz. An efficient coding approach to the debate on grounded cognition. Synthese, 195(12):5245–5269, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1815-7;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1815-7.pdf.

Bickle:2018:CEN

[1817] John Bickle and Aaron Kostko. Connection experiments in neurobiology. Synthese, 195(12):5271–5295, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1838-0;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1838-
0.pdf.

Gadsby:2018:AAA

[1818] Stephen Gadsby and Daniel Williams. Action, affordances, and anorexia: body representation and basic cognition. Synthese, 195(12):5297– 5317, December 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1843-3;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1843-3.pdf.

Elber-Dorozko:2018:MKW

[1819] Lotem Elber-Dorozko. Manipulation is key: on why non-mechanistic explanations in the cognitive sciences also describe relations of manipulation and control. Synthese, 195(12):5319–5337, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-018-01901-3;   http://link.springer.com/content/pdf/
10.1007/s11229-018-01901-3.pdf.

Buckner:2018:EMT

[1820] Cameron Buckner. Empiricism without magic: transformational abstraction in deep convolutional neural networks. Synthese, 195(12):5339–5372, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-018-01949-1;   http://link.springer.com/content/pdf/
10.1007/s11229-018-01949-1.pdf.

Haueis:2018:BCM

[1821] Philipp Haueis. Beyond cognitive myopia: a patchwork approach to the concept of neural function. Synthese, 195(12):5373–5402, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-018-01991-z;   http://link.springer.com/content/pdf/
10.1007/s11229-018-01991-z.pdf.

Morgan:2018:MAU

[1822] Alex Morgan. Mindless accuracy: on the ubiquity of content in nature. Synthese, 195(12):5403–5429, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-018-02011-w;   http://link.springer.
com/content/pdf/10.1007/s11229-018-02011-w.pdf.

Karaca:2018:LLH

[1823] Koray Karaca. Lessons from the Large Hadron Collider for model-based experimentation: the concept of a model of data acquisition and the scope of the hierarchy of models. Synthese, 195(12):5431–5452, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

017-1453-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1453-5.pdf.

Plakias:2018:RMM

[1824] Alexandra Plakias. The response model of moral disgust. Synthese, 195(12):5453–5472, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1455-3;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1455-3.pdf.

Dougherty:2018:WIE

[1825] John Dougherty. What inductive explanations could not be. Synthese, 195(12):5473–5483, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1457-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1457-1.pdf.

Pulcini:2018:PCL

[1826] Gabriele Pulcini and Achille C. Varzi. Paraconsistency in classical logic. Synthese, 195(12):5485–5496, December 2018. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1458-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1458-
0.pdf.

Goldwater:2018:PSC

[1827] Jonah Goldwater. Physicalism and the sortalist conception of objects. Synthese, 195(12):5497–5519, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1459-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1459-
z.pdf.

Griffin:2018:BB

[1828] Nicholas Griffin. Brandom and the brutes. Synthese, 195(12):5521– 5547, December 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1460-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1460-6.pdf.

Gray:2018:LRP

[1829] Aidan Gray. Lexical-rule predicativism about names. Synthese, 195(12): 5549–5569, December 2018. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1462-4;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1462-4.pdf.

Thicke:2018:ME

[1830] Michael Thicke. Market epistemology. Synthese, 195(12):5571–5594, December 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1464-2;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1464-2.pdf.

Muller:2019:IRP

[1831] Thomas Müller, Antje Rumberg, and Verena Wagner. An introduction to real possibilities, indeterminism, and free will: three contingencies of the debate. Synthese, 196(1):1–10, January 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1842-4;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1842-
4.pdf.

Placek:2019:LDT

[1832] Tomasz Placek. Laplace’s demon tries on Aristotle’s cloak: on two approaches to determinism. Synthese, 196(1):11–30, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1158-1;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1158-1.pdf.

Broersen:2019:ANE

[1833] Jan Broersen. Agents necessitating effects in Newtonian time and space: from power and opportunity to effectivity. Synthese, 196(1):31–68, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1769-9;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1769-9.pdf.

Ohrstrom:2019:CDP

[1834] Peter Øhrstrøm. A critical discussion of Prior’s philosophical and tense-logical analysis of the ideas of indeterminism and human freedom. Synthese, 196(1):69–85, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1149-2;

 http://link.springer.com/content/pdf/10.1007/s11229-016-1149-
2.pdf.

Haji:2019:PCF

[1835] Ishtiyaque Haji. A paradox concerning Frankfurt examples. Synthese, 196(1):87–103, January 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1025-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1025-0.pdf.

Mayr:2019:APA

[1836] Erasmus Mayr. Alternative possibilities and asymmetry. Synthese, 196 (1):105–125, January 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1503-z;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1503-z.pdf.

Ekstrom:2019:TPE

[1837] Laura W. Ekstrom. Toward a plausible event-causal indeterminist account of free will. Synthese, 196(1):127–144, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1143-8;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1143-8.pdf.

Kane:2019:CTF

[1838] Robert Kane. The complex tapestry of free will: striving will, indeterminism and volitional streams. Synthese, 196(1):145–160, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1046-8;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1046-8.pdf.

vanMiltenburg:2019:LPP

[1839] Niels van Miltenburg and Dawa Ometto. The libertarian predicament: a plea for action theory. Synthese, 196(1):161–178, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1148-3;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1148-3.pdf.

Groff:2019:SFW

[1840] Ruth Groff. Sublating the free will problematic: powers, agency and causal determination. Synthese, 196(1):179–200, January 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1124-y;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1124-
y.pdf.

Vetter:2019:AD

[1841] Barbara Vetter. Are abilities dispositions? Synthese, 196(1):201– 220, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1152-7;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1152-7.pdf.

Jonsson:2019:UAC

[1842] Martin L. Jönsson and Tomoji Shogenji. A unified account of the conjunction fallacy by coherence. Synthese, 196(1):221–237, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1467-z;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1467-z.pdf.

Maunu:2019:PII

[1843] Ari Maunu. The principle of the indiscernibility of identicals requires no restrictions. Synthese, 196(1):239–246, January 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1468-y;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1468-
y.pdf.

Barrantes:2019:ORE

[1844] Manuel Barrantes. Optimal representations and the Enhanced Indispensability Argument. Synthese, 196(1):247–263, January 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1470-4;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1470-
4.pdf.

Mundici:2019:FCP

[1845] Daniele Mundici. De Finetti coherence and the product law for independent events. Synthese, 196(1):265–271, January 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1471-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1471-
3.pdf.

Vidal:2019:DFS

[1846] Mathieu Vidal and Denis Perrin. A default-free solution to the imperfective paradox. Synthese, 196(1):273–297, January 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1474-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1474-
0.pdf.

Gao:2019:MPR

[1847] Shan Gao. The measurement problem revisited. Synthese, 196(1): 299–311, January 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1476-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1476-y.pdf. See reply [2081].

Levy:2019:DDD

[1848] Neil Levy. Due deference to denialism: explaining ordinary people’s rejection of established scientific findings. Synthese, 196(1):313–327, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1477-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1477-x.pdf.

Azzano:2019:QRP

[1849] Lorenzo Azzano. The question of realism for powers. Synthese, 196 (1):329–354, January 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1478-9;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1478-9.pdf.

Weaver:2019:HIS

[1850] Sara Weaver. The harms of ignoring the social nature of science. Synthese, 196(1):355–375, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1479-8;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1479-8.pdf.

Arfini:2019:OCV

[1851] Selene Arfini, Tommaso Bertolotti, and Lorenzo Magnani. Online communities as virtual cognitive niches. Synthese, 196(1):377–397, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1482-0;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1482-0.pdf.

Eva:2019:SCA

[1852] Benjamin Eva. A-symmetric confirmation and anthropic skepticism. Synthese, 196(1):399–412, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1486-9;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1486-9.pdf.

Bruner:2019:MDA

[1853] Justin P. Bruner. Minority (dis)advantage in population games. Synthese, 196(1):413–427, January 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1487-8;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1487-8.pdf.

Thrasher:2019:CRS

[1854] John Thrasher. Constructivism, representation, and stability: pathdependence in public reason theories of justice. Synthese, 196(1): 429–450, January 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1488-7;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1488-7.pdf.

Rowbottom:2019:SRW

[1855] Darrell P. Rowbottom. Scientific realism: what it is, the contemporary debate, and new directions. Synthese, 196(2):451–484, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1484-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1484-y.pdf.

Wolff:2019:NQS

[1856] Johanna Wolff. Naturalistic quietism or scientific realism? Synthese, 196 (2):485–498, February 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0873-3;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0873-3.pdf.

Nanay:2019:ERS

[1857] Bence Nanay. Entity realism and singularist semirealism. Synthese, 196 (2):499–517, February 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1179-9;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1179-9.pdf.

Elgin:2019:NRO

[1858] Catherine Z. Elgin. Nominalism, realism and objectivity. Synthese, 196 (2):519–534, February 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1114-0;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1114-0.pdf.

Frost-Arnold:2019:SHM

[1859] Greg Frost-Arnold. Should a historically motivated anti-realist be a Stanfordite? Synthese, 196(2):535–551, February 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1050-z;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1050-
z.pdf.

Wray:2019:DTR

[1860] K. Brad Wray. Discarded theories: the role of changing interests. Synthese, 196(2):553–569, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1058-4;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1058-4.pdf.

Vickers:2019:TRS

[1861] Peter Vickers. Towards a realistic success-to-truth inference for scientific realism. Synthese, 196(2):571–585, February 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1150-9;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1150-
9.pdf.

Asay:2019:GLD

[1862] Jamin Asay. Going local: a defense of methodological localism about scientific realism. Synthese, 196(2):587–609, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1072-6;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1072-6.pdf.

Saatsi:2019:WTP

[1863] Juha Saatsi. What is theoretical progress of science? Synthese, 196 (2):611–631, February 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1118-9;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1118-9.pdf.

Jaque:2019:MRU

[1864] Claudia Picazo Jaque. Are mental representations underdeterminacyfree? Synthese, 196(2):633–654, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1494-9;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1494-9.pdf.

Zahle:2019:WMI

[1865] Julie Zahle and Harold Kincaid. Why be a methodological individualist? Synthese, 196(2):655–675, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1523-8;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1523-
8.pdf.

Aydede:2019:EPT

[1866] Murat Aydede. Is the experience of pain transparent? Synthese, 196 (2):677–708, February 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1528-3;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1528-3.pdf.

Surovell:2019:SEE

[1867] Jonathan Reid Surovell. Stance empiricism and epistemic reason. Synthese, 196(2):709–733, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1539-0;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1539-0.pdf.

Bonevac:2019:FCR

[1868] Daniel Bonevac. Free choice reasons. Synthese, 196(2):735–760, February 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1540-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1540-7.pdf.

Bschir:2019:ISN

[1869] Karim Bschir, Simon Lohse, and Hasok Chang. Introduction: systematicity, the nature of science? Synthese, 196(3):761–773, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1685-z;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1685-z.pdf.

Mantzavinos:2019:NSD

[1870] C. Mantzavinos. The nature of science. A dialogue. Synthese, 196 (3):775–793, March 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1154-5;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1154-5.pdf.

Green:2019:SCS

[1871] Sara Green. Science and common sense: perspectives from philosophy and science education. Synthese, 196(3):795–818, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1276-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1276-9.pdf.

Wray:2019:SCT

[1872] K. Brad Wray. Systematicity and the Continuity Thesis. Synthese, 196(3):819–832, March 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1088-y;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1088-y.pdf.

Lyons:2019:STM

[1873] Timothy D. Lyons. Systematicity theory meets Socratic scientific realism: the systematic quest for truth. Synthese, 196(3):833–861, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

017-1561-2;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1561-2.pdf.

Bird:2019:SKB

[1874] Alexander Bird. Systematicity, knowledge, and bias. how systematicity made clinical medicine a science. Synthese, 196(3):863–879, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1342-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1342-y.pdf.

Oreskes:2019:SNS

[1875] Naomi Oreskes. Systematicity is necessary but not sufficient: on the problem of facsimile science. Synthese, 196(3):881–905, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1481-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1481-1.pdf.

Hoyningen-Huene:2019:R

[1876] Paul Hoyningen-Huene. Replies. Synthese, 196(3):907–928, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1741-8;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1741-8.pdf.

Schirn:2019:FPG

[1877] Matthias Schirn. Frege’s philosophy of geometry. Synthese, 196(3): 929–971, March 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1489-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1489-6.pdf.

Grinbaum:2019:EMP

[1878] Alexei Grinbaum. The effectiveness of mathematics in physics of the unknown. Synthese, 196(3):973–989, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1490-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1490-
0.pdf.

Ginnobili:2019:WDN

[1879] Santiago Ginnobili and Daniel Blanco. Wallace’s and Darwin’s natural selection theories. Synthese, 196(3):991–1017, March 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1491-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1491-
z.pdf.

Georgi:2019:PRT

[1880] Geoff Georgi. Propositions, representation, and truth. Synthese, 196 (3):1019–1043, March 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1492-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1492-y.pdf.

Fujimoto:2019:DBA

[1881] Kentaro Fujimoto. Deflationism beyond arithmetic. Synthese, 196(3): 1045–1069, March 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1495-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1495-8.pdf.

Barnett:2019:TDS

[1882] Zach Barnett. Tolerance and the distributed sorites. Synthese, 196 (3):1071–1077, March 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1499-4;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1499-4.pdf.

Vassend:2019:COE

[1883] Olav B. Vassend. Confirmation and the ordinal equivalence thesis. Synthese, 196(3):1079–1095, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1500-2;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1500-2.pdf.

Tesic:2019:CGN

[1884] Marko Tesić. Confirmation and the generalized Nagel–Schaffner model of reduction: a Bayesian analysis. Synthese, 196(3):1097–1129, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1501-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1501-1.pdf.

Hudetz:2019:SVT

[1885] Laurenz Hudetz. The semantic view of theories and higher-order languages. Synthese, 196(3):1131–1149, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1502-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1502-
0.pdf.

Johnson:2019:IIM

[1886] Casey Rebecca Johnson. Investigating illocutionary monism. Synthese, 196(3):1151–1165, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1508-7;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1508-7.pdf.

Hirvela:2019:GSH

[1887] Jaakko Hirvelä. Global safety: how to deal with necessary truths. Synthese, 196(3):1167–1186, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1511-z;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1511-z.pdf.

Haug:2019:NMC

[1888] Matthew C. Haug. No microphysical causation? no problem: selective causal skepticism and the structure of completeness-based arguments for physicalism. Synthese, 196(3):1187–1208, March 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1519-4;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1519-4.pdf.

Garcia-Carpintero:2019:IPV

[1889] Manuel García-Carpintero and Bjørn Jespersen. Introduction: primitivism versus reductionism about the problem of the unity of the proposition. Synthese, 196(4):1209–1224, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1727-6;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1727-
6.pdf.

Eklund:2019:RUF

[1890] Matti Eklund. Regress, unity, facts, and propositions. Synthese, 196 (4):1225–1247, April 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1155-4;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1155-4.pdf.

Duzi:2019:ISP

[1891] Marie Duzí. If structured propositions are logical procedures then how are procedures individuated? Synthese, 196(4):1249–1283, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1595-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1595-5.pdf.

Jespersen:2019:AP

[1892] Bjørn Jespersen. Anatomy of a proposition. Synthese, 196(4):1285–1324, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1512-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1512-y.pdf.

Gaskin:2019:UPL

[1893] Richard Gaskin. From the unity of the proposition to linguistic idealism. Synthese, 196(4):1325–1342, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-016-1081-5;
 http://link.springer.com/content/pdf/10.1007/s11229-016-1081-
5.pdf.

King:2019:PFG

[1894] Jeffrey C. King. On propositions and fineness of grain (again!). Synthese, 196(4):1343–1367, April 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1291-x;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1291-x.pdf.

Soames:2019:PCA

[1895] Scott Soames. Propositions as cognitive acts. Synthese, 196(4): 1369–1383, April 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1168-z;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1168-z.pdf.

Hanks:2019:C

[1896] Peter Hanks. On cancellation. Synthese, 196(4):1385–1402, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1260-4;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1260-4.pdf.

Recanati:2019:FC

[1897] François Recanati. Force cancellation. Synthese, 196(4):1403–1424, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1223-9;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1223-9.pdf.

Pickel:2019:UTT

[1898] Bryan Pickel. Unity through truth. Synthese, 196(4):1425–1452, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1279-6;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1279-6.pdf.

Speaks:2019:ATA

[1899] Jeff Speaks. Act theories and the attitudes. Synthese, 196(4):1453–1473, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1219-5;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1219-5.pdf.

Ostertag:2019:SPL

[1900] Gary Ostertag. Structured propositions and the logical form of predication. Synthese, 196(4):1475–1499, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1420-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1420-
1.pdf.

Pagin:2019:GAA

[1901] Peter Pagin. A general argument against structured propositions. Synthese, 196(4):1501–1528, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-016-1244-4;   http://link.springer.com/
content/pdf/10.1007/s11229-016-1244-4.pdf.

Keller:2019:WPS

[1902] Lorraine Juliano Keller. What propositional structure could not be. Synthese, 196(4):1529–1553, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1585-7;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1585-7.pdf.

Zakkou:2019:DRC

[1903] Julia Zakkou. Denial and retraction: a challenge for theories of taste predicates. Synthese, 196(4):1555–1573, April 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1520-y;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1520-
y.pdf.

Crawford:2019:BBD

[1904] Lindsay Crawford. Believing the best: on doxastic partiality in friendship. Synthese, 196(4):1575–1593, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1521-x;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1521-
x.pdf.

McNulty:2019:CCK

[1905] Michael Bennett McNulty. Continuity of change in Kant’s dynamics. Synthese, 196(4):1595–1622, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1527-4;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1527-4.pdf.

Castro:2019:IID

[1906] Clinton Castro and Casey Hart. The imprecise impermissivist’s dilemma. Synthese, 196(4):1623–1640, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1530-9;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1530-9.pdf.

Cote-Bouchard:2019:OIC

[1907] Charles Côté-Bouchard. ‘Ought’ implies ‘can’ against epistemic deontologism: beyond doxastic involuntarism. Synthese, 196(4):1641–1656, April 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

017-1531-8;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1531-8.pdf.

Fletcher:2019:IIS

[1908] Samuel C. Fletcher, Patricia Palacios, Laura Ruetsche, and Elay Shech. Infinite idealizations in science: an introduction. Synthese, 196(5):1657– 1669, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-018-02069-6;   http://link.springer.com/content/pdf/
10.1007/s11229-018-02069-6.pdf.

Nefdt:2019:IFL

[1909] Ryan M. Nefdt. Infinity and the foundations of linguistics. Synthese, 196(5):1671–1711, May 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1574-x;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1574-x.pdf.

Strevens:2019:SAI

[1910] Michael Strevens. The structure of asymptotic idealization. Synthese, 196(5):1713–1731, May 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1646-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1646-y.pdf.

Bianchi:2019:CFI

[1911] Silvia De Bianchi. Combining finite and infinite elements: Why do we use infinite idealizations in engineering? Synthese, 196(5):1733–1748, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1864-y;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1864-y.pdf.

Fletcher:2019:MAN

[1912] Samuel C. Fletcher. Minimal approximations and Norton’s dome. Synthese, 196(5):1749–1760, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-018-1676-0;   http://link.springer.com/
content/pdf/10.1007/s11229-018-1676-0.pdf.

Valente:2019:PRP

[1913] Giovanni Valente. On the paradox of reversible processes in thermodynamics. Synthese, 196(5):1761–1781, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1560-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1560-
3.pdf.

Robertson:2019:SSE

[1914] Katie Robertson. Stars and steam engines: To what extent do thermodynamics and statistical mechanics apply to self-gravitating systems? Synthese, 196(5):1783–1808, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-018-02032-5;   http://link.springer.
com/content/pdf/10.1007/s11229-018-02032-5.pdf.

Wayne:2019:PPE

[1915] Andrew Wayne. Point-particle explanations: the case of gravitational waves. Synthese, 196(5):1809–1829, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1638-y;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1638-
y.pdf.

Feintzeig:2019:DDI

[1916] Benjamin H. Feintzeig. Deduction and definability in infinite statistical systems. Synthese, 196(5):1831–1861, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1497-6;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1497-

6.pdf. See correction [2080].

vanWierst:2019:PPT

[1917] Pauline van Wierst. The paradox of phase transitions in the light of constructive mathematics. Synthese, 196(5):1863–1884, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1637-z;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1637-z.pdf.

Liu:2019:IIC

[1918] Chuang Liu. Infinite idealization and contextual realism. Synthese, 196(5):1885–1918, May 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1767-y;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1767-y.pdf.

Bangu:2019:DSD

[1919] Sorin Bangu. Discontinuities and singularities, data and phenomena: for referentialism. Synthese, 196(5):1919–1937, May 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1747-2;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1747-
2.pdf.

Baron:2019:ILE

[1920] Sam Baron. Infinite lies and explanatory ties: idealization in phase transitions. Synthese, 196(5):1939–1961, May 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1678-y;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1678-
y.pdf.

Shech:2019:IIE

[1921] Elay Shech. Infinitesimal idealization, easy road nominalism, and fractional quantum statistics. Synthese, 196(5):1963–1990, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1680-4;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1680-4.pdf.

Earman:2019:RIA

[1922] John Earman. The role of idealizations in the Aharonov–Bohm effect. Synthese, 196(5):1991–2019, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1522-9;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1522-9.pdf.

Pinillos:2019:KPA

[1923] N. Ángel Pinillos. Knowledge and the permissibility of action. Synthese, 196(5):2021–2043, May 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1567-9;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1567-9.pdf.

Zieba:2019:NRA

[1924] Paweł Jakub Zieba. Naïve realism about unconscious perception. Synthese, 196(5):2045–2073, May 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1570-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1570-1.pdf.

Madison:2019:OMT

[1925] B. J. C. Madison. Is open-mindedness truth-conducive? Synthese, 196 (5):2075–2087, May 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1571-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1571-0.pdf.

Tarasov:2019:SRW

[1926] Leonid Tarasov. Semantic relativism and ways of knowing. Synthese, 196(5):2089–2109, May 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1572-z;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1572-z.pdf.

Tekin:2019:BMB

[1927] Serife Tekin and Edouard Machery. Beyond mind-body dualism: embracing pluralism in psychiatric research- introduction to the special issue, ”psychiatry and its philosophy”. Synthese, 196(6):2111–2115, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1836-2;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1836-2.pdf.

Lemoine:2019:NRP

[1928] Maël Lemoine. On the neurobiological redefinition of psychiatric symptoms: elimination, reduction, or what? Synthese, 196(6):2117–2133, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1270-2;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1270-2.pdf.

Robins:2019:CCM

[1929] Sarah K. Robins. Confabulation and constructive memory. Synthese, 196 (6):2135–2151, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1315-1;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1315-1.pdf.

Stinson:2019:ABP

[1930] Catherine Stinson. The absent body in psychiatric diagnosis, treatment, and research. Synthese, 196(6):2153–2176, June 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1507-8;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1507-
8.pdf.

Tabb:2019:PPA

[1931] Kathryn Tabb. Philosophy of psychiatry after diagnostic kinds. Synthese, 196(6):2177–2195, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1659-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1659-6.pdf.

Tekin:2019:MSS

[1932] Serife Tekin. The missing self in scientific psychiatry. Synthese, 196 (6):2197–2215, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1324-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1324-0.pdf.

Hoffman:2019:CIP

[1933] Ginger A. Hoffman. Collectively ill: a preliminary case that groups can have psychiatric disorders. Synthese, 196(6):2217–2241, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1379-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1379-y.pdf.

Dentith:2019:CTB

[1934] M R. X. Dentith. Conspiracy theories on the basis of the evidence. Synthese, 196(6):2243–2261, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1532-7;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1532-7.pdf.

Rosa:2019:RR

[1935] Luis Rosa. Reasoning without regress. Synthese, 196(6):2263–2278, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1535-4;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1535-4.pdf.

Fallis:2019:TFA

[1936] Don Fallis and Peter J. Lewis. Toward a formal analysis of deceptive signaling. Synthese, 196(6):2279–2303, June 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1536-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1536-
3.pdf.

Boesch:2019:MEA

[1937] Brandon Boesch. The means-end account of scientific, representational actions. Synthese, 196(6):2305–2322, June 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1537-2;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1537-
2.pdf.

Weinberger:2019:MME

[1938] Naftali Weinberger. Mechanisms without mechanistic explanation. Synthese, 196(6):2323–2340, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1538-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1538-1.pdf.

Eronen:2019:RRL

[1939] Markus I. Eronen. Robust realism for the life sciences. Synthese, 196 (6):2341–2354, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1542-5;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1542-5.pdf.

Cahen:2019:NAR

[1940] Arnon Cahen. Nonconceptual apprehension and the reason-giving character of perception. Synthese, 196(6):2355–2383, June 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1543-4;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1543-
4.pdf.

Copeland:2019:SSD

[1941] Samantha Copeland. On serendipity in science: discovery at the intersection of chance and wisdom. Synthese, 196(6):2385–2406, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1544-3;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1544-3.pdf.

Jantzen:2019:KPL

[1942] Benjamin C. Jantzen. Kinds of process and the levels of selection. Synthese, 196(6):2407–2433, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1546-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1546-1.pdf.

Almotahari:2019:SDD

[1943] Mahrad Almotahari. Semantic deflationism deflated. Synthese, 196 (6):2435–2454, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1547-0;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1547-0.pdf.

Vasudevan:2019:BIE

[1944] Anubav Vasudevan. Biased information and the exchange paradox. Synthese, 196(6):2455–2485, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1550-5;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1550-5.pdf.

Scharp:2019:IM

[1945] Kevin Scharp. On the indeterminacy of the meter. Synthese, 196 (6):2487–2517, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1552-3;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1552-3.pdf.

Surovell:2019:ESP

[1946] Jonathan Surovell. Empirical significance, predictive power, and explication. Synthese, 196(6):2519–2539, June 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1554-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1554-
1.pdf.

Whittle:2019:CSR

[1947] Bruno Whittle. Correction to: Self-referential propositions. Synthese, 196(6):2541, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1714-y;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1714-y.pdf. See [1573].

vanDitmarsch:2019:CDL

[1948] Hans van Ditmarsch. Correction to: Dynamics of lying. Synthese, 196(6):2543, June 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1888-3;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1888-3.pdf. See [771].

Koslow:2019:MNE

[1949] Arnold Koslow. The modality and non-extensionality of the quantifiers. Synthese, 196(7):2545–2554, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-014-0539-6;   http://link.springer.com/
content/pdf/10.1007/s11229-014-0539-6.pdf.

Cook:2019:PPA

[1950] Roy T. Cook. Possible predicates and actual properties. Synthese, 196 (7):2555–2582, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-014-0592-1;   http://link.springer.com/content/
pdf/10.1007/s11229-014-0592-1.pdf.

Novaes:2019:AAF

[1951] Catarina Dutilh Novaes. Axiomatizations of arithmetic and the first-order/second-order divide. Synthese, 196(7):2583–2597, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
014-0636-6;   http://link.springer.com/content/pdf/10.1007/
s11229-014-0636-6.pdf.

Azzouni:2019:CML

[1952] Jody Azzouni. The challenge of many logics: a new approach to evaluating the role of ideology in Quinean commitment. Synthese, 196 (7):2599–2619, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0657-9;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0657-9.pdf.

Paseau:2019:MIR

[1953] A. C. Paseau. A measure of inferential-role preservation. Synthese, 196(7):2621–2642, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0705-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0705-5.pdf.

Hale:2019:SOL

[1954] Bob Hale. Second-order logic: properties, semantics, and existential commitments. Synthese, 196(7):2643–2669, July 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0764-7;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0764-
7.pdf.

Cohen:2019:MGA

[1955] Liron Cohen and Arnon Avron. The middle ground-ancestral logic. Synthese, 196(7):2671–2693, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-015-0784-3;   http://link.springer.com/
content/pdf/10.1007/s11229-015-0784-3.pdf.

Kyriacou:2019:EDM

[1956] Christos Kyriacou. Evolutionary debunking: the Milvian Bridge destabilized. Synthese, 196(7):2695–2713, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1555-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1555-
0.pdf.

Larvor:2019:EGK

[1957] Brendan Larvor. From Euclidean geometry to knots and nets. Synthese, 196(7):2715–2736, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1558-x;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1558-x.pdf.

Pettigrew:2019:AIA

[1958] Richard Pettigrew. Aggregating incoherent agents who disagree. Synthese, 196(7):2737–2776, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1613-7;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1613-7.pdf.

Norton:2019:PTM

[1959] James Norton and Kristie Miller. A psychologistic theory of metaphysical explanation. Synthese, 196(7):2777–2802, July 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1566-x;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1566-
x.pdf.

Ventura:2019:ASP

[1960] Rafael Ventura. Ambiguous signals, partial beliefs, and propositional content. Synthese, 196(7):2803–2820, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1580-z;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1580-
z.pdf.

Westra:2019:STM

[1961] Evan Westra. Stereotypes, theory of mind, and the action-prediction hierarchy. Synthese, 196(7):2821–2846, July 2019. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/s11229-017-1575-9;   http://link.
springer.com/content/pdf/10.1007/s11229-017-1575-9.pdf.

Kiverstein:2019:FGN

[1962] Julian Kiverstein, Mark Miller, and Erik Rietveld. The feeling of grip: novelty, error dynamics, and the predictive brain. Synthese, 196(7): 2847–2869, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1583-9;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1583-9.pdf.

Shaw:2019:BCP

[1963] Kegan J. Shaw. The bifurcated conception of perceptual knowledge: a new solution to the basis problem for epistemological disjunctivism. Synthese, 196(7):2871–2884, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1590-x;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1590-
x.pdf.

Lundgren:2019:DSI

[1964] Björn Lundgren. Does semantic information need to be truthful? Synthese, 196(7):2885–2906, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1587-5;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1587-5.pdf.

Riccardi:2019:PPA

[1965] Mattia Riccardi. Perceptual presence: an attentional account. Synthese, 196(7):2907–2926, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1588-4;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1588-4.pdf.

vanDitmarsch:2019:AET

[1966] Hans van Ditmarsch, Sophia Knight, and Aybüke Özgün. Announcement as effort on topological spaces. Synthese, 196(7):2927–2969, July 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1592-8;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1592-8.pdf.

Roski:2019:BKG

[1967] Stefan Roski. Bolzano and Kim on grounding and unification. Synthese, 196(7):2971–2999, July 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1593-7;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1593-7.pdf.

Claveau:2019:VET

[1968] François Claveau and Olivier Grenier. The variety-of-evidence thesis: a Bayesian exploration of its surprising failures. Synthese, 196(8): 3001–3028, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1607-5;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1607-5.pdf.

Mayo-Wilson:2019:CIP

[1969] Conor Mayo-Wilson. Causal identifiability and piecemeal experimentation. Synthese, 196(8):3029–3065, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1826-4;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1826-
4.pdf.

Heesen:2019:VMT

[1970] Remco Heesen, Liam Kofi Bright, and Andrew Zucker. Vindicating methodological triangulation. Synthese, 196(8):3067–3081, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1294-7;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1294-7.pdf.

Wuthrich:2019:PEA

[1971] Nicolas Wüthrich and Katie Steele. The problem of evaluating automated large-scale evidence aggregators. Synthese, 196(8):3083–3102, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1627-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1627-1.pdf.

Reiss:2019:AEV

[1972] Julian Reiss. Against external validity. Synthese, 196(8):3103–3121, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1796-6;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1796-6.pdf.

Frank:2019:ESQ

[1973] David M. Frank. Ethics of the scientist qua policy advisor: inductive risk, uncertainty, and catastrophe in climate economics. Synthese, 196 (8):3123–3138, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1617-3;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1617-3.pdf.

Vieland:2019:NEA

[1974] Veronica J. Vieland and Hasok Chang. No evidence amalgamation without evidence measurement. Synthese, 196(8):3139–3161, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1666-7;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1666-7.pdf.

Fletcher:2019:EAS

[1975] Samuel C. Fletcher, Jürgen Landes, and Roland Poellinger. Evidence amalgamation in the sciences: an introduction. Synthese, 196(8):

3163–3188, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1840-6;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1840-6.pdf.

Holman:2019:DMA

[1976] Bennett Holman. In defense of meta-analysis. Synthese, 196(8): 3189–3211, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1690-2;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1690-2.pdf.

Danks:2019:AED

[1977] David Danks and Sergey Plis. Amalgamating evidence of dynamics. Synthese, 196(8):3213–3230, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1568-8;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1568-8.pdf.

Baetu:2019:PEM

[1978] Tudor M. Baetu. On pain experience, multidisciplinary integration and the level-laden conception of science. Synthese, 196(8):3231–3250, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1429-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1429-5.pdf.

Wilde:2019:ERW

[1979] Michael Wilde and Veli-Pekka Parkkinen. Extrapolation and the Russo– Williamson thesis. Synthese, 196(8):3251–3262, August 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1573-y;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1573-
y.pdf.

Kao:2019:UBJ

[1980] Molly Kao. Unification beyond justification: a strategy for theory development. Synthese, 196(8):3263–3278, August 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1515-8;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1515-
8.pdf.

Bertolaso:2019:EAP

[1981] Marta Bertolaso and Fabio Sterpetti. Evidence amalgamation, plausibility, and cancer research. Synthese, 196(8):3279–3317, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1591-9;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1591-9.pdf.

Kalkman:2019:NPD

[1982] David Kalkman. New problems for defining animal communication in informational terms. Synthese, 196(8):3319–3336, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1598-2;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1598-2.pdf.

Zerilli:2019:MRC

[1983] John Zerilli. Multiple realization and the commensurability of taxonomies. Synthese, 196(8):3337–3353, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1599-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1599-
1.pdf.

Thorpe:2019:RIS

[1984] Joshua Rowan Thorpe. Radical interpretation, scepticism, and the possibility of shared error. Synthese, 196(8):3355–3368, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1600-z;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1600-z.pdf.

Deasy:2019:TAA

[1985] Daniel Deasy. The triviality argument against presentism. Synthese, 196 (8):3369–3388, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1601-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1601-y.pdf.

Kim:2019:BCO

[1986] Namjoong Kim. Bad company objection to Joongol Kim’s adverbial theory of numbers. Synthese, 196(8):3389–3407, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1602-x;   http://link.springer.com/content/pdf/10.1007/

s11229-017-1602-x.pdf. See [572, 2075] and correction [2075].

Dimmock:2019:KBE

[1987] Paul Dimmock. Knowledge, belief, and egocentric bias. Synthese, 196 (8):3409–3432, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1603-9;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1603-9.pdf.

Elliott:2019:IWP

[1988] Edward Elliott. Impossible worlds and partial belief. Synthese, 196 (8):3433–3458, August 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1604-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1604-8.pdf.

Burnston:2019:CCP

[1989] Daniel C. Burnston. Correction to: Cognitive penetration and the cognition–perception interface. Synthese, 196(8):3459, August 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-018-01967-z;   http://link.springer.com/content/pdf/
10.1007/s11229-018-01967-z.pdf. See [1506].

Demeter:2019:UAM

[1990] Tamás Demeter and Eric Schliesser. The uses and abuses of mathematics in early modern philosophy: introduction. Synthese, 196(9): 3461–3464, September 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1670-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1670-y.pdf.

Wilson:2019:WIL

[1991] Mark Wilson. What I’ve learned from the early moderns. Synthese, 196(9):3465–3481, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1361-8;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1361-8.pdf.

Nelson:2019:DLU

[1992] Alan Nelson. Descartes on the limited usefulness of mathematics. Synthese, 196(9):3483–3504, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1328-9;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1328-9.pdf.

Domski:2019:IMM

[1993] Mary Domski. Imagination, metaphysics, mathematics: Descartes’s arguments for the Vortex Hypothesis. Synthese, 196(9):3505–3526, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1533-6;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1533-6.pdf.

Peterman:2019:EVS

[1994] Alison Peterman. Empress vs. Spider-Man: Margaret Cavendish on pure and applied mathematics. Synthese, 196(9):3527–3549, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1504-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1504-y.pdf.

Shaheen:2019:PND

[1995] Jonathan L. Shaheen. Part of nature and division in Margaret Cavendish’s materialism. Synthese, 196(9):3551–3575, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1326-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1326-y.pdf.

Zemplen:2019:DCA

[1996] Gábor Áron Zemplén. Diagrammatic carriers and the acceptance of Newton’s optical theory. Synthese, 196(9):3577–3593, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1356-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1356-5.pdf.

Ducheyne:2019:CMI

[1997] Steffen Ducheyne. Constraining (mathematical) imagination by experience: Nieuwentijt and van Musschenbroek on the abuses of mathematics. Synthese, 196(9):3595–3613, September 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1392-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1392-
1.pdf.

Demeter:2019:HSC

[1998] Tamás Demeter. Hume on the social construction of mathematical knowledge. Synthese, 196(9):3615–3631, September 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1655-x;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1655-
x.pdf.

Wolfe:2019:VAM

[1999] Charles T. Wolfe. Vital anti-mathematicism and the ontology of the emerging life sciences: from Mandeville to Diderot. Synthese, 196(9): 3633–3654, September 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1350-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1350-y.pdf.

Peterson:2019:PAD

[2000] Jared Peterson. A puzzle about desire. Synthese, 196(9):3655–3676, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1606-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1606-6.pdf.

Sandgren:2019:MTI

[2001] Alexander Sandgren. A metarepresentational theory of intentional identity. Synthese, 196(9):3677–3695, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1609-3;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1609-
3.pdf.

Juhl:2019:EIH

[2002] Cory Juhl and Brian Knab. On the empirical inaccessibility of higher-level modality and its significance for cosmological finetuning. Synthese, 196(9):3697–3710, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1610-x;

 http://link.springer.com/content/pdf/10.1007/s11229-017-1610-
x.pdf.

Koberinski:2019:DEC

[2003] Adam Koberinski, Lucas Dunlap, and William L. Harper. Do the EPR correlations pose a problem for causal decision theory? Synthese, 196(9): 3711–3722, September 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1611-9;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1611-9.pdf.

Longenecker:2019:EAR

[2004] Michael Tze-Sung Longenecker. Existentialism, aliens and referentially unrestricted worlds. Synthese, 196(9):3723–3738, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1618-2;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1618-2.pdf.

Casati:2019:NMI

[2005] Filippo Casati and Naoya Fujikawa. Nothingness, Meinongianism and inconsistent mereology. Synthese, 196(9):3739–3772, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1619-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1619-1.pdf.

Backes:2019:BPC

[2006] Marvin Backes. A bitter pill for closure. Synthese, 196(9):3773– 3787, September 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1620-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1620-8.pdf.

Carnielli:2019:EAP

[2007] Walter Carnielli and Abilio Rodrigues. An epistemic approach to paraconsistency: a logic of evidence and truth. Synthese, 196(9):3789– 3813, September 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1621-7;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1621-7.pdf.

Gibbs:2019:BB

[2008] Cameron Gibbs. Basing for the Bayesian. Synthese, 196(9):3815– 3840, September 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1622-6;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1622-6.pdf.

Polonioli:2019:PMB

[2009] Andrea Polonioli. A plea for minimally biased naturalistic philosophy. Synthese, 196(9):3841–3867, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1628-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1628-
0.pdf.

Oakley:2019:RAA

[2010] Tim Oakley. The reductio argument against epistemic infinitism. Synthese, 196(9):3869–3887, September 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1629-z;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1629-z.pdf.

Cunningham:2019:BNR

[2011] J. J. Cunningham. Is believing for a normative reason a composite condition? Synthese, 196(9):3889–3910, September 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1630-6;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1630-
6.pdf.

Bhakthavatsalam:2019:SRU

[2012] Sindhuja Bhakthavatsalam and Ian James Kidd. Science, realism, and unconceived alternatives: introduction to the special issue on unconceived alternatives. Synthese, 196(10):3911–3913, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-019-02353-z;   http://link.springer.com/content/pdf/
10.1007/s11229-019-02353-z.pdf.

Stanford:2019:UAC

[2013] P. Kyle Stanford. Unconceived alternatives and conservatism in science: the impact of professionalization, peer-review, and Big Science. Synthese, 196(10):3915–3932, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0856-4;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0856-
4.pdf.

Ruhmkorff:2019:UAC

[2014] Samuel Ruhmkorff. Unconceived alternatives and the cathedral problem. Synthese, 196(10):3933–3945, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-015-0947-2;
 http://link.springer.com/content/pdf/10.1007/s11229-015-0947-
2.pdf.

Rowbottom:2019:EAU

[2015] Darrell P. Rowbottom. Extending the argument from unconceived alternatives: observations, models, predictions, explanations, methods, instruments, experiments, and values. Synthese, 196(10):3947–3959, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1132-y;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1132-y.pdf.

Bonilla:2019:RVA

[2016] Jesús P. Zamora Bonilla. Realism versus anti-realism: philosophical problem or scientific concern? Synthese, 196(10):3961–3977, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0988-6;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0988-6.pdf.

Saatsi:2019:HIO

[2017] Juha Saatsi. Historical inductions, old and new. Synthese, 196(10): 3979–3993, October 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-015-0855-5;   http://link.springer.com/content/
pdf/10.1007/s11229-015-0855-5.pdf.

Kinney:2019:IEG

[2018] David Kinney. Inductive explanation and Garber-style solutions to the problem of old evidence. Synthese, 196(10):3995–4009, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-

017-1634-2;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1634-2.pdf.

Carter:2019:EFD

[2019] Jessica Carter. Exploring the fruitfulness of diagrams in mathematics. Synthese, 196(10):4011–4032, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1635-1;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1635-
1.pdf.

Skipper:2019:GDB

[2020] Mattias Skipper and Asbjørn Steglich-Petersen. Group disagreement: a belief aggregation perspective. Synthese, 196(10):4033–4058, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1636-0;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1636-0.pdf.

Hughes:2019:DE

[2021] Nick Hughes. Dilemmic epistemology. Synthese, 196(10):4059–4090, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1639-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1639-x.pdf.

Bartunek:2019:TI

[2022] Nicoletta Bartunek. Truth in the investigations. Synthese, 196(10): 4091–4111, October 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1643-1;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1643-1.pdf.

Davies:2019:TTS

[2023] James E. Davies. Towards a theory of singular thought about abstract mathematical objects. Synthese, 196(10):4113–4136, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1644-0;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1644-0.pdf.

Wenmackers:2019:SWP

[2024] Sylvia Wenmackers. The Snow White problem. Synthese, 196(10): 4137–4153, October 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1647-x;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1647-x.pdf.

Hommen:2019:OCF

[2025] David Hommen. Ontological commitments of frame-based knowledge representations. Synthese, 196(10):4155–4183, October 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1649-8;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1649-

8.pdf. See [2076].

Mazzola:2019:GRC

[2026] Claudio Mazzola. Generalised Reichenbachian common cause systems. Synthese, 196(10):4185–4209, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-017-1650-2;   http://link.springer.
com/content/pdf/10.1007/s11229-017-1650-2.pdf.

Nissan-Rozen:2019:PAA

[2027] Ittay Nissan-Rozen and Levi Spectre. A pragmatic argument against equal weighting. Synthese, 196(10):4211–4227, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1651-1;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1651-1.pdf.

Bowker:2019:SBM

[2028] Mark Bowker. Saying a bundle: meaning, intention, and underdetermination. Synthese, 196(10):4229–4252, October 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1652-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1652-
0.pdf.

Binney:2019:MPM

[2029] Nicholas Binney. Meno’s paradox and medicine. Synthese, 196(10): 4253–4278, October 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1654-y;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1654-y.pdf.

Millar:2019:SO

[2030] Becky Millar. Smelling objects. Synthese, 196(10):4279–4303, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1657-8;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1657-8.pdf.

Douskos:2019:SSI

[2031] Christos Douskos. The spontaneousness of skill and the impulsivity of habit. Synthese, 196(10):4305–4328, October 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1658-7;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1658-
7.pdf.

Peterson:2019:APR

[2032] Clayton Peterson. Accommodation, prediction and replication: model selection in scale construction. Synthese, 196(10):4329–4350, October 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1660-0;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1660-0.pdf.

Holman:2019:MKS

[2033] Bennett Holman, Sven Bernecker, and Luciana Garbayo. Medical knowledge in a social world: Introduction to the special issue. Synthese, 196(11):4351–4361, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1779-7;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1779-
7.pdf.

Holman:2019:PD

[2034] Bennett Holman. Philosophers on drugs. Synthese, 196(11):4363– 4390, November 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1642-2;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1642-2.pdf.

Robinson:2019:FEN

[2035] Mark D. Robinson. Financializing epistemic norms in contemporary biomedical innovation. Synthese, 196(11):4391–4407, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1704-0;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1704-0.pdf.

Kukla:2019:IER

[2036] Rebecca Kukla. Infertility, epistemic risk, and disease definitions. Synthese, 196(11):4409–4428, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1405-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1405-
0.pdf.

Merrick:2019:IDC

[2037] Teri Merrick. From ‘intersex’ to ‘DSD’: a case of epistemic injustice. Synthese, 196(11):4429–4447, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1327-x;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1327-
x.pdf.

Ben-Moshe:2019:IMM

[2038] Nir Ben-Moshe. The internal morality of medicine: a constructivist approach. Synthese, 196(11):4449–4467, November 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1466-0;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1466-
0.pdf.

vanBaalen:2019:ITC

[2039] Sophie van Baalen and Annamaria Carusi. Implicit trust in clinical decision-making by multidisciplinary teams. Synthese, 196(11):4469– 4492, November 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1475-z;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1475-z.pdf.

Brun:2019:LEL

[2040] Georg Brun. Logical expressivism, logical theory and the critique of inferences. Synthese, 196(11):4493–4509, November 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1662-y;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1662-
y.pdf.

Boespflug:2019:WRW

[2041] Mark Boespflug. Why Reid was no dogmatist. Synthese, 196(11): 4511–4525, November 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1663-x;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1663-x.pdf.

Vernazzani:2019:SSE

[2042] Alfredo Vernazzani. The structure of sensorimotor explanation. Synthese, 196(11):4527–4553, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1664-9;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1664-9.pdf.

Montevil:2019:PSN

[2043] Maël Montévil. Possibility spaces and the notion of novelty: from music to biology. Synthese, 196(11):4555–4581, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1668-5;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1668-5.pdf.

Lawler:2019:UWK

[2044] Insa Lawler. Understanding why, knowing why, and cognitive achievements. Synthese, 196(11):4583–4603, November 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1672-9;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1672-
9.pdf.

Tee:2019:CRM

[2045] Sim-Hui Tee. Constructing reality with models. Synthese, 196(11): 4605–4622, November 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1673-8;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1673-8.pdf.

Halimi:2019:SMM

[2046] Brice Halimi. Settings and misunderstandings in mathematics. Synthese, 196(11):4623–4656, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1671-x;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1671-x.pdf.

Asghari:2019:EAH

[2047] Amir Asghari. Equivalence: an attempt at a history of the idea. Synthese, 196(11):4657–4677, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1674-2;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1674-
2.pdf.

Simion:2019:PEW

[2048] Mona Simion. A puzzle for epistemic WAMs. Synthese, 196(11):4679– 4689, November 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1681-3;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1681-3.pdf.

Petersen:2019:CCN

[2049] Esben Nedenskov Petersen. A case for a certainty norm of assertion. Synthese, 196(11):4691–4710, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1682-2;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1682-
2.pdf.

Baxendale:2019:MCR

[2050] Matthew Baxendale. Mapping the continuum of research strategies. Synthese, 196(11):4711–4733, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1683-1;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1683-
1.pdf.

Rice:2019:ESP

[2051] Collin Rice, Yasha Rohwer, and André Ariew. Explanatory schema and the process of model building. Synthese, 196(11):4735–4757, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1686-y;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1686-y.pdf.

Hlobil:2019:FNV

[2052] Ulf Hlobil. Faithfulness for naive validity. Synthese, 196(11):4759– 4774, November 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1687-x;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1687-x.pdf.

Lennertz:2019:MBA

[2053] Benjamin Lennertz. Might-beliefs and asymmetric disagreement. Synthese, 196(11):4775–4805, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-018-1688-9;   http://link.springer.
com/content/pdf/10.1007/s11229-018-1688-9.pdf.

Ciardelli:2019:CQI

[2054] Ivano Ciardelli. Correction to: Questions as information types. Synthese, 196(11):4807, November 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-019-02321-7;   http://link.springer.
com/content/pdf/10.1007/s11229-019-02321-7.pdf. See [1592].

Salice:2019:TAG

[2055] Alessandro Salice, John Michael, and András Szigeti. Thinking (about) groups: a special issue of Synthese. Synthese, 196(12):4809–4812, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1597-3;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1597-3.pdf.

Seemann:2019:RTJ

[2056] Axel Seemann. Reminiscing together: joint experiences, epistemic groups, and sense of self. Synthese, 196(12):4813–4828, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
016-1156-3;   http://link.springer.com/content/pdf/10.1007/
s11229-016-1156-3.pdf.

Thomasson:2019:OSG

[2057] Amie L. Thomasson. The ontology of social groups. Synthese, 196(12): 4829–4845, December 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-016-1185-y;   http://link.springer.com/content/
pdf/10.1007/s11229-016-1185-y.pdf.

Leon:2019:ESE

[2058] Felipe León, Thomas Szanto, and Dan Zahavi. Emotional sharing and the extended mind. Synthese, 196(12):4847–4867, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1351-x;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1351-x.pdf.

Rovane:2019:GAS

[2059] Carol Rovane. Is group agency a social phenomenon? Synthese, 196(12): 4869–4898, December 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-017-1384-1;   http://link.springer.com/content/
pdf/10.1007/s11229-017-1384-1.pdf.

Epstein:2019:WSG

[2060] Brian Epstein. What are social groups? Their metaphysics and how to classify them. Synthese, 196(12):4899–4932, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
017-1387-y;   http://link.springer.com/content/pdf/10.1007/
s11229-017-1387-y.pdf.

Michaelian:2019:CMT

[2061] Kourken Michaelian and John Sutton. Collective mental time travel: remembering the past and imagining the future together. Synthese, 196(12):4933–4960, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/s11229-017-1449-1;   http://link.springer.com/
content/pdf/10.1007/s11229-017-1449-1.pdf.

Wahlberg:2019:WSS

[2062] Tobias Hansson Wahlberg. Why the social sciences are irreducible. Synthese, 196(12):4961–4987, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-017-1472-2;
 http://link.springer.com/content/pdf/10.1007/s11229-017-1472-
2.pdf.

Carter:2019:KVS

[2063] J. Adam Carter and Robin McKenna. Kornblith versus Sosa on grades of knowledge. Synthese, 196(12):4989–5007, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1689-8;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1689-8.pdf.

Pittard:2019:FDL

[2064] John Pittard. Fundamental disagreements and the limits of instrumentalism. Synthese, 196(12):5009–5038, December 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1691-1;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1691-
1.pdf.

Verhaegh:2019:SOQ

[2065] Sander Verhaegh. Sign and object: Quine’s forgotten book project. Synthese, 196(12):5039–5060, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/s11229-018-1693-z;   http://link.springer.
com/content/pdf/10.1007/s11229-018-1693-z.pdf.

Friend:2019:CPC

[2066] Toby Friend. Can parts cause their wholes? Synthese, 196(12):5061– 5082, December 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1694-y;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1694-y.pdf.

Baltimore:2019:EVM

[2067] Joseph A. Baltimore. Expanding the vector model for dispositionalist approaches to causation. Synthese, 196(12):5083–5098, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1695-x;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1695-x.pdf.

Stoutenburg:2019:DEP

[2068] Gregory Stoutenburg. In defense of an epistemic probability account of luck. Synthese, 196(12):5099–5113, December 2019. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1699-6;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1699-
6.pdf.

Gallegos:2019:MSE

[2069] Sergio A. Gallegos. Models as signs: extending Kralemann and Lattman’s proposal on modeling models within Peirce’s theory of signs. Synthese, 196(12):5115–5136, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/s11229-018-1700-4;
 http://link.springer.com/content/pdf/10.1007/s11229-018-1700-
4.pdf.

Zylstra:2019:EG

[2070] Justin Zylstra. The essence of grounding. Synthese, 196(12):5137– 5152, December 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1701-3;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1701-3.pdf.

Fileva:2019:JAA

[2071] Iskra Fileva and Linda A. W. Brakel. Just another article on Moore’s paradox, but we don’t believe that. Synthese, 196(12):5153–5167, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1703-1;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1703-1.pdf.

DeHaro:2019:HFD

[2072] Sebastian De Haro. The heuristic function of duality. Synthese, 196(12): 5169–5203, December 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1708-9;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1708-9.pdf.

Billon:2019:PH

[2073] Alexandre Billon. Paradoxical hypodoxes. Synthese, 196(12):5205– 5229, December 2019. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-018-1711-1;   http://link.springer.com/content/
pdf/10.1007/s11229-018-1711-1.pdf.

Bruineberg:2019:GEI

[2074] Jelle Bruineberg, Anthony Chemero, and Erik Rietveld. General ecological information supports engagement with affordances for ‘higher’ cognition. Synthese, 196(12):5231–5251, December 2019. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1716-9;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1716-9.pdf.

Kim:2020:CBC

[2075] Namjoong Kim. Correction to: Bad company objection to Joongol Kim’s adverbial theory of numbers. Synthese, 197(3):1379, March 2020. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1738-3;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1738-3.pdf. See [1986].

Hommen:2020:COC

[2076] David Hommen. Correction to: Ontological commitments of framebased knowledge representations. Synthese, 197(4):1839–1840, April 2020. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
018-1772-1;   http://link.springer.com/content/pdf/10.1007/
s11229-018-1772-1.pdf. See [2025].

Cousins:2020:EJL

[2077] Robert D. Cousins. Erratum to: The Jeffreys–Lindley paradox and discovery criteria in high energy physics. Synthese, 197(8):3643, August 2020. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
015-0687-3;   http://link.springer.com/content/pdf/10.1007/
s11229-015-0687-3.pdf. See [1349].

Miller:2020:CHE

[2078] Mark Miller and Andy Clark. Correction to: Happily entangled: prediction, emotion, and the embodied mind. Synthese, 197(8):3645, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/

s11229-019-02260-3;   http://link.springer.com/content/pdf/
10.1007/s11229-019-02260-3.pdf. See [1697].

Fisher:2020:CLA

[2079] A. R. J. Fisher. Correction to: On Lewis against magic: a study of method in metaphysics. Synthese, 197(11):4743, November 2020. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-019-02323-5;   http://link.springer.com/content/pdf/
10.1007/s11229-019-02323-5.pdf. See [1687].

Feintzeig:2020:CDD

[2080] Benjamin H. Feintzeig. Correction to: Deduction and definability in infinite statistical systems. Synthese, 197(12):5539–5540, December 2020. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-018-01968-y;   http://link.springer.com/content/pdf/
10.1007/s11229-018-01968-y.pdf. See [1916].

Oldofredi:2021:SRM

[2081] Andrea Oldofredi. Some remarks on the mentalistic reformulation of the measurement problem: a reply to S. Gao. Synthese, 198(2):1217–1233, February 2021. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL http://link.springer.com/article/10.1007/
s11229-019-02101-3;   http://link.springer.com/content/pdf/
10.1007/s11229-019-02101-3.pdf. See [1847].

Pedersen:2021:CNR

[2082] Nikolaj Jang Lee Linding Pedersen. Correction to: Non-rational action in the face of disagreement: an argument against (strong) nonconformism. Synthese, 198(6):5951, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01964-2. See [1713].

Khudairi:2022:CGC

[2083] Hasen Khudairi. Correction to: Grounding, conceivability, and the mindbody problem. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03640-y. See [1619].

Finlay:2022:CWO

[2084] Stephen Finlay. Correction to: What ought probably means, and why you can’t detach it. Synthese, 200(3):??, June 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03605-1.

Baras:2019:HCI

[2085] Dan Baras. How close are impossible worlds? A critique of Brogaard and Salerno’s account of counterpossibles. Dialectica: International Review of Philosophy of Knowledge, 73(3):315–329, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).